A Complete Bibliography of Publications in Computer Networks (Amsterdam, Netherlands: 2000–2009)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

29 July 2017 Version 1.19

Title word cross-reference

1 [1183]. ^{2} [262, 313, 1277, 1320]. ^{3} [1986]. _{2} [1868]. _{ELLULAR} [732]. d [876, 1211]. H_{∞} [1335]. ∞ [1868]. K [1183, 1683, 1717, 1825, 1884]. ⇔ [87]. N [1868].

* [1832].

-based [1764]. -dominating [876]. -hop [876, 1211]. -Neigh [1825]. -target [1683].

1 [430].

2 [276, 277, 887, 888]. 2000 [155, 222, 569]. 2001 [509].

3.5G [1499]. 324M [304]. 3G [304, 1122, 1306, 1533, 1630, 1636, 1965]. 3G-324M [304]. 3G/WiMax [1965]. 3rd [1071].

4 [680]. 4G [1556].

802.11 [278, 547, 622, 1009, 1319, 1329, 1351, 1380, 1545, 1552, 1605, 1647, 1983]. 802.11-Based [622, 839]. 802.11-compliant [1351, 1545]. 802.11-operated [1319]. 802.11a [548]. 802.11b [1360, 1437, 1788]. 802.11b/g [1788]. 802.11e [844, 1116, 1244, 1353, 1378, 1379, 1404, 1419]. 802.11s [1381]. 802.15.4 [1437, 1685, 1711, 1713, 1998]. 802.16 [969, 1434, 1547]. 802.16-based [1358].

802.16d [1803]. 802.17 [1179].

ABR [118–120, 134, 190, 293]. ABridges [1581]. abstractable [1260]. abstraction [307]. abusive [681]. AC [1852]. academic [21, 164, 329]. accelerating [1510, 1678, 1739]. accelerator [336, 1823]. Access [35, 37, 53, 64, 102, 129, 322, 361, 371, 379, 391, 498, 520, 522, 606, 636, 668, 681, 686, 690, 797, 841, 863, 870, 872, 939, 953, 968, 1021, 1030, 1138, 1141, 1232, 1273, 1292, 1335, 1353, 1386, 1398, 1416, 1418, 1427, 1431, 1434, 1454, 1475, 1517, 1545, 1554, 1678, 1745, 1759, 1772, 1774, 1803, 1805, 1833, 1845, 1852, 1896, 1924, 1972]. access/cognitive [1141]. accessibility [87]. accountability [1225]. accountable [746]. accounting [1952]. accumulation [765]. accumulation-based [765]. accuracy [823, 860, 921, 1999]. Accurate [312, 574, 1552, 1707, 1843]. achieve [168]. Achieving [468, 472, 483, 1146, 1329, 1433, 1578]. Acknowledgement [207, 362, 541, 727, 864]. acknowledgment [1109]. acoustic [1714]. acquisition [1069]. across [446, 772, 842, 1423]. action [397, 1100]. Active [226– 229, 231, 232, 237, 240, 422, 613, 658, 662, 861, 1113, 1162, 1164, 1168, 1184, 1229, 1233, 1629]. activities [1620]. actor [1727, 1905, 1982]. Actuator [1251]. ad [202, 268, 281, 338, 361, 375, 431, 507, 511, 534, 537, 552, 622, 719, 759, 796, 813, 833, 876, 926, 939, 1063, 1090, 1111, 1127, 1205, 1211, 1214, 1244, 1283, 1285, 1287, 1304, 1326, 1380, 1384, 1390, 1395, 1424, 1452, 1459, 1475, 1482, 1522, 1535, 1538, 1549, 1550, 1562, 1567, 1573, 1587, 1602, 1624, 1627, 1642, 1645, 1653, 1705, 1718, 1766, 1883, 1899, 1901, 1912, 1951, 1953, 1963, 1968]. ad-hoc [507, 1127, 1535, 1602, 1705, 1718, 1951]. adaptation [351, 862, 937, 972, 1066, 1249, 1329, 1392, 1647]. Adapting [375]. Adaptive [15, 77, 148, 183, 191, 247, 271, 300, 342, 343, 358, 368, 481, 576, 603, 609, 612, 646, 662, 680, 702, 840, 901, 917, 927, 954, 972, 998, 1073, 1086, 1144, 1179, 1248, 1283, 1295, 1316, 1329, 1331, 1354, 1443, 1463, 1551, 1618, 1639, 1704, 1705, 1710, 1737, 1828, 1883, 1903, 1962, 1995, 2000]. Adaptive-level [15]. added [1484]. additive [869, 1206, 1823]. address [493, 715, 787, 1063, 1445, 1494, 1561]. Addressing [535, 1591, 1982]. adjusting [937]. adjustment [1651]. ADMIRE [77]. Admission [4, 114, 132, 193, 239, 309, 364, 376, 439, 520, 597, 619, 682, 701, 777, 897, 1111, 1276, 1402, 1434, 1446, 1470, 1518, 1621, 1729, 1763, 1795, 1868, 1909]. ADMs [1648]. ADSL [174, 701, 1759]. Advance [294, 1665, 1743]. advanced [212, 329, 1658]. Advances [143, 455, 798, 1368, 1371]. adversarial [1502]. advertisement [280]. advertisement-based [280]. advertising [100, 1154]. AF [1000]. AF-based [1000]. against [907, 1151, 1294, 1394, 1445, 1468, 2000]. Agent [46, 161, 162, 232, 290, 372, 610, 820, 1646, 1764]. Agent-based [46, 290, 372, 1646]. agents [45, 144, 816, 963, 1190, 1228, 1375, 1760]. aggregatable [1079]. aggregate [684, 920]. aggregated [1112, 1707]. aggregation [169, 285, 480, 593, 811, 1109, 1206, 1451, 1546, 1650, 1798, 1849, 1926, 1947]. aggressiveness [1468]. Agile [1023]. agreements [772, 825, 881]. AHP [264]. AID [1490]. aided [803]. AIMD [616, 849, 869, 943, 1119, 1456, 1503]. AIMD-based [616]. AIMD/RED [1503]. airtime [1690]. alarm [1194]. alarms [145]. alerts [1272]. algebraic [595]. Algorithm [78, 129, 192, 242, 298, 316, 405, 410, 427, 514, 643, 656, 657, 762, 834, 861, 898, 901, 944, 961, 1081, 1091, 1114, 1120, 1149, 1176, 1192, 1211, 1240, 1355, 1361, 1426, 1430, 1471, 1491, 1518, 1575, 1576, 1608, 1633, 1645, 1705, 1711, 1737, 1761, 1887, 1942, 1994]. Algorithmic [1104]. Algorithms [132, 141, 349, 356, 376, 424, 492, 503, 536, 579, 593, 618, 722, 803, 876, 963, 977, 984, 1019, 1032, 1076, 1146, 1181, 1203, 1295, 1325, 1336, 1369, 1422, 1476, 1618, 1662, 1732, 1754, 1779, 1784, 1798, 1848, 1849, 1955, 1960]. Aligned [880]. all-optical [32, 37, 185, 549, 1663, 1666, 1793, 1833]. All-Photonic [1023]. All-to-all [1634]. Alleviating [1758]. allocating [1537]. allocation [30, 120, 141, 275, 367, 441, 445, 448, 538, 612, 613, 662, 715, 745, 775, 781, 854, 959, 1002, 1084, 1146, 1171, 1204, 1402, 1404, 1408, 1434, 1558, 1651, 1679, 1690, 1763, 1795, 1892, 1907]. allocations [468, 483, 1224]. Aloha [760]. along [1231]. alpha [909]. alpha-beta [909]. alternate [305, 1486]. alternative [326]. alternatives [212]. altitude [1446]. always [1981]. always-on [1981]. ambients [14]. amplification [1675]. amplifier [1675]. AMRST [646]. Amsterdam [2002]. analogies [1564]. analogy [1919]. Analysing [757]. Analysis [5, 9– 11, 13, 33, 73, 74, 111, 123, 130, 145, 169, 186, 210, 238, 241, 269, 302, 313, 384, 393, 397, 403, 430, 444, 450, 482, 489, 540, 590, 656, 693, 697, 841, 849, 881, 882, 888, 904, 913, 914, 927, 943, 944, 964, 969, 974, 983, 1012, 1031, 1051, 1073, 1088, 1116, 1128, 1129, 1139, 1193, 1211, 1221, 1274, 1288, 1299, 1306, 1361, 1370, 1378, 1390, 1431, 1437, 1456, 1458, 1465, 1466, 1469, 1520, 1565, 1582, 1607, 1612, 1619, 1656, 1663, 1686, 1720, 1756, 1775, 1793, 1814, 1825, 1828, 1860, 1875, 1913, 1927, 1931, 1969, 1973, 1998, 2001]. analysis [246, 988, 1380, 1419, 1450, 1478, 1741, 1781, 1804, 1910]. analytic [963, 1353, 1869]. Analytical [317, 318, 463, 530, 725, 1081, 1216, 1397, 1449, 1524, 1801, 1943]. Analyzing [50, 79, 574, 850, 1155]. anchor [1332]. ANEMA [1809]. animation [474–478]. Annotation [61]. Annotation-based [61]. annotations [419, 1735]. Annotea [419]. Announced [522]. anomalies [1720]. Anomaly [149, 723, 1444, 1472, 1546, 1843].

Anomaly-based [149, 723]. anonymity [669, 1598, 1777]. anonymization [787]. anonymous [490, 649, 746]. ANQL [1168]. answers [1884]. ant [963]. ant-like [963]. antennas [1645, 1850]. anti [189, 1460, 1490]. anti-DoS [1490]. anti-jamming [189]. anti-phishing [1460]. Anticipatory [1393]. AntNet [1361]. ants [1507]. Anycasting [537, 1771]. Anycasting-based [537]. AoA [1999]. APDU [252]. API [181, 252, 685]. APOSN [1333]. appliance [90]. Application [10, 181, 186, 227, 418, 458, 475, 638, 640, 818, 900, 903, 915, 932, 949, 1002, 1192, 1281, 1423, 1425, 1436, 1578, 1673, 1736, 1782, 1838, 1868, 1907, 1914]. application-aware [1673]. application-layer [1782]. application-specific [1914]. applications [83, 255, 415, 424, 437, 438, 456, 491, 509, 520, 559, 586, 687, 724, 731, 834, 852, 867, 1053, 1058, 1078, 1101, 1161, 1224, 1257, 1284, 1298, 1310, 1468, 1504, 1520, 1527, 1601, 1615, 1628, 1651, 1702, 1760, 1773, 1835]. Applying [219, 371, 1072, 1248, 1774]. approach [24, 25, 43, 74, 100, 182, 218, 233, 240, 267, 326, 365, 372, 374, 401, 409, 460, 468, 483, 615, 619, 620, 625, 687, 699, 719, 721, 742, 783, 788, 814, 825, 903, 931, 963, 995, 1000, 1022, 1068, 1072, 1102, 1106, 1150, 1170, 1178, 1212, 1247, 1260, 1314, 1335, 1356, 1408, 1434, 1476, 1483, 1500, 1530, 1533, 1630, 1663, 1725, 1742, 1752, 1771, 1773, 1776, 1799, 1836, 1892, 1907, 1937]. approaches [63, 1407, 1662, 1696, 1748, 1929]. Approaching [1631]. appropriate [1344]. Approximate [1612, 1875]. approximating [1415]. approximation [117, 744, 859]. APSD [1379]. AQM [1144, 1335, 1456, 1515]. AQoSM [1010]. arbitrary [117, 536]. architectural [527, 570]. Architecture [12, 35, 76, 150, 194, 195, 206, 211, 248, 315, 336, 347, 406, 415, 435, 526, 529, 665, 676, 796, 835, 845, 872, 891, 922, 973, 999, 1078, 1273, 1299, 1327, 1373, 1486, 1514, 1519, 1561, 1659, 1809, 1818, 1839, 1896, 1933, 1935].

Architectures [156, 212, 311, 357, 436, 530, 687, 782, 1130, 1288, 1660, 1673]. archive [94]. Area [35, 133, 200, 547, 649, 761, 938, 969, 1328, 1342, 1363, 1389, 1542, 1583, 1994]. areas [32, 341]. ARQ [1498, 1590]. ARQ-based [1590]. arrayed [516]. arrayed-waveguide [516]. arrival [540, 575]. arrivals [456, 675]. art [608, 694, 1502, 1990]. artifacts [201, 1603]. artificial [114]. ARTP [1330]. AS-level [700, 1735]. ASON [1247]. ASON/GMPLS [1247]. Aspect [515]. Aspect-oriented [515]. aspects [156]. Assessing [1700]. Assessment [660, 932, 1542, 1615]. assigning [618]. assignment [31, 836, 1005, 1149, 1642, 1662, 1692, 1848, 1912, 1923, 1958]. assignments [335, 1937]. assisted [630, 851, 1846]. associated [1203]. Association [1495, 1780]. Associative [1348]. assurance [1000]. Assured [517]. asymmetric [357, 1443]. asymmetry [1395]. asymptotically [1919]. Asynchronous [1107, 1171, 1246, 1766]. ATM [2, 4, 11, 18, 19, 34, 110, 112, 114– 116, 119, 122, 124, 125, 128, 132, 137, 138, 174, 188, 190, 267, 369, 385, 408, 1284]. ATM-based [385]. attack [1151, 1294, 1384, 1462]. attacks [694, 802, 902, 930, 931, 934, 1200, 1394, 1403, 1450, 1540, 1599, 1600, 1965, 1997, 2000]. attribute [1773]. attributes [75]. auction [778, 1537]. auctions [1596]. audio [172, 1714]. audio-sensor [1714]. audio-visual [172]. audit [1472]. augmented [136, 201]. authenticated [1685]. Authentication [254, 669, 721, 904, 980, 1042, 1191, 1287, 1371, 1372, 1646, 1695]. Author [224, 272, 330, 380, 433, 484, 486, 542, 599, 652, 707, 768, 827, 884, 945]. authority [320, 371, 1783]. auto [1593]. auto-configuration [1593]. autoconfiguration [1063]. autocorrelated [1617]. autocorrelation [457]. automata [742]. Automated [51, 647, 1247, 1316].

Automatic [101, 159, 265, 491, 1317].

Automating [82]. automation [1062, 1978]. automotive [1867]. Autonomic [1398, 1725, 1809, 1894, 1896, 1898, 1900]. autonomization [1895]. Autonomous [144, 736, 820, 1362, 1738]. availability [326, 1154, 1747, 1947]. Availability-aware [1947]. available [1028, 1054, 1129, 1594, 1617, 1698, 1846, 1966]. available-bandwidth [1966]. AVCOL [1947]. avoidance [616, 1822, 1878, 1924]. Avoiding [394, 610]. AVoN [84]. aware [268, 438, 450, 511, 551, 553, 642, 809, 826, 831, 841, 923, 970, 974, 1047, 1090, 1152, 1184, 1196, 1314, 1330, 1355, 1423, 1434, 1480, 1549, 1595, 1641, 1662, 1665, 1673, 1682, 1690, 1747, 1772, 1779, 1848, 1859, 1889, 1947, 1958, 1987]. AXL [84].

back [198]. backbone

[443, 908, 1194, 1400, 1562, 1970].

background [26, 1340]. backoff [278, 1329]. backup [1026, 1879, 1886]. backward [1924]. backward-compatible [1924]. backwards [90]. balance [1631]. balancing [122, 450, 853, 1074, 1305, 1387, 1557, 1652, 1784]. balking [1873]. band [1015, 1416]. Bandwidth [102, 118, 141, 310, 312, 364, 368, 383, 389, 390, 471, 483, 538, 550, 579, 596, 606, 662, 680, 705, 735, 774, 775, 784, 797, 804, 811, 832, 910, 920, 959, 1000, 1002, 1021, 1028, 1047, 1054, 1084, 1111, 1128, 1129, 1138, 1146, 1195, 1204, 1268, 1284, 1305, 1337, 1345, 1392, 1405, 1408, 1434, 1440, 1467, 1537, 1558, 1565, 1572, 1617, 1640, 1645, 1678, 1698, 1763, 1833, 1846, 1892, 1962, 1966]. Bandwidth-delay [579, 1345]. bandwidth-guaranteed [784, 1645]. bandwidth-limited [389]. bandwidth-on-demand [383]. Banyan [11]. Bargaining [1204]. base [511, 859, 1547]. Based [10, 17, 25, 46, 61, 114, 172, 194, 204, 217, 260, 266, 280, 290, 313, 323, 344, 372, 429, 447, 515, 517, 519, 537, 662, 666, 671, 674, 682, 697, 723, 765, 787, 788, 823, 826, 839, 862, 871, 874, 877, 898, 903, 969, 991, 1000, 1030, 1032, 1129, 1130, 1133, 1137, 1144, 1199, 1207, 1211, 1232, 1243, 1302, 1342, 1343, 1374, 1384, 1394, 1404, 1449, 1459, 1509, 1515, 1529, 1561, 1575, 1583, 1591, 1598, 1607, 1621, 1632, 1645, 1649, 1653, 1664, 1665, 1695, 1698, 1705, 1769, 1778, 1818, 1871, 1878, 1905, 1929, 1933, 1936, 1943, 1945, 1953, 1962, 1963, 1978]. based [2, 119, 126, 129, 137, 149, 184, 213, 233, 267, 311, 374, 385, 389, 402, 405, 427, 428, 476, 509, 550, 616, 622, 645, 660, 664, 681, 766, 804, 845, 879, 880, 892, 904, 924, 925, 931, 934, 971, 1041, 1050, 1073, 1084, 1091, 1092, 1105, 1148, 1164, 1166, 1244, 1260, 1285, 1298, 1303, 1312, 1313, 1323, 1346, 1358, 1375, 1442, 1470, 1476, 1494, 1507, 1508, 1564, 1572, 1574, 1600, 1646, 1647, 1668, 1670, 1710, 1725, 1728, 1750, 1760, 1764, 1765, 1768, 1820, 1833, 1836, 1841, 1843, 1861, 1863, 1883, 1886, 1903, 1909, 1911, 1922, 1923, 1941, 1942, 1951, 1959, 1994]. based [309, 364, 477, 642, 820, 860, 1054, 1590, 1753, 1847, 1924, 1934]. Basestation [508]. Basic [990, 1770]. Batch [301]. battlefield [152]. battlefield-intelligence [152]. Bayesian [1663]. be [1988]. Beacon [1430]. bearer [279]. behavior [70, 104, 677, 830, 908, 1775, 1872, 1947]. behaviors [501, 1101]. behaviour [113, 580]. being [306]. benchmark [23]. benchmarking [525]. Benefits [1154]. best [19, 673, 1564, 1827]. best-effort [19, 673, 1827]. beta [909]. between [98, 104, 299, 365, 390, 408, 547, 568, 621, 718, 802, 1038, 1265, 1723, 1749, 1829]. Beyond [43, 212, 283, 1213, 1242, 1527]. BGP [714, 736, 899, 960, 1134, 1200, 1208, 1649, 1938]. BGP-RCN [899]. bi [69, 636, 1614]. bi-directional [636, 1614]. Bi-Layer [69]. bias [463]. biased [1235]. big [1721]. billing [243, 919, 1646]. Binary [5, 1269, 1699]. Binding [289]. bio [1900]. bio-inspired [1900]. biologically [1514]. biologically-inspired [1514]. BiSNET [1514]. bit [344]. BitTorrent [1817]. Black [1989]. Black-box [1989]. bloated [1760]. block [1491]. blocking [581, 612, 733, 995, 1415]. Blogistan [1046]. Bloom [1637]. BLT [69]. Bluetooth [508, 509, 547, 719, 720, 1360, 1480, 1726]. BMAPs [675]. board [383, 385, 454, 465, 473, 479, 488, 497, 506, 513, 523, 532, 544, 555, 564, 573, 582, 592, 601, 607, 617, 627, 635, 644, 654, 661, 670, 678, 689, 698, 709, 716, 728, 739, 750, 758, 770, 780, 789, 799, 807, 819, 829, 837, 848, 856, 865, 875, 886, 894, 905, 916, 928, 936, 947, 956, 965, 975, 985, 993, 1013, 1024, 1034, 1043, 1052, 1061, 1070, 1082, 1093, 1110, 1125, 1140, 1157, 1169, 1188, 1202, 1218, 1254, 1266, 1289, 1307, 1318, 1338, 1347, 1367, 1385, 1412, 1435]. Board [1457, 1473, 1488, 1505, 1521, 1532, 1560, 1570, 1588, 1597, 1610, 1625, 1635, 1644, 1657, 1676, 1687, 1701, 1716, 1730, 1746, 1757, 1767, 1785, 1796, 1810, 1821, 1834, 1844, 1864, 1876, 1893, 1908, 1921, 1939, 1949, 1964, 1975, 1984, 1991]. Bolch [1865]. Bone [235]. Bookmarks [1511]. both [306]. bottleneck [471, 535, 1354, 1800, 1804]. bottlenecks [1306]. bound [855, 941, 1006]. Bounded [18, 355, 1925]. bounds [313, 480, 803, 822, 1067, 1108, 1408]. box [248, 1989]. Brahms [1946]. Branch [1003]. Braving [534]. Bridge [320, 1386]. Bringing [63, 1938]. Broadband [9, 129, 183, 282, 298, 382, 386, 387, 606, 795, 953, 970, 1377, 1418, 1431, 1434]. Broadcast [13, 189, 534, 1296, 1555]. broadcasting [1553, 1559, 1765, 1891]. Broker [1641]. Broker-placement [1641]. brokers [310, 538]. browsing [94, 103]. BSD [1210]. bucket [2]. buffer [126, 130, 271, 429, 441, 462, 577, 775, 782, 1146, 1230, 1330, 1352, 1366, 1456, 1929, 1961, 1983]. buffer-aware [1330]. buffer-based [126]. buffer-less [1929]. buffered [117, 781, 1148, 1491]. buffering [1941]. bufferless [1885]. buffers [188]. Building [350, 416, 1162, 1604]. bulk [1601, 1609, 1977]. bullet [950]. bundles [598, 1118]. Burst [111, 376, 459, 492, 498, 692, 766, 853, 940, 1020, 1091, 1221, 1267, 1270, 1301, 1569, 1671, 1673, 1798]. Burst-Contention [1569]. burst-level [111]. burstiness [1828]. Bursty [597, 1468, 1925]. bus [764]. Business [71, 284, 292, 432, 589, 825]. business-oriented [825]. business-to-business [589]. byte [562]. bytecode [1974]. Byzantine [1042, 1946].

C [732]. cable [606, 1759]. CAC [115, 1216]. cache [323, 613, 651, 665, 1561, 1745]. cache-based [1561]. caches [15, 377, 378, 933]. Caching [91– 93, 535, 559, 613, 667, 880, 1413, 1458, 1599]. caching/prefetching [1458]. calculus [14, 941, 1072, 1133, 1778, 1878, 1925]. calculus-based [1778]. Call [132, 196, 210, 257, 364, 365, 556, 563, 591, 619, 733, 1209, 1263, 1265, 1297, 1446]. calling [84]. calls [1700]. campus [326, 1581, 1706, 1722, 1816, 1969]. campus-wide [326, 1722]. Can [1341, 1867]. candidate [659, 1030]. capabilities [891]. capability [1463]. capacities [428]. Capacity [275, 445, 510, 634, 854, 915, 1016, 1069, 1075, 1306, 1327, 1431, 1539, 1548, 1728, 1851]. capture [81]. capturing [700]. CAPWAP [1794]. CARD [251, 256, 980, 1030, 1370, 1374, 1375]. card-based [1375]. cards [249, 250, 253–255, 1368, 1369, 1371, 1376]. carrier [1393, 1647]. carrier-grade [1393]. Cartesian [139]. Casals [729, 730, 740]. Case [161, 209, 307, 388, 756, 763, 907, 989, 1320, 1425, 1454, 1612]. cases [221, 1248, 1288]. catalogs [43]. Categorizing [1036]. category [1236]. category-to-wavelength [1236]. cause [899, 1656]. caused [552]. CCSDS [1478]. CD [184]. CDMA [376, 556, 597, 826, 961, 1069, 1446, 1674, 1882, 1934]. CDMA-based [1934]. CDMA2000 [786]. CeBIT [509]. Cell [112, 311, 428, 823, 1417, 1583, 1724, 1745]. cell-based [311, 823]. cell-by-cell [1583]. cells [18, 618]. Cellular [345, 364, 447, 581, 611, 742, 761, 805, 826, 835, 838, 840, 846, 942, 972, 998, 1051, 1271, 1282, 1297, 1327, 1328, 1446, 1453, 1533, 1613, 1646, 1654, 1764, 1851]. center [1990]. centers [1789, 1985, 1987]. centralised [281]. centralized [1855, 1993]. centric [83, 1401, 1772, 1819]. Centroid [1902]. Certificate [1750]. Certificate-based [1750]. certificates [746, 1122]. Certification [320, 371, 1783]. chains [73]. challenge [904]. challenge/response [904]. Challenges [608, 753, 1513, 1592, 1658, 1990]. change [1102, 1966, 1978]. Changing [8, 912, 1722]. Channel [334, 342, 417, 612, 649, 745, 794, 812, 836, 873, 927, 961, 998, 1069, 1112, 1282, 1331, 1353, 1404, 1558, 1586, 1591, 1642, 1845, 1891, 1937, 1958, 1977]. channels [262, 846, 1578, 1829, 1936]. Characteristics [1008, 1048, 1501, 1816, 1847]. characterization [675, 906, 941, 997, 1083, 1203, 1238, 1693]. Characterizing [714, 1837]. charge [260]. charging [776]. charts [990]. checking [1089]. Chilean [1097]. chip [1855]. choice [1544]. Chord [1277]. churn [1197]. CINSim [1874]. circuit [1181]. CISS [1298]. CitiTime [214]. class [115, 188, 193, 252, 295, 367, 405, 472, 519, 705, 1280, 1331, 1392, 1871, 1980]. class-level [193]. classes [577, 774]. Classification [694, 810, 1092, 1121, 1219, 1256, 1511, 1536, 1714, 1759, 1835, 1836, 1838–1841, 1863, 1956]. classifier [1843]. classifiers [594]. classroom [163]. click [1098]. click-sequences [1098]. client [1458]. client-side [1458]. Clock [762, 1124]. clocking [1467]. Clos [695]. Clos-network [695]. close [605]. closed [765, 1552]. closed-form [1552]. closed-loop [765]. clothing [200]. cluster [642, 994, 1287, 1303, 1386, 1430]. cluster-based [642, 1303]. cluster-tree [994]. clustered [655]. Clustering [762, 1211, 1298, 1576, 1653, 1758, 1779]. clusters [328, 450]. CNPGSDP [1199]. co [1313]. co-processor-based [1313]. Coalescing [1967]. Coarse [712]. Coarse-grain [712]. coax [246]. CoCONet [1670]. code [59, 1160, 1692]. codes [373, 818, 1674]. coding [1478, 1704, 1931, 2000]. CODIS [893]. Coexistence [574, 690, 1360]. Cognitive [321, 1141, 1589–1591, 1593– 1595, 1679, 1866, 1880]. cognizant [1334]. collaboration [508, 1285]. collaborative [91, 95, 195, 1511, 1755]. collapse [394]. collecting [79]. collection [1858]. collision [184, 1025, 1670, 1924]. collision-free [1670]. collusion [1648]. color [1575]. color-theory-based [1575]. Colored [1272, 1364, 1409]. Combating [1760]. combination [26, 1329]. Combinatorial [504]. combined [504, 1206]. combining [530]. commerce [40, 41, 255, 289, 292, 589, 746, 1504]. commercial [46]. commitment [344]. commodity [261]. common [964]. communicating [269]. Communication [3, 5, 20, 29, 134, 172, 195, 203, 262, 337, 355, 373, 423, 476, 554, 595, 614, 647, 655, 657, 669, 720, 749, 764, 792, 795, 1062, 1108, 1118, 1235, 1240, 1259, 1328, 1543, 1583, 1591, 1689, 1706, 1726, 1731, 1808, 1867, 1880, 1898, 1925, 1976]. communications [194, 200, 259, 283, 511, 649, 808, 815, 816, 818, 843, 870, 1417, 1420, 1448, 1506, 1510, 1805]. communities [1814, 1817]. community [80, 86, 94, 164, 173, 1811, 1815, 1818]. community-oriented [86]. Comparative [302, 444, 778]. Comparing [779, 1096]. Comparison [116, 280, 299, 370, 562, 633, 686, 1028, 1073, 1083, 1332, 1442]. comparisons [313]. compatible [184, 1924]. compensation [1751]. competing [1119]. competition [1917]. Competitive [554, 610]. compiler [244]. complete [1935]. Complex [1731]. complexity [1114, 1124, 1715, 1737, 1938]. compliance [313]. compliant [844, 1351, 1545]. Component [148, 756, 1164]. component-based [1164]. components [179, 568, 992]. composable [1147]. composing [1742]. composition [588, 1163, 1899]. compound [456]. comprehensive [1812, 1926, 1928]. compressed [958]. compression [1175, 1715]. computation [444, 910, 1240, 1432, 1879, 1886]. computational [1114]. computations [1231]. Computer [8, 201, 453, 466, 638, 1450, 1589, 1658, 1731, 1865, 2001]. computer-augmented [201]. Computing [105, 198, 204, 205, 237, 264, 469, 687, 948, 995, 1423, 1548, 1761]. concast [1166]. concast-based [1166]. concept [277, 366, 1045]. Concepts [21, 75, 159]. concerns [149]. condition [397, 1100]. conditioner [879]. conditioning [374]. conditions [342]. conducting [1732]. Conference [319, 2002]. config [1992]. Configurable [673, 821, 1629]. configuration [604, 631, 685, 817, 972, 1116, 1226, 1559, 1593, 1794, 1809, 1909]. configurations [800, 1393]. Configuring [505, 1581, 1624]. conflict [594, 1121, 1732]. conflict-free [1121, 1732]. conflicts [1263]. Conformance [881, 1139]. congested [1208]. Congestion [9, 10, 19, 183, 187, 265, 293, 363, 394, 430, 467, 499, 501, 502, 580, 616, 648, 697, 763, 778, 896, 944, 964, 1033, 1064, 1172, 1270, 1321, 1335, 1346, 1401, 1421, 1463, 1497, 1616, 1629, 1632, 1740, 1778, 1806, 1827, 1854, 1878, 1911, 1958, 1962]. Congestion-aware [1958]. Conjugate [1072]. Connected [876, 1683, 1727]. Connection [4, 114, 239, 456, 559, 650, 1078, 1236, 1585]. connection-oriented [1078, 1236]. connections [318, 357, 663, 693, 964, 1552, 1743]. connectivity [1551]. connectors [568]. conservation [1690]. conserve [1912]. considerations [31, 93, 527, 801, 1718]. considered [496]. Considering [691, 1583, 1671]. Consistency [142, 567, 1060]. Consistent [1356]. constrained [579, 786, 1284, 1432, 1633, 1900]. constraint [413, 804]. constraints [452, 895, 1076, 1227, 1637, 1779, 1897, 1980]. construct [623]. Constructing [352, 451]. construction [1177, 1409, 1945]. consumption [552, 720, 1573, 1685, 1987]. container [1670]. container-based [1670]. containment [1223, 1311, 1316]. Content [61, 378, 448, 598, 645, 854, 893, 978, 1001, 1066, 1252, 1323, 1403, 1427, 1507, 1586, 1784, 1811, 1817, 1841, 1887, 1971]. content-based [1323, 1507]. content-distribution [1887]. content-sharing [1817]. contention [1569, 1710, 1924, 1930]. Contents [712, 986]. contest [23, 28]. Context [71, 81, 210, 402, 1012, 1184, 1772]. Context-aware [1184, 1772]. Context-based [402]. continuity [895]. Continuous [83, 306, 738, 1064, 1891]. Contour [1182]. Contract [286]. Contract-driven [286]. contracts [1180]. contractual [589]. Control [4, 10, 16, 53, 114, 119, 121, 132, 141, 167, 190, 193, 240, 265, 293, 298, 300, 365, 369, 376, 430, 431, 439, 483, 499, 510, 520, 553, 571, 616, 619, 641, 648, 662, 681, 682, 684, 685, 697, 755, 763, 777, 786, 896, 939, 944, 1000, 1033, 1064, 1120, 1138, 1150, 1172, 1232, 1265, 1270, 1276, 1292, 1320, 1321, 1330, 1398, 1401, 1402, 1416, 1421, 1427, 1429, 1434, 1446, 1463, 1470, 1489, 1504, 1518, 1525, 1572, 1573, 1577, 1591, 1621, 1629, 1632, 1645, 1672, 1682, 1729, 1740, 1778, 1795, 1806, 1825, 1832, 1841, 1852, 1861, 1909, 1953, 1962, 1977, 1993, 1996]. control [19, 115, 129, 130, 183, 239, 267, 302, 309, 361, 363, 364, 467, 501, 502, 504, 548, 554, 597, 636, 656, 701, 897, 983, 1111, 1230, 1263, 1273, 1324, 1335, 1346, 1356, 1454, 1497, 1533, 1616, 1763, 1805, 1854, 1882, 1911, 1924, 1980]. control-theoretical [483]. controllability [1829]. controlled [85, 1057, 1083, 1530, 1727, 1780, 1840, 1916]. controller [2, 6, 964, 1144, 1270, 1827, 1910]. controllers [4]. Controlling [738, 1396]. controls [748, 767, 1414]. convenient [102]. conventional [1159]. convergecast [1580]. converged [213, 845, 1530]. convergence [208, 899, 1031, 1555, 1608, 1993]. conversations [1705]. conversion [171, 1666, 1830]. conversions [995]. converter [38, 87]. converters [185, 1227, 1274]. convex [1806]. convolution [1455]. cooperation [1749]. Cooperative [967, 978, 1039, 1434, 1586, 1613, 1802]. Coordinate [814]. Coordinated [1296]. coordinates [1902]. coordination [756, 1619]. CORBA [140, 1895]. core [483, 1554, 1670, 1733]. core-stateless [483]. CORNet [1672]. Corral [1160]. correcting [247]. correction [1267]. Correctness [190]. Correlation [366, 540, 731, 1194, 1563]. correlations [741, 1008, 1612]. corresponding [1520]. Corrigendum [2001]. Cost [102, 117, 123, 355, 676, 976, 1005, 1045, 1170, 1176, 1190, 1248, 1252, 1323, 1469, 1761]. Cost-effective [117, 676, 1170, 1469]. cost-efficient [1252]. cost-minimization [1176]. Cost-optimal [1045]. costs [135, 655]. Counteracting [1584]. counting [805]. Coupon [306]. Coverage [570, 859, 1362, 1683, 1699, 1706, 1727, 1853, 1861, 1942]. covering [581]. CPU [1628]. crawling [95]. CREAM [585]. CREAting [585]. creation [214, 286]. credit [129]. credit-based [129]. criteria [116]. critical [496, 992]. Cross [60, 332, 890, 957, 1018, 1498, 1590, 1611, 1704, 1840, 1857]. Cross-layer [890, 957, 1498, 1590, 1611, 1704, 1857]. cross-media [60]. cross-path [332]. cross-producting [1840]. Cross-sharing [1018]. crossbar [781, 1148]. crossings [1650]. crowds [1177]. Cryptanalysis [649]. Cryptographic [254, 1203, 1369, 1535]. Cryptography [253, 787, 821, 1454]. cryptography-based [787]. CSMA [184]. CSMA/CD [184]. cumulative [92]. cumulus [603, 882]. CuPIDS [1313]. Current [249]. curve [643]. customers [1868, 1873]. Customizable [236]. Customized [89]. cycle [1634, 1710]. cycles [216].

DAGs [958]. danger [1315]. DARPA [146]. Data [29, 77, 85, 90, 92, 140, 145, 191, 208, 246, 357, 376, 421, 430, 593, 634, 797, 846, 933, 968, 1051, 1074, 1085, 1099, 1215, 1222, 1237, 1271, 1315, 1342, 1472, 1495, 1510, 1538, 1574, 1762, 1789, 1819, 1849, 1853, 1858, 1891, 1923, 1926, 1941, 1985, 1987, 1990]. data-centric [1819]. Data-gathering [634]. data-rate [1237]. data-transmission [1853]. data-transporting [1510]. database [64, 83]. database-centric [83]. databases [761, 1636]. dataflow [1102]. DCCP [1751]. DDoS [694, 1540]. de-allocation [745]. Deadline [897, 1574]. Deadline-based [1574]. DEAPspace [202]. decentralised [812]. Decentralized [1112, 1195, 1572, 1650, 1913, 2001]. decision [5, 815, 1663, 1914]. decomposition [995, 1871]. decrease [869, 1806, 1962]. Decreasing [1938]. decrement [407]. deep [1478]. defence [1314]. Defending [1151, 1445]. defense [694, 1315]. defenses [1599]. deferred [850]. Deficit [192]. defined [1708]. deflection [1019]. Delay [313, 334, 355, 391, 480, 482, 579, 640, 797, 834, 877, 941, 943, 1006, 1031, 1153, 1171, 1345, 1356, 1400, 1428, 1449, 1465, 1477, 1601, 1878, 1930, 1980]. Delay-based [877, 1449, 1878]. delay-bounded [355]. delay-sensitive [834]. delay-throughput [334]. delays [785, 915, 1503]. delegation [180]. delivering [1230]. Delivery [598, 893, 978, 1001, 1056, 1396, 1478, 1574]. Demand [375, 383, 676, 804, 836, 847, 1106, 1231, 1326, 1447, 1496, 1812, 1817, 1927]. demands [832, 1426]. Demographically [609]. demonstration [999]. denial [150, 902, 934, 974, 1089, 1314, 1403, 1450]. denial-of-service [150, 902]. dense [1573, 1733, 1942]. dense-core [1733]. density [464, 1647]. departure [1798]. dependable [268]. dependence [463, 906, 911, 912]. dependencies [1885]. dependent [430, 455, 461, 961, 1051, 1079]. deployable [468]. deployed [1853]. Deployment [235, 438, 1164, 1363, 1681, 1831, 1957]. deployments [1724]. derivation [20, 647]. Deriving [1250]. Describing [568, 1359]. Description [638, 1050, 1734]. descriptions [97]. Design [11, 44, 49, 53, 201, 218, 282, 287, 309, 340, 360, 418, 423, 476, 504, 520, 529, 557, 565, 570, 576, 613, 625, 639, 656, 677, 737, 879, 888, 938, 960, 992, 1023, 1045, 1063, 1076, 1131, 1175, 1178, 1197, 1241, 1295, 1300, 1335, 1358, 1360, 1378, 1389, 1591, 1622, 1643, 1674, 1704, 1713, 1753, 1783, 1910, 1935, 1978]. designed [1190]. designer [104]. Designing [58, 741, 1504, 1776, 1960]. desynchronized [1491]. detect [1192]. Detecting [27, 754, 1151, 1782, 1997]. Detection [23–26, 28, 143, 144, 146– 150, 152, 182, 280, 470, 594, 723, 810, 901, 902, 931, 1085, 1102, 1259, 1309, 1311, 1322, 1444, 1462, 1466, 1472, 1540, 1603, 1650, 1799, 1800, 1828, 1841, 1843, 1917, 1955, 1965, 1966]. Detector [1637]. detectors [1312]. deterioration [1139]. Determining [800, 1344]. deterministic [332, 1399, 1867]. developed [262]. Developing [426, 664, 992]. Development [217, 251, 266, 415, 595, 991, 1095, 1880]. developments [1369]. device [1771]. devices [63, 202, 977, 1516]. DFA [1310]. DHA [1764]. DHT [1298, 1768, 1820]. DHT-based [1298, 1768]. Diagnosis [749, 1085, 1182, 1306]. diagnostic [913]. diagonal [1092]. diagonal-based [1092]. diagrams [5, 223, 1261]. dialup [1759]. diameter [1604, 1736]. Did [168]. Diff [1010, 1827]. Diff-Serv [1010, 1827]. different [1683]. differential [1153, 1774]. Differentiated [12, 170, 305, 440, 441, 470, 472, 499, 682– 684, 726, 782, 879, 922, 1001, 1029, 1068, 1280, 1302, 1305, 1470, 1669, 1674, 1699, 1773, 1826]. differentiated-services [1470]. differentiation [437, 704, 1210, 1356, 1396, 1449, 1647, 1911]. Diffie [1685]. diffserv [191, 374, 551, 604, 685, 699, 981, 1000, 1402, 1669]. Diffusion [1861]. Diffusion-based [1861]. digital [41, 165, 301, 1442]. dimensional [1121, 1461, 1509, 1955]. dimensioning [38, 1023, 1281, 1801]. direct [720]. directed [302, 1485]. direction [1243, 1566]. directional [636, 1614, 1850]. Directions [249, 751]. directory [371]. discard [917]. discipline [405, 662, 941, 1774, 1915]. disciplines [263]. Discovering [104]. discovery [407, 1030, 1107, 1199, 1205, 1246, 1486, 1508, 1542, 1677, 1927]. discrete [36, 540]. discrete-time [36, 540]. Disjoint [895, 1364, 1409, 1634]. Dispatching [695]. dispersal [270]. disruption [1899]. dissemination [46, 788, 1495, 1586, 1613, 1941, 1943]. distance [500, 1243, 1339, 1346, 1461, 1697, 1968]. Distinguishing [802]. distortion [1207]. Distributed [155, 161, 163, 179, 205, 244, 264, 351, 388, 432, 500, 526, 611, 633, 665, 666, 767, 833, 872, 876, 898, 902, 923, 959, 996, 1025, 1104, 1120, 1135, 1150, 1151, 1154, 1201, 1257, 1265, 1315, 1353, 1373, 1379, 1393, 1403, 1409, 1413, 1483, 1486, 1496, 1546, 1550, 1580, 1593, 1595, 1618, 1636, 1662, 1714, 1803, 1822, 1832, 1884, 1932, 1947, 1982, 1995]. Distribution [44, 51, 312, 323, 427, 645, 854, 855, 920, 955, 976, 997, 1053, 1077, 1127, 1268, 1288, 1403, 1461, 1484, 1493, 1523, 1784, 1811, 1833, 1887]. Distribution-based [1833]. distributions [1117, 1455]. diverse [1696]. diversity [1605]. division [335]. DMTP [1396]. DNS [535, 645, 1115]. DNS-based [645]. do [97, 324]. DOC.COM [288]. document [56, 409, 490]. documents [53, 1102]. domain [1, 440, 688, 883, 984, 1115, 1180, 1196, 1278, 1518, 1591, 1781, 1993]. domains [1222, 1423, 1820]. dominating [876]. don’t [1357]. doped [1675]. DoS-resistant [1087]. Double [391, 1475, 1684]. Downlink [334, 968, 1774]. download [1887]. Downloading [949]. Dr [1865]. driven [218, 286, 477, 529, 786, 809, 909, 1608, 1771]. drivers [436]. DropTail [1341]. DS [556]. DS-CDMA [556]. DSA [301]. DSA-type [301]. DSM [1098]. DSM-PLW [1098]. DSMCast [699]. DTNs [1819]. DTSMA [1803]. Dual [405, 1072, 1764]. dual-class [405]. duration [664]. durations [1127]. during [509]. dust [817]. duty [1710]. DVB [891, 892, 999, 1704]. DVB-RCS [999, 1704]. DVB-S [891]. DVB-T [892]. Dynamic [2, 19, 30, 45, 65, 87, 229, 235, 278, 279, 285, 348, 372, 429, 503, 522, 604, 620, 685, 712, 726, 745, 748, 783, 813, 824, 826, 853, 895, 920, 923, 1002, 1020, 1054, 1060, 1065, 1067, 1099, 1113, 1141, 1177, 1209, 1212, 1240, 1286, 1354, 1362, 1406, 1408, 1546, 1559, 1567, 1569, 1596, 1628, 1661, 1686, 1710, 1734]. dynamical [619]. Dynamically [402, 673, 998, 1160, 1433]. Dynamics [17, 241, 395, 400, 1094, 1095, 1097, 1103, 1238, 1585, 1888].

e-business [432]. e-commerce

[255, 589, 746, 1504]. E-LOTOS [307].

e-mail [1841]. E-markets [287]. E-Model [1174]. earliest [897]. early [470]. Earth [790–792, 795, 797, 798, 1888]. Earth-fixed [1888]. earth-science [797]. EASES [1655]. easily [451]. ECC [1685]. Ecological [625]. economic [779, 1663, 1907]. economics [258, 771, 1654]. EDCA [1116, 1419]. Edge [309, 666, 830, 924]. Edge-based [309, 924]. Edge-to-edge [666]. edged [237]. editing [171]. Editorial [7, 108, 109, 197, 319, 353, 454, 465, 473, 479, 533, 592, 617, 627, 635, 644, 654, 661, 670, 678, 689, 698, 709, 716, 717, 728, 739, 750, 758, 770, 780, 789, 799, 807, 819, 829, 837, 848, 856, 857, 865, 875, 886, 894, 905, 916, 928, 936, 947, 956, 965, 975, 985, 993, 1013, 1024, 1034, 1043, 1044, 1052, 1061, 1070, 1082, 1093, 1110, 1125, 1140, 1157, 1169, 1188, 1202, 1218, 1254, 1266, 1289, 1290, 1307, 1318, 1338, 1347, 1367, 1385, 1412, 1435, 1457, 1473, 1488, 1571, 1797]. Editorial [160, 274, 284, 308, 382, 395, 411, 435, 455, 474, 488, 497, 506, 513, 523, 532, 544, 555, 564, 573, 582, 601, 607, 966, 1158, 1255, 1505, 1521, 1532, 1541, 1553, 1560, 1570, 1588, 1597, 1610, 1625, 1635, 1644, 1657, 1676, 1687, 1701, 1716, 1730, 1746, 1757, 1767, 1785, 1786, 1796, 1810, 1821, 1834, 1844, 1864, 1876, 1893, 1894, 1908, 1921, 1939, 1940, 1949, 1964, 1975, 1984, 1991]. Editorial-autonomic [1894]. Editorials [1865]. education [163]. EEMC [1576]. EFCM [964]. effect [74, 463, 775, 950, 952, 1671]. Effective [16, 117, 288, 421, 676, 917, 1170, 1469]. Effectiveness [335, 892]. effects [1077, 1724, 1788, 1983]. efficiency [616, 942, 1132, 1365, 1523, 1565]. Efficient [51, 78, 99, 191, 369, 424, 453, 490, 514, 548, 576, 593, 603, 624, 632, 657, 668, 673, 711, 722, 726, 744, 761, 800, 801, 834, 926, 977, 980, 1002, 1021, 1058, 1102, 1134, 1195, 1199, 1219, 1229, 1231, 1233, 1240, 1252, 1283, 1286, 1298, 1303, 1326, 1387, 1407, 1440, 1443, 1477, 1533, 1540, 1542, 1549, 1562, 1575, 1576, 1593, 1614, 1642, 1655, 1715, 1753, 1765, 1780, 1831, 1838, 1882, 1892, 1897, 1916, 1981]. efficiently [1209]. effort [19, 673, 1827]. EFSM [1824]. eHIP [1303]. elastic [738, 1249, 1301]. electric [1062]. Electronic [40, 43, 46, 164, 284, 285, 288, 289, 292, 423, 448, 918]. elements [1147, 1215, 1742]. elimination [1139]. Elliott [927]. Elsevier [1589]. embedded [199, 755, 821]. embedding [1634]. emergency [874, 877]. Emerging [212, 1418]. empirical [1306, 1544]. enable [200, 650, 798]. enabled [84, 753, 777, 1012, 1430, 1533, 1673]. enablers [436]. Enabling [417, 420, 1527, 1660, 1778]. encapsulation [888, 889]. encoded [680]. encoding [99, 658]. encounter [605]. encryption [721]. End [50, 169, 181, 195, 313, 352, 432, 704, 772, 874, 915, 941, 1057, 1064, 1076, 1108, 1129, 1145, 1264, 1284, 1401, 1582, 1609, 1617, 1742, 1747, 1877, 1897, 1980]. End-host [1057]. end-point [1401]. End-to-end [50, 169, 181, 195, 313, 352, 432, 704, 772, 874, 915, 941, 1064, 1076, 1108, 1129, 1145, 1284, 1582, 1617, 1742, 1747, 1877, 1897, 1980]. Energy [389, 510, 548, 552, 642, 801, 1222, 1303, 1423, 1443, 1477, 1542, 1573, 1575, 1576, 1682, 1685, 1690, 1715, 1822, 1882, 1897, 1912, 1916, 1930, 1987]. Energy-aware [642, 1423, 1682]. Energy-conservation [1690]. Energy-efficient [548, 1303, 1443, 1542, 1576, 1882, 1897, 1916]. energy-limited [389]. enforcement [317]. engine [140, 1096]. engineered [1149]. Engineering [213, 219, 442–444, 455, 481, 674, 845, 924, 1247, 1481, 1686, 1694, 1737, 1970]. engines [76, 77, 400, 401, 1095]. enhance [550, 1145, 1879]. Enhanced [312, 373, 746, 968, 979, 1353, 1486, 1601, 1713, 1790]. enhancement [1327, 1355, 1374, 1431, 1483]. enhancements [957]. Enhancing [87, 720, 1145, 1405, 1637, 1706, 1747, 1856, 1858]. ensure [1145]. entanglement [1262]. enterprise [1607]. enterprises [286]. entertainment [948]. entropy [100, 2000]. Envelope [910]. environment [112, 172, 174, 213, 228, 251, 266, 360, 379, 384, 388, 754, 794, 809, 1122, 1165, 1284, 1358, 1675]. environments [178, 195, 244, 264, 432, 469, 1041, 1100, 1406, 1478, 1483, 1518, 1606, 1749, 1771, 1772]. ephemeral [350, 1685]. Epidemic [1083, 1410, 1598, 1883]. Epidemic-based [1883]. equal [976]. equal-cost-multi-paths [976]. equalization [1430]. equation [517, 1346, 1366]. equation-based [517, 1346]. EquiCast [1948]. Equilibrium [316, 1775]. equivalence [1089]. equivalence-checking [1089]. equivalent [910]. erasure [247, 373, 1478]. erbium [1675]. erbium-doped [1675]. ergodic [1872]. Error [793, 1185, 1267, 1918]. error-prone [793]. errors [1187]. establishment [759, 1154]. estimate [804]. estimates [1922]. Estimating [920, 1522]. estimation [429, 464, 530, 538, 612, 640, 814, 913, 915, 1037, 1054, 1124, 1128, 1388, 1422, 1626, 1698, 1707, 1846, 1863, 1966]. estimators [463]. ETEN [793]. eternal [1762]. Ethernet [1029, 1153, 1531, 1581, 1640, 1759, 1790, 1856, 1915]. Ethernet-over-SONET [1153]. Eulogies [1865]. Eulogy [730]. Evaluating [1736]. Evaluation [110, 120, 146, 263, 264, 275, 304, 318, 326, 339, 360, 377, 388, 459, 530, 685, 732, 782, 787, 879, 964, 969, 971, 1030, 1131, 1136, 1137, 1174, 1175, 1293, 1294, 1324, 1333, 1360, 1411, 1422, 1429, 1499, 1558, 1643, 1789, 1812, 1854, 1867, 1874]. evaluations [1819]. even [1950]. Event [397, 477, 1100, 1655, 1923]. event-based [1923]. Event-condition-action [397, 1100]. event-driven [477]. everyday [201]. evidence [911]. Evolution [72, 399, 583, 968, 1906, 1990]. evolutionary [1512, 1906]. evolving [1099]. example [566, 1511]. exceedance [1863]. excess [550]. exchange [99, 490, 820, 1685, 1750]. execution [478]. exemplified [262]. Existing [1444]. Exogenous [1152].

Exogenous-loss [1152]. expectations [104, 337, 753]. Expected [765, 1766]. expedited [1006]. experience [94, 201, 509]. Experiences [166, 327, 1858, 1992]. experiment [509]. Experimental [163, 797, 1411, 1724, 1854]. experiments [76, 469, 874]. experts [70]. Explicit [81, 120, 187, 793, 1170, 1172, 1343, 1345, 1749, 1827, 1854]. explicit-rate [120]. exploitation [688]. Exploiting [311, 348, 357, 616, 763, 1501, 1749, 1944]. exploration [874, 1313]. explorations [1201]. exponent [1388]. extended [1482, 1663]. Extending [420, 922, 1160, 1346, 1421, 1594]. Extensible [71, 347, 451, 1050]. extension [181, 431, 1751]. extensions [156, 1262, 1383, 1520]. extraction [421].

Fabric [11, 200, 453, 1281]. fabrics [311, 1148]. factors [50]. FAD [1877]. fading [1788]. Failure [1090, 1197, 1466, 1640]. failure-tolerant [1090]. failures [1017, 1085, 1200, 1696]. Fair [383, 468, 483, 490, 643, 662, 668, 820, 852, 919, 920, 961, 1002, 1021, 1320, 1375, 1392, 1679, 1740, 1750, 1756, 1774, 1915, 1941]. fair-share [1941]. Fairness [517, 616, 725, 766, 937, 939, 1091, 1146, 1152, 1414, 1479, 1672, 1826, 1944, 1954, 1977]. false [1758]. family [1372]. fan [407]. fan-out [407]. Fano [898]. Fast [64, 594, 672, 742, 1027, 1121, 1176, 1195, 1269, 1284, 1339, 1349, 1383, 1534, 1606, 1623, 1640, 1761, 1945]. fast-rerouting [1195]. FastTrack [1059]. Fault [721, 749, 1017, 1042, 1194, 1223, 1228, 1770, 1776, 1824, 1830]. fault-tolerant [1017, 1770, 1776]. faults [1824]. FCFS [1808]. FDTs [299]. feasible [603, 1055, 1491]. Feature [22–26, 28, 496, 751– 755, 1256, 1257, 1259, 1261, 1262, 1264, 1265]. feature-rich [1259]. featured [1260]. features [27, 651, 704, 902]. featuring [891].

FEC [112, 390, 393, 997, 1077]. fed [459]. Federal [320]. federated [178]. Feedback [141, 1083, 1343, 1345, 1404, 1503, 1846, 1910]. Feedback-assisted [1846]. Fermat [1968]. fewest [1829]. Fi [734]. fiber [246, 1005, 1293, 1675, 1793]. fiber/coax [246]. fidelity [451]. field [874, 1755]. FIFO [480, 577, 1302, 1502]. FIFO-based [1302]. Fifth [28]. file [1011, 1288, 1478, 1517, 1527, 1609, 1613, 1815]. file-sharing [1815]. filter [1393, 1536]. filtering [95, 365, 1086, 1966, 2000]. filters [950, 1121, 1637]. Finding [242, 1733, 1884]. fingerprinting [878, 1782]. Finite [462, 1568, 1853, 1873, 1983]. Finite-source [1873]. Firewall [1300, 1429, 1979]. FireWire [307]. First [23, 327, 897, 1438]. fixed [335, 690, 1280, 1888]. fixing [297]. FLAME [1167]. flash [1177]. flat [260]. flat-rate [260]. flaw [297]. FLC [267]. FLC-based [267]. Flexible [3, 294, 526, 540, 796, 870, 1163, 1166, 1167, 1527, 1839]. flooding [552, 1105, 1407]. flooding-based [1105]. Flow [6, 16, 119, 129, 130, 302, 439, 440, 504, 554, 580, 609, 684, 786, 901, 1008, 1088, 1120, 1212, 1280, 1354, 1587, 1607, 1643, 1684, 1769, 1996]. Flow-level [1354]. flows [170, 296, 471, 539, 550, 671, 738, 910, 1081, 1119, 1123, 1335, 1414, 1980]. fluctuations [1181]. Fluid [671, 1088, 1132, 1242, 1769]. fluid-based [671]. Fluid-flow [1088]. FluNet [1534]. Focused [62, 1313]. focussing [326]. foe [817]. Fokker [1366]. forecast [496]. foreground [26]. foreground/background [26]. forest [713]. form [1552]. Formal [252, 289, 299, 337, 569, 639, 1568]. format [60, 99, 171]. formation [1998]. forms [416]. formula [1552]. formulation [1699]. forward [247, 1267]. forwarding [277, 350, 517, 632, 1006, 1269, 1526, 1567, 1579, 1638, 1933]. foundations [186]. fountain [1442]. fP2P [1818]. Fractional [458]. fragmentations [1692]. frame [266, 664, 917]. frame-size-based [664]. Framework [131, 154, 179, 288, 317, 318, 418, 445, 518, 546, 603, 638, 683, 800, 821, 840, 932, 962, 1099, 1136, 1137, 1163, 1175, 1213, 1241, 1298, 1305, 1345, 1384, 1394, 1402, 1550, 1606, 1801, 1837, 1869, 1898, 1971]. frameworks [1677]. free [516, 1106, 1121, 1326, 1432, 1584, 1670, 1694, 1712, 1732]. Freenet [806]. frequency [873, 1422]. freshness [667]. friend [817]. friendliness [1503]. friendly [306, 374, 501, 603, 862]. friends [1988]. front [187]. full [417, 1860]. Fully [647]. function [677]. functionality [265, 870, 1159]. functions [123, 367, 464, 1784]. Fundamental [527, 1288, 1618, 1925]. further [1372]. Future [153, 283, 436, 575, 792, 863, 868, 1078, 1351, 1481]. Fuzzy [4, 879, 1356, 1515, 1827, 1951].

G [173, 430, 1788, 1868]. G-WiN [173]. G.805 [1664]. gains [1605]. Game [344, 760, 953, 1068, 1150, 1193, 1434, 1442, 1741, 1748, 1906, 1972]. game-theoretic [1068, 1748]. games [950, 952, 954, 1655]. gaming [951, 1487]. gamma [1455]. gap [104]. gateway [1551, 1774, 1992]. gateways [431]. gathering [593, 634, 1222]. Gaussian [460, 1281]. Gbit [173]. Gbit/s [173]. GBN [927]. general [460, 489]. Generalized [461, 462, 471, 643, 742, 941, 1017, 1084, 1204, 1321, 1456, 1925]. generated [402]. Generating [572, 1253]. Generation [12, 29, 41, 88, 214, 283, 317, 377, 453, 522, 725, 815, 822, 977, 989, 1141, 1216, 1259, 1260, 1513, 1555, 1952, 1972, 1992, 2002]. generation/dynamic [1141]. Generic [502, 605, 973, 1203, 1213, 1261, 1952]. Genesis [1135]. genetic [290, 316, 503, 1295, 1476, 1608]. GEO [387]. geocast [537]. Geographic [860, 874, 1187, 1214, 1459, 1485, 1567, 1968]. geographical [32, 1538]. geographically [1232]. Geometrical [1919]. Georoy [1355]. geostationary [383, 385, 609, 847, 1558]. German [173]. GFR [369]. Gilbert [927]. Global [130, 710, 1152, 1350, 1490, 1650, 1820]. Global-ISP [1350]. GLR [1214]. GMPLS [1247, 1621, 1662, 1665, 1666, 1668]. GMPLS-based [1621, 1665, 1668]. Gnutella [1528, 1585]. go [949]. goal [1272]. goals [42]. good [937]. goodputs [137]. GoS [1276]. gossip [1941–1943, 1945]. gossip-based [1941–1943, 1945]. GPRS [157, 315, 1186, 1192]. grade [1393]. grain [712]. Granular [1275]. Granularity [32]. Graph [68, 519, 1017, 1245, 1696, 1847]. graph-based [1847]. Graphic [478]. graphs [1733]. grating [516]. gray [1989]. gray-box [1989]. greedy [1334]. grid [1500, 1651]. GridJet [1510]. Grids [436]. groomed [733]. Grooming [503, 726, 783, 982, 1212, 1661]. Group [3, 706, 720, 1016, 1199, 1240, 1420, 1448, 1476, 1493, 1523, 1636, 1696, 1805]. group-based [1199]. group-to-tree [1420]. grouping [1166]. groups [1039, 1118, 1240, 1879]. growth [398, 714, 1512]. GS [631]. GSM [157, 428]. GTS [1711]. guarantee [1006]. Guaranteed [120, 263, 298, 386, 480, 621, 784, 901, 1078, 1180, 1223, 1252, 1405, 1645, 1808]. guarantees [332, 844, 938, 984, 1302, 1433]. guard [8, 612, 826]. Guest [160, 274, 284, 308, 382, 435, 455, 474, 966, 1158, 1255, 1541, 1786]. Guided [402, 1426]. Guidelines [1675]. Gunter [1865].

H [1820]. H-P2PSIP [1820]. H.264 [997]. H.323 [172, 718]. HACK [404]. hand [280]. hand-offs [280]. handle [357]. Handling [1177, 1264, 1395]. Handoff [345, 512, 737, 805, 815, 841, 1051, 1123, 1282, 1643]. handoffs [1323]. handover [581, 974, 1184, 1533, 1888, 1914]. handovers [1383]. HAP [1446]. hardware [348]. Harnessing [1348]. harvesting [1215]. hash [1784]. hashing [91]. HB [1372].

HB-family [1372]. HB-MP [1372]. HCCA [1378]. Header [1050, 1108, 1175]. Header-size [1108]. healing [370, 631, 1898]. heavy [462, 464, 1892]. heavy-tailed [462, 464]. Hellman [1685]. helping [1538]. here [949]. HERO [1790]. heterogeneity [1277]. heterogeneous [342, 560, 605, 767, 842, 892, 1123, 1288, 1299, 1395, 1439, 1530, 1554, 1578, 1595, 1616, 1699, 1742, 1749, 1755, 1772, 1851, 1859, 1962, 1970, 1987]. Heuristic [185, 1229, 1849, 1886, 1892, 1960]. heuristics [1699, 1736]. HFC [17]. hidden [1272, 1380, 1719]. hide [918]. hiding [1055]. Hierarchical [135, 356, 357, 372, 643, 667, 761, 805, 1022, 1029, 1332, 1511, 1523, 1557, 1594, 1801, 1805, 1820, 1918, 1933]. hierarchically [655]. hierarchically-clustered [655]. hierarchies [1805]. hierarchy [425, 1427]. High [102, 128, 170, 300, 324, 348, 391, 451, 643, 648, 797, 872, 901, 944, 968, 1148, 1160, 1237, 1311, 1317, 1340–1343, 1345, 1346, 1446, 1616, 1647, 1684, 1697, 1790, 1839, 1934, 1994, 1999]. high-bandwidth [102]. high-fidelity [451]. high-interaction [1311]. High-performance [170, 348, 1148, 1317]. high-rate [1994]. high-speed [300, 324, 901, 1160, 1340– 1342, 1346, 1616, 1790, 1839]. High-Throughput [872]. higher [1830]. higher-level [1830]. highest [1556, 1884]. highest-ranked [1884]. Highly [526, 759, 1163, 1231, 1313, 1839]. hinted [280]. hinted- [280]. HiperLAN [276, 277]. HiperLAN/2 [276, 277]. HiperLAN2 [275]. histogram [1973]. history [259]. hitlist [1445]. Hitting [205]. HN [1818]. hoc [202, 268, 281, 338, 361, 375, 431, 507, 511, 534, 537, 552, 622, 719, 759, 796, 813, 833, 876, 926, 939, 1063, 1090, 1111, 1127, 1205, 1211, 1214, 1244, 1283, 1285, 1287, 1304, 1326, 1380, 1384, 1390, 1395, 1424, 1452, 1459, 1475, 1482, 1522, 1535, 1538, 1549, 1550, 1562, 1567, 1573, 1587, 1602, 1624, 1627, 1642, 1645, 1653, 1705, 1718, 1766, 1883, 1899, 1901, 1912, 1951, 1953, 1963, 1968]. Hogwarts [198]. home [545–547, 578, 1228, 1439, 1764]. homes [1992]. homogeneous [1288, 1707]. homomorphisms [1744]. honeypots [1311]. Hop [6, 303, 361, 510, 622, 703, 833, 871, 876, 939, 967, 1111, 1190, 1211, 1244, 1432, 1461, 1627, 1765, 1781]. Hop-by-hop [6, 303, 703, 1432]. hop-limited [1765]. HORNET [35]. hose [1337, 1529, 1754]. hose-model [1529]. host [1057, 1312, 1749]. host-based [1312]. hosted [87]. hosting [1628, 1913, 2001]. hosts [860, 1390, 1678]. hot [243, 509, 1794]. hot-spots [1794]. hotspots [734, 970, 1586]. hour [1311]. HSDPA [968, 1868]. HTML [87, 1102]. HTML/XML [1102]. HTTP [69, 88, 723, 1009, 1310, 1508]. HTTP-based [1508]. hub [1021, 1469]. Hunter [1782]. Hurst [913]. Hybrid [27, 158, 211, 246, 248, 384, 516, 676, 788, 847, 862, 869, 978, 1122, 1303, 1336, 1534, 1577, 1882, 1892, 1937]. hypermedia [85]. hypothesis [749].

I-Views [86]. IBGP [1226]. IC [1374]. IC-card-based [1374]. ID [388]. identification [1373, 1742, 1838, 1842]. Identifying [1258, 1501]. Identity [291, 817, 1285]. Identity-based [1285]. IEEE [278, 547, 548, 622, 969, 1009, 1179, 1319, 1329, 1351, 1353, 1360, 1378, 1380, 1381, 1404, 1419, 1434, 1437, 1545, 1547, 1552, 1605, 1647, 1685, 1711, 1713, 1788, 1803, 1983, 1998]. IEEE802.11 [737]. IGM [826]. IGMP [1852]. IGMP-AC [1852]. IIN [209]. IMA [186]. image [1185, 1715]. Imaging [103]. immune [1315]. Impact [125, 345, 359, 432, 580, 1340, 1386, 1452, 1482, 1524, 1545, 1563, 1622, 1659, 1675, 1704, 1758, 1807, 1813, 1936]. impairment [1662]. impairment-aware [1662]. impairments [1848]. impatient [1873]. imperative [639]. implement [244]. Implementation [44, 53, 170, 309, 310, 418, 557, 737, 888, 977, 1063, 1376, 1389, 1516, 1713, 1794, 1935]. implementations [1959]. implementing [232, 809]. implications [232, 741, 911, 1048, 1245, 1741, 1816]. implicit [701]. implied [135]. imposed [1203]. Improve [404, 806, 1115, 1179, 1944]. Improved [75, 851, 1826, 1950]. improvement [510, 611, 766, 1498, 1697]. improvements [1878]. Improving [137, 147, 187, 220, 265, 437, 658, 684, 691, 783, 811, 823, 843, 860, 896, 899, 1000, 1025, 1112, 1144, 1272, 1365, 1403, 1436, 1464, 1538, 1639, 1749, 1968, 1977]. iMST [1645]. IMT [154, 155]. IMT-2000 [154, 155]. IN-Internet [211]. inaccurate [824]. Incentive [1953]. Incentive-based [1953]. incentives [779, 1041, 1440]. incident [152]. incident-response [152]. incoming [1566]. incomplete [1472]. inconsistencies [1322]. inconsistency [1085]. incorporates [1426]. Increase [869, 1697, 1823, 1962]. Increasing [1086, 1523]. incremental [749]. incrementally [468]. independent [1138, 1954]. Index [47, 48, 106, 107, 175, 176, 224, 225, 272, 273, 330, 331, 380, 381, 433, 434, 484– 487, 542, 543, 599, 600, 652, 653, 707, 708, 768, 769, 827, 828, 884, 885, 945, 946, 1004, 1234]. indexing [64, 1011]. individual [1287]. indoor [509]. industrial [764]. industry [337]. inelastic [1414]. inference [463, 1129, 1877]. influence [50, 313]. Information [46, 79, 270, 322, 575, 760, 766, 824, 874, 880, 981, 1313, 1404, 1451, 1509, 1755, 1819, 1943, 1947, 1997]. infrastructural [534]. infrastructure [124, 173, 200, 227, 341, 419, 942, 1058, 1394, 1506, 1877, 1983]. infrastructures [232, 743, 1811]. initiation [951]. innovating [1769]. Innovations [1506]. Input [311, 460, 462, 575, 1049, 1491, 1855]. input-buffered [1491]. Input-queued [311, 575, 1049]. Insensitive [1074]. insight [1998]. inspired [1514, 1900]. installation [229]. integer [1699]. integrate [1209, 1299]. Integrated [4, 251, 315, 333, 360, 405, 620, 702, 719, 824, 840, 843, 1051, 1178, 1270, 1284, 1489, 1646, 1764]. integrated-routing [620]. Integrating [49, 57, 1133]. Integration [42, 158, 159, 208, 246, 578, 838, 1049, 1251]. intelligence [152]. Intelligent [2, 183, 200, 208, 816, 1667]. Inter [1, 148, 387, 984, 1278, 1417, 1493, 1724, 1781]. inter-cell [1417, 1724]. inter-component [148]. Inter-domain [1, 984, 1278, 1781]. inter-group [1493]. inter-satellite [387]. interacting [1316]. interaction [23–26, 28, 223, 228, 339, 365, 496, 515, 751, 752, 1259, 1262, 1296, 1311, 1723]. Interactions [22, 27, 28, 568, 753– 755, 1205, 1256–1258, 1264, 1265, 1927]. interactive [437, 757, 954, 1249, 1615]. Interactivity [1479, 1483]. intercepting [1700]. interchange [60]. interconnection [773, 891, 1820, 1874]. interconnectivity [736]. interdependence [1927]. interdomain [1191]. interest [1944]. interface [42, 625, 1579, 1613, 1772]. interface-specific [1579]. interfaces [650, 1281]. interference [826, 1114, 1437, 1619, 1724, 1779, 1788, 1822]. interfering [120]. interleaved [997]. intermediaries [425]. Intermediary [743]. intermediate [1456]. internal [941]. International [28, 679, 795, 1071, 1686, 2002]. Internet [12, 21, 29, 43, 63, 70, 86, 89, 102, 103, 128, 140, 171, 180, 206, 208, 211–213, 235, 259, 262, 314, 317, 327, 379, 394, 424, 435, 436, 439, 442, 482, 509, 603, 608, 681, 700, 710, 711, 713, 736, 747, 754, 762, 771, 777, 860, 866– 868, 887, 908, 911–913, 949, 1008, 1012, 1048, 1078, 1180, 1189, 1200, 1201, 1209, 1224, 1232, 1252, 1257, 1359, 1400, 1414, 1470, 1481, 1484, 1485, 1492, 1515, 1526, 1534, 1540, 1551, 1558, 1561, 1599, 1719, 1737, 1806, 1910]. Internet-PSTN [213]. Internet-scale [206]. Interoperability [20, 547, 647].

Interoperable [215]. InterPlaNetary [608, 866]. interpretation [1878]. interpreter [1611, 1974]. Interrupt [1967]. intervals [184]. Interworking [718, 839, 1554]. Intra [56, 883, 1180, 1493, 1993]. intra-document [56]. Intra-domain [883, 1180, 1993]. Introducing [1611, 1668]. Introduction [566, 571]. intruder [1272]. intruders [151]. Intrusion [143, 144, 146– 150, 152, 723, 1303, 1309, 1312, 1316, 1955]. Intrusion-detection [152]. IntServ [315]. IntServ/RSVP [315]. invariants [903, 1739]. investigating [1239]. Investigation [21, 463, 1868]. invocation [1518]. IP [69, 126, 137, 138, 215, 248, 276, 280, 341, 360, 384, 393, 407, 408, 440, 443, 446, 451, 604, 611, 620, 650, 668, 672, 675, 679, 701, 732, 787, 824, 825, 830, 838, 845, 853, 863, 870, 888, 892, 900, 907, 957, 973, 1019, 1026, 1028, 1047, 1056, 1063, 1071, 1075, 1076, 1170, 1213, 1220, 1228, 1242, 1247, 1278, 1284, 1286, 1349, 1393, 1438, 1453, 1518, 1530, 1554, 1559, 1566, 1577, 1621, 1669, 1739, 1809, 1818, 1826, 1957, 1969]. IP-based [892, 1818]. IP-in-IP [650]. IP-over-WDM [620, 668, 824, 853, 1026]. IP/MPLS [1247, 1284, 1826]. IP/MPLS-over-WDM [1826]. IP/WDM [1669]. IPsec [1203]. IPTV [1814, 1920]. IPv6 [295, 447, 1079, 1145, 1156, 1332]. IRIS [1862]. isolation [1462]. ISP [772, 1350, 1741]. Issue [1091, 1094, 1553, 1589, 1658, 1709, 1786, 1865]. issues [38, 44, 167, 282, 300, 340, 839, 948, 1990]. Italia [174]. Italian [164]. iterative [1294]. itineraries [1190]. ITU [565, 1664]. ITU-T [565, 1664]. IXP [1974].

jamming [189]. Java [59, 251]. JavaCard [252]. JavaML [59]. Jellyfish [1733]. Jetmax [1616]. jitter [313, 482]. JML [252]. job [97, 1987]. Joint [520, 675, 997, 1146, 1297, 1587, 1647, 1679, 1912]. jointly [1585]. journal [466, 1658].

Key [449, 855, 926, 1042, 1118, 1240, 1427, 1493, 1523, 1685, 1805]. keying [1251, 1304]. Keys [412]. keyword [57, 147]. keywords [43]. knowledge [420, 1978]. known [105]. knows [1564].

label [392, 442, 784, 1349, 1649, 1856, 1933]. label-switched [784]. LambdaGrids [1440]. LAN [456, 838, 839, 997, 1280, 1531, 1724, 1759, 1995]. Language [55, 58, 59, 233, 588, 935, 991, 1050, 1168, 1262, 1264]. language-based [233]. languages [565, 987]. LANS [275, 809, 844, 878, 1025, 1329, 1379, 1404, 1892]. large [32, 92, 151, 453, 500, 630, 649, 665, 926, 954, 1134, 1135, 1214, 1240, 1403, 1420, 1474, 1483, 1492, 1540, 1576, 1626, 1652, 1739, 1780, 1858, 1875, 1919, 1932, 1947, 1982]. large-scale [453, 500, 630, 649, 1134, 1135, 1483, 1540, 1576, 1626, 1652, 1739, 1780, 1932]. latencies [536]. Latency [18, 91, 270, 409, 432, 539, 621, 952, 1222, 1626, 1641, 1822]. latency-aware [1641]. Latency-Energy [1822]. Latency-sensitive [91]. latent [1348]. latest [1444]. Layer [69, 406, 686, 811, 890, 900, 957, 1026, 1273, 1277, 1425, 1477, 1498, 1542, 1590, 1593, 1611, 1622, 1639, 1664, 1704, 1726, 1736, 1782, 1841, 1848, 1857, 1907]. layer-2 [1593]. layer-3 [1841]. Layered [514, 1162, 1857, 1861]. layering [432, 1159]. layout [245, 1950]. leaf [1585]. leaky [2]. Learning [132, 329, 566, 878, 1099, 1310, 1463, 1566, 1832, 1929]. learning-based [1929]. learnt [1397]. leave [1523]. Ledge [1978]. Legendre [1072]. LEO [386, 1238, 1496, 1888]. less [1929]. let [1988]. Level [15, 111, 162, 193, 227, 390, 439, 556, 580, 650, 700, 772, 825, 881, 958, 1083, 1133, 1271, 1354, 1430, 1536, 1576, 1735, 1752, 1830, 1877]. Leveraging [1079, 1131]. Lévy [458]. LIAD [1962]. libraries [165]. life [216, 1163]. lifecycle [589]. lifetime [593, 1183, 1208, 1522, 1543, 1996]. lifetimes [1430]. lightpath [895, 1026, 1154]. Lightweight [128, 371, 888, 981, 1161, 1205, 1225, 1283, 1372, 1508]. like [323, 363, 963]. limitations [172, 704]. limited [185, 389, 671, 706, 1109, 1274, 1362, 1666, 1765]. limits [575]. line [146, 810, 1138, 1227]. Linear [641, 767, 1409]. Link [73–75, 123, 387, 390, 643, 677, 686, 766, 824, 972, 1017, 1020, 1149, 1181, 1207, 1329, 1354, 1639, 1792, 1863, 1877, 1879]. link-layer [686]. link-level [390, 1877]. link-sharing [643]. link-structure [74]. links [314, 404, 554, 1047, 1501, 1684, 1718, 1788, 1800]. Linux [1479]. LIPS [1225]. List [1308]. Lite [1462]. live [1789, 1813]. lived [663, 693, 1552]. Load [122, 271, 450, 846, 853, 1074, 1305, 1343, 1354, 1387, 1557, 1573, 1631, 1652, 1784, 1792]. Load-balancing [122, 1387, 1784]. loaded [1892]. loads [559]. local [507, 938, 950, 1205, 1223, 1426, 1648]. local-scope [1205]. locality [1944]. localization [633, 1388, 1391, 1723, 1999]. Locating [1165]. Location [268, 611, 711, 761, 809, 814, 822, 841, 860, 878, 1055, 1176, 1186, 1187, 1243, 1328, 1355, 1459, 1583, 1636, 1723, 1753, 1764, 1777, 1922]. location-aware [809, 841, 1355]. Location-based [1753]. location-hiding [1055]. locations [56]. Locator [297]. Logarithmic [1697, 1962]. logic [1515]. logic-based [1515]. login [1374]. Long [455, 461, 463, 664, 785, 906, 911, 912, 1339, 1346, 1552, 1697, 1976]. long-distance [1339, 1346]. long-duration [664]. long-lived [1552]. Long-range [455, 461, 463, 912]. longest [1438]. look [911]. lookup [672, 958, 1438, 1561]. Loop [765, 1106, 1326, 1432, 1694]. Loop-free [1106, 1326, 1432, 1694]. loops [1226, 1526, 1579]. Loss [127, 499, 738, 804, 997, 1083, 1119, 1152, 1180, 1267, 1302, 1334, 1388, 1415, 1569, 1632, 1647, 1793, 1863, 1877, 1929]. Loss-cognizant [1334]. Loss-resilient [1632]. lossy [404, 1501, 1977]. LOTOS [25, 244, 307, 477, 478, 764]. LOTOS-based [25]. Low [102, 355, 391, 889, 1124, 1151, 1190, 1294, 1311, 1604, 1678, 1715, 1737, 1850]. low- [1311]. low-bandwidth [1678]. Low-complexity [1124, 1715, 1737]. Low-cost [102, 355, 1190]. low-diameter [1604]. low-rate [1151, 1294, 1850]. lower [855, 1108]. LRU [1612]. lunar [874]. LVMSR [514].

M [262, 430, 1320, 1868, 1873]. M/G/1 [430]. MAC [184, 516, 673, 1244, 1324, 1431, 1477, 1642, 1682, 1710, 1803, 1850, 1881, 1916, 1959]. Machine [1397]. Machine-learnt [1397]. machines [1312, 1568, 1956, 1989]. MAI [597]. mail [1841]. Maintenance [14, 1060, 1277, 1362, 1567]. malformed [1394]. Malicious [1040]. MAN [1531, 1945]. managed [1628, 1904]. Management [34, 161, 162, 165, 167, 177– 181, 230, 235, 240, 271, 279, 291, 316, 356, 358, 360, 372, 385, 389, 429, 441, 449, 512, 577, 596, 602, 606, 611, 625, 642, 662, 681, 688, 712, 719, 766, 767, 782, 800, 822, 826, 833, 861, 863, 890, 926, 1035, 1037, 1152, 1164, 1177, 1179, 1186, 1233, 1243, 1328, 1341, 1358, 1360, 1363, 1427, 1439, 1447, 1477, 1495, 1523, 1764, 1772, 1794, 1809, 1900, 1903, 1914, 1933, 1988, 1989]. Managing [816, 825, 1265, 1514, 1897, 1986]. MANETs [1199, 1454, 1897]. Manhattan [561]. manufacturing [72]. many [296, 1118, 1235, 1376]. many-to-one [1235, 1376]. Mapping [408, 1138]. maps [989, 1182]. margin [826]. mark [187]. mark-front [187]. marker [981]. market [44, 220]. marketing [101]. marketplaces [285, 288]. markets [287]. marking [517, 551, 1081, 1220, 1278, 1286, 1399, 1827]. Markov [73, 1719]. Markovian [36, 849, 1963]. Markovian-based [1963]. markup [59]. MARQ [630]. Martian [873]. masking [1824]. massive [1784]. massively [1201, 1655]. master [1386]. matching [1420, 1491, 1536]. MATE [481]. Mathematical [776, 882, 1862]. mature [1722]. max [1120, 1616]. Maximal [1183]. Maximizing [1105, 1605, 1727, 1851]. maximum [549, 593, 1320]. May [2002]. MBMS [1533]. McPAD [1843]. MCR [134]. MCR-weighted [134]. MDP [127, 794]. mean [262]. means [402, 409, 1719, 1837]. measure [1231]. Measurement [121, 321, 428, 500, 666, 787, 804, 860, 1008, 1028, 1059, 1130, 1484, 1526, 1572, 1603, 1684, 1815, 1839]. Measurement-based [428, 666, 787, 804, 860, 1130, 1572]. measurements [762, 791, 1063, 1582, 1732, 1816]. Measuring [723, 830, 1986]. mechanism [183, 188, 271, 278, 306, 404, 407, 917, 1151, 1270, 1374, 1537, 1621, 1703, 1862]. Mechanisms [275, 328, 685, 694, 704, 753, 849, 877, 1379, 1677, 1745, 1888]. Media [60, 325, 352, 676, 738, 847, 1064, 1123, 1253, 1330, 1524, 1773, 1857]. media-oriented [1773]. Mediacups [201]. Medium [203, 361, 939, 1292, 1416, 1924]. meeting [337]. MEHROM [1080]. membership [1946]. memetic [618]. Memex [94]. Memory [15, 374, 729, 1637, 1865, 1869]. memory-based [374]. memoryless [1108]. Mental [344]. mesh [445, 726, 858, 891, 895, 1018, 1171, 1198, 1381, 1387, 1441, 1489, 1495, 1639, 1680, 1681, 1779, 1783, 1803, 1935–1937, 1958, 1960]. Meshed [609, 632, 1669]. message [262, 990, 1394, 1396, 1938]. messages [552]. messaging [1516]. meta [77, 85]. metadata [585, 1100]. method [119, 438, 889, 898, 1209, 1252, 1336, 1366, 1415, 1999]. Methodologies [287]. methodology [132, 218, 451, 529, 914, 1721]. Methods [280, 337, 664, 686, 1096, 1217, 1558, 1680, 1800]. metric [426]. metrics [408, 1206]. Metro [516, 668, 776, 1640, 1672]. metropolitan [35, 341, 969, 982, 1531]. MF [889]. MF-TDMA [889]. mice [1242]. MicroISPs [102]. middleboxes [1841]. Middleware [319, 329, 821, 1389, 1486, 1513, 1514, 1519, 1520, 1542]. middleware-layer [1542]. Military [152, 808, 816]. Millau [99]. millennium [222, 223]. MIMD [944]. MIMO [1718, 1845]. min [1120, 1616]. mine [94]. minimal [1692]. minimization [1176, 1648, 1822]. minimize [1987]. Minimizing [655, 1153, 1918]. Minimum [765, 859, 1114, 1573, 1761, 1899]. Mining [96, 145, 401, 1098, 1099]. MIPv6 [1383, 1411]. misaligned [1572]. misbehaving [1401]. missions [792, 868]. Mistreatment [1413]. Mistreatment-resilient [1413]. mitigate [1600]. Mitigating [1579]. mitigation [846, 1781, 1788]. mix [912]. Mixed [1075, 1319, 1686]. Mixed-mode [1319]. mixtures [464]. MLP [4]. MMORPG [1193]. MobDHop [1211]. Mobile [14, 54, 161, 179, 194, 243, 268, 280, 292, 304, 358, 372, 431, 446, 447, 507, 511, 537, 552, 581, 611, 618, 619, 630, 660, 669, 687, 703, 746, 786, 811, 820, 833, 845, 863, 919, 926, 957, 963, 1051, 1080, 1123, 1127, 1156, 1176, 1211, 1215, 1228, 1287, 1296, 1304, 1332, 1384, 1390, 1423, 1424, 1452, 1493, 1516, 1535, 1549, 1555, 1562, 1567, 1575, 1583, 1602, 1606, 1613, 1643, 1646, 1653, 1749, 1760, 1766, 1818, 1825, 1851, 1868, 1883, 1899, 1901, 1951, 1963, 1968]. mobile-to-mobile [1613]. mobiles [316]. Mobility [155, 156, 339, 341, 356, 359, 360, 364, 366, 375, 384, 406, 446, 630, 732, 842, 970, 1080, 1211, 1332, 1358, 1362, 1390, 1417, 1453, 1482, 1530, 1549, 1649, 1653, 1734, 1779, 1914, 1933]. Mobility-assisted [630]. Mobility-aware [970, 1779]. Mobility-based [364, 1211]. MobiSEC [1935]. mode [1030, 1319, 1379, 1386]. Model [27, 77, 209, 210, 289, 294, 326, 333, 399, 413, 414, 456, 574, 603, 651, 671, 776, 806, 809, 896, 909, 931, 1009, 1066, 1081, 1084, 1129, 1155, 1174, 1175, 1275, 1315, 1337, 1449, 1485, 1512, 1529, 1664, 1754, 1769, 1807, 1862, 1943, 1951]. Model-based [1129]. model-driven [809]. Modeling [36, 58, 131, 263, 366, 450, 455, 458, 470, 615, 663, 675, 741, 873, 906, 962, 1101, 1137, 1208, 1216, 1253, 1333, 1384, 1410, 1429, 1548, 1583, 1719, 1865, 1869, 1870, 1874, 1885, 1963, 1969, 1971, 1983, 1986]. Modelling [110, 113, 531, 718, 963, 1126, 1242]. models [26, 36, 116, 261, 292, 349, 424, 491, 520, 609, 664, 988, 1073, 1142, 1242, 1260, 1397, 1422, 1502, 1528, 1680, 1719, 1815, 1816, 1824, 1875, 1973]. modem [1759]. moderation [112]. modern [1315, 1835]. modes [562, 1995]. modifications [209]. Modular [239, 529]. Molecular [1976]. moment [1221]. monitor [1091, 1957]. monitor-based [1091]. monitoring [576, 666, 900, 1112, 1167, 1182, 1190, 1194, 1306, 1400, 1499, 1607, 1620, 1760, 1802, 1828]. most [460, 1040, 1536]. Mother [1564]. motion [458]. move [949, 1531]. movement [280]. MoVES [1550]. MP [1372]. MPEG [317, 664, 680, 887, 888]. MPEG-2 [887, 888]. MPEG-4 [680]. MPEG-trace [664]. MPLS [392, 442, 444, 558, 637, 674, 832, 845, 1027, 1194, 1247, 1284, 1529, 1622, 1694, 1886, 1890, 1970]. MPLS-based [674, 845, 1529]. MPLS-over-WDM [1826]. MPOA [168]. mRDP [1508]. MSC [216, 223, 567, 568]. MSC-2000 [223, 567]. MSCs [572]. MSL [414]. MSOCKS [406]. MTCP [363]. Multi [117, 127, 140, 188, 237, 239, 295, 314, 361, 367, 379, 442, 472, 504, 510, 515, 610, 618, 622, 705, 745, 802, 833, 835, 841, 849, 863, 871, 876, 931, 939, 967, 976, 1002, 1026, 1090, 1111, 1190, 1244, 1280, 1293, 1327, 1331, 1337, 1386, 1392, 1430, 1461, 1518, 1542, 1576, 1586, 1587, 1622, 1627, 1642, 1649, 1664, 1703, 1737, 1770, 1773, 1781, 1793, 1855, 1890, 1958, 1980]. multi-access [841, 863]. multi-agent [610]. multi-application [1002]. multi-attribute [1773]. multi-buffered [117]. multi-channel [745, 1586, 1642, 1958]. multi-chip [1855]. multi-class [188, 295, 367, 472, 705, 1331, 1392]. multi-cluster [1386]. multi-domain [1518]. multi-edged [237]. multi-fiber [1293, 1793]. multi-hop [361, 510, 622, 833, 871, 876, 939, 967, 1111, 1190, 1244, 1627, 1781]. multi-hop-distance [1461]. Multi-layer [1026, 1622, 1664]. multi-level [1430, 1576]. multi-model [931]. multi-operator [379]. multi-organisational [515]. Multi-path [1090, 1337, 1703, 1737, 1890]. multi-population [618]. multi-power [835, 1327]. multi-protocol [442, 1649]. multi-radio [1542]. multi-rate [127, 1587]. multi-ring [504]. multi-search [140]. multi-service [239, 314, 1280]. multi-source [802]. multi-state [849]. multi-topology [1770]. multiaccess [343]. multicarrier [1590]. Multicast [3, 268, 310, 324, 338, 350, 363, 407, 425, 427, 449, 495, 499, 514, 536, 613, 636, 646, 658, 683, 703, 715, 831, 855, 862, 890, 900, 925, 982, 1003, 1007, 1010, 1018, 1030, 1033, 1048, 1049, 1056– 1058, 1136, 1170, 1172, 1227, 1237, 1321, 1383, 1406, 1421, 1425, 1476, 1493, 1523, 1555, 1602, 1629, 1726, 1736, 1783, 1852, 1855, 1857, 1883, 1897, 1948, 1968, 1981]. multicast-streams [427]. Multicast/broadcast [1555]. multicasting [355, 389, 699, 922, 994, 1083, 1198, 1283, 1565, 1780]. multiclass [663]. multicriteria [815]. multidimensional [1725]. multidomain [310]. multifiber [1005]. multifractal [615, 910]. Multifractality [907]. multihomed [1643]. multihoming [1079, 1436]. Multihop [281, 1462, 1548, 1551, 1654]. multilateral [291]. Multilevel [1693]. Multimedia [12, 43, 92, 93, 325, 333, 368, 376, 512, 619, 971, 1007, 1137, 1224, 1244, 1268, 1271, 1291, 1554, 1611, 1618, 1709, 1710, 1712, 1713, 1820, 1891, 1896, 1911]. multipath [481, 632, 1364, 1365, 1409, 1634]. Multiplayer [950, 951, 1655]. multiple [35, 39, 242, 301, 391, 516, 650, 772, 774, 872, 1079, 1190, 1423, 1475, 1745, 1752, 1784, 1803, 1804, 1824, 1830, 1843, 1924, 1936]. multiple-access [35]. multiple-bottleneck [1804]. multiple-cell [1745]. multiple-round [1924]. multiplexed [33]. multiplexing [335, 673, 1622]. multiplicative [869, 1806]. multipoint [134, 355]. multipoint-to-point [134]. multiprocessor [526]. multiprotocol [1349]. multirate [896, 1756]. Multiservice [193, 443, 679, 845, 1071, 1276, 1446]. multisignatures [1109]. multistage [781]. multithread [1769]. multiuser [1605]. Multiway [672, 1269]. MuON [1598]. MuPAC [835, 1327]. MuSeQoR [1090]. music [95, 949]. Mutual [1437, 1598, 1695].

names [1115, 1196]. nano [1789].

nanonetworks [1689, 1976]. NASA [791]. Nash [1204]. native [1160]. nature [1564]. navigation [82, 84, 402]. NBM [1003]. NCTUns [557]. near [67, 1902]. near-shortest [1902]. near-uniform [67]. Nearly [577]. need [329]. needed [125]. negotiating [1898]. negotiation [148, 287, 288, 290, 1209, 1773]. negotiations [423]. Neigh [1825]. neighbor [1631]. Neighborhood [1653, 1901]. neighboring [1997]. nested [1453]. Net [1272]. NetFlow [1607]. Netherlands [2002]. NetMine [1837]. NetPDL [1050]. nets [1250]. Network [34, 124, 133, 141, 159, 166, 177, 199, 200, 229, 232, 236, 238, 260, 281, 321, 326, 341, 344, 359, 372, 375, 447, 458, 470, 480, 519, 524, 525, 529, 530, 561, 596, 604, 614, 625, 638, 666, 668, 763, 791, 810, 870, 874, 906, 909, 915, 938, 941, 953, 963, 969, 1020, 1027, 1033, 1072, 1078, 1087, 1126, 1133, 1135, 1161, 1167, 1274, 1370, 1398, 1426, 1445, 1453, 1465, 1472, 1499, 1529, 1556, 1594, 1622, 1649, 1659, 1664, 1672, 1681, 1686, 1693, 1722, 1739, 1742, 1752, 1778, 1802, 1809, 1837, 1870, 1878, 1933, 1936, 1950, 1969, 1972, 2000]. network [4, 21, 35, 113, 237, 403, 451, 516, 520, 526– 528, 531, 557, 578, 606, 643, 650, 695, 733, 809, 814, 832, 841, 872, 891, 893, 901, 914, 988, 1047, 1076, 1079, 1084, 1128, 1178, 1184, 1212, 1222, 1344, 1358, 1363, 1371, 1386, 1388, 1391, 1430, 1433, 1493, 1530, 1531, 1555, 1670, 1685, 1691, 1725, 1729, 1741, 1759, 1760, 1765, 1772, 1775, 1816, 1820, 1884, 1931, 1944, 1974]. network-calculus-based [1133]. network-wide [1765]. networked [951, 1479, 1978]. Networking [200, 202, 227, 258, 274, 319, 382, 453, 790, 798, 871, 948, 1165, 1542, 1658]. Networks [31, 32, 112, 114, 122, 132, 268, 278, 298, 300, 310, 333, 346, 354, 383, 392, 430, 431, 436, 466, 536, 581, 628, 629, 646, 655, 679, 779, 811, 826, 842, 858, 863, 944, 959, 979, 981, 995, 1000, 1015, 1027, 1028, 1065, 1071, 1123, 1141, 1149, 1168, 1205, 1242, 1251, 1291, 1304, 1325, 1341, 1387, 1417, 1431, 1439, 1443, 1469, 1480, 1514, 1549, 1550, 1556, 1562, 1575, 1589, 1598, 1607, 1645, 1649, 1669, 1671, 1680, 1682, 1699, 1709, 1723, 1726, 1727, 1745, 1791, 1811, 1848, 1865, 1890, 1914, 1919, 1929, 1962, 1963, 1972, 1996, 1999, 2001]. networks [17, 29, 128, 158, 191, 226, 304, 312, 340, 386, 429, 505, 507, 546, 614, 630, 632, 634, 648, 663, 674, 682, 684, 685, 725, 747, 774, 797, 825, 873, 887, 902, 939, 967, 972, 1001, 1010, 1011, 1016, 1017, 1037, 1062, 1063, 1086, 1106, 1108, 1127, 1132, 1152, 1155, 1198, 1211, 1215, 1229, 1243, 1271, 1295, 1297, 1334, 1342, 1348, 1400, 1408, 1482, 1502, 1512, 1525, 1535, 1537, 1551, 1559, 1567, 1580, 1619, 1621, 1633, 1639, 1641, 1665, 1686, 1732, 1740, 1744, 1747, 1795, 1807– 1809, 1818, 1827, 1866, 1880, 1887, 1905, 1917, 1930, 1953, 1998]. networks [9, 117, 119, 127, 141, 147, 193, 265, 358, 367, 368, 374, 405, 445, 500, 518, 537, 558, 593, 619, 633, 691, 692, 748, 761, 785, 786, 793, 803, 823, 824, 835, 838, 846, 853, 869, 881, 892, 904, 1006, 1060, 1074, 1091, 1133, 1139, 1177, 1179, 1181, 1183, 1197, 1246, 1270, 1283, 1292, 1301, 1339, 1351, 1382, 1384, 1390, 1403, 1407, 1462, 1489, 1495, 1498, 1501, 1513, 1522, 1576, 1581, 1583, 1584, 1593, 1602, 1624, 1627, 1631, 1642, 1653, 1654, 1659, 1697, 1703, 1705, 1731, 1739, 1766, 1776, 1779, 1783, 1801, 1803, 1850, 1859, 1871, 1881, 1896, 1899, 1916, 1925, 1937, 1965, 1982]. networks [126, 138, 153, 183, 208, 293, 305, 349, 382, 385, 428, 483, 510, 522, 545, 551, 554, 556, 609, 618, 637, 645, 656, 665, 702, 726, 759, 772, 784, 801, 845, 876, 923, 955, 957, 982, 999, 1005, 1014, 1019, 1022, 1090, 1145, 1173, 1182, 1214, 1217, 1227, 1244, 1260, 1267, 1285, 1306, 1326, 1327, 1332, 1345, 1362, 1381, 1392, 1402, 1424, 1434, 1452, 1464, 1533, 1546, 1557, 1573, 1613, 1630, 1662, 1674, 1711, 1714, 1715, 1748, 1754, 1755, 1763, 1764, 1777, 1825, 1826, 1833, 1835, 1851, 1862, 1872, 1874, 1883, 1904, 1915, 1923, 1926, 1935, 1951, 1958, 1994]. networks [13, 30, 129, 154, 185, 192, 230, 233, 234, 237, 239, 335, 356, 446, 492, 504, 552, 622, 660, 686, 705, 719, 724, 731, 745, 766, 776, 796, 812, 813, 840, 847, 888, 924, 962, 998, 1002, 1018, 1045, 1055, 1069, 1085, 1113, 1163, 1164, 1171, 1176, 1192, 1221, 1238, 1241, 1247, 1273, 1287, 1303, 1315, 1324, 1380, 1393, 1395, 1411, 1441, 1450, 1461, 1548, 1553–1555, 1564, 1569, 1578, 1587, 1636, 1643, 1646, 1666, 1668, 1679, 1710, 1712, 1718, 1728, 1734, 1743, 1758, 1830, 1858, 1882, 1885, 1886, 1889, 1907, 1942, 1952, 1959– 1961, 1968, 1970, 1980]. networks [2, 8, 110, 246, 262, 282, 355, 361, 364, 370, 389, 391, 495, 498, 508, 547, 611, 612, 620, 631, 642, 822, 833, 843, 895, 921, 926, 942, 1023, 1026, 1029, 1075, 1111, 1185, 1194, 1196, 1208, 1212, 1282, 1296, 1305, 1335, 1346, 1405, 1415, 1471, 1475, 1543, 1577, 1589, 1590, 1595, 1616, 1658, 1661, 1673, 1683, 1713, 1753, 1804, 1822, 1849, 1861, 1901, 1911, 1912, 1922, 1931, 1934, 1955, 1957, 1997]. neural [114, 147, 239, 1578]. newbies [70]. NewReno [241]. news [1927]. newspaper [245]. Next [12, 88, 214, 317, 453, 522, 725, 815, 822, 1003, 1141, 1216, 1272, 1513, 1555, 1952, 1972, 1992, 2002]. next-generation [214, 453]. NGN [1520]. NICE [1039]. NICs [1790]. Ninja [206]. Ninth [2002]. no [10, 1026, 1227]. nodal [9]. node [335, 561, 891, 1274, 1682, 1887]. node-selection [1887]. NodeOS [1159]. nodes [830, 1229, 1538, 1643, 1872]. nomadic [204]. Non [130, 387, 574, 609, 1005, 1015, 1242, 1375, 1434, 1545, 1802, 1872]. non-cooperative [1434, 1802]. non-ergodic [1872]. non-GEO [387]. non-geostationary [609]. non-repudiation [1375]. non-stationary [1242]. Non-TCP [574]. non-uniform [130, 1005, 1015, 1545]. nonblocking [503]. Nonlinear [651, 983, 1233, 1836]. normal [1231]. normality [723]. notation [566, 571]. note [977, 1688]. Notes [222]. notification [187, 793, 899, 1569]. notion [915]. Novel [267, 401, 611, 1179, 1214, 1283, 1299, 1343, 1407, 1569, 1621, 1630, 1695, 1890, 1902, 1920, 1929, 1942]. NSDF [638]. ntop [166]. number [549, 1274, 1751]. numbers [1106].

Object [103, 155, 651, 1298, 1389, 1955]. object-oriented [155]. objective [168]. objectives [1572]. objects [250, 325, 352]. OBS [1471, 1569, 1672, 1929]. OBS/OPS [1471]. Observation [791, 795, 1683]. occlusion [1712]. occlusion-free [1712]. occupancy [1366]. odd [1950]. ODP [142, 452]. OFDM [969, 1591]. OFDM-based [969, 1591]. OFDMA [1382, 1619, 1931]. off [146, 909]. off-line [146]. offer [1830]. offline [1124]. offs [280, 1038]. offset [766]. Olga [729, 730, 740]. OMNI [1058]. OMP [410]. OMS [1668]. on-board [383, 385]. On-demand [676, 847, 1106, 1231, 1326, 1496]. On-line [1138, 1227]. one [623, 1235, 1376, 1461, 1509]. one-dimensional [1461, 1509]. one-time [623]. Online [705, 954, 1067, 1124, 1443, 1655]. Open [85, 419, 475, 663, 839, 1794]. OpenCAPWAP [1794]. operated [1133, 1319]. Operating [1159, 1866]. operation [286, 476, 548, 1241, 1333, 1523, 1790]. operational [1192, 1465]. operator [379]. opportunistic [518, 803, 1365, 1845]. opportunities [1527, 1658]. OPS [1471]. Optical [29, 31–35, 37–39, 185, 327, 492, 495, 498, 549, 561, 688, 692, 702, 733, 747, 766, 853, 923, 940, 962, 995, 1014, 1016, 1017, 1019, 1020, 1022, 1027, 1029, 1091, 1179, 1221, 1236, 1267, 1270, 1274, 1293, 1301, 1489, 1621, 1658, 1659, 1661, 1663, 1665, 1666, 1668, 1670, 1671, 1673– 1675, 1793, 1798, 1826, 1833, 1848, 1915, 1934]. Optics [1919]. Optimal [270, 298, 335, 356, 367, 376, 392, 448, 561, 577, 623, 785, 921, 1033, 1045, 1064, 1116, 1130, 1153, 1222, 1224, 1329, 1433, 1580, 1711, 1763, 1847, 1849, 1887, 1919, 1996]. optimisation [397, 927]. optimised [1681]. Optimising [118]. Optimization [124, 245, 368, 449, 637, 651, 660, 742, 854, 1027, 1066, 1084, 1156, 1295, 1331, 1577, 1583, 1587, 1590, 1671, 1680, 1718, 1754, 1791, 1809, 1818, 1909]. optimized [974, 1270, 1857, 1930]. Optimizing [343, 773, 1007, 1173, 1325, 1896]. options [1976]. orbit [1873]. orchestrating [1500]. order [192, 1938]. ordered [940]. ordering [604, 1806]. organisational [515]. organising [1894, 1907]. organization [634]. organized [1905]. organizing [282]. oriented [86, 155, 415, 467, 515, 825, 1078, 1236, 1771, 1773]. orienting [1712]. Origin [1191, 1225]. origins [911]. orthogonal [1674]. OSA [1881]. oscillations [1, 1226]. OSPF [410, 788, 924]. OSPF-based [788]. OSPF-OMP [410]. other [149, 1032]. outdoor [1788]. Outline [281]. output [188]. overbooking [1471]. Overcoming [1829]. overflow [805, 1352, 1415]. Overhead [801, 889, 1277, 1484]. overheads [1038, 1203]. overlap [1766]. Overlay [124, 234, 235, 683, 831, 934, 955, 1053– 1055, 1058–1060, 1087, 1130, 1152, 1177, 1196, 1239, 1241, 1288, 1363, 1411, 1466, 1565, 1643, 1698, 1732, 1741, 1747, 1758, 1781, 1791, 1820, 1847, 1887, 1914, 1945]. overlay-based [934, 1698]. overlays [925, 1237]. Overload [553, 1504, 1832]. overview [645, 1444, 1926]. OVSF [1692].

P2P [1036, 1037, 1100, 1173, 1177, 1197, 1325, 1454, 1509, 1598, 1620, 1631, 1758, 1812, 1815, 1818, 1842, 1904]. P2P-based [1818]. P2PSIP [1820]. P3P [935]. pacing [1551, 1639]. Packet [35, 38, 39, 141, 157, 192, 312, 321, 332, 334, 343, 391, 405, 429, 489, 517, 551, 561, 562, 575, 577, 594, 609, 660, 668, 675, 688, 742, 747, 776, 959, 961, 968, 972, 1006, 1049, 1050, 1077, 1081, 1092, 1121, 1128, 1133, 1160, 1174, 1219, 1225, 1236, 1269, 1275, 1286, 1293, 1297, 1393, 1399, 1465, 1536, 1546, 1556, 1638, 1705, 1793, 1807, 1808, 1833, 1839, 1840, 1885, 1933, 2000]. packet-based [660, 1705]. Packet-E-Model [1174]. packet-level [1133]. packet-over-WDM [35]. packet-pair [1128]. Packet-switched [192, 429, 609, 668, 747, 972, 1556, 1808, 1833]. packet-switching [1885]. Packets [137, 1566, 1675, 1967]. page [105]. pages [64, 66]. paging [973, 1170]. pair [1128, 1275]. pairings [1750]. Palm [1127]. paper [325]. Papers [196, 257, 319, 563, 591, 679, 1071]. PAPI [322]. paradigm [115, 263, 405, 725, 1350, 1689, 1769, 1902]. paradoxes [610]. Parallel [554, 1084, 1135, 1142, 1237, 1491, 1550, 1793, 1887]. parameter [2, 682, 913]. Parameters [121, 323, 788, 927, 1353, 1860]. parcels [1118]. Pareto [459, 1455, 1860]. Paris [776]. parity [658]. Parlay [213]. Parlay-based [213]. partial [637, 760, 981, 1706, 1751]. particle [1084]. partitioning [1328, 1336]. pass [1098]. Passive [34, 391, 903, 1029, 1306, 1425, 1720, 1915, 1934]. Passport [52]. password [1460]. Past [1351]. Patching [1268]. Path [6, 39, 73, 242, 332, 370, 392, 444, 460, 637, 831, 849, 1031, 1090, 1098, 1127, 1153, 1337, 1344, 1388, 1402, 1432, 1451, 1486, 1614, 1633, 1694, 1698, 1703, 1737, 1765, 1768, 1853, 1879, 1886, 1890, 1902]. Path-aware [831]. paths [352, 659, 784, 943, 976, 1284, 1742]. pattern [130, 218, 997]. patterns [992, 1098, 1452]. payload [1843]. payload-based [1843]. PCF [548]. PCN [1909]. PCN-based [1909]. PCS [194, 356, 612, 1243]. PDAs [62]. PDE [1769]. PDP [1205]. Peer [711, 724, 779, 846, 1011, 1035, 1038– 1042, 1060, 1105, 1155, 1217, 1246, 1298, 1348, 1355, 1407, 1487, 1527, 1584, 1604, 1641, 1655, 1677, 1768, 1813, 1814, 1941, 1971]. Peer-to-peer [724, 779, 1011, 1035, 1038, 1040– 1042, 1060, 1105, 1155, 1217, 1246, 1298, 1355, 1407, 1487, 1527, 1584, 1604, 1641, 1655, 1677, 1768, 1813, 1814, 1941, 1971]. peering [1741]. peers [1040]. PEN [199]. peninsula [1245]. people [203]. per-application [1914]. per-class [1280]. Per-flow [440]. per-session [332]. perceived [49, 1524]. perception [950]. Performance [33, 38, 50, 110, 116, 126, 169, 190, 219, 243, 246, 275, 280, 300, 318, 321, 334, 384, 388, 404, 408, 432, 447, 459, 489, 491, 530, 531, 535, 574, 575, 685, 691, 720, 734, 747, 782, 803, 822, 849, 938, 952, 969– 972, 974, 978, 979, 1012, 1019, 1020, 1025, 1028, 1030, 1051, 1077, 1136, 1137, 1144, 1197, 1211, 1221, 1270, 1274, 1288, 1293, 1299, 1317, 1324, 1333, 1340, 1361, 1378, 1380, 1410, 1429, 1456, 1458, 1478, 1482, 1498, 1499, 1545, 1558, 1619, 1724, 1747, 1789, 1807, 1819, 1860, 1865, 1867, 1869, 1878, 1936, 1953, 1973, 1998]. performance [115, 147, 170, 186, 263, 296, 302, 348, 358, 437, 450, 461, 783, 791, 806, 843, 850, 869, 879, 904, 957, 962, 964, 988, 1067, 1079, 1112, 1115, 1145, 1148, 1179, 1329, 1331, 1332, 1386, 1408, 1452, 1464, 1546, 1637, 1749, 1784, 1813, 1874, 1977]. perimeter [629]. periodic [722]. permit [1225]. Persistent [1, 1226, 1322, 1526].

Personal [156, 1994]. Personalised [1517, 1681]. personalization [1398]. personalized [60, 970]. perspective [382, 398, 953, 1193, 1906]. perspectives [569]. Pervasive [198, 202, 469]. Petri [1272]. Petri-Net [1272]. Phantom [16]. phase [1483, 1585]. phenomenon [1139, 1245]. phishing [1460]. Photo [740]. Photonic [1023]. Photonics [1660]. physical [1647, 1848]. physics [398, 1564]. picture [1721]. Piggybacking [1115]. PKI [320, 746]. place [44, 814]. Placement [133, 185, 561, 614, 925, 1641, 1682]. Planck [1366]. plane [685]. planetary [874]. planning [359, 558, 761, 1075, 1422, 1426, 1680, 1978]. platform [51, 584, 999, 1363, 1446, 1839, 1914]. platforms [155, 1628]. playing [352]. playout [1705, 1813]. PLR [1886]. PLR-based [1886]. plugins [348]. PLW [1098]. PMM [1030]. PNNI [123, 136]. point [134, 322, 355, 1332, 1401, 1465, 1866]. point-to-multipoint [355]. point-to-point [1465]. points [1138, 1968]. Poisson [459, 1860]. Poissonian [1275]. poker [344]. polices [734]. Policies [54, 141, 378, 741, 752, 771, 773, 1041, 1263, 1619, 1763]. Policy [392, 613, 753, 1200, 1735, 1781, 1900, 1979]. Policy-constrained [1900]. Policy-enabled [753]. Polish [327]. Pollen [203]. Pollution [1599, 1971]. Polynomial [859]. pool [619, 1988]. POOSL [531]. popularity [1117]. population [618]. port [137]. port-based [137]. Portable [214, 417, 977, 1279]. Portal [87]. portals [80]. positioning [823, 1752]. possession [1020]. posteriori [37]. Potential [172, 510, 616, 1150]. Power [31, 510, 511, 548, 622, 720, 725, 734, 833, 835, 1150, 1324, 1327, 1379, 1395, 1477, 1525, 1544, 1624, 1647, 1648, 1752, 1766, 1874, 1882, 1995]. Power-aware [511]. power-level [1752]. Power-saving [622, 734, 1624, 1766]. power/rate [1882]. Practical [166, 930, 1194, 1278, 1360, 1400, 1626, 1931, 1974]. practice [715]. Pragmatic [365, 586]. pRAM [114]. Pre [192, 549, 589, 680, 1231]. pre- [1231]. pre-contractual [589]. pre-encoded [680]. Pre-order [192]. pre-routed [549]. prebuffering [1920]. Precomputation [883]. predefined [1735]. Predicate [1250]. Predicate/Transition [1250]. Predicate/Transition-nets [1250]. predictability [403, 1474]. Predicting [262, 471, 1272, 1366]. prediction [73, 596, 1653, 1792, 1962]. Predictive [182, 364, 528, 550, 597, 1230, 1282]. Preemptive [877]. preference [877, 935]. preferences [918]. Preferential [539, 1485]. Prefetching [92, 409, 1458]. Prefix [93, 787, 880, 1269, 1438]. prefix-caching [93]. Prefix-preserving [787]. prefixes [1079]. Premium [825]. prepaid [243]. presence [1187, 1380, 1558, 1685, 1802, 1970]. present [1351]. Presenting [97]. preserving [787, 1739, 1777]. Preventing [1226]. prevention [1352, 1603]. priced [1433]. prices [261]. Pricing [19, 45, 101, 259, 560, 603, 705, 771, 774– 778, 882, 1001, 1068, 1180, 1334, 1440, 1630, 1681, 1773]. primary [1866]. prime [977]. Principles [1146]. priori [37]. Priorities [521]. prioritization [1282]. prioritized [782]. Priority [10, 188, 472, 681, 744, 1795, 1868, 1871, 1911, 1980]. Priority-based [1911]. Privacy [149, 1519, 1744, 1777, 1889]. Privacy-aware [1889]. Private [236, 340, 596, 674, 1075, 1373, 1577, 1744, 1754]. Proactive [92, 535, 648, 1988]. proactively [692]. Probabilistic [270, 449, 472, 749, 1011, 1037, 1286, 1384, 1799, 1927, 2000]. probabilities [581, 733, 995]. Probability [464, 522, 612, 901, 1415, 1461, 1929]. probability-guaranteed [901]. probable [460]. Probe [174, 1470]. Probe-based [1470]. probes [576]. probing [1128]. problem [132, 427, 558, 715, 938, 1476, 1591, 1648, 1683, 1760]. problems [149, 361, 1104, 1829]. procedures [384, 732]. process [14, 152, 228, 459, 540, 910, 1798, 1998]. processes [462, 1900]. Processing [57, 383, 385, 526, 802, 1160, 1717, 1755]. processor [53, 461, 462, 525, 529–531, 941, 1313]. processors [524, 527, 528, 1974]. product [43, 1345]. producting [1840]. production [491]. products [797]. PROFIBUS [299, 1417]. profiles [605, 832]. Profiling [1842]. profitable [1241]. Program [1262]. Programmable [230, 233, 346, 348, 349, 505, 754, 1103, 1165, 1974]. programming [424, 588, 1699, 1905]. programs [54]. progress [325]. prohibition [1303]. project [216, 396]. promising [1928]. Prompt [637]. prone [793]. propagation [1441]. Properties [42, 292, 937, 1770, 1823, 1853]. proportion [1428]. proportional [659, 1280, 1356]. proposal [863, 1469]. proposed [839]. proprietary [1788]. Protecting [934, 1018, 1310, 1312, 1394, 1468, 1898]. protection [42, 620, 629, 895, 974, 997, 1016, 1026, 1309, 1313, 1519, 1614, 1633, 1667, 1668, 1694, 1890, 1960]. Protocol [3, 13, 52, 93, 103, 134, 278, 338, 361, 371, 391, 442, 469, 474, 477, 490, 493, 516, 522, 537, 576, 646, 649, 669, 673, 814, 888, 951, 960, 973, 994, 1003, 1069, 1123, 1197, 1199, 1205, 1250, 1279, 1283, 1330, 1343, 1423, 1460, 1463, 1478, 1480, 1494, 1508, 1510, 1549, 1561, 1624, 1649, 1655, 1710, 1726, 1750, 1803, 1825, 1850, 1916, 1943]. protocol-adaptive [576]. Protocols [5, 20, 131, 167, 254, 287, 360, 373, 426, 476, 498, 501, 547, 595, 622, 639, 647, 668, 687, 764, 833, 867, 869, 918, 939, 943, 990, 1012, 1031, 1273, 1339, 1340, 1372, 1429, 1442, 1452, 1602, 1623, 1639, 1673, 1718, 1881, 1906]. prototype [199, 874]. provide [193, 877, 1041, 1428]. provider [1079]. provider-dependent [1079]. providers [322]. Providing [102, 332, 405, 547, 669, 772, 844, 1244, 1301]. Provision [231, 386, 781]. Provisioning [248, 605, 683, 705, 774, 922, 955, 970, 1010, 1047, 1147, 1337, 1351, 1392, 1539, 1554, 1628, 1795, 1981]. ProWGen [377]. proxies [1103, 1131]. proximity [873, 1631]. proxy [93, 136, 377, 691, 925, 1145, 1230]. proxy-based [925]. Proxy-PAR [136]. PSTN [213]. public [1042]. publish [1944]. publish-subscribe [1944]. Publisher [1688]. publishing [60]. purely [704]. purposes [989, 1492]. push [970, 1030]. push-mode-multicast [1030]. Pushing [422].

Q [313, 546, 1080, 1320]. Q-M [1320].

Q-MEHROM [1080]. Q-Soft [546]. QNS [1875]. QoE [1582, 1896]. QoS [181, 193, 236, 242, 248, 279, 295, 304, 326, 405, 426, 440, 450, 452, 494, 519, 546, 550, 553, 556, 606, 624, 657, 665, 673, 679, 688, 775, 777, 788, 826, 840, 845, 883, 891, 898, 923, 974, 984, 999, 1010, 1012, 1047, 1067, 1071, 1076, 1078, 1090, 1113, 1139, 1181, 1213, 1227, 1231, 1236, 1244, 1252, 1297, 1301, 1336, 1351, 1354, 1382, 1392, 1423, 1432–1434, 1436, 1439, 1449, 1486, 1554, 1556, 1563, 1577, 1578, 1582, 1595, 1608, 1617, 1630, 1672, 1710, 1771, 1774, 1859, 1959, 1980]. QoS-aware [450, 553, 826, 923, 974, 1047, 1434, 1595, 1859]. QoS-driven [1608, 1771]. QoS-enabled [777]. QoS-IP [679, 1071]. QoS-MAC [1710]. QoS-relevant [1617]. QoS-sensitive [673]. QoS-sensitivity [1563]. QoSbox [1210]. QoSockets [181]. QUAD [1922]. Quadrant [1922]. Quadrant-based [1922]. quadratic [1695]. Qualitative [1201]. Quality [49, 128, 231, 276, 303, 314, 328, 332, 339, 343, 439, 781, 834, 842, 844, 1138, 1147, 1173, 1174, 1272, 1408, 1424, 1451, 1524, 1600, 1615, 1674, 1773, 1795]. Quality-of-Service [1451]. Quantile [1400]. quantiles [464]. quantitative [633, 1210]. quasi [847, 1539]. quasi-geostationary [847]. quasi-static [1539]. queries [630, 1717, 1802]. Query [55, 57, 76, 416, 1168, 1631]. Querying [83, 140, 587, 1065]. Queue [240, 459, 462, 640, 662, 677, 804, 861, 1144, 1179, 1233, 1366, 1534, 1860, 1868, 1873]. queue-based [1144]. queued [311, 575, 1049, 1855]. Queueing [10, 137, 430, 460, 852, 914, 1009, 1129, 1320, 1428, 1450, 1871]. queues [656, 1281, 1341]. Queuing [393, 663, 1073, 1774, 1808]. Quick [1608]. quorum [1766]. quorumcast [657]. quota [681]. quota-based [681].

radii [859]. Radio [89, 157, 343, 489, 686, 826, 873, 972, 1141, 1480, 1542, 1590, 1591, 1593, 1594, 1613, 1679, 1726, 1866]. radios [1936]. Rainbow [852]. ramifications [1095]. random [470, 1242, 1325, 1390, 1544, 1617, 1946]. randomization [1445]. randomized [242, 1768, 1955]. randomizing [1032]. randomly [1853]. range [185, 431, 455, 461, 463, 672, 906, 911, 912, 1631, 1860, 1912, 1976]. ranges [516, 1853]. rank [78, 1117]. ranked [1884]. rankings [1096]. rapid [214, 415]. Rappel [1944]. Rate [2, 112, 119, 120, 127, 190, 191, 260, 298, 300, 429, 441, 468, 480, 520, 597, 621, 656, 662, 697, 738, 765, 937, 1006, 1151, 1172, 1207, 1237, 1249, 1270, 1294, 1320, 1330, 1331, 1344, 1354, 1543, 1587, 1618, 1823, 1850, 1877, 1882, 1911, 1980, 1994]. rate-based [119, 697]. Rate-distortion [1207]. Rate-lifetime [1543]. Rates [125, 968]. RBF [902]. RCN [899]. RCS [999, 1704]. RDF [71, 418–420, 1100]. re [445, 1027, 1251]. re-allocation [445]. re-keying [1251]. re-optimization [1027]. reachability [900]. reachable [626]. reacting [1101]. reactive [422, 1923, 1988]. reading [1606]. Real [121, 184, 245, 302, 467, 477, 804, 897, 902, 952, 992, 1058, 1078, 1163, 1378, 1417, 1497, 1515, 1574, 1592, 1596, 1686, 1705, 1773, 1808, 1966]. real-life [1163]. Real-time [121, 184, 302, 477, 804, 897, 902, 952, 992, 1058, 1078, 1378, 1417, 1497, 1515, 1574, 1592, 1596, 1705, 1773, 1808, 1966]. Realistic [1136, 1253, 1441, 1807]. reality [198, 753]. Realizing [950]. RealSecure [388]. reasoning [1799]. Receiver [467, 786, 862, 1401, 1852]. receiver-based [862]. Receiver-centric [1401]. receiver-driven [786]. receiver-oriented [467]. receivers [1080]. recommendation [1951]. recommending [95]. reconfigurable [54, 348, 687]. reconfiguration [1016]. reconstruction [1317]. records [535]. recovery [706, 1083, 1197, 1267, 1317, 1322, 1466, 1623, 1633, 1640, 1899]. recurrence [1836]. recurrence-based [1836]. RED [562, 641, 677, 693, 983, 1088, 1233, 1503]. redirection [996]. redistribution [1783]. Redland [418]. reduce [539, 559, 1428, 1929]. reduced [1114]. reducible [1260]. Reducing [409, 552, 720, 1181, 1248, 1277]. reduction [494, 519, 1600]. redundancy [1830]. redundant [1471]. REFACING [1725]. reflection [1226]. Refreshment [378]. regarding [1675]. regimes [1534]. region [1331]. registration [1443, 1583, 1636]. regression [1792]. regular [335, 1693]. reinforcement [1929]. Reinforcing [626]. rekeying [706]. related [1115]. relations [96, 98]. Relationship [621, 1735]. relationships [1744]. relative [1922]. relay [1931]. relaying [1482]. Relevance [1201]. relevant [1617]. Reliability [264, 811, 890, 1038, 1669, 1751, 1897, 1932]. Reliable [51, 363, 373, 613, 658, 759, 812, 949, 994, 1083, 1296, 1543, 1702, 1703, 1883]. relocation [1727]. Remote [930, 980, 1105, 1374, 1800]. remotely [1230]. removal [1124]. rendezvous [1459]. rendezvous-based [1459]. renegotiating [125]. Reno [318]. Reordered [1967]. Repacking [836]. repairing [1960]. replacement [741]. Replacing [817]. replanning [428]. replica [712]. replicated [1232]. replication [356, 712, 1862]. reply [126]. reports [416]. repositioning [979]. repositories [422]. repository [66, 165]. representation [5, 99, 417, 420, 423, 424]. representations [1310]. representative [700, 1922]. Representing [713]. repudiation [1375]. Reputation [1035–1038, 1041]. Reputation-based [1041]. reputations [105]. request [126, 275]. request-reply [126]. requests [549, 615, 1326, 1774]. requirement [1258]. Requirements [242, 292, 525, 560, 566, 839, 990, 1516]. reroute [558, 1349]. Rerouting [1195, 1661]. rescheduling [492]. rescue [1127]. Research [436, 608, 751, 1217, 1602, 1702]. reservation [13, 294, 364, 368, 619, 920, 998, 1069, 1111, 1282, 1471]. reservation-pool [619]. reservations [850, 1080, 1665, 1743, 1845]. residential [1436, 1992]. residues [1695]. resilience [1403]. resiliency [1200]. Resilient [959, 1185, 1413, 1487, 1604, 1632, 1729, 1909, 1946]. resistant [150, 1087]. resolution [493, 630, 1025, 1257, 1494]. Resource [30, 97, 125, 230, 297, 316, 343, 367, 385, 389, 407, 538, 685, 767, 775, 781, 826, 920, 1064, 1080, 1107, 1209, 1246, 1651, 1677, 1696, 1707, 1754, 1791, 1795, 1817, 1859, 1907, 1988, 1989]. resources [78, 614, 786, 816, 1146, 1165, 1382, 1986]. response [152, 262, 874, 904, 1675]. responsive [1231]. restorable [1529]. Restoration [370, 784, 962, 1666, 1830]. restricted [1868]. restrictive [1206]. restructuring [1480]. results [1096, 1201, 1564, 1704]. retransmit [1623]. retrial [1873]. Retrieval [667, 1509, 1819]. retrospective [1099]. reuse [218, 1978]. reuse-driven [218]. Revealing [361]. revenue [1073]. revenue-based [1073]. revenues [1086, 1433]. review [496, 1370]. reviewers [207, 362, 541, 727, 864, 1308]. Revisiting [1143]. revolution [583]. RFID [1373, 1606, 1652, 1695]. rich [1259]. riding [1584]. right [324, 1041]. rigorous [1500]. ring [35, 498, 504, 959, 982, 1469, 1668, 1672]. rings [36, 37, 503, 549, 673, 783, 1021]. Risk [895, 1180, 1384, 1696, 1700, 1879]. Risk-based [1384]. risk-disjoint [895]. Risks [52]. RL [1994]. RL-based [1994]. road [250, 581]. roaming [1646]. robin [192, 312]. robot [1799]. Robust [56, 206, 926, 997, 1175, 1217, 1278, 1304, 1540, 1763, 1802, 1846]. robustness [1089, 1109]. Role [138, 1923, 1937]. root [899, 1656]. RoQ [1600]. Round [192, 312, 785, 1924]. round-trip [785]. Route [1, 375, 1226, 1326, 1567, 1671, 1818, 1900]. route-demand [375]. routed [549, 1491]. router [311, 347, 1030, 1345, 1600, 1663]. router-based [1600]. routers [348, 1210, 1335, 1790, 1974]. routes [626, 1359, 1781]. Routing [3, 31, 39, 122, 123, 127, 135, 136, 139, 167, 215, 305, 338, 375, 387, 426, 494, 495, 519, 554, 609, 610, 632, 646, 659, 665, 714, 736, 784, 801, 824, 832, 834, 876, 883, 898, 923, 963, 984, 1003, 1005, 1017, 1019, 1022, 1031, 1057, 1067, 1090, 1106, 1114, 1130, 1147, 1156, 1181, 1187, 1191, 1198, 1200, 1214, 1227, 1251, 1326, 1336, 1337, 1364, 1387, 1409, 1424, 1432, 1452, 1459, 1476, 1489, 1496, 1507, 1549, 1562, 1575, 1602, 1627, 1634, 1662, 1696, 1718, 1743, 1761, 1770, 1777, 1783, 1790, 1791, 1848, 1849, 1890, 1909, 1919, 1923, 1982, 1993]. routing [1, 30, 76, 303, 410, 504, 534, 550, 579, 611, 620, 624, 636, 657, 660, 703, 796, 1054, 1223, 1231, 1305, 1323, 1410, 1451, 1538, 1587, 1768, 1889, 1902, 1912, 1918, 1968]. RPR [1179]. RQL [587]. RSVP [315, 438]. RSVP-aware [438]. RT [262, 835]. RT-MuPAC [835]. RTID [145]. rule [1905]. rule-based [1905]. rules [397, 422, 1100]. Running [90]. s [108, 109, 173, 445, 891, 1873]. S-SLSP [445]. safety [992, 1323]. safety-critical [992]. SafetyNet [233]. SALSA [74]. SAM [1642]. SAM-MAC [1642]. samples [1717]. Sampling [67, 1189, 1400, 1492, 1684, 1721, 1946]. sand [810]. Sandpiper [1989]. satellite [158, 293, 357, 382, 383, 385– 387, 609, 646, 793, 847, 890, 999, 1138, 1238, 1496, 1558, 1704, 1888]. satellite/terrestrial [158]. SATIP6 [999]. Satisfying [549]. Saturated [1419]. saturation [1545]. save [1379]. saving [510, 622, 734, 1379, 1624, 1766, 1995]. Scalability [1235, 1459, 1745, 1839]. Scalable [51, 309, 363, 372, 426, 594, 631, 636, 641, 672, 699, 704, 788, 845, 920, 978, 1010, 1135, 1155, 1173, 1190, 1195, 1251, 1373, 1448, 1481, 1581, 1616, 1789, 1840, 1948, 1981]. scale [206, 453, 500, 630, 649, 954, 1006, 1134, 1135, 1403, 1420, 1483, 1540, 1576, 1626, 1652, 1739, 1780, 1932, 1982]. Scaled [744]. scaling [680, 908, 1048, 1336]. scatternet [719]. scenarios [477, 1683, 1686]. Schedule [692, 1774]. scheduler [1378]. schedulers [312, 313, 1428, 1795, 1855]. Scheduling [18, 188, 192, 342, 343, 376, 405, 472, 518, 528, 536, 702, 722, 766, 782, 803, 897, 940, 941, 961, 1029, 1049, 1143, 1150, 1183, 1271, 1365, 1382, 1443, 1491, 1547, 1574, 1580, 1601, 1609, 1619, 1679, 1682, 1705, 1711, 1732, 1743, 1756, 1793, 1813, 1882, 1891, 1915, 1931, 1955, 1987, 1994]. schema [414, 420, 423, 1509, 1838]. schema-based [1509]. scheme [16, 18, 126, 190, 344, 369, 429, 538, 611, 737, 745, 759, 765, 776, 787, 805, 815, 826, 862, 882, 937, 974, 980, 998, 1025, 1051, 1064, 1170, 1179, 1194, 1195, 1214, 1220, 1233, 1278, 1302, 1358, 1404, 1477, 1497, 1523, 1639, 1640, 1642, 1695, 1715, 1764, 1833, 1859, 1890, 1891, 1920, 1924, 1952, 2000]. schemes [122, 137, 193, 279, 310, 370, 393, 512, 660, 695, 697, 778, 859, 927, 978, 1027, 1067, 1171, 1332, 1349, 1360, 1416, 1623, 1661, 1819]. SCI [302]. science [792, 797, 798]. sciences [790]. scientific [173]. SCM [33, 391]. scope [1205]. SCPS [1624]. SCTP [1142]. SDH [124]. SDL [216–222, 565, 569, 570, 696, 992]. SDL-2000 [222, 569]. SDL-based [217]. Seamless [446, 1383]. Search [57, 70, 76, 77, 81, 140, 400, 401, 427, 494, 1092, 1095, 1096, 1155, 1217, 1245, 1325, 1348, 1407, 1426, 1438, 1648, 1873]. searches [1269]. searching [62, 1011]. SearchPad [81]. second [41, 679, 911]. secondary [1592, 1866]. secret [1783]. sectorized [1645]. Secure [180, 493, 629, 919, 1109, 1118, 1145, 1285, 1376, 1389, 1406, 1448, 1493, 1504, 1523, 1628, 1655, 1753, 1805, 1926, 1934]. Securely [933]. Securing [351, 1520]. Security [54, 166, 167, 232, 291, 340, 432, 638, 639, 724, 746, 787, 929, 932, 1035, 1090, 1145, 1156, 1235, 1251, 1279, 1374, 1520, 1527, 1725, 1738, 1935]. security-aware [1090]. seeing [1721]. segment [1667]. segmentation [664, 686]. Selected [177, 319, 679, 1071]. selecting [1229]. selection [147, 221, 279, 659, 812, 954, 1153, 1236, 1332, 1556, 1608, 1698, 1772, 1866, 1887]. selections [645]. selective [632, 1170]. Self [113, 282, 370, 631, 763, 911, 937, 1018, 1131, 1189, 1304, 1309, 1314, 1352, 1467, 1514, 1581, 1624, 1628, 1651, 1712, 1809, 1832, 1894, 1897, 1898, 1903–1905, 1907, 1910, 1960, 2000]. Self-* [1832]. Self-adaptive [1903, 2000]. self-adjusting [937]. Self-adjustment [1651]. self-aware [1314]. self-clocking [1467]. self-configuration [631, 1809]. self-configuring [1581, 1624]. self-healing [370, 631, 1898]. self-keying [1304]. self-learning [1832]. Self-managed [1628, 1904]. Self-managing [1514, 1897]. self-optimization [1809]. self-organising [1894, 1907]. self-organized [1905]. self-organizing [282]. Self-orienting [1712]. Self-prevention [1352]. self-protecting [1898]. self-protection [1309]. self-repairing [1960]. self-sharing [1018]. self-similar [113, 1189]. Self-similarity [763, 911, 1131]. self-tuning [1910]. selfish [1948]. Selfishness [1648]. Semantic [80, 583–587, 589, 1508, 1509]. semantics [569, 1261, 1348]. SemTalk [586]. sender [636, 862]. senders [1080]. sending [1032, 1344]. sense [1475]. sensing [200, 1647, 1942]. sensitive [91, 673, 834, 836, 1171, 1249, 1305, 1573, 1601, 1640, 1711]. Sensitivity [1328, 1563]. Sensor [354, 593, 628–634, 642, 731, 801, 810, 812, 814, 921, 979, 1065, 1085, 1182, 1183, 1185, 1215, 1222, 1229, 1251, 1291, 1292, 1295, 1303, 1324, 1362, 1388, 1391, 1430, 1443, 1461, 1514, 1525, 1543, 1573, 1575, 1576, 1580, 1682, 1683, 1691, 1699, 1709, 1710, 1712–1715, 1723, 1727, 1728, 1748, 1752, 1753, 1755, 1777, 1807, 1822, 1849, 1850, 1861, 1882, 1889, 1905, 1911, 1916, 1922, 1923, 1926, 1930, 1942, 1955, 1982, 1996–1999]. sensors [655, 1853]. separation [517, 1275, 1775]. sequence [558, 990, 1106, 1751]. sequences [572, 1098]. Sequential [618]. series [1617]. Serv [1010, 1827]. server [49, 336, 450, 621, 645, 748, 851, 917, 954, 1143, 1224, 1253, 1749, 1973]. server-side [917]. servers [15, 133, 461, 480, 553, 934, 1230, 1268, 1294]. Service [119, 134, 140, 150, 156, 157, 162, 212, 213, 229, 231, 236, 239, 243, 263, 276, 285, 294, 297, 303, 314, 328, 332, 334, 339, 365, 371, 405, 415, 417, 437, 439, 489, 529, 537, 560, 588, 604, 605, 643, 704, 726, 754, 772, 774, 781, 825, 834, 842, 844, 881, 902, 934, 961, 974, 1051, 1089, 1138, 1147, 1163, 1166, 1180, 1199, 1210, 1236, 1239, 1241, 1250, 1256, 1280, 1305, 1314, 1403, 1408, 1424, 1449–1451, 1490, 1518, 1542, 1608, 1646, 1665, 1674, 1677, 1706, 1713, 1771, 1791, 1795, 1871, 1899, 1911, 1920, 1927]. service-aware [1665]. service-dependent [1051]. service-differentiated [726]. service-driven [529]. service-oriented [415, 1771]. Services [12, 19, 63, 90, 120, 159, 170, 206, 211, 214, 226, 231, 260, 298, 305, 329, 337, 350, 365, 369, 386, 405, 422, 440, 441, 470, 472, 590, 602, 603, 605, 658, 682–684, 757, 765, 782, 809, 879, 892, 897, 922, 970, 971, 1001, 1012, 1029, 1068, 1162, 1164, 1232, 1264, 1265, 1316, 1405, 1433, 1447, 1459, 1470, 1484, 1500, 1513, 1517, 1557, 1773, 1820, 1831, 1898, 1981]. Serving [276, 933]. Session [332, 371, 389, 951, 1208, 1232, 1903]. session-based [389]. sessions [505, 982, 1018, 1130, 1354, 1423]. set [572]. sets [876, 1065, 1287]. setup [392, 1284]. shaping [115, 191, 1892]. share [1274, 1941]. shared [130, 334, 419, 1016, 1633, 1667, 1696, 1793, 1879, 1986]. shared-per-wavelength [1793]. shared-segment [1667]. sharing [38, 86, 379, 461, 462, 471, 550, 643, 735, 846, 941, 1018, 1021, 1026, 1195, 1331, 1354, 1381, 1467, 1527, 1536, 1594, 1595, 1679, 1815, 1817, 1859]. SherLOCK [1637]. Short [539, 663, 693, 1046, 1545, 1612]. short-lived [663]. Short-term [1545]. Shortest [1765, 1768, 1902]. Shortest-path [1765, 1768]. side [917, 1458, 1788]. side-effects [1788]. signal [802, 1755]. Signaling [10, 128, 265, 360, 387, 559, 701, 822, 1209, 1213, 1528, 1787, 1965]. signalling [1643]. signature [1655]. signatures [301, 623, 851, 1750]. signon [52]. SIL [1155]. silk [1159]. SilkRoute [98]. similar [113, 1189]. similarity [763, 911, 1131]. Simple [16, 43, 278, 305, 393, 520, 917, 1233, 1302, 1917]. simulated [794]. simulating [1359]. Simulation [17, 21, 266, 313, 377, 410, 447, 498, 530, 590, 693, 960, 1054, 1126, 1132, 1133, 1135, 1137, 1319, 1441, 1492, 1550, 1686, 1734, 1739, 1778, 1862, 1907]. simulation-based [17, 266, 1137]. simulations [1134]. simulator [557, 1534]. simulators [451]. simultaneous [650, 1020]. simultaneously [1428]. SingBroad [1891]. single [52, 802, 1021, 1098, 1366, 1460, 1469, 1891]. single-hub [1021, 1469]. Single-pass [1098]. singleton [880]. Sink [979]. SIP [718, 971, 1265, 1394, 1673]. SIP-based [971, 1394]. SIP-enabled [1673]. site [72, 104]. sites [58]. size [664, 677, 1077, 1108]. sizeable [262]. sizes [675, 1456, 1766]. sketch [1762]. skew [1124]. Skype [1582, 1787]. SLA [178]. SLAs [443]. slave [720, 1386]. slave-to-slave [720]. slice [1863]. slice-based [1863]. sliding [785]. sliding-window [785]. Slot [1171]. slotted [36, 37, 760, 1833]. SLSP [445]. Small [545, 806, 908, 977, 1048, 1177, 1961]. Small-buffer [1961]. Small-time [908]. Small-world [806, 1048, 1177]. Smart [248–250, 253–256, 817, 980, 1368– 1371, 1375, 1376, 1992]. Smooth [247, 737]. smoothness [738, 937]. SMP [1628]. SMS [578]. snapshot [208]. snooping [665]. Snow [1159]. social [1037, 1512, 1744]. socket [1352]. sockets [181]. soft [345, 546, 1123]. soft-handoff [1123]. softswitch [559]. software [21, 45, 546, 570, 1258, 1790, 1869]. Sojourn [9]. Sojourn-time [9]. solution [248, 558, 1204, 1481, 1547, 1772, 1995]. solutions [41, 122, 839, 1401, 1444, 1448, 1788]. solve [1546]. some [1006]. Someone [1620]. SONET [503, 783, 1153]. SONET/WDM [503]. Sorting [1950, 1967]. source [59, 135, 683, 802, 909, 1175, 1225, 1638, 1794, 1873]. sources [118]. SPA [1877]. Space [384, 494, 792, 794–796, 867– 871, 874, 1092, 1445, 1478, 1509, 1585, 1856]. Space-based [871, 874]. space-terrestrial [869]. Spacecraft [872]. SPAD [1486]. spam [1396, 1841]. spanning [1761]. Spare [445, 1016, 1539]. sparse [495, 1501]. spatial [1838]. Spatio [731, 1987]. Spatio-temporal [731, 1987]. spawning [230]. Special [1094, 1553, 1589, 1658, 1709, 1786, 1865]. Specific [79, 683, 1536, 1579, 1914]. specification [252, 299, 588, 764]. specifications [159, 221, 244, 567, 595, 696, 990, 1250]. specified [367]. specify [452]. specifying [27]. spectral [516, 1800]. Spectrum [690, 813, 913, 1141, 1381, 1592, 1594– 1596, 1679]. speech [1704]. Speed [300, 324, 648, 836, 901, 944, 968, 1160, 1340– 1343, 1346, 1616, 1684, 1697, 1790, 1839, 1934]. speed-sensitive [836]. speedup [1220]. splitting [495, 933]. spot [205]. spots [509, 1794]. SPP [1460]. spread [51, 690, 1803]. spurious [691, 1192]. Squeal [55]. SRAMT [862]. SRD [117]. SRRG [1696]. SS7 [265]. SS7-signaling [265]. SSL [933]. SSVP [1497]. Stability [13, 641, 1502, 1503, 1627, 1653, 1804, 1838]. stability-based [1653]. stabilization [1223]. stabilizing [656]. stable [677, 978, 1144, 1872]. Stacked [11]. Stacked-Banyan [11]. stage [1536]. standard [421, 867]. Standards [256]. star [391]. start [241, 1609]. start-up [241]. StarTrek [198]. State [350, 430, 608, 694, 766, 824, 849, 981, 1181, 1207, 1322, 1502, 1568, 1760, 1990]. State-dependent [430]. state-of-the-art [608, 694]. Stateless [440, 483, 1493, 1831]. states [961]. static [122, 1212, 1462, 1539, 1606, 1795]. station [795, 859, 1547]. stationary [1242]. stations [266, 511]. Statistical [182, 263, 398, 735, 878, 897, 902, 1528, 1622, 1782, 1800]. Status [569]. Steiner [1736]. Steiner-tree [1736]. step [639, 1270, 1372]. Steps [491]. Stepwise [1941]. still [1357]. Stochastic [74, 261, 366, 399, 1128, 1336, 1512, 1806, 1867, 1875]. stochastically [1925]. storage [854]. stored [917, 1230, 1252]. storm [534]. strategies [37, 519, 560, 712, 785, 1001, 1119, 1384, 1692]. strategy [187, 632, 952, 1017, 1504, 1690, 1755]. stratified [1721]. Stream [34, 306, 550, 1064]. stream-based [550]. streaming [12, 86, 325, 676, 680, 847, 917, 1098, 1224, 1252, 1253, 1330, 1481, 1497, 1524, 1557, 1601, 1698, 1789, 1812, 1813, 1971, 1981]. streamline [1671]. streams [93, 126, 427, 1099, 1173, 1331, 1611, 1618, 1970]. street [561]. strict [1674]. Strictly [503]. striping [650]. structure [68, 74, 457, 597, 1097, 1762]. Structured [55, 1197, 1300]. structures [1053, 1563]. Structuring [1196]. studies [797, 1137]. Study [17, 161, 209, 307, 388, 410, 447, 498, 664, 734, 736, 745, 756, 778, 921, 962, 1008, 1020, 1054, 1059, 1328, 1353, 1484, 1524, 1526, 1544, 1545, 1724, 1745, 1815, 1953]. studying [1585]. Sub [1017, 1633]. Sub-graph [1017]. sub-path [1633]. subcarrier [33]. SubCast [1982]. subflows [1142]. Subject [225, 242, 273, 331, 381, 434, 485, 487, 543, 600, 653, 708, 769, 828, 885, 946, 1004, 1234]. Submission [165]. subnets [1190]. subnetworks [1872]. subscribe [1944]. subscriber [1122]. subscription [1681]. subsystem [821]. successful [509]. succinct [713]. suitability [210]. suite [20, 647]. sunshine [1762]. supplemental [960]. supply [1817]. Support [81, 288, 291, 314, 329, 341, 355, 440, 443, 534, 546, 551, 589, 701, 792, 795, 892, 1029, 1080, 1122, 1171, 1236, 1424, 1487, 1574, 1630, 1646, 1649, 1672, 1743, 1792, 1809, 1855, 1897, 1956]. Supporting [54, 1271, 1335, 1383, 1417, 1511, 1601, 1713, 1815]. surface [873]. surrogates [417]. surveillance [1183, 1807]. Survey [354, 735, 858, 957, 1033, 1062, 1141, 1217, 1291, 1349, 1369, 1416, 1418, 1564, 1623, 1677, 1691, 1748, 1848, 1881, 1928, 1959]. survivability [1826, 1870]. survivable [1087, 1633]. survive [1341]. sweet [205]. SweetBait [1311]. switch [11, 38, 116, 332, 575, 1281, 1491, 1793]. switched [188, 192, 388, 392, 429, 432, 609, 668, 692, 747, 784, 972, 1020, 1221, 1267, 1301, 1556, 1649, 1671, 1808, 1833]. switches [130, 188, 618, 688, 695, 781, 1049, 1293, 1855]. switching [39, 311, 442, 453, 492, 498, 561, 742, 766, 853, 940, 1015, 1091, 1148, 1236, 1270, 1274, 1349, 1660, 1673, 1798, 1856, 1885]. Sybil [1997]. Symbiotic [1249]. symbol [1376]. symmetric [357]. Synchronization [512, 762, 921]. synchronous [24]. synergy [1553]. synthesis [1549, 1972]. synthetic [377]. System [10, 14, 79, 86, 101, 163, 164, 171, 204, 214, 290, 322, 325, 342, 358, 432, 565, 578, 638, 690, 736, 791, 805, 919, 927, 934, 972, 983, 1062, 1135, 1225, 1264, 1276, 1303, 1313, 1458, 1591, 1607, 1724, 1812, 1828, 1843, 1860, 1905, 1982]. systematic [721, 1136]. Systems [22, 29, 33, 143, 161, 172, 177, 189, 206, 219, 269, 284, 323, 343, 345, 376, 387, 388, 460, 475, 512, 515, 667, 669, 690, 749, 755, 808, 815, 821, 847, 988, 992, 1035, 1036, 1042, 1051, 1104, 1105, 1159, 1216, 1258, 1259, 1299, 1315, 1316, 1323, 1328, 1353, 1360, 1377, 1446, 1479, 1496, 1503, 1509, 1599, 1614, 1652, 1660, 1692, 1695, 1704, 1738, 1766, 1768, 1769, 1832, 1868, 1888, 1894, 1898, 1932, 1944, 1947, 1971, 1978].

T [565, 892, 1664, 1945]. T-Man [1945]. tâtonnement [778]. table [714]. tables [736, 958]. Tabu [427, 1220]. tabu-search [427]. Tackling [1103, 1420]. tactical [194, 815, 818, 1862]. tag [1373, 1606]. tailed [462, 464]. tailored [97, 1705]. tandem [733, 1281]. TAP [584]. target [810, 1183, 1683]. targeted [100]. Targeting [677, 1841]. targets [270, 1714]. tariffing [603]. TARP [1480, 1494]. task [1065]. taxonomies [1507]. Taxonomy [1036, 1258, 1623]. TCAMs [1219]. TCDMA [342]. TCP [17, 69, 170, 241, 296, 306, 318, 363, 369, 374, 390, 394, 404, 437, 451, 456, 457, 467, 471, 501, 551, 574, 640, 641, 656, 663, 671, 691, 693, 725, 747, 763, 786, 830, 843, 862, 896, 907, 917, 964, 983, 1032, 1081, 1088, 1133, 1145, 1151, 1179, 1192, 1242, 1335, 1340, 1357, 1397, 1405, 1421, 1442, 1456, 1464, 1467, 1474, 1503, 1551, 1552, 1569, 1656, 1697, 1720, 1749, 1778, 1878, 1917, 1956, 1962, 1969, 1977, 1983]. TCP-aware [551]. TCP-friendliness [1503]. TCP-friendly [374, 501]. TCP-like [363]. TCP-operated [1133]. TCP-Real [467]. TCP/AQM [1456]. TCP/IP [69, 451, 907, 1969]. TCP/RED [641]. TDD [342]. TDD/TCDMA [342]. TDM [536]. TDM/WDM [536]. TDMA [889, 1645, 1822, 1882, 1892, 1936]. TDMA-based [1645]. TDMA/CDMA [1882]. technical [186]. technique [1236, 1600]. Techniques [219, 220, 444, 449, 972, 1028, 1391, 1444, 1478, 1525, 1535]. techno [1663]. techno-economic [1663]. technological [1444]. technologies [250, 421, 798, 808, 872, 1418, 1660]. technology [161, 162, 232, 466, 868, 1113, 1184, 1654, 1928]. telecom [174, 261]. telecommunication [159, 475, 756, 1553]. Telecommunications [22, 358, 987, 1447]. telecooperation [289]. telephony [138, 212, 214, 215, 559, 754, 1209]. teleteaching [172]. template [1260]. template-based [1260]. temporal [595, 597, 731, 1756, 1838, 1987]. Temporally [373]. TERENA [319]. term [64, 1545, 1612]. terminal [155]. terminals [1380, 1559]. terms [64]. terrestrial [158, 357, 384, 646, 686, 869]. terrestrial-satellite [357]. Test [20, 159, 221, 571, 572, 647, 989, 1248, 1259]. Testability [269, 696]. testers [1829]. Testing [24, 159, 570, 571, 794, 903, 932, 1568]. TFRC [306, 1524]. their [42, 464, 1053, 1095, 1537]. Theorem [1127]. theoretic [240, 983, 1000, 1068, 1434, 1442, 1741, 1748]. theoretical [313, 483]. theory [715, 731, 852, 878, 1575, 1663]. thermal [1987]. thermal-aware [1987]. third [283]. thread [1161]. three [426, 1955]. three-dimensional [1955]. three-metric [426]. Threshold [1454, 1650, 1909]. thresholds [10, 667]. Throughput [334, 391, 697, 811, 872, 896, 1116, 1142, 1397, 1419, 1422, 1474, 1548, 1552, 1627, 1639, 1756]. Ticket [1494]. Ticket-based [1494]. tier [1287]. tight [443]. Time [9, 36, 121, 184, 220, 245, 262, 302, 325, 477, 540, 567, 623, 671, 688, 744, 804, 832, 859, 897, 902, 908, 921, 950, 952, 992, 1058, 1078, 1249, 1378, 1404, 1409, 1417, 1497, 1515, 1574, 1580, 1592, 1593, 1596, 1617, 1633, 1705, 1711, 1773, 1803, 1808, 1938, 1966]. Time-efficient [1593]. time-limited [671]. time-optimal [1580]. time-sensitive [1711]. time-series [1617]. time-spread [1803]. time-to-market [220]. time-varying [832]. timed [405, 1568, 1824]. timeout [691, 901]. timeouts [1192]. timers [1357]. times [1032, 1609]. timing [930, 1824]. TINA [156]. TKC [74]. TMCP [1726]. today [212]. token [405]. tolerance [721, 1228]. tolerant [1017, 1042, 1090, 1770, 1776]. Tool [111, 262, 377, 475, 991]. Tool-based [991]. toolkit [1656]. tools [159, 219, 220, 238, 906, 914]. top [1717]. top- [1717]. topic [79]. topic-specific [79]. Topics [177, 1377]. Topological [1023, 1238, 1336, 1932]. topologically [1196]. topologies [335, 700, 1048, 1239, 1492, 1604, 1693, 1735, 1847, 1904, 1922]. Topology [124, 407, 891, 1201, 1206, 1485, 1489, 1573, 1585, 1645, 1682, 1747, 1770, 1825, 1888, 1917, 1945, 1998]. topology-aware [1747]. total [859]. tours [402]. TR [1758]. TR-clustering [1758]. trace [664]. traceback [1220, 1278, 1286]. traceroute [1603]. traces [478, 741]. Tracing [69]. tracking [611, 660, 732, 810, 1038, 1176, 1636, 1901]. Trade [1038]. Trade-offs [1038]. tradeoff [390, 1477, 1543, 1627, 1996]. tradeoffs [1618]. trading [98, 1572]. Traffic [34, 111, 117, 121, 128, 137, 169, 191, 238, 263, 293, 333, 342, 343, 369, 374, 379, 442, 444, 455, 457, 458, 481, 482, 536, 538, 574, 576, 596, 609, 673, 674, 685, 702, 723, 726, 748, 763, 783, 803, 877, 906, 941, 976, 1009, 1075, 1083, 1131, 1137, 1149, 1152, 1181, 1189, 1193, 1242, 1249, 1271, 1275, 1340, 1365, 1392, 1414, 1426, 1458, 1468, 1480, 1481, 1515, 1528, 1539, 1558, 1582, 1633, 1685, 1686, 1694, 1707, 1719, 1721, 1737, 1802, 1807, 1814, 1835, 1837, 1842, 1851, 1863, 1871, 1892, 1925, 1956, 1957, 1963, 1969, 1970, 1983, 1994]. traffic [113, 115, 120, 130, 267, 295, 403, 461, 495, 503, 597, 664, 675, 719, 733, 845, 879, 907– 909, 912–914, 924, 953, 1171, 1212, 1244, 1247, 1305, 1452, 1661, 1741, 1816, 1836, 1839]. traffic-aware [1480]. traffic-engineered [1149]. traffic-groomed [733]. traffic-transmission [343]. Trajectory [431]. transaction [1279]. transactional [1869]. transactions [292, 918, 1711]. transceiver [389]. transceiver-limited [389]. transcoding [61, 1007]. transfer [128, 131, 409, 1012, 1237, 1342, 1474]. transfers [1609]. transform [1072, 1591]. transformation [1696]. Transient [202, 501, 581, 1579, 1675, 1875]. TranSim [1739]. transition [241]. Transition-nets [1250]. Transmission [317, 343, 725, 862, 887, 967, 1033, 1064, 1324, 1376, 1525, 1538, 1590, 1690, 1703, 1853, 1977]. transmission-strategy-aware [1690]. transmissions [184]. transmit [548, 1647]. Transparent [438, 1895]. Transport [30, 191, 406, 469, 604, 793, 797, 811, 917, 1019, 1022, 1083, 1123, 1185, 1249, 1343, 1429, 1536, 1623, 1639, 1906, 1916, 1954]. transport-controlled [1916]. Transport-independent [1954]. transporting [1510]. trapping [151]. traversal [1098]. treatment [539]. tree [576, 994, 1337, 1420, 1438, 1736]. trees [672, 1007, 1018, 1364, 1409, 1650, 1761, 1765, 1960]. Trends [1099, 1444, 1880]. Trial [1704]. trials [1702]. TRIDENT [1247]. triggers [737]. trip [785]. troubleshooting [1979]. trust [759, 817, 1035, 1036, 1898, 1951]. trustworthiness [1725]. TRW [1637]. TSpaces [205]. Tspec [317]. Tstat [830].

TTCN [571]. TTCN-3 [571]. Tuning [278, 536, 725, 1353, 1910]. Tunnel [1782]. tunneling [650]. tunnelling [1161]. tunnels [394, 1782]. tuple [1092]. turbo [818]. tutorial [1990]. Two [63, 115, 238, 299, 370, 577, 778, 1121, 1221, 1270, 1273, 1277, 1287, 1428, 1483, 1536, 1726, 1871, 1955]. two-class [115, 1871]. two-dimensional [1121]. two-layer [1273, 1277, 1726]. two-phase [1483]. two-step [1270]. Two-tier [1287]. type [301, 430, 463]. types [1759].

U [1379]. U-APSD [1379]. ubiquitous [200, 1517, 1659, 1706, 1771]. UBR [137]. UDP [747, 1342]. UDP-based [1342]. UDT [1342]. ultra [949, 1416, 1585]. ultrawideband [690]. UML [217, 452, 987, 988]. UMTS [334, 359, 686, 839, 968, 971, 1178, 1192, 1582]. UMTS/GPRS [1192]. uncertain [1426]. uncollaborative [1947]. undeniable [1646]. underlying [1510]. Understanding [157, 296, 1055, 1222, 1603, 1787]. underutilized [1344]. unequal [997]. unfair [681]. unfairness [1143]. unicast [350, 706, 1033, 1049, 1130]. unidirectional [503]. unified [413, 1613, 1827]. Uniform [67, 130, 1005, 1015, 1545]. unintrusive [100]. Universal [358, 1056, 1453]. universality [1117]. UNIX [438]. unplanned [1724]. unreliable [1872]. unstructured [1011, 1598]. untraceability [669]. untrusted [933]. update [1207]. updates [672]. updating [749]. Uplink [556, 1178, 1382]. Upper [1271]. Upper-level [1271]. urban [1441]. URL [67]. usage [2, 260, 401, 1268, 1592, 1722, 1856]. usage-based [260]. Use [71, 132, 201, 650, 963, 989, 1326, 1362, 1917]. User [42, 49, 85, 104, 366, 367, 566, 580, 603, 669, 909, 918, 952, 980, 1101, 1524, 1646, 1772, 1775, 1927]. user-centric [1772]. user-friendly [603]. user-perceived [49, 1524]. user-specified [367]. users [660, 1334, 1537, 1866, 1948]. Using [4, 5, 14, 73, 94, 103, 114, 135, 144, 147, 152, 165, 203, 220, 232, 235, 264, 307, 316, 394, 407, 417, 422, 437, 442, 446, 452, 503, 516, 517, 531, 586, 595, 598, 651, 658, 675, 704, 737, 764, 802, 806, 812, 816, 867, 876, 896, 902, 980, 988, 1000, 1045, 1055, 1065, 1089, 1092, 1099, 1106, 1113, 1129, 1170, 1184, 1196, 1219, 1230, 1267, 1268, 1270, 1295, 1296, 1311, 1315, 1364, 1411, 1440, 1445, 1498, 1507, 1509, 1536, 1634, 1650, 1693, 1698, 1704, 1754, 1783, 1800, 1845, 1850, 1879, 1890, 1937, 1966, 1973, 1997, 1999]. using [24, 620, 1231]. utility [367, 1120, 1392, 1740, 1903, 1996]. utility-based [1903]. utility-fair [1392]. utility-lifetime [1996]. utilization [343, 643, 677, 813, 1064, 1440]. Utilize [1830]. Utilizing [217]. UWB [1431, 1728, 1850]. UWB-based [1728].

v [325]. valid [1566]. Validation [999, 1099]. value [1484]. value-added [1484]. variable [366, 943, 1674]. variable-weight [1674]. variance [463]. variance-type [463]. variant [1340]. variations [1736]. varying [184, 832]. VBR [112, 1863]. Vector [64, 1031, 1792, 1956]. Vegas [640, 896]. vehicular [1550]. VERA [347]. verification [159, 639, 718, 990, 1638, 1723, 1979]. verify [1089]. verifying [301]. version [1482, 1663]. versus [37, 1015, 1337, 1397, 1988]. vertical [841]. via [26, 87, 548, 811, 999, 1086, 1277, 1306, 1365, 1524, 1554, 1690, 1727, 1747]. Viceroy [1355]. video [86, 112, 171, 317, 425, 514, 680, 917, 997, 1173, 1230, 1497, 1557, 1698, 1789, 1812, 1863, 1927, 1971, 1994]. video-on-demand [1812]. videophones [304]. VIDOS [171]. View [476, 1932]. View-based [476]. Viewpoint [142]. viewpoints [1712]. Views [86]. Vinci [415]. violations [1781]. VIPSec [1708]. Virtual [6, 44, 195, 236, 237, 286, 335, 340, 511, 596, 656, 674, 804, 1075, 1181, 1312, 1343, 1483, 1562, 1577, 1754, 1755, 1845, 1888, 1902, 1989]. virtual-backbone [1562]. virtualized [1985, 1987]. Virtuosity [230]. visual [172]. visualization [75, 477, 478, 906, 914]. visualizing [625]. VLAN [1856]. VLAN-label [1856]. VM [1986]. VoD [126, 1859]. Voice [84, 208, 246, 357, 393, 482, 508, 701, 757, 835, 1051, 1615, 1705]. voice-enabled [84]. Voice/data [208, 1051]. VoIP [210, 1045, 1174, 1582, 1700, 1702– 1704, 1706, 1707, 1831, 1928]. Volume [330, 331, 380, 381, 433, 434, 484, 485, 599, 600, 652, 653, 707, 708, 768, 769, 827, 828, 884, 885, 1234]. Volumes [486, 487]. voting [164]. Vox [87]. VoxML [87]. VPN [295, 1045, 1337, 1529]. VPNs [1694]. VQ [804]. vs [1018, 1037, 1379, 1479]. vulnerabilities [1401, 1520].

W [1446]. W-CDMA [1446]. W3C [414]. waiting [744]. walk [1046, 1325]. walking [1390]. walks [1544]. WAN [456]. wanted [1040]. WAP [63, 578, 903]. war [1620]. watermarking [425]. waveguide [516]. Wavelength [31, 38, 185, 335, 688, 895, 995, 1005, 1154, 1198, 1227, 1236, 1274, 1491, 1662, 1666, 1793, 1830, 1848]. wavelength-continuity [895]. wavelength-routed [1491]. wavelength-routing [1198]. wavelet [913]. WCDMA [304, 1276, 1692]. WDM [13, 30, 32, 35–38, 355, 391, 445, 492, 498, 503, 516, 536, 620, 668, 673, 702, 726, 766, 824, 853, 895, 923, 962, 982, 995, 1005, 1015, 1017, 1021, 1026, 1212, 1227, 1469, 1614, 1633, 1661, 1669, 1675, 1776, 1826, 1830]. WDM/SCM [391]. Web [15, 49, 50, 55, 58, 61, 62, 64– 66, 68, 70, 72, 76, 79– 85, 87, 90, 92, 94, 95, 99, 101, 104, 105, 204, 245, 297, 328, 336, 377, 378, 395, 398– 400, 402, 409, 415, 419– 421, 450, 491, 515, 539, 553, 583– 588, 590, 615, 651, 712, 741, 743, 748, 929, 931, 932, 934, 978, 996, 1066, 1094, 1095, 1097– 1099, 1101, 1103, 1131, 1143, 1245, 1256, 1310, 1458, 1506, 1517, 1608, 1628, 1678, 1799, 1828, 1832, 1903, 1973, 2002]. Web [75, 78, 91, 96, 100, 401, 589, 598, 1510]. Web-based [204, 515, 931]. Web-collaborative [95]. Web-hosted [87]. Web-server [450, 748]. WebAccel [1678]. WebBase [66]. WebML [58]. website [1117]. WebSOS [934]. WebVCR [82]. WebVigiL [1102]. weight [1149, 1674]. Weighted [134, 312, 383, 609, 765, 1320]. well [1357]. Westwood [1088, 1697]. WF [313, 1320]. WFBoD [383]. WHIM [425]. Wi [734]. Wi-Fi [734]. Wide [15, 133, 326, 743, 996, 1342, 1363, 1389, 1416, 1722, 1765, 2002]. wide-area [133, 1363, 1389]. wideband [342]. WiFi [1794, 1801]. Wikipedia [1913, 2001]. will [97]. wily [151]. WiMAX [1801, 1959, 1965]. WiN [173]. window [697, 785, 1032, 1632, 1697, 1710]. window-based [697, 1632]. WinPEPSY [1875]. WinPEPSY-QNS [1875]. wired [847, 1086, 1417, 1464, 1501]. wired/wireless [847, 1086, 1417, 1501]. Wireless [200, 274, 281, 283, 316, 333, 342, 354, 512, 574, 593, 628, 632, 634, 649, 655, 690, 691, 725, 803, 810, 811, 822, 823, 838, 839, 842, 844, 858, 863, 878, 892, 938, 939, 949, 967, 970, 979, 1025, 1086, 1141, 1205, 1215, 1229, 1280, 1291, 1295, 1377, 1387, 1391, 1398, 1404, 1417, 1431, 1443, 1459, 1462, 1489, 1498, 1514, 1522, 1525, 1550, 1551, 1575, 1576, 1580, 1589, 1624, 1627, 1639, 1659, 1680– 1682, 1691, 1705, 1709, 1722–1724, 1727, 1745, 1752, 1783, 1803, 1807, 1822, 1850, 1857, 1916, 1919, 1930, 1936, 1937, 1962, 1982, 1996, 1998]. wireless [153, 155, 275, 282, 314, 338, 361, 364, 368, 389, 509–511, 518, 522, 631, 633, 669, 731, 745, 786, 793, 801, 814, 843, 847, 904, 921, 942, 997, 1085, 1090, 1145, 1171, 1176, 1185, 1214, 1216, 1222, 1244, 1251, 1273, 1282, 1285, 1292, 1296, 1299, 1303, 1324, 1329, 1335, 1379, 1381, 1388, 1392, 1418, 1434, 1464, 1475, 1493, 1495, 1501, 1543, 1546, 1548, 1557, 1564, 1590, 1595, 1636, 1643, 1679, 1683, 1703, 1710, 1712–1715, 1728, 1748, 1753, 1755, 1759, 1763, 1772, 1777, 1779, 1825, 1849, 1882, 1911, 1912, 1922, 1923, 1926, 1935, 1942, 1958, 1965, 1969, 1977, 1994, 1995, 1997]. Wireless [1861]. Wireless-LAN [839]. within [1613]. without [995, 1830, 1877]. WLAN [447, 840, 846, 1122, 1422, 1477, 1586, 1646, 1647, 1706, 1764, 1983]. WLAN-3G [1122]. WLAN-based [447]. WLAN/cellular [1646]. WLANs [521, 548, 1009, 1319, 1378, 1545, 1552, 1605, 1690, 1756, 1780]. WLESS [266]. WLESS-frame [266]. work [325, 1105, 1357]. Workbench [21]. workload [377, 1913, 1973, 2001]. workloads [1253]. Workshop [28, 679, 1071]. World [15, 256, 743, 806, 996, 1048, 1177, 2002]. World-Wide [996]. worm [1311]. wormhole [1462]. worms [1445]. Worp [1462]. Worst [1320]. Worst-case [1320]. WPANs [720]. written [1250]. WSANs [1251]. WTMS [79]. WWW [165]. WWW9 [2002].

X [235]. X-Bone [235]. XL [588]. XML [53, 57, 98, 99, 397, 412–414, 416, 421– 424, 588, 1050, 1102, 1516, 1517]. XML-based [1050]. XPref [935]. XQForms [416]. XSLT [97]. Xyleme [396].

Yellow [861]. YouTube [1816].

Zero [1311, 1992]. Zero-hour [1311]. Zipf [323]. Zipf-like [323]. zone [1583]. zonebased [1583].

References

Varadhan:2000:PRO

[1] Kannan Varadhan, Ramesh Govindan, and Deborah Estrin. Persistent route oscillations in inter-domain routing. Computer Networks (Amsterdam, Netherlands: 1999), 32 (1):1–16, January 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/25/27/article.pdf.

Yaghmaee:2000:IUP

[2] Mohamad Hossien Yaghmaee, Mostafa Safavi, and Mohamad Bagher Menhaj. An intelligent usage parameter controller based on dynamic rate leaky bucket for ATM networks. Computer Networks (Amsterdam, Netherlands: 1999), 32(1):17–34, January 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/25/28/article.pdf.

Aggarwal:2000:FMR

[3] Sudhir Aggarwal, Sanjoy Paul, Daniel Massey, and Daniela Caldararu. A flexible multicast routing protocol for group communication. Computer Networks (Amsterdam, Netherlands: 1999), 32(1):35–60, January 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/25/29/article.pdf.

Ng:2000:CAC

[4] Nelson O. L. Ng and C. K. Tham.

Connection admission control of ATM network using integrated MLP and fuzzy controllers. Computer Networks (Amsterdam, Netherlands: 1999), 32 (1):61–79, January 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/25/30/article.pdf.

Sisto:2000:UBD

[5] R. Sisto. Using binary decision diagrams for representation and analysis of communication protocols. Computer Networks (Amsterdam, Netherlands: 1999), 32(1):81–98, January 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/25/31/article.pdf.

Zhang:2000:HHF

[6] Hongyi Zhang, Oliver W. Yang, and H. Mouftah. A hop-by-hop flow controller for a virtual path. Computer Networks (Amsterdam, Netherlands: 1999), 32(1):99–119, January 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/25/32/article.pdf.

Akyildiz:2000:E

[7] Ian F. Akyildiz and Harry Rudin. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 32 (1):vi, January 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/25/26/article.pdf.

Enslow:2000:CGC

[8] Philip H. Enslow, Jr. Changing the guard at Computer Networks. Computer Networks (Amsterdam, Netherlands: 1999), 32(1):v, January 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/25/25/article.pdf.

Lau:2000:STA

[9] Wing Cheong Lau and Sanqi Li. Sojourn-time analysis on nodal congestion in broadband networks. Computer Networks (Amsterdam, Netherlands: 1999), 32(2):125–147, February 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/27/25/article.pdf.

Choi:2000:APQ

[10] Bong Dae Choi, Sung Ho Choi, Bara Kim, and Dan Keun Sung. Analysis of priority queueing system based on thresholds and its application to signaling system no. 7 with congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 32 (2):149–170, February 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/27/26/article.pdf.

Kraimeche:2000:DAS

[11] B. Kraimeche. Design and analysis of the Stacked-Banyan ATM switch fabric. Computer Networks (Amsterdam, Netherlands: 1999), 32(2): 171–184, February 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/27/27/article.pdf.

Hou:2000:DSA

[12] Yiwei Thomas Hou et al. A differentiated services architecture for multimedia streaming in next generation Internet. Computer Networks (Amsterdam, Netherlands: 1999), 32 (2):185–209, February 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/27/28/article.pdf.

Sivaraman:2000:RPB

[13] Vijay Sivaraman and George N. Rouskas. A reservation protocol for broadcast WDM networks and stability analysis. Computer Networks (Amsterdam, Netherlands: 1999), 32 (2):211–227, February 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/29/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/15/
22/45/27/29/article.pdf.

Ando:2000:MMS

[14] Toshihiko Ando, Kaoru Takahashi, Yasushi Kato, and Norio Shiratori. Maintenance of mobile system ambients using a process calculus. Computer Networks (Amsterdam, Netherlands: 1999), 32(2):229–256, February 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/27/30/article.pdf.

Chang:2000:ALM

[15] Da-Wei Chang, Hao-Ren Ke, and Ruei-Chuan Chang. Adaptive-level memory caches on World Wide Web servers. Computer Networks (Amsterdam, Netherlands: 1999), 32(3): 261–275, March 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/28/25/article.pdf.

Afek:2000:PSE

[16] Yehuda Afek, Yishay Mansour, and Zvi Ostfeld. Phantom: a simple and effective flow control scheme. Computer Networks (Amsterdam, Netherlands: 1999), 32(3):277–305, March 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/28/26/article.pdf.

Elloumi:2000:SBS

[17] Omar Elloumi, Nada Golmie, Hossam Afifi, and David Su. A simulationbased study of TCP dynamics over HFC networks. Computer Networks (Amsterdam, Netherlands: 1999), 32 (3):307–323, March 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/28/27/article.pdf.

Dolev:2000:BLS

[18] Shlomi Dolev and Alexander Kesselman. Bounded latency scheduling scheme for ATM cells. Computer Networks (Amsterdam, Netherlands: 1999), 32(3):325–331, March 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/28/28/article.pdf.

Peha:2000:DPC

[19] Jon M. Peha. Dynamic pricing and congestion control for best-effort ATM services. Computer Networks (Amsterdam, Netherlands: 1999), 32 (3):333–345, March 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/28/29/article.pdf.

Kang:2000:ITS

[20] Sungwon Kang, Jaehwi Shin, and Myungchul Kim. Interoperability test suite derivation for communication protocols. Computer Networks (Amsterdam, Netherlands: 1999), 32 (3):347–364, March 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/28/30/article.pdf.

Pullen:2000:NWN

[21] J. Mark Pullen. The Network Workbench: network simulation software for academic investigation of Internet concepts. Computer Networks (Amsterdam, Netherlands: 1999), 32 (3):365–378, March 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/28/31/article.pdf.

Bouma:2000:FIT

[22] L. G. Bouma, Nancy Griffeth, and Kristofer Kimbler. Feature interactions in telecommunications systems. Computer Networks (Amsterdam, Netherlands: 1999), 32 (4):383–387, April 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/29/25/article.pdf.

Griffeth:2000:FIB

[23] Nancy Griffeth, Ralph Blumenthal, Jean-Charles Gregoire, and Tadashi Ohta. A feature interaction benchmark for the first feature interaction detection contest. Computer Networks (Amsterdam, Netherlands: 1999), 32(4):389–418, April 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/29/26/article.pdf.

duBousquet:2000:FID

[24] L. du Bousquet, F. Ouabdesselam, J.- L. Richier, and N. Zuanon. Feature interaction detection using a synchronous approach and testing. Computer Networks (Amsterdam, Netherlands: 1999), 32(4):419–431, April 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/29/27/article.pdf.

Fu:2000:FID

[25] Q. Fu, P. Harnois, L. Logrippo, and J. Sincennes. Feature interaction detection: a LOTOS-based approach. Computer Networks (Amsterdam, Netherlands: 1999), 32 (4):433–448, April 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/29/28/article.pdf.

Hall:2000:FCI

[26] Robert J. Hall. Feature combination and interaction detection via foreground/background models. Computer Networks (Amsterdam, Netherlands: 1999), 32(4):449–469, April 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/29/29/article.pdf.

Siddiqi:2000:HMS

[27] Saheem Siddiqi and Joanne M. Atlee. A hybrid model for specifying features and detecting interactions. Computer Networks (Amsterdam, Netherlands: 1999), 32(4):471–485, April 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/29/30/article.pdf.

Griffeth:2000:FID

[28] Nancy Griffeth, Ralph Blumenthal, Jean-Charles Gregoire, and Tadashi Ohta. Feature interaction detection contest of the Fifth International Workshop on Feature Interactions. Computer Networks (Amsterdam, Netherlands: 1999), 32 (4):487–510, April 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/29/31/article.pdf.

Gagnaire:2000:ONN

[29] Maurice Gagnaire and Roberto Sabella. Optical networks for new generation Internet and data communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 32 (5):515–517, May 15, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/30/25/article.pdf.

Spath:2000:DRR

[30] Jan Späth. Dynamic routing and resource allocation in WDM transport networks. Computer Networks (Amsterdam, Netherlands: 1999), 32 (5):519–538, May 15, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/30/26/article.pdf.

Ali:2000:RWA

[31] Maher Ali, Byrav Ramamurthy, and Jitender S. Deogun. Routing and wavelength assignment with power considerations in optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 32(5):539–555, May 15, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/30/27/article.pdf.

Binetti:2000:GAO

[32] S. Binetti, G. Notaro, and R. Sabella. Granularity in all-optical WDM networks for large geographical areas. Computer Networks (Amsterdam, Netherlands: 1999), 32(5): 557–562, May 15, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/30/28/article.pdf.

Betti:2000:APS

[33] S. Betti, E. Bravi, and M. Giaconi. Analysis of the performance of subcarrier multiplexed (SCM) optical systems. Computer Networks (Amsterdam, Netherlands: 1999), 32 (5):563–569, May 15, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/30/29/article.pdf.

Gagnaire:2000:STM

[34] Maurice Gagnaire and Sašo Stojanovski. Stream traffic management over an ATM passive optical network. Computer Networks (Amsterdam, Netherlands: 1999), 32(5): 571–586, May 15, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/30/30/article.pdf.

White:2000:AHP

[35] I. M. White et al. The architecture of HORNET: A packet-over-WDM multiple-access optical metropolitan area ring network. Computer Networks (Amsterdam, Netherlands: 1999), 32 (5):587–598, May 15, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/30/31/article.pdf.

Marsan:2000:MSW

[36] M. Ajmone Marsan, E. Leonardi, M. Meo, and F. Neri. Modeling slotted WDM rings with discrete-time Markovian models. Computer Networks (Amsterdam, Netherlands: 1999), 32 (5):599–615, May 15, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/30/32/article.pdf.

Bianco:2000:PVP

[37] A. Bianco, V. Distefano, A. Fumagalli, E. Leonardi, and F. Neri. A posteriori versus a priori access strategies in slotted all-optical WDM rings. Computer Networks (Amsterdam, Netherlands: 1999), 32(5):617–631, May 15, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/30/33/article.pdf.

Eramo:2000:WCS

[38] Vincenzo Eramo and Marco Listanti. Wavelength converter sharing in a WDM optical packet switch: dimensioning and performance issues. Computer Networks (Amsterdam, Netherlands: 1999), 32(5): 633–651, May 15, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/30/34/article.pdf.

Castanon:2000:OPS

[39] Gerardo Castañón, Lubo Tancevski, and Lakshman Tamil. Optical packet switching with multiple path routing. Computer Networks (Amsterdam, Netherlands: 1999), 32(5): 653–662, May 15, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/30/35/article.pdf.

Field:2000:EC

[40] Simon Field and Michael Waidner. Electronic Commerce. Computer Networks (Amsterdam, Netherlands: 1999), 32(6):667–668, May 30, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/31/25/article.pdf.

Mott:2000:SGD

[41] Steve Mott. The second generation of digital commerce solutions. Computer Networks (Amsterdam, Netherlands: 1999), 32(6):669–683, May 30, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/31/26/article.pdf.

Wolf:2000:PPG

[42] Gritta Wolf and Andreas Pfitzmann. Properties of protection goals and their integration into a user interface. Computer Networks (Amsterdam, Netherlands: 1999), 32(6): 685–700, May 30, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/31/27/article.pdf.

Stanoevska-Slabeva:2000:IEP

[43] Katarina Stanoevska-Slabeva and Beat Schmid. Internet electronic product catalogs: an approach beyond simple keywords and multimedia. Computer Networks (Amsterdam, Netherlands: 1999), 32(6):701–715, May 30, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/31/28/article.pdf.

Hoffner:2000:DID

[44] Yigal Hoffner, Christian Facciorusso, Simon Field, and Andreas Schade. Distribution issues in the design and implementation of a virtual market place. Computer Networks (Amsterdam, Netherlands: 1999), 32(6): 717–730, May 30, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        45/25/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/45/31/29/article.pdf.

Kephart:2000:DPS

[45] Jeffrey O. Kephart, James E. Hanson, and Amy R. Greenwald. Dynamic pricing by software agents. Computer Networks (Amsterdam, Netherlands: 1999), 32(6):731–752, May 30, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/31/30/article.pdf.

Konstantas:2000:ABC

[46] Dimitri Konstantas and Jean-Henry Morin. Agent-based commercial dissemination of electronic information. Computer Networks (Amsterdam, Netherlands: 1999), 32(6): 753–765, May 30, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/31/31/article.pdf.

Anonymous:2000:Ia

[47] Anonymous.  Index.  Computer

Networks (Amsterdam, Netherlands: 1999), 32(6):771–773, May 30, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/31/32/article.pdf.

Anonymous:2000:Ib

[48] Anonymous.  Index.  Computer

Networks (Amsterdam, Netherlands: 1999), 32(6):775–776, May 30, 2000.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
45/25/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/45/31/33/article.pdf.

Bhatti:2000:IUP

[49] Nina Bhatti, Anna Bouch, and Allan Kuchinsky. Integrating user-perceived quality into Web server design. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):1–16, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/27/article.pdf;   http:
//www9.org/w9cdrom/92/92.html.

Krishnamurthy:2000:AFI

[50] Balachander Krishnamurthy and Craig E. Wills. Analyzing factors that influence end-to-end Web performance. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):17–32, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/28/article.pdf;   http:
//www9.org/w9cdrom/371/371.html.

Rodriguez:2000:SSP

[51] Pablo Rodriguez and Sandeep Sibal. SPREAD: Scalable platform for reliable and efficient automated distribution. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):33–49, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/29/article.pdf;   http:
//www9.org/w9cdrom/411/411.html.

Kormann:2000:RPS

[52] David P. Kormann and Aviel D. Rubin. Risks of the Passport single signon protocol. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):51–58, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/30/article.pdf;   http:
//www9.org/w9cdrom/114/114.html.

Damiani:2000:DIA

[53] Ernesto Damiani, Sabrina De Capitani di Vimercati, Stefano Paraboschi, and Pierangela Samarati. Design and implementation of an access control processor for XML documents. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):59–75, June 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/31/article.pdf;   http:
//www9.org/w9cdrom/419/419.html.

Hashii:2000:SRS

[54] B. Hashii, S. Malabarba, R. Pandey, and M. Bishop. Supporting reconfigurable security policies for mobile programs. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):77–93, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/32/article.pdf;   http:
                                        //www9.org/w9cdrom/345/345.html.

Spertus:2000:SSQ

[55] Ellen Spertus and Lynn Andrea Stein. Squeal: a structured query language for the Web. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):95–103, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/33/article.pdf;   http:
                                        //www9.org/w9cdrom/222/222.html.

Phelps:2000:RID

[56] Thomas A. Phelps and Robert Wilensky. Robust intra-document locations. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):105–118, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/34/article.pdf;   http:
                                        //www9.org/w9cdrom/312/312.html.

Florescu:2000:IKS

[57] Daniela Florescu, Donald Kossmann, and Ioana Manolescu. Integrating keyword search into XML query processing. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):119–135, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/35/article.pdf;   http:
//www9.org/w9cdrom/324/324.html.

Ceri:2000:WML

[58] Stefano Ceri, Piero Fraternali, and Aldo Bongio. Web Modeling Language (WebML): a modeling language for designing Web sites. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):137–157, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/36/article.pdf;   http:
//www9.org/w9cdrom/177/177.html.

Badros:2000:JML

[59] Greg J. Badros. JavaML: a markup language for Java source code. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):159–177, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/37/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/37/article.pdf;   http:
//www9.org/w9cdrom/342/342.html.

vanAmstel:2000:IFC

[60] Patrick van Amstel, Pim van der Eijk, Evert Haasdijk, and David Kuilman. An interchange format for cross-media personalized publishing. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):179–195, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/38/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/38/article.pdf;   http:
                                        //www9.org/w9cdrom/263/263.html.

Hori:2000:ABW

[61] Masahiro Hori, Goh Kondoh, Kouichi Ono, Shin ichi Hirose, and Sandeep Singhal. Annotation-based Web content transcoding. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):197–211, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/39/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/39/article.pdf;   http:
                                        //www9.org/w9cdrom/169/169.html.

Buyukkokten:2000:FWS

[62] Orkut Buyukkokten, Hector Garcia-Molina, and Andreas Paepcke. Focused Web searching with PDAs. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):213–230, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/40/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/40/article.pdf;   http:
                                        //www9.org/w9cdrom/195/195.html.

Kaasinen:2000:TAB

[63] Eija Kaasinen, Matti Aaltonen, Juha Kolari, Suvi Melakoski, and Timo Laakko. Two approaches to bringing Internet services to WAP devices. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):231–246, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/41/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/41/article.pdf;   http:
//www9.org/w9cdrom/228/228.html.

Stata:2000:TVD

[64] Raymie Stata, Krishna Bharat, and Farzin Maghoul. The Term Vector Database: fast access to indexing terms for Web pages. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):247–255, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/42/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/42/article.pdf;   http:
//www9.org/w9cdrom/159/159.html.

Brewington:2000:HDW

[65] Brian E. Brewington and George Cybenko. How dynamic is the Web? Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):257–276, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/43/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/43/article.pdf;   http:
//www9.org/w9cdrom/264/264.html.

Hirai:2000:WRW

[66] Jun Hirai, Sriram Raghavan, Hector Garcia-Molina, and Andreas Paepcke. WebBase: a repository of Web pages. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):277–293, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/44/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/44/article.pdf;   http:
//www9.org/w9cdrom/296/296.html.

Henzinger:2000:NUU

[67] Monika R. Henzinger, Allan Heydon, Michael Mitzenmacher, and Marc Najork. On near-uniform URL sampling. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):295–308, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/45/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/45/article.pdf.

Broder:2000:GSWb

[68] Andrei Broder, Ravi Kumar, Farzin Maghoul, Prabhakar Raghavan, Sridhar Rajagopalan, Raymie Stata, Andrew Tomkins, and Janet Wiener. Graph structure in the Web. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):309–320, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/46/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/46/article.pdf;   http:
//www9.org/w9cdrom/160/160.html.

Feldmann:2000:BBL

[69] Anja Feldmann. BLT: Bi-Layer Tracing of HTTP and TCP/IP. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):321–335, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/47/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/47/article.pdf;   http:
//www9.org/w9cdrom/367/367.html.

Holscher:2000:WSBb

[70] Christoph Hölscher and Gerhard Strube. Web search behavior of Internet experts and newbies. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):337–346, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/48/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/48/article.pdf;   http:
//www9.org/w9cdrom/81/81.html.

Forsberg:2000:EUR

[71] Kerstin Forsberg and Lars Dannstedt. Extensible use of RDF in a business context. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):347–364, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/49/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/49/article.pdf;   http:
//www9.org/w9cdrom/323/323.html.

Fowler:2000:EMW

[72] Shannon L. Fowler, Anne-Marie J. Novack, and Michael J. Stillings. The evolution of a manufacturing Web site. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):365–376, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/50/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/50/article.pdf;   http:
//www9.org/w9cdrom/346/346.html.

Sarukkai:2000:LPP

[73] Ramesh R. Sarukkai. Link prediction and path analysis using Markov chains. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):377–386, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/51/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/51/article.pdf;   http:
//www9.org/w9cdrom/68/68.html.

Lempel:2000:SAL

[74] R. Lempel and S. Moran. The stochastic approach for link-structure analysis (SALSA) and the TKC effect. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):387–401, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/52/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/52/article.pdf;   http:
//www9.org/w9cdrom/175/175.html.

Weinreich:2000:CIV

[75] Harald Weinreich and Winfried Lamersdorf. Concepts for improved visualization of Web link attributes. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):403–416, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/53/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/53/article.pdf;   http:
//www9.org/w9cdrom/319/319.html.

Sugiura:2000:QRW

[76] Atsushi Sugiura and Oren Etzioni. Query routing for Web search engines: architecture and experiments. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):417–429, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/54/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/54/article.pdf;   http:
//www9.org/w9cdrom/139/139.html.

Huang:2000:AAD

[77] Lieming Huang, Matthias Hemmje, and Erich J. Neuhold. ADMIRE: an adaptive data model for meta search engines. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):431–448, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/55/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/55/article.pdf;   http:
//www9.org/w9cdrom/165/165.html;

                                         http://www9.org/w9cdrom/374/374.
                                        html.

Zhang:2000:EAR

[78] Dell Zhang and Yisheng Dong. An efficient algorithm to rank Web resources. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):449–455, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/56/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/56/article.pdf;   http:
                                        //www9.org/w9cdrom/251/251.html.

Mukherjea:2000:WSC

[79] Sougata Mukherjea. WTMS: a system for collecting and analyzing topicspecific Web information. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):457–471, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/57/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/57/article.pdf;   http:
                                        //www9.org/w9cdrom/293/293.html.

Staab:2000:SCW

[80] S. Staab, J. Angele, S. Decker, M. Erdmann, A. Hotho, A. Mädche, H.-P. Schnurr, and R. Studer. Semantic community Web portals. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):473–491, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/58/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/

22/48/25/58/article.pdf;   http:
//www9.org/w9cdrom/134/134.html.

Bharat:2000:SEC

[81] Krishna Bharat. SearchPad: explicit capture of search context to support Web search. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):493–501, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/59/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/59/article.pdf;   http:
//www9.org/w9cdrom/173/173.html.

Anupam:2000:AWN

[82] Vinod Anupam, Juliana Freire, Bharat Kumar, and Daniel Lieuwen. Automating Web navigation with the WebVCR. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):503–517, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/60/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/60/article.pdf;   http:
//www9.org/w9cdrom/208/208.html.

Shafer:2000:CQD

[83] John C. Shafer and Rakesh Agrawal. Continuous querying in databasecentric Web applications. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):519–531, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/61/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/

                                        22/48/25/61/article.pdf;   http:
                                        //www9.org/w9cdrom/378/378.html.

Rollins:2000:ACA

[84] Sami Rollins and Neel Sundaresan. AVoN calling: AXL for voice-enabled Web navigation. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):533–551, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/62/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/62/article.pdf;   http:
                                        //www9.org/w9cdrom/55/55.html.

Gronbaek:2000:OHU

[85] Kaj Grønbæk, Lennert Sloth, and Niels Olof Bouvin. Open hypermedia as user controlled meta data for the Web. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):553–566, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/63/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/63/article.pdf;   http:
                                        //www9.org/w9cdrom/183/183.html.

Pan:2000:VCO

[86] Pengkai Pan and Glorianna Davenport. I-views: a community-oriented system for sharing streaming video on the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):567–581, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/64/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/15/
22/48/25/64/article.pdf;   http:
//www9.org/w9cdrom/199/199.html.

Goose:2000:EWA

[87] Stuart Goose, Mike Newman, Claus Schmidt, and Laurent Hue. Enhancing Web accessibility via the Vox Portal and a Web-hosted dynamic HTML ⇔ VoxML converter. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):583–592, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/65/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/65/article.pdf;   http:
//www9.org/w9cdrom/354/354.html.

Spreitzer:2000:HNG

[88] Mike Spreitzer and Bill Janssen. HTTP ‘Next Generation’. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):593–607, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/66/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/66/article.pdf;   http:
//www9.org/w9cdrom/60/60.html.

Krishnan:2000:CIR

[89] Venky Krishnan and S. Grace Chang. Customized Internet radio. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):609–618, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/67/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/

                                        22/48/25/67/article.pdf;   http:
                                        //www9.org/w9cdrom/353/353.html.

Huang:2000:RWB

[90] Andrew C. Huang, Benjamin C. Ling, John J. Barton, and Armando Fox. Running the Web backwards: appliance data services. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):619–631, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/68/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/68/article.pdf.

Wu:2000:LSH

[91] Kun-Lung Wu and Philip S. Yu. Latency-sensitive hashing for collaborative Web caching. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):633–644, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/69/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/69/article.pdf;   http:
                                        //www9.org/w9cdrom/301/301.html.

Jung:2000:PWC

[92] Jaeyeon Jung, Dongman Lee, and Kilnam Chon. Proactive Web caching with cumulative prefetching for large multimedia data. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):645–655, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/70/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/

22/48/25/70/article.pdf;   http:
//www9.org/w9cdrom/315/315.html.

Gruber:2000:PCP

[93] Stephane Gruber, Jennifer Rexford, and Andrea Basso. Protocol considerations for a prefix-caching proxy for multimedia streams. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):657–668, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/71/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/71/article.pdf;   http:
//www9.org/w9cdrom/349/349.html.

Chakrabarti:2000:UMA

[94] Soumen Chakrabarti, Sandeep Srivastava, Mallela Subramanyam, and Mitul Tiwari. Using Memex to archive and mine community Web browsing experience. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):669–684, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/72/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/72/article.pdf;   http:
//www9.org/w9cdrom/98/98.html.

Cohen:2000:WCF

[95] William W. Cohen and Wei Fan. Webcollaborative filtering: recommending music by crawling the Web. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):685–698, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/

                                        48/25/73/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/73/article.pdf;   http:
                                        //www9.org/w9cdrom/266/266.html.

Sundaresan:2000:MWR

[96] Neel Sundaresan and Jeonghee Yi. Mining the Web for relations. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):699–711, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/15/22/48/25/74/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/15/22/48/25/74/article.pdf;
                                        http://www9.org/w9cdrom/363/363.
                                        html.

Cawsey:2000:PTR

[97] Alison Cawsey. Presenting tailored resource descriptions: will XSLT do the job? Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):713–722, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/75/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/75/article.pdf;   http:
                                        //www9.org/w9cdrom/119/119.html.

Fernandez:2000:STBb

[98] Mary Fernández, Wang-Chiew Tan, and Dan Suciu. SilkRoute: trading between relations and XML. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):723–745, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/76/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/15/
22/48/25/76/article.pdf;   http:
//www9.org/w9cdrom/202/202.html.

Girardot:2000:MEF

[99] Marc Girardot and Neel Sundaresan. Millau: an encoding format for efficient representation and exchange of XML over the Web. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):747–765, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/77/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/77/article.pdf;   http:
//www9.org/w9cdrom/154/154.html.

Tomlin:2000:EAU

[100] John A. Tomlin. An entropy approach to unintrusive targeted advertising on the Web. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):767–774, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/78/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/78/article.pdf;   http:
//www9.org/w9cdrom/214/214.html.

Abe:2000:WMS

[101] Naoki Abe and Tomonari Kamba. A Web marketing system with automatic pricing. Computer Networks (Amsterdam, Netherlands: 1999), 33 (1–6):775–788, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/79/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/79/article.pdf;   http:
                                        //www9.org/w9cdrom/217/217.html.

Brustoloni:2000:MPC

[102] José Brustoloni and Juan Garay. MicroISPs: providing convenient and low-cost high-bandwidth Internet access. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):789–802, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/80/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/80/article.pdf;   http:
                                        //www9.org/w9cdrom/249/249.html.

Martinez:2000:OBU

[103] Kirk Martinez, Steve Perry, and John Cupitt. Object browsing using the Internet Imaging Protocol. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):803–810, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/81/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/81/article.pdf;   http:
                                        //www9.org/w9cdrom/122/122.html.

Nakayama:2000:DGB

[104] Takehiro Nakayama, Hiroki Kato, and Yohei Yamane. Discovering the gap between Web site designers’ expectations and users’ behavior. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):811–822, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/

48/25/82/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/82/article.pdf;   http:
//www9.org/w9cdrom/213/213.html.

Rafiei:2000:WPK

[105] Davood Rafiei and Alberto O. Mendelzon. What is this page known for? computing Web page reputations. Computer Networks (Amsterdam, Netherlands: 1999), 33(1– 6):823–835, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/83/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/83/article.pdf;   http:
//www9.org/w9cdrom/368/368.html.

Anonymous:2000:Ic

[106] Anonymous.  Index.  Computer

Networks (Amsterdam, Netherlands: 1999), 33(1–6):837–839, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/84/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/84/article.pdf.

Anonymous:2000:Id

[107] Anonymous.  Index.  Computer

Networks (Amsterdam, Netherlands: 1999), 33(1–6):841–843, June 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
48/25/85/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/48/25/85/article.pdf.

Anonymous:2000:Ea

[108] Anonymous. Editorial(s). Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):v–vi, June 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/25/article.pdf.

Anonymous:2000:Eb

[109] Anonymous. Editorial(s). Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):vii–viii, June 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        48/25/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/48/25/26/article.pdf.

Kouvatsos:2000:PME

[110] Demetres Kouvatsos. Performance Modelling and Evaluation of ATM Networks. Computer Networks (Amsterdam, Netherlands: 1999), 34(1):1–2, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/25/25/article.pdf.

Evangelatos:2000:BLT

[111] A. Evangelatos, Th. Rapsomanikis, N. Mitrou, and G. Stassinopoulos. A burst-level traffic analysis tool. Computer Networks (Amsterdam, Netherlands: 1999), 34(1):3–22, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/25/26/article.pdf.

Alheraish:2000:CRM

[112] A. Alheraish and D. G. Smith. Cell rate moderation of VBR video under an FEC environment in ATM networks. Computer Networks (Amsterdam, Netherlands: 1999), 34 (1):23–35, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/25/27/article.pdf.

Stathis:2000:MSS

[113] Christos Stathis and Basil Maglaris. Modelling the self-similar behaviour of network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 34(1):37–47, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/25/28/article.pdf.

Balestrieri:2000:ACA

[114] Francesco Balestrieri, Pantelis L. Panteli, Vasillios Dionissopoulos, and Trevor G. Clarkson. ATM connection admission control using pRAM based artificial neural networks. Computer Networks (Amsterdam, Netherlands: 1999), 34(1):49–63, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/25/29/article.pdf.

Soldatos:2000:CTS

[115] J. Soldatos, E. Vayias, and N. Mitrou. CAC and traffic shaping for performance control in ATM: the twoclass paradigm. Computer Networks (Amsterdam, Netherlands: 1999), 34(1):65–83, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/25/30/article.pdf.

Becker:2000:PCC

[116] M. Becker and A.-L. Beylot. Performance comparison criteria for ATM switch models. Computer Networks (Amsterdam, Netherlands: 1999), 34(1):85–95, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/25/31/article.pdf.

Kouvatsos:2000:CEA

[117] Demetres D. Kouvatsos, Irfan U. Awan, Rod J. Fretwell, and George Dimakopoulos. A cost-effective approximation for SRD traffic in arbitrary multi-buffered networks. Computer Networks (Amsterdam, Netherlands: 1999), 34(1):97–113, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/

49/25/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/25/32/article.pdf.

Bhabuta:2000:OBA

[118] Madhu D. K. Bhabuta and Peter G. Harrison. Optimising bandwidth of ABR sources. Computer Networks (Amsterdam, Netherlands: 1999), 34 (1):115–128, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/25/33/article.pdf.

Lengliz:2000:RBF

[119] Ilhem Lengliz and Farouk Kamoun. A rate-based flow control method for ABR service in ATM networks. Computer Networks (Amsterdam, Netherlands: 1999), 34(1):129–138, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/25/34/article.pdf.

Vojnovi:2000:EAE

[120] Milan Vojnovi and Nikola Ro i. An evaluation of the ABR explicit-rate allocation interfering with the guaranteed services traffic. Computer Networks (Amsterdam, Netherlands: 1999), 34(1):139–155, July 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/25/35/article.pdf.

Naudts:2000:TRT

[121] Jan Naudts. Towards real-time measurement of traffic control parameters. Computer Networks (Amsterdam, Netherlands: 1999), 34 (1):157–167, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/25/36/article.pdf.

Casetti:2000:LBS

[122] C. Casetti, R. Lo Cigno, and M. Mellia. Load-balancing solutions for static routing schemes in ATM networks. Computer Networks (Amsterdam, Netherlands: 1999), 34 (1):169–180, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/25/37/article.pdf.

Magi:2000:ALC

[123] Ádám Magi, Áron Szentesi, and Balázs Szviatovszki. Analysis of link cost functions for PNNI routing. Computer Networks (Amsterdam, Netherlands: 1999), 34(1):181–197, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/38/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/25/38/article.pdf.

Do:2000:TOO

[124] Tien Van Do, Thong T. Nguyen, Hung T. Tran, Gábor Kalvach, and Balázs Varga. Topology optimization of an overlay ATM network in an SDH infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 34(1):199–210, July 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/39/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/25/39/article.pdf.

Denteneer:2000:IRN

[125] D. Denteneer, V. Pronk, J. M. Griffiths, and L. G. Cuthbert. Impact of the resource needed for renegotiating ATM rates. Computer Networks (Amsterdam, Netherlands: 1999), 34 (1):211–225, July 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/40/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/25/40/article.pdf.

Poon:2000:PBB

[126] S. M. Poon, Song Jie, B. S. Lee, and C. K. Yeo. Performance of bufferbased request-reply scheme for VoD streams over IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 34(2):229–240, August 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/27/25/article.pdf.

Hwang:2000:MRM

[127] Ren-Hung Hwang, James F. Kurose, and Don Towsley. MDP routing for multi-rate loss networks. Computer Networks (Amsterdam, Netherlands: 1999), 34(2):241–261, August 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/27/26/article.pdf.

Guillemin:2000:LSA

[128] Fabrice Guillemin, Olivier Dugeon, Jacqueline Boyer, and Christophe Mangin. Lightweight signaling in ATM networks for high quality transfer of Internet traffic. Computer Networks (Amsterdam, Netherlands: 1999), 34 (2):263–278, August 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/27/27/article.pdf.

Song:2000:CBF

[129] Shaowen Song. A credit-based flow control algorithm for broadband access networks. Computer Networks (Amsterdam, Netherlands: 1999), 34 (2):279–295, August 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/27/28/article.pdf.

Saleh:2000:ASB

[130] Mahmoud Saleh and Mohammed Atiquzzaman. Analysis of shared buffer switches under non-uniform traffic pattern and global flow control. Computer Networks (Amsterdam, Netherlands: 1999), 34(2):297–315, August 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/27/29/article.pdf.

Herrmann:2000:FMT

[131] Peter Herrmann and Heiko Krumm. A framework for modeling transfer protocols. Computer Networks (Amsterdam, Netherlands: 1999), 34(2): 317–337, August 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/27/30/article.pdf.

Atlasis:2000:ULA

[132] Antonios F. Atlasis, Nikolaos H. Loukas, and Athanasios V. Vasilakos. The use of learning algorithms in ATM networks call admission control problem: a methodology. Computer Networks (Amsterdam, Netherlands: 1999), 34(3):341–353, September 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/29/24/article.pdf.

Al-Fawzan:2000:PNS

[133] M. A. Al-Fawzan and Fahad Hoymany. Placement of network servers in a widearea network. Computer Networks (Amsterdam, Netherlands: 1999), 34 (3):355–361, September 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/29/25/article.pdf.

Habra:2000:MWP

[134] Shatha K. Habra and Ahmed E. Kamal. An MCR-weighted protocol for multipoint-to-point communication over ABR service. Computer Networks (Amsterdam, Netherlands: 1999), 34 (3):363–378, September 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/29/26/article.pdf.

Montgomery:2000:HSR

[135] Michael Montgomery and Gustavo de Veciana. Hierarchical source routing using implied costs. Computer Networks (Amsterdam, Netherlands: 1999), 34(3):379–397, September 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/25/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/29/27/article.pdf.

Haas:2000:PAR

[136] Robert Haas, Patrick Droz, and Daniel Bauer. PNNI Augmented Routing (PAR) and Proxy-PAR. Computer Networks (Amsterdam, Netherlands: 1999), 34(3):399–418, September 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/

49/25/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/29/28/article.pdf.

Lai:2000:IGI

[137] Wei Kuang Lai and Mu-Rung Shiu. Improving goodputs of IP packets under ATM UBR traffic with port-based queueing schemes. Computer Networks (Amsterdam, Netherlands: 1999), 34 (3):419–437, September 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/29/29/article.pdf.

Mainwaring:2000:RAT

[138] Keith Mainwaring. A role for ATM in telephony and IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 34(3):439–454, September 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/29/30/article.pdf.

Hughes:2000:CR

[139] Larry Hughes, Omid Banyasad, and Evan Hughes. Cartesian routing. Computer Networks (Amsterdam, Netherlands: 1999), 34(3): 455–466, September 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/29/31/article.pdf.

Chang:2000:NMS

[140] Yue-Shan Chang, Shyan-Ming Yuan, and Winston Lo. A new multisearch engine for querying data through an Internet search service on CORBA. Computer Networks (Amsterdam, Netherlands: 1999), 34 (3):467–480, September 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/29/32/article.pdf.

Hou:2000:NBA

[141] Y. Thomas Hou, Bo Li, Shivendra S. Panwar, and Henry Tzeng. On network bandwidth allocation policies and feedback control algorithms for packet networks. Computer Networks (Amsterdam, Netherlands: 1999), 34 (3):481–501, September 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/29/33/article.pdf.

Boiten:2000:VCO

[142] Eerke Boiten, Howard Bowman, John Derrick, Peter Linington, and Maarten Steen. Viewpoint consistency in ODP. Computer Networks (Amsterdam, Netherlands: 1999), 34(3): 503–537, September 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/29/34/article.pdf.

Frincke:2000:RAI

[143] Deborah A. Frincke and Ming-Yuh Huang. Recent advances in intrusion detection systems. Computer Networks (Amsterdam, Netherlands: 1999), 34(4):541–545, October 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/30/24/article.pdf.

Spafford:2000:IDU

[144] Eugene H. Spafford and Diego Zamboni. Intrusion detection using autonomous agents. Computer Networks (Amsterdam, Netherlands: 1999), 34 (4):547–570, October 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/30/25/article.pdf.

Manganaris:2000:DMA

[145] Stefanos Manganaris, Marvin Christensen, Dan Zerkle, and Keith Hermiz. A data mining analysis of RTID alarms. Computer Networks (Amsterdam, Netherlands: 1999), 34(4): 571–577, October 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/30/26/article.pdf.

Lippmann:2000:DLI

[146] Richard Lippmann, Joshua W. Haines, David J. Fried, Jonathan Korba, and Kumar Das. The 1999 DARPA off-line intrusion detection evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 34(4): 579–595, October 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/30/27/article.pdf.

Lippmann:2000:IID

[147] Richard P. Lippmann and Robert K. Cunningham. Improving intrusion detection performance using keyword selection and neural networks. Computer Networks (Amsterdam, Netherlands: 1999), 34(4):597–603, October 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/30/28/article.pdf.

Feiertag:2000:IDI

[148] Richard Feiertag et al. Intrusion detection inter-component adaptive negotiation. Computer Networks (Amsterdam, Netherlands: 1999), 34(4): 605–621, October 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/30/29/article.pdf.

Lundin:2000:ABI

[149] Emilie Lundin and Erland Jonsson. Anomaly-based intrusion detection: privacy concerns and other problems. Computer Networks (Amsterdam, Netherlands: 1999), 34(4): 623–640, October 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/30/30/article.pdf.

Mell:2000:DSR

[150] Peter Mell, Donald Marks, and Mark McLarnon. A denial-of-service resistant intrusion detection architecture. Computer Networks (Amsterdam, Netherlands: 1999), 34(4): 641–658, October 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/30/31/article.pdf.

Mansfield:2000:TTW

[151] Glenn Mansfield, Kohei Ohta, Y. Takei, N. Kato, and Y. Nemoto. Towards trapping wily intruders in the large. Computer Networks (Amsterdam, Netherlands: 1999), 34(4): 659–670, October 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/25/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/30/32/article.pdf.

Yuill:2000:IDI

[152] J. Yuill et al. Intrusion-detection for incident-response, using a military battlefield-intelligence process. Computer Networks (Amsterdam, Netherlands: 1999), 34(4):671–697, October 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/15/22/49/25/33/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/15/22/49/30/33/article.pdf.

Logrippo:2000:FWN

[153] Luigi Logrippo and John Visser. Future wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 34 (5):701–703, November ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/31/24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/31/24/article.pdf.

Schmidts:2000:FIN

[154] Harald Schmidts and John Visser. Framework for IMT-2000 networks. Computer Networks (Amsterdam, Netherlands: 1999), 34(5):705–715, November ??, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/31/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/31/25/article.pdf.

Samou:2000:IWT

[155] Jean Claude Samou, Drissa Houatra, Philippe Godlewski, and Elie Najm. IMT-2000 wireless terminal mobility over object-oriented distributed platforms. Computer Networks (Amsterdam, Netherlands: 1999), 34(5): 717–743, November ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/31/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/31/26/article.pdf.

Lombardo:2000:PMA

[156] Alfio Lombardo, Ferdinando Lucidi, Marina Samarotto, and Sebastiano Trigila. Personal mobility aspects in service architectures: extensions to TINA. Computer Networks (Amsterdam, Netherlands: 1999), 34(5): 745–762, November ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/31/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/31/27/article.pdf.

Ghribi:2000:UGG

[157] Brahim Ghribi and Luigi Logrippo. Understanding GPRS: the GSM packet radio service. Computer Networks (Amsterdam, Netherlands: 1999), 34 (5):763–779, November ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/31/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/31/28/article.pdf.

Daoud:2000:HST

[158] Fawzi Daoud. Hybrid satellite/terrestrial networks integration. Computer Networks (Amsterdam, Netherlands: 1999), 34(5):781–797, November ??, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/15/22/49/31/29/abstract.html;

                                         http://www.elsevier.nl/gej-ng/
                                        10/15/22/49/31/29/article.pdf.

Maggiore:2000:NIT

[159] Giulio Maggiore, Giulio Brusasco, and Marco Vecchiato. Network integration testing: concepts, test specifications and tools for automatic telecommunication services verification. Computer Networks (Amsterdam, Netherlands: 1999), 34(5):799– 819, November ??, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/31/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/31/30/article.pdf.

Allocchio:2000:GE

[160] Claudio Allocchio. Guest editorial. Computer Networks (Amsterdam, Netherlands: 1999), 34(6):821– 822, December ??, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/32/24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/49/32/24/article.pdf.

Raibulet:2000:MAT

[161] Claudia Raibulet and Claudio Demartini. Mobile agent technology for the management of distributed systems — a case study. Computer Networks (Amsterdam, Netherlands: 1999), 34(6): 823–830, December ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        49/32/25/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/15/
22/49/32/25/article.pdf.

Bissel:2000:SLM

[162] Torsten Bissel, Manfred Bogen, Christian Bonkowski, and Volker Hadamschek. Service level management with agent technology. Computer Networks (Amsterdam, Netherlands: 1999), 34(6):831–841, December ??, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/49/
32/26/abstract.html;   http://www.
elsevier.nl/gej-ng/10/15/22/49/
32/26/article.pdf.

Orbaek:2000:ESD

[163] Peter Ørbæk. An experimental system for distributed classroom education. Computer Networks (Amsterdam, Netherlands: 1999), 34(6): 843–850, December ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/32/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/32/27/article.pdf.

Bonetti:2000:IAC

[164] Pierluigi Bonetti, Stefano Ravaioli, and Simone Piergallini. The Italian academic community’s electronic voting system. Computer Networks (Amsterdam, Netherlands: 1999), 34(6): 851–860, December ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/32/28/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/15/
22/49/32/28/article.pdf.

Karvounarakis:2000:SRM

[165] Gregory Karvounarakis and Sarantos Kapidakis. Submission and repository management of digital libraries, using WWW. Computer Networks (Amsterdam, Netherlands: 1999), 34(6): 861–872, December ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/32/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/32/29/article.pdf.

Deri:2000:PNS

[166] Luca Deri and Stefano Suin. Practical network security: experiences with ntop. Computer Networks (Amsterdam, Netherlands: 1999), 34(6): 873–880, December ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/32/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/32/30/article.pdf.

Baltatu:2000:SIC

[167] Madalina Baltatu, Antonio Lioy, Fabio Maino, and Daniele Mazzocchi. Security issues in control, management and routing protocols. Computer Networks (Amsterdam, Netherlands: 1999), 34 (6):881–894, December ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/32/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/32/31/article.pdf.

Hommes:2000:DMA

[168] Ferdinand Hommes, Eva Pless, and Lothar Zier. Did MPOA achieve its objective? Computer Networks (Amsterdam, Netherlands: 1999), 34(6): 895–904, December ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/32/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/32/32/article.pdf.

Ferrari:2000:EEP

[169] Tiziana Ferrari. End-to-end performance analysis with traffic aggregation. Computer Networks (Amsterdam, Netherlands: 1999), 34(6): 905–914, December ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/32/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/32/33/article.pdf.

Sander:2000:DSI

[170] Volker Sander, Ian Foster, Alain Roy, and Linda Winkler. A differentiated services implementation for high-performance TCP flows. Computer Networks (Amsterdam, Netherlands: 1999), 34(6):915–929, December ??, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/32/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/32/34/article.pdf.

Boudier:2000:VSV

[171] Thomas Boudier and David M. Shotton. VIDOS, a system for video editing and format conversion over the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 34(6): 931–944, December ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/32/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/32/35/article.pdf.

Balaouras:2000:PLT

[172] Pantelis Balaouras, Ioannis Stavrakakis, and Lazaros Merakos. Potential and limitations of a teleteaching environment based on H.323 audio-visual communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 34(6):945–958, December ??, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/32/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/32/36/article.pdf.

Hoffmann:2000:GWG

[173] Gertraud Hoffmann. G-WiN — the Gbit/s infrastructure for the German scientific community. Computer Networks (Amsterdam, Netherlands: 1999), 34(6):959–964, December ??, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/49/
32/37/abstract.html;   http://www.
elsevier.nl/gej-ng/10/15/22/49/
32/37/article.pdf.

Milanovic:2000:AAP

[174] Stanislav Milanovic and Alessandro Maglianella. ATM over ADSL probe in Telecom Italia environment. Computer Networks (Amsterdam, Netherlands: 1999), 34(6): 965–980, December ??, 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
49/32/38/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/49/32/38/article.pdf.

Anonymous:2000:Ie

[175] Anonymous.  Index.  Computer

Networks (Amsterdam, Netherlands: 1999), 34(6):981–983, December ??, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/15/22/49/32/39/abstract.html;
http://www.elsevier.nl/gej-ng/
10/15/22/49/32/39/article.pdf.

Anonymous:2000:If

[176] Anonymous.  Index.  Computer

Networks (Amsterdam, Netherlands: 1999), 34(6):984–986, December ??, 2000. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/15/22/49/32/40/abstract.html;
http://www.elsevier.nl/gej-ng/
10/15/22/49/32/40/article.pdf.

Lupu:2001:STN

[177] Emil Lupu, Subrata Mazumdar, and Rolf Stadler. Selected topics in network and systems management. Computer Networks (Amsterdam, Netherlands: 1999), 35(1):1–4, January ??, 2001.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/24/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/24/24/article.pdf.

Bhoj:2001:SMF

[178] P. Bhoj, S. Singhal, and S. Chutani. SLA management in federated environments. Computer Networks (Amsterdam, Netherlands: 1999), 35(1): 5–24, January ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/24/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/24/25/article.pdf.

Feridun:2001:FDM

[179] M. Feridun and J. Krause. A framework for distributed management with mobile components. Computer Networks (Amsterdam, Netherlands: 1999), 35(1):25–38, January ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/24/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/24/26/article.pdf.

Schonwalder:2001:SIM

[180] Jürgen Schönwälder and Jürgen Quittek. Secure Internet management by delegation. Computer Networks (Amsterdam, Netherlands: 1999), 35 (1):39–56, January ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/24/27/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/15/
22/52/24/27/article.pdf.

Florissi:2001:QNE

[181] P. G. S. Florissi, Y. Yemini, and D. Florissi. QoSockets: a new extension to the sockets API for end-to-end application QoS management. Computer Networks (Amsterdam, Netherlands: 1999), 35(1):57–76, January ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/24/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/24/28/article.pdf.

Hellerstein:2001:SAP

[182] Joseph L. Hellerstein, Fan Zhang, and Perwez Shahabuddin. A statistical approach to predictive detection. Computer Networks (Amsterdam, Netherlands: 1999), 35(1): 77–95, January ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/24/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/24/29/article.pdf.

Polykalas:2001:ACC

[183] Spyros E. Polykalas, George T. Kolyvas, and Iakovos S. Venieris. An adaptive congestion control mechanism for intelligent broadband networks. Computer Networks (Amsterdam, Netherlands: 1999), 35(2– 3):97–115, February ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/27/24/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/27/24/article.pdf.

Sharon:2001:CCC

[184] Oran Sharon and Michael Spratt. A CSMA/CD compatible MAC for realtime transmissions based on varying collision intervals. Computer Networks (Amsterdam, Netherlands: 1999), 35 (2–3):117–142, February ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        52/27/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/27/25/article.pdf.

Venugopal:2001:HPL

[185] K. R. Venugopal, M. Shiva Kumar, and P. Sreenivasa Kumar. A heuristic for placement of limited range wavelength converters in all-optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 35(2–3): 143–163, February ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        52/27/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/27/26/article.pdf.

Postigo-Boix:2001:ITF

[186] Marcos Postigo-Boix, Joan García-Haro, and Mónica Aguilar-Igartua. IMA: technical foundations, application and performance analysis. Computer Networks (Amsterdam, Netherlands: 1999), 35(2–3): 165–183, February ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/

52/27/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/27/27/article.pdf.

Liu:2001:IEC

[187] Chunlei Liu and Raj Jain. Improving explicit congestion notification with the mark-front strategy. Computer Networks (Amsterdam, Netherlands: 1999), 35(2–3): 185–201, February ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/27/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/27/28/article.pdf.

Li:2001:SPS

[188] Z. G. Li, X. J. Yuan, C. Y. Wen, and B. H. Soong. A switched priority scheduling mechanism for ATM switches with multi-class output buffers. Computer Networks (Amsterdam, Netherlands: 1999), 35(2– 3):203–221, February ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/27/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/27/29/article.pdf.

Desmedt:2001:BAJ

[189] Yvo Desmedt et al. Broadcast antijamming systems. Computer Networks (Amsterdam, Netherlands: 1999), 35 (2–3):223–236, February ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/27/30/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/15/
22/52/27/30/article.pdf.

Lee:2001:CPA

[190] David Lee, K. K. Ramakrishnan, and W. Melody Moh. Correctness and performance of the ATM ABR rate control scheme. Computer Networks (Amsterdam, Netherlands: 1999), 35(2– 3):237–261, February ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/27/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/27/31/article.pdf.

DeCnodder:2001:RAS

[191] Stefaan De Cnodder, Omar Elloumi, and Kenny Pauwels. Rate adaptive shaping for the efficient transport of data traffic in diffserv networks. Computer Networks (Amsterdam, Netherlands: 1999), 35(2–3): 263–285, February ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/27/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/27/32/article.pdf.

Tsao:2001:POD

[192] Shih-Chiang Tsao and Ying-Dar Lin. Pre-order Deficit Round Robin: a new scheduling algorithm for packet-switched networks. Computer Networks (Amsterdam, Netherlands: 1999), 35(2–3):287– 305, February ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/

52/27/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/27/33/article.pdf.

Kalyanasundaram:2001:ACS

[193] Suresh Kalyanasundaram, Edwin K. P. Chong, and Ness B. Shroff. Admission control schemes to provide class-level QoS in multiservice networks. Computer Networks (Amsterdam, Netherlands: 1999), 35(2–3): 307–326, February ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/27/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/27/34/article.pdf.

Cayirci:2001:PBA

[194] Erdal Cayirci and Cem Ersoy. A PCS based architecture for tactical mobile communications. Computer Networks (Amsterdam, Netherlands: 1999), 35 (2–3):327–350, February ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/27/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/27/35/article.pdf.

Shirmohammadi:2001:EEC

[195] Shervin Shirmohammadi and Nicolas D. Georganas. An end-to-end communication architecture for collaborative virtual environments. Computer Networks (Amsterdam, Netherlands: 1999), 35(2–3):351– 367, February ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/

                                        52/27/36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/27/36/article.pdf.

Anonymous:2001:CPa

[196] Anonymous. Call for Papers. Computer Networks (Amsterdam, Netherlands: 1999), 35(2–3):369– 370, February ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        52/27/37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/27/37/article.pdf.

Anonymous:2001:E

[197] Anonymous. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 35(4):371, March ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        52/30/24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/30/24/article.pdf.

Husemann:2001:PCH

[198] Dirk Husemann. Pervasive computing: Hogwarts, StarTrek, reality and back. Computer Networks (Amsterdam, Netherlands: 1999), 35 (4):373–375, March ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        52/30/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/30/25/article.pdf.

Jones:2001:PEN

[199] Alan Jones and Andrew Hopper. The prototype embedded network (PEN). Computer Networks (Amsterdam, Netherlands: 1999), 35(4): 377–390, March ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/30/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/30/26/article.pdf.

Hum:2001:FAN

[200] Alex P. J. Hum. Fabric area network — a new wireless communications infrastructure to enable ubiquitous networking and sensing on intelligent clothing. Computer Networks (Amsterdam, Netherlands: 1999), 35 (4):391–399, March ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/30/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/30/27/article.pdf.

Beigl:2001:MED

[201] Michael Beigl, Hans-W. Gellersen, and Albrecht Schmidt. Mediacups: experience with design and use of computer-augmented everyday artifacts. Computer Networks (Amsterdam, Netherlands: 1999), 35(4): 401–409, March ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/30/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/30/28/article.pdf.

Hermann:2001:DTA

[202] Reto Hermann et al. DEAPspace — Transient ad hoc networking of pervasive devices. Computer Networks (Amsterdam, Netherlands: 1999), 35 (4):411–428, March ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        52/30/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/30/29/article.pdf.

Glance:2001:PUP

[203] Natalie Glance, Dave Snowdon, and Jean-Luc Meunier. Pollen: using people as a communication medium. Computer Networks (Amsterdam, Netherlands: 1999), 35(4): 429–442, March ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        52/30/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/30/30/article.pdf.

Kindberg:2001:WBN

[204] Tim Kindberg and John Barton. A Web-based nomadic computing system. Computer Networks (Amsterdam, Netherlands: 1999), 35(4): 443–456, March ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        52/30/31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/30/31/article.pdf.

Lehman:2001:HDC

[205] Tobin J. Lehman et al. Hitting the distributed computing sweet spot with TSpaces. Computer Networks (Amsterdam, Netherlands: 1999), 35 (4):457–472, March ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/30/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/30/32/article.pdf.

Gribble:2001:NAR

[206] Steven D. Gribble et al. The Ninja architecture for robust Internet-scale systems and services. Computer Networks (Amsterdam, Netherlands: 1999), 35 (4):473–497, March ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/30/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/30/33/article.pdf.

Anonymous:2001:AR

[207] Anonymous. Acknowledgement to reviewers. Computer Networks (Amsterdam, Netherlands: 1999), 35(4): 499–502, March ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/30/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/30/34/article.pdf.

Magedanz:2001:VDI

[208] Thomas Magedanz and Mikhail Smirnov. Voice/data integration — a snapshot of intelligent networks and Internet convergence. Computer Networks (Amsterdam, Netherlands: 1999), 35(5):503–505, April ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/31/24/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/31/24/article.pdf.

Zhu:2001:IMM

[209] Xiaomin Zhu, Jianxin Liao, and Junliang Chen. IIN model: modifications and case study. Computer Networks (Amsterdam, Netherlands: 1999), 35 (5):507–519, April ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        52/31/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/31/25/article.pdf.

Chapron:2001:ACM

[210] Jean-Emmanuel Chapron and Bruno Chatras. An analysis of the IN call model suitability in the context of VoIP. Computer Networks (Amsterdam, Netherlands: 1999), 35(5): 521–535, April ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        52/31/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/31/26/article.pdf.

Licciardi:2001:AIH

[211] C. A. Licciardi, G. Canal, A. Andreetto, and P. Lago. An architecture for IN-Internet hybrid services. Computer Networks (Amsterdam, Netherlands: 1999), 35(5): 537–549, April ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        52/31/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/31/27/article.pdf.

Glitho:2001:EAT

[212] Roch H. Glitho. Emerging alternatives to today’s advanced service architectures for Internet telephony: IN and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 35 (5):551–563, April ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/31/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/31/28/article.pdf.

Perdikeas:2001:PBS

[213] Menelaos K. Perdikeas and Iakovos S. Venieris. Parlay-based service engineering in a converged Internet-PSTN environment. Computer Networks (Amsterdam, Netherlands: 1999), 35 (5):565–578, April ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/31/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/31/29/article.pdf.

Anjum:2001:CSR

[214] Farooq Anjum, Francesco Caruso, Ravi Jain, Paolo Missier, and Adalberto Zordan. CitiTime: a system for rapid creation of portable next-generation telephony services. Computer Networks (Amsterdam, Netherlands: 1999), 35 (5):579–595, April ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
52/31/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/52/31/30/article.pdf.

Kantola:2001:IRI

[215] Raimo Kantola, Jose Costa Requena, and Nicklas Beijar. Interoperable routing for IN and IP Telephony. Computer Networks (Amsterdam, Netherlands: 1999), 35(5):597–609, April ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        52/31/31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/52/31/31/article.pdf.

Dssouli:2001:MSP

[216] Rachida Dssouli and Yair Lahav. MSC and SDL in project life cycles. Computer Networks (Amsterdam, Netherlands: 1999), 35(6):611–612, May ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.nl/gej%
                                        2Dng/10/15/22/52/32/25/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej%2Dng/10/15/22/52/32/25/article.
                                        pdf.

Andersson:2001:UUS

[217] Michael Andersson, Anders Ek, and Niklas Landin. Utilizing UML in SDL-based development. Computer Networks (Amsterdam, Netherlands: 1999), 35(6):613–625, May ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.nl/gej%
                                        2Dng/10/15/22/52/32/26/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej%2Dng/10/15/22/52/32/26/article.
                                        pdf.

Geppert:2001:SPA

[218] Birgit Geppert and Frank Rößler. The SDL pattern approach — a reusedriven SDL design methodology. Computer Networks (Amsterdam, Netherlands: 1999), 35(6):627–645, May ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/52/32/27/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/52/32/27/article.
pdf.

Hintelmann:2001:ATT

[219] Jörg Hintelmann, Richard Hofmann, Frank Lemmen, Andreas Mitschele-Thiel, and Bruno Müller-Clostermann. Applying techniques and tools for the performance engineering of SDL systems. Computer Networks (Amsterdam, Netherlands: 1999), 35 (6):647–665, May ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.nl/gej%2Dng/
10/15/22/52/32/28/abstract.html;
http://www.elsevier.nl/gej%2Dng/
10/15/22/52/32/28/article.pdf.

Mansurov:2001:ITM

[220] Nikolai N. Mansurov and Robert L. Probert. Improving time-to-market using SDL tools and techniques. Computer Networks (Amsterdam, Netherlands: 1999), 35(6):667–691, May ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/52/32/29/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/52/32/29/article.
pdf.

Bourhfir:2001:TCS

[221] C. Bourhfir, E. Aboulhamid, F. Khendek, and R. Dssouli. Test cases selection from SDL specifications. Computer Networks (Amsterdam, Netherlands: 1999), 35(6):693–708, May ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/52/32/30/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/52/32/30/article.
pdf.

Reed:2001:NSN

[222] Rick Reed. Notes on SDL-2000 for the new millennium. Computer Networks (Amsterdam, Netherlands: 1999), 35(6):709–720, May ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/52/32/31/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/52/32/31/article.
pdf.

Haugen:2001:MID

[223] Ø. Haugen. MSC-2000 interaction diagrams for the new millennium. Computer Networks (Amsterdam, Netherlands: 1999), 35(6):721–732, May ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/52/32/32/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/52/32/32/article.
pdf.

Anonymous:2001:AIa

[224] Anonymous. Author Index. Computer Networks (Amsterdam, Netherlands: 1999), 35(6):733–735, May ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/52/32/33/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/52/32/33/article.
pdf.

Anonymous:2001:SIa

[225] Anonymous. Subject Index. Computer Networks (Amsterdam, Netherlands: 1999), 35(6):737–739, May ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/52/32/34/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/52/32/34/article.
pdf.

Crowcroft:2001:ANS

[226] J. Crowcroft et al. Active networks and services. Computer Networks (Amsterdam, Netherlands: 1999), 36(1):1–3, June ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/25/25/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/25/25/article.
pdf.

Ghosh:2001:IAL

[227] Atanu Ghosh, Michael Fry, and Glen MacLarty. An infrastructure for application level active networking. Computer Networks (Amsterdam, Netherlands: 1999), 36(1):5–20, June ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/25/26/abstract.
html;   http://www.elsevier.nl/

                                        gej%2Dng/10/15/22/61/25/26/article.
                                        pdf.

Kornblum:2001:API

[228] Jessica A. Kornblum, Danny Raz, and Yuval Shavitt. The active process interaction with its environment. Computer Networks (Amsterdam, Netherlands: 1999), 36(1):21–34, June ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.nl/gej%
                                        2Dng/10/15/22/61/25/27/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej%2Dng/10/15/22/61/25/27/article.
                                        pdf.

Fernando:2001:DNS

[229] Amila Fernando, Darren Williams, Alan Fekete, and Bob Kummerfeld. Dynamic network service installation in an active network. Computer Networks (Amsterdam, Netherlands: 1999), 36(1):35–48, June ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.nl/gej%
                                        2Dng/10/15/22/61/25/28/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej%2Dng/10/15/22/61/25/28/article.
                                        pdf.

Villela:2001:VPR

[230] Daniel Villela, Andrew T. Campbell, and John Vicente. Virtuosity: Programmable resource management for spawning networks. Computer Networks (Amsterdam, Netherlands: 1999), 36(1):49–73, June ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.nl/gej%
                                        2Dng/10/15/22/61/25/29/abstract.

html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/25/29/article.
pdf.

Marshall:2001:PQS

[231] Ian W. Marshall and Chris Roadknight. Provision of quality of service for active services. Computer Networks (Amsterdam, Netherlands: 1999), 36(1):75–85, June ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/25/30/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/25/30/article.
pdf.

Karnouskos:2001:SII

[232] Stamatis Karnouskos. Security implications of implementing active network infrastructures using agent technology. Computer Networks (Amsterdam, Netherlands: 1999), 36 (1):87–100, June ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.nl/gej%2Dng/
10/15/22/61/25/31/abstract.html;
http://www.elsevier.nl/gej%2Dng/
10/15/22/61/25/31/article.pdf.

Wakeman:2001:SLB

[233] Ian Wakeman, Alan Jeffrey, Tim Owen, and Damyan Pepper. SafetyNet: A language-based approach to programmable networks. Computer Networks (Amsterdam, Netherlands: 1999), 36(1):101–114, June ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/25/32/abstract.

html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/25/32/article.
pdf.

Touch:2001:ON

[234] Joe Touch. Overlay networks. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):115–116, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/25/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/25/article.
pdf.

Touch:2001:DIO

[235] Joe Touch. Dynamic Internet overlay deployment and management using the X-Bone. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):117–135, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/26/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/26/article.
pdf.

Lim:2001:CVP

[236] L. Keng Lim et al. Customizable virtual private network service with QoS. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):137–151, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/27/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/27/article.
pdf.

Su:2001:VAN

[237] Gong Su and Yechiam Yemini. Virtual active networks: towards multiedged network computing. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):153–168, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/28/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/28/article.
pdf.

Ayedemir:2001:TTN

[238] M. Ayedemir et al. Two tools for network traffic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):169–179, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/29/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/29/article.
pdf.

Soh:2001:MNN

[239] Wee-Seng Soh and Chen-Khong Tham. Modular neural networks for multi-service connection admission control. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):181–202, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/30/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/30/article.
pdf.

Aweya:2001:CTA

[240] James Aweya, Michel Ouellette, and Delfin Y. Montuno. A control theoretic approach to active queue management. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):203–235, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/31/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/31/article.
pdf.

Joo:2001:AST

[241] Changhee Joo and Saewoong Bahk. Analysis of start-up transition dynamics of TCP NewReno. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):237–250, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/32/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/32/article.
pdf.

Korkmaz:2001:RAF

[242] Turgay Korkmaz and Marwan Krunz. A randomized algorithm for finding a path subject to multiple QoS requirements. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):251–268, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/33/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/33/article.
pdf.

Chang:2001:PHB

[243] Ming-Feng Chang, Yi-Bing Lin, and Wei-Zu Yang. Performance of hot billing mobile prepaid service. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):269–290, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/34/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/34/article.
pdf.

Yasumoto:2001:CIL

[244] Keiichi Yasumoto, Teruo Higashino, and Kenichi Taniguchi. A compiler to implement LOTOS specifications in distributed environments. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):291–310, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/35/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/35/article.
pdf.

Gonzalez:2001:OWN

[245] J. González et al. Optimization of Web newspaper layout in real time. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):311–321, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/36/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/36/article.
pdf.

Xiao:2001:PAV

[246] Xiaojun Xiao, Yong Huat Chew, Winston K. G. Seah, and Chi Chung Ko. Performance analysis for voice and data integration in hybrid fiber/coax networks. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):323–341, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/37/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/37/article.
pdf.

Dolev:2001:SAF

[247] Shlomi Dolev, Boris Fitingof, Avraham Melkman, and Olga Tubman. Smooth and adaptive forward erasure correcting. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):343–355, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/38/abstract.
html;   http://www.elsevier.nl/
gej%2Dng/10/15/22/61/27/38/article.
pdf.

Yener:2001:SBA

[248] Bülent Yener, Gong Su, and Eran Gabber. Smart box architecture: a hybrid solution for IP QoS provisioning. Computer Networks (Amsterdam, Netherlands: 1999), 36(2–3):357–375, July ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.nl/gej%
2Dng/10/15/22/61/27/39/abstract.
html;   http://www.elsevier.nl/

gej%2Dng/10/15/22/61/27/39/article.
pdf.

Domingo-Ferrer:2001:CDS

[249] Josep Domingo-Ferrer and Pieter Hartel. Current directions in smart cards. Computer Networks (Amsterdam, Netherlands: 1999), 36(4): 377–379, July 16, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/28/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/28/25/article.pdf.

Praca:2001:SCS

[250] Denis Praca and Claude Barral. From smart cards to smart objects: the road to new smart technologies. Computer Networks (Amsterdam, Netherlands: 1999), 36(4):381–389, July 16, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/28/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/28/26/article.pdf.

Attali:2001:IDE

[251] Isabelle Attali, Denis Caromel, Carine Courbis, Ludovic Henrio, and Henrik Nilsson. An integrated development environment for Java Card. Computer Networks (Amsterdam, Netherlands: 1999), 36(4):391–405, July 16, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/28/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/28/27/article.pdf.

Poll:2001:FSJ

[252] Erik Poll, Joachim van den Berg, and Bart Jacobs. Formal specification of the JavaCard API in JML: the APDU class. Computer Networks (Amsterdam, Netherlands: 1999), 36 (4):407–421, July 16, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        61/28/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/61/28/28/article.pdf.

Borst:2001:CSC

[253] Johan Borst, Bart Preneel, and Vincent Rijmen. Cryptography on smart cards. Computer Networks (Amsterdam, Netherlands: 1999), 36(4): 423–435, July 16, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        61/28/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/61/28/29/article.pdf.

Guillou:2001:CAP

[254] L. C. Guillou, M. Ugon, and J-J. Quisquater. Cryptographic authentication protocols for smart cards. Computer Networks (Amsterdam, Netherlands: 1999), 36(4): 437–451, July 16, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        61/28/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/61/28/30/article.pdf.

MRaihi:2001:CAS

[255] David M’Raïhi and Moti Yung. E-commerce applications of smart cards. Computer Networks (Amsterdam, Netherlands: 1999), 36(4): 453–472, July 16, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/28/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/28/31/article.pdf.

Husemann:2001:SSC

[256] Dirk Husemann. Standards in the smart card world. Computer Networks (Amsterdam, Netherlands: 1999), 36 (4):473–487, July 16, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/28/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/28/32/article.pdf.

Anonymous:2001:CPb

[257] Anonymous. Call for Papers. Computer Networks (Amsterdam, Netherlands: 1999), 36(4):489–490, July 16, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/28/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/28/33/article.pdf.

Dermler:2001:EN

[258] Gabriel Dermler and Beat Liver. The economics of networking. Computer Networks (Amsterdam, Netherlands: 1999), 36(5–6):491–492, August ??, 2001. CODEN ????

ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/61/
29/25/abstract.html;   http://www.
elsevier.nl/gej-ng/10/15/22/61/
29/25/article.pdf.

Odlyzko:2001:IPH

[259] Andrew Odlyzko. Internet pricing and the history of communications. Computer Networks (Amsterdam, Netherlands: 1999), 36(5– 6):493–517, August ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/29/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/29/26/article.pdf.

Altmann:2001:HCN

[260] Jörn Altmann and Karyen Chu. How to charge for network services — flat-rate or usage-based? Computer Networks (Amsterdam, Netherlands: 1999), 36 (5–6):519–531, August ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/29/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/29/27/article.pdf.

Kenyon:2001:SMT

[261] Chris Kenyon and Giorgos Cheliotis. Stochastic models for telecom commodity prices. Computer Networks (Amsterdam, Netherlands: 1999), 36 (5–6):533–555, August ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/29/28/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/15/
22/61/29/28/article.pdf.

Wong:2001:MTD

[262] Allan K. Y. Wong, Tharam S. Dillon, Wilfred W. K. Lin, and May T. W. Ip. M^{2}RT: A tool developed for predicting the mean message response time of communication channels in sizeable networks exemplified by the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 36 (5–6):557–577, August ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/29/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/29/29/article.pdf.

Lombardo:2001:STM

[263] Alfio Lombardo, Giacomo Morabito, and Giovanni Schembra. Statistical traffic modeling and guaranteed service disciplines: a performance evaluation paradigm. Computer Networks (Amsterdam, Netherlands: 1999), 36 (5–6):579–595, August ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/29/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/29/30/article.pdf.

Fahmy:2001:RED

[264] Hossam Mahmoud Ahmad Fahmy. Reliability evaluation in distributed computing environments using the AHP. Computer Networks (Amsterdam, Netherlands: 1999), 36(5– 6):597–615, August ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        61/29/31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/61/29/31/article.pdf.

Koster:2001:IAC

[265] Gerta Köster. Improving the automatic congestion control functionality in SS7-signaling networks. Computer Networks (Amsterdam, Netherlands: 1999), 36(5–6):617– 624, August ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        61/29/32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/61/29/32/article.pdf.

Bellotti:2001:WFS

[266] F. Bellotti, A. De Gloria, D. Grosso, and L. Noli. WLESS-frame: a simulation-based development environment for 802.11 stations. Computer Networks (Amsterdam, Netherlands: 1999), 36(5–6):625–641, August ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/15/22/61/
                                        29/33/abstract.html;   http://www.
                                        elsevier.nl/gej-ng/10/15/22/61/
                                        29/33/article.pdf.

Yaghmaee:2001:NFB

[267] Mohammad Hossien Yaghmaee, Mohammad Bagher Menhaj, and Mostafa Safavi. A novel FLC-based approach for ATM traffic control. Computer Networks (Amsterdam, Netherlands: 1999), 36(5–6):643– 658, August ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/29/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/29/34/article.pdf.

Basagni:2001:LAD

[268] Stefano Basagni, Imrich Chlamtac, and Violet R. Syrotiuk. Location aware, dependable multicast for mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 36(5– 6):659–670, August ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/29/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/29/35/article.pdf.

Drira:2001:TAC

[269] Khalil Drira, Pierre Azéma, and Pierre de Saqui Sannes. Testability analysis in communicating systems. Computer Networks (Amsterdam, Netherlands: 1999), 36(5– 6):671–693, August ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
61/29/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/61/29/36/article.pdf.

Nakayama:2001:OID

[270] Marvin K. Nakayama and Bülent Yener. Optimal information dispersal for probabilistic latency targets. Computer Networks (Amsterdam, Netherlands: 1999), 36(5–6): 695–707, August ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        61/29/37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/61/29/37/article.pdf.

Aweya:2001:LAM

[271] James Aweya, Michel Ouellette, Delfin Y. Montuno, and Alan Chapman. A load adaptive mechanism for buffer management. Computer Networks (Amsterdam, Netherlands: 1999), 36(5–6):709–728, August ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/15/22/61/
                                        29/38/abstract.html;   http://www.
                                        elsevier.nl/gej-ng/10/15/22/61/
                                        29/38/article.pdf.

Anonymous:2001:AIb

[272] Anonymous. Author Index. Computer Networks (Amsterdam, Netherlands: 1999), 36(5–6):729–731, August ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/15/22/61/
                                        29/39/abstract.html;   http://www.
                                        elsevier.nl/gej-ng/10/15/22/61/
                                        29/39/article.pdf.

Anonymous:2001:SIb

[273] Anonymous. Subject Index. Computer Networks (Amsterdam, Netherlands: 1999), 36(5–6):733–734, August ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/15/22/61/
                                        29/40/abstract.html;   http://www.

elsevier.nl/gej-ng/10/15/22/61/
29/40/article.pdf.

Grillo:2001:WN

[274] Davide Grillo. Guest editorial: Wireless networking. Computer Networks (Amsterdam, Netherlands: 1999), 37 (1):1–4, September ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/27/27/abstract.html.

Lenzini:2001:PEC

[275] L. Lenzini and E. Mingozzi. Performance evaluation of capacity request and allocation mechanisms for HiperLAN2 wireless LANs. Computer Networks (Amsterdam, Netherlands: 1999), 37(1):5–15, September ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/27/28/abstract.html.

Kadelka:2001:SIQ

[276] Arndt Kadelka and Arno Masella. Serving IP quality of service with HiperLAN/2. Computer Networks (Amsterdam, Netherlands: 1999), 37 (1):17–24, September ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/27/29/abstract.html.

Esseling:2001:FCH

[277] Norbert Esseling, Harbinder S. Vandra, and Bernhard Walke. A forwarding concept for HiperLAN/2. Computer Networks (Amsterdam, Netherlands: 1999), 37(1):25–32, September ??, 2001. CODEN ????

ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
27/30/abstract.html.

Bruno:2001:SPD

[278] Raffaele Bruno, Marco Conti, and Enrico Gregori. A simple protocol for the dynamic tuning of the backoff mechanism in IEEE 802.11 networks. Computer Networks (Amsterdam, Netherlands: 1999), 37(1): 33–44, September ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/27/31/abstract.html.

Dunlop:2001:QMD

[279] John Dunlop, Gwenaël Le Bodic, James Irvine, and Demessie Girma. QoS management with dynamic bearer selection schemes. Computer Networks (Amsterdam, Netherlands: 1999), 37 (1):45–53, September ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/27/32/abstract.html.

Fikouras:2001:PCH

[280] N. A. Fikouras and C. Görg. Performance comparison of hinted- and advertisement-based movement detection methods for mobile IP handoffs. Computer Networks (Amsterdam, Netherlands: 1999), 37(1):55– 62, September ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
27/33/abstract.html.

Habetha:2001:OCM

[281] Jörg Habetha and Martin Nadler. Outline of a centralised multihop ad hoc wireless network. Computer Networks (Amsterdam, Netherlands: 1999), 37 (1):63–71, September ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/27/34/abstract.html.

Xu:2001:DIS

[282] Bangnan Xu and Bernhard Walke. Design issues of self-organizing broadband wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 37 (1):73–81, September ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/27/35/abstract.html.

Aretz:2001:FWC

[283] Kurt Aretz, Martin Haardt, Walter Konhäuser, and Werner Mohr. The future of wireless communications beyond the third generation. Computer Networks (Amsterdam, Netherlands: 1999), 37(1):83– 92, September ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
27/36/abstract.html.

Lamersdorf:2001:EBS

[284] Winfried Lamersdorf, Andreas Bartelt, Dietrich Fahrenholtz, and Mark Tuan Tu. Guest editorial: Electronic business systems. Computer Networks (Amsterdam, Netherlands: 1999), 37 (2):93–94, October ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/33/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/67/33/27/article.pdf.

Piccinelli:2001:DSA

[285] Giacomo Piccinelli, Giuliano Di Vitantonio, and Leonid Mokrushin. Dynamic service aggregation in electronic marketplaces. Computer Networks (Amsterdam, Netherlands: 1999), 37 (2):95–109, October ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/33/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/67/33/28/article.pdf.

Hoffner:2001:CDC

[286] Yigal Hoffner, Simon Field, Paul Grefen, and Heiko Ludwig. Contractdriven creation and operation of virtual enterprises. Computer Networks (Amsterdam, Netherlands: 1999), 37 (2):111–136, October ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/33/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/67/33/29/article.pdf.

Bichler:2001:MDN

[287] Martin Bichler and Arie Segev. Methodologies for the design of negotiation protocols on E-markets. Computer Networks (Amsterdam, Netherlands: 1999), 37(2):137–152, October ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
33/30/abstract.html;   http://www.
elsevier.nl/gej-ng/10/15/22/67/
33/30/article.pdf.

Schoop:2001:DCF

[288] Mareike Schoop and Christoph Quix. DOC.COM: a framework for effective negotiation support in electronic marketplaces. Computer Networks (Amsterdam, Netherlands: 1999), 37(2): 153–170, October ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/33/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/67/33/31/article.pdf.

Grimm:2001:BTF

[289] Rüdiger Grimm and Peter Ochsenschläger. Binding telecooperation — a formal model for electronic commerce. Computer Networks (Amsterdam, Netherlands: 1999), 37(2): 171–193, October ??, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/33/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/15/
22/67/33/32/article.pdf.

Choi:2001:GAB

[290] Samuel P. M. Choi, Jiming Liu, and Sheung-Ping Chan. A genetic agent-based negotiation system. Computer Networks (Amsterdam, Netherlands: 1999), 37(2):195– 204, October ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/

                                        67/33/33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/67/33/33/article.pdf.

Clauss:2001:IMS

[291] Sebastian Clauß and Marit Köhntopp. Identity management and its support of multilateral security. Computer Networks (Amsterdam, Netherlands: 1999), 37(2):205–219, October ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/15/22/67/
                                        33/34/abstract.html;   http://www.
                                        elsevier.nl/gej-ng/10/15/22/67/
                                        33/34/article.pdf.

Tsalgatidou:2001:BMT

[292] Aphrodite Tsalgatidou and Evaggelia Pitoura. Business models and transactions in mobile electronic commerce: requirements and properties. Computer Networks (Amsterdam, Netherlands: 1999), 37(2):221– 236, October ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/33/35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/15/
                                        22/67/33/35/article.pdf.

Morabito:2001:CCA

[293] Giacomo Morabito and Sergio Palazzo. Congestion control for ABR traffic in satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 37 (3–4):237–249, November 5, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/34/27/abstract.html.

Chen:2001:FSM

[294] Y. T. Chen and K. H. Lee. A flexible service model for advance reservation. Computer Networks (Amsterdam, Netherlands: 1999), 37(3–4): 251–262, November 5, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/34/28/abstract.html.

Zhang:2001:ITM

[295] Liren Zhang and Li Zheng. IPv6 traffic with multi-class QoS in VPN. Computer Networks (Amsterdam, Netherlands: 1999), 37(3–4):263– 275, November 5, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
34/29/abstract.html.

Qiu:2001:UPM

[296] Lili Qiu, Yin Zhang, and Srinivasan Keshav. Understanding the performance of many TCP flows. Computer Networks (Amsterdam, Netherlands: 1999), 37(3–4):277– 306, November 5, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
34/30/abstract.html.

Evans:2001:RLS

[297] M. P. Evans and S. M. Furnell. The Resource Locator Service: fixing a flaw in the web. Computer Networks (Amsterdam, Netherlands: 1999), 37(3– 4):307–330, November 5, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/34/31/abstract.html.

Bruni:2001:ORC

[298] Carlo Bruni and Caterina Scoglio. An optimal rate control algorithm for guaranteed services in broadband networks. Computer Networks (Amsterdam, Netherlands: 1999), 37(3–4): 331–344, November 5, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/34/32/abstract.html.

Marino:2001:PFS

[299] P. Mariño, J. Nogueira, C. Sigüenza, F. Poza, and M. Domínguez. The PROFIBUS formal specification: a comparison between two FDTs. Computer Networks (Amsterdam, Netherlands: 1999), 37(3–4):345– 362, November 5, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/15/22/67/
                                        34/33/abstract.html.

Abdelaziz:2001:ARC

[300] Mohamed Abdelaziz and Ioannis Stavrakakis. Adaptive rate control in high-speed networks: performance issues. Computer Networks (Amsterdam, Netherlands: 1999), 37(3–4): 363–382, November 5, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/34/34/abstract.html.

Shao:2001:BVM

[301] Zuhua Shao. Batch verifying multiple DSA-type digital signatures.

Computer Networks (Amsterdam, Netherlands: 1999), 37(3–4):383– 389, November 5, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
34/35/abstract.html.

Todd:2001:CPA

[302] Robert W. Todd, Matthew C. Chidester, and Alan D. George. Comparative performance analysis of directed flow control for real-time SCI. Computer Networks (Amsterdam, Netherlands: 1999), 37(3–4): 391–406, November 5, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/34/36/abstract.html.

VanMieghem:2001:HHQ

[303] Piet Van Mieghem, Hans De Neve, and Fernando Kuipers. Hop-by-hop quality of service routing. Computer Networks (Amsterdam, Netherlands: 1999), 37 (3–4):407–423, November 5, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/34/37/abstract.html.

Curcio:2001:QEM

[304] Igor D. D. Curcio, Ville Lappalainen, and Miraj-E-Mostafa. QoS evaluation of 3G-324M mobile videophones over WCDMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 37(3–4):425– 445, November 5, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
34/38/abstract.html.

Patek:2001:SAR

[305] Stephen D. Patek, Raja Venkateswaran, and Jörg Liebeherr. Simple alternate routing for differentiated services networks. Computer Networks (Amsterdam, Netherlands: 1999), 37(3– 4):447–466, November 5, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/34/39/abstract.html.

Zhao:2001:CTM

[306] Yongxiang Zhao and Chang jia Chen. Coupon TFRC: a mechanism being friendly to both TCP and continuous stream. Computer Networks (Amsterdam, Netherlands: 1999), 37(3– 4):467–479, November 5, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/34/40/abstract.html.

Shankland:2001:CSA

[307] Carron Shankland and Alberto Verdejo. A case study in abstraction using E-LOTOS and the FireWire. Computer Networks (Amsterdam, Netherlands: 1999), 37(3–4):481– 502, November 5, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/15/22/67/
                                        34/41/abstract.html.

Marsan:2001:GE

[308] Marco Ajmone Marsan and Andrea Bianco. Guest editorial. Computer Networks (Amsterdam, Netherlands: 1999), 37(5):503–505, November 22, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
35/27/abstract.html.

Yuan:2001:DIS

[309] Ping Yuan, Julie Schlembach, Anders Skoe, and Edward Knightly. Design and implementation of scalable edge-based admission control. Computer Networks (Amsterdam, Netherlands: 1999), 37(5):507–518, November 22, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
35/28/abstract.html.

Borella:2001:ISM

[310] Andrea Borella, Giovanni Cancellieri, Elena Pagani, and Gian Paolo Rossi. Implementation schemes for multicast bandwidth brokers in multidomain networks. Computer Networks (Amsterdam, Netherlands: 1999), 37(5): 519–540, November 22, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/35/29/abstract.html.

Marsan:2001:IQR

[311] M. Ajmone Marsan, A. Bianco, P. Giaccone, E. Leonardi, and F. Neri. Input-queued router architectures exploiting cell-based switching fabrics. Computer Networks (Amsterdam, Netherlands: 1999), 37(5):541– 559, November 22, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
35/30/abstract.html.

Francini:2001:EWR

[312] A. Francini, F. M. Chiussi, R. T. Clancy, K. D. Drucker, and N. E. Idirene. Enhanced weighted round robin schedulers for accurate bandwidth distribution in packet networks. Computer Networks (Amsterdam, Netherlands: 1999), 37(5): 561–578, November 22, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/35/31/abstract.html.

Ciulli:2001:ASW

[313] Nicola Ciulli and Stefano Giordano. Analysis and simulation of WF^{2}Q + based schedulers: comparisons, compliance with theoretical bounds and influence on end-to-end delay jitter. Computer Networks (Amsterdam, Netherlands: 1999), 37(5):579– 599, November 22, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/15/22/67/
                                        35/32/abstract.html.

Xylomenos:2001:QSS

[314] George Xylomenos and George C. Polyzos. Quality of service support over multi-service wireless Internet links. Computer Networks (Amsterdam, Netherlands: 1999), 37(5): 601–615, November 22, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/35/33/abstract.html.

Priggouris:2001:GRI

[315] G. Priggouris, S. Hadjiefthymiades, and L. Merakos. GPRS + IntServ/RSVP: an integrated architecture. Computer Networks (Amsterdam, Netherlands: 1999), 37(5): 617–629, November 22, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/35/34/abstract.html.

Moustafa:2001:WRM

[316] Mohamed Moustafa, Ibrahim Habib, and Mahmoud Naghshineh. Wireless resource management using genetic algorithm for mobiles equilibrium. Computer Networks (Amsterdam, Netherlands: 1999), 37(5):631– 643, November 22, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
35/35/abstract.html.

Lombardo:2001:TEM

[317] A. Lombardo and G. Schembra. Tspec enforcement for MPEG video transmission over the next generation Internet: an analytical framework. Computer Networks (Amsterdam, Netherlands: 1999), 37(5): 645–667, November 22, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/35/36/abstract.html.

Casetti:2001:AFP

[318] Claudio Casetti and Michela Meo. An analytical framework for the performance evaluation of TCP Reno connections. Computer Networks (Amsterdam, Netherlands: 1999), 37(5): 669–682, November 22, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/35/37/abstract.html.

Anonymous:2001:ENM

[319] Anonymous. Editorial: Networking middleware: selected papers from the TERENA Networking Conference 2001. Computer Networks (Amsterdam, Netherlands: 1999), 37(6):683– 684, December 15, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/15/22/67/
                                        36/27/abstract.html.

Alterman:2001:UFP

[320] Peter Alterman. The US Federal PKI and the Federal Bridge Certification Authority. Computer Networks (Amsterdam, Netherlands: 1999), 37(6): 685–690, December 15, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/36/28/abstract.html.

Gelenbe:2001:MPC

[321] Erol Gelenbe, Ricardo Lent, and Zhiguang Xu. Measurement and performance of a cognitive packet network. Computer Networks (Amsterdam, Netherlands: 1999), 37(6): 691–701, December 15, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/36/29/abstract.html.

Castro-Rojo:2001:PSP

[322] Rodrigo Castro-Rojo and Diego R. López. The PAPI system: point of access to providers of information. Computer Networks (Amsterdam, Netherlands: 1999), 37(6):703– 710, December 15, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
36/30/abstract.html.

Dolgikh:2001:PCS

[323] Dmitry G. Dolgikh and Andrei M. Sukhov. Parameters of cache systems based on a Zipf-like distribution. Computer Networks (Amsterdam, Netherlands: 1999), 37(6):711– 716, December 15, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/67/
36/31/abstract.html.

Dorries:2001:HDH

[324] Gundula Dörries and Lothar Zier. How to do high-speed multicast right! Computer Networks (Amsterdam, Netherlands: 1999), 37(6): 717–728, December 15, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/36/32/abstract.html.

Feustel:2001:MOT

[325] Björn Feustel and Thomas C. Schmidt. Media objects in time — a multimedia streaming system — work in progress paper v 1.5. Computer Networks (Amsterdam, Netherlands: 1999), 37(6): 729–737, December 15, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/36/33/abstract.html.

Griera:2001:QEM

[326] Martí Griera, Maribel Jiménez, and Juan Antonio Martínez. QOS evaluation model for a campus-wide network: an alternative approach focussing on availability. Computer Networks (Amsterdam, Netherlands: 1999), 37(6): 739–746, December 15, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/36/34/abstract.html.

Binczewski:2001:FEP

[327] Artur Binczewski et al. First experiences with the Polish Optical Internet. Computer Networks (Amsterdam, Netherlands: 1999), 37(6): 747–759, December 15, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/15/22/
                                        67/36/35/abstract.html.

Cardellini:2001:MQS

[328] Valeria Cardellini, Emiliano Casalicchio, Michele Colajanni, and Salvatore Tucci. Mechanisms for quality of service in Web clusters. Computer Networks (Amsterdam, Netherlands: 1999), 37(6):761–771, December 15, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/15/22/67/
                                        36/36/abstract.html.

Konstantopoulos:2001:NAM

[329] M. Konstantopoulos, T. Spyrou, and J. Darzentas. The need for academic middleware to support advanced learning services. Computer Networks (Amsterdam, Netherlands: 1999), 37(6):

773–781, December 15, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/36/37/abstract.html.

Anonymous:2001:AIV

[330] Anonymous. Author index: Volume 37. Computer Networks (Amsterdam, Netherlands: 1999), 37(6): 783–785, December 15, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/36/38/abstract.html.

Anonymous:2001:SIV

[331] Anonymous. Subject index: Volume 37. Computer Networks (Amsterdam, Netherlands: 1999), 37(6): 787–789, December 15, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
67/36/39/abstract.html.

Chan:2002:PPS

[332] M. C. Chan, P. P. To, and T. T. Lee. Providing per-session deterministic quality of service guarantees in cross-path packet switch. Computer Networks (Amsterdam, Netherlands: 1999), 38(1):1–23, January 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/27/27/abstract.html.

Antunes:2002:ITM

[333] Nelson Antunes, António Pacheco, and Rui Rocha. An integrated traffic model for multimedia wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 38(1):

25–41, January 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
27/28/abstract.html.

Borgonovo:2002:PSU

[334] Flaminio Borgonovo, Antonio Capone, Matteo Cesana, and Luigi Fratta. Packet service in UMTS: delaythroughput performance of the downlink shared channel. Computer Networks (Amsterdam, Netherlands: 1999), 38(1):43–59, January 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/27/29/abstract.html.

Siu:2002:EON

[335] Fai Siu and Rocky K. C. Chang. Effectiveness of optimal node assignments in wavelength division multiplexing networks with fixed regular virtual topologies. Computer Networks (Amsterdam, Netherlands: 1999), 38 (1):61–74, January 15, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/27/30/abstract.html.

Song:2002:AWS

[336] Junehwa Song, Arun Iyengar, Eric Levy-Abegnoli, and Daniel Dias. Architecture of a Web server accelerator. Computer Networks (Amsterdam, Netherlands: 1999), 38(1): 75–97, January 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
27/31/abstract.html.

Dietrich:2002:FMC

[337] F. Dietrich and J.-P. Hubaux. Formal methods for communication services: meeting the industry expectations. Computer Networks (Amsterdam, Netherlands: 1999), 38(1): 99–120, January 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
27/32/abstract.html.

Lee:2002:NWA

[338] Seungjoon Lee and Chongkwon Kim. A new wireless ad hoc multicast routing protocol. Computer Networks (Amsterdam, Netherlands: 1999), 38 (2):121–135, February 5, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/33/27/abstract.html.

Manner:2002:EMQ

[339] Jukka Manner et al. Evaluation of mobility and quality of service interaction. Computer Networks (Amsterdam, Netherlands: 1999), 38(2): 137–163, February 5, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/33/28/abstract.html.

Cheung:2002:VPN

[340] K. H. Cheung and J. Mišić. On virtual private networks security design issues. Computer Networks (Amsterdam, Netherlands: 1999), 38(2): 165–179, February 5, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/33/29/abstract.html.

Hristea:2002:NII

[341] Cristina Hristea and Fouad Tobagi. A network infrastructure for IP mobility support in metropolitan areas. Computer Networks (Amsterdam, Netherlands: 1999), 38(2):181– 206, February 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
33/30/abstract.html.

Ferracioli:2002:CAS

[342] Mirko Ferracioli, Velio Tralli, and Roberto Verdone. Channel adaptive scheduling for a wideband TDD/TCDMA wireless system under heterogeneous traffic conditions. Computer Networks (Amsterdam, Netherlands: 1999), 38(2):207–223, February 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
33/31/abstract.html.

Baiocchi:2002:ORR

[343] Andrea Baiocchi, Francesca Cuomo, and Cristina Martello. Optimizing the radio resource utilization of multiaccess systems with a traffictransmission quality adaptive packet scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 38(2): 225–246, February 5, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/33/32/abstract.html.

Chou:2002:MPG

[344] Jue-Sam Chou and Yi-Shiung Yeh. Mental poker game based on a bit commitment scheme through network. Computer Networks (Amsterdam, Netherlands: 1999), 38(2): 247–255, February 5, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/33/33/abstract.html.

Hegde:2002:ISH

[345] Nidhi Hegde and Khosrow Sohraby. On the impact of soft handoff in cellular systems. Computer Networks (Amsterdam, Netherlands: 1999), 38 (2):257–271, February 5, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/33/34/abstract.html.

Campbell:2002:PN

[346] A. T. Campbell, David Wetherall, and Raj Yavatkar. Programmable networks. Computer Networks (Amsterdam, Netherlands: 1999), 38(3): 273–275, February 21, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/34/27/abstract.html.

Karlin:2002:VER

[347] Scott Karlin and Larry Peterson. VERA: an extensible router architecture. Computer Networks (Amsterdam, Netherlands: 1999), 38(3): 277–293, February 21, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/34/28/abstract.html.

Taylor:2002:DHP

[348] David E. Taylor, Jonathan S. Turner, John W. Lockwood, and Edson L. Horta. Dynamic hardware plugins: exploiting reconfigurable hardware for high-performance programmable routers. Computer Networks (Amsterdam, Netherlands: 1999), 38(3): 295–310, February 21, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/34/29/abstract.html.

Raz:2002:NMA

[349] Danny Raz and Yuval Shavitt. New models and algorithms for programmable networks. Computer Networks (Amsterdam, Netherlands: 1999), 38(3):311–326, February 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
34/30/abstract.html.

Wen:2002:BMS

[350] Su Wen, James Griffioen, and Kenneth L. Calvert. Building multicast services from unicast forwarding and ephemeral state. Computer Networks (Amsterdam, Netherlands: 1999), 38 (3):327–345, February 21, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/34/31/abstract.html.

Li:2002:SDA

[351] Jun Li, Mark Yarvis, and Peter Reiher. Securing distributed adaptation. Computer Networks (Amsterdam, Netherlands: 1999), 38(3):347– 371, February 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
34/32/abstract.html.

Nakao:2002:CEE

[352] Akihiro Nakao, Larry Peterson, and Andy Bavier. Constructing endto-end paths for playing media objects. Computer Networks (Amsterdam, Netherlands: 1999), 38(3):373– 389, February 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
34/33/abstract.html.

Anonymous:2002:E

[353] Anonymous. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 38(4):391–392, March 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/35/27/abstract.html.

Akyildiz:2002:WSN

[354] I. F. Akyildiz, W. Su, Y. Sankarasubramaniam, and E. Cayirci. Wireless sensor networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 38(4):393–422, March 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/35/28/abstract.html.

Znati:2002:LCD

[355] Taieb F. Znati, Tawfig Alrabiah, and Rami Melhem. Low-cost, delaybounded point-to-multipoint communication to support multicasting over WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 38 (4):423–445, March 15, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/35/29/abstract.html.

Tian:2002:ORA

[356] Qing Tian and Donald C. Cox. Optimal replication algorithms for hierarchical mobility management in PCS networks. Computer Networks (Amsterdam, Netherlands: 1999), 38(4): 447–459, March 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
35/30/abstract.html.

Iera:2002:EHT

[357] Antonio Iera, Antonella Molinaro, and Giuseppe Araniti. Exploiting hierarchical terrestrial-satellite architectures to handle voice, symmetric data, and asymmetric data connections. Computer Networks (Amsterdam, Netherlands: 1999), 38(4): 461–475, March 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
35/31/abstract.html.

Lindemann:2002:APM

[358] Christoph Lindemann, Marco Lohmann, and Axel Thümmler. Adaptive performance management for universal mobile telecommunications system networks. Computer Networks (Amsterdam, Netherlands: 1999), 38(4): 477–496, March 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
35/32/abstract.html.

Litjens:2002:IMU

[359] Remco Litjens. The impact of mobility on UMTS network planning. Computer Networks (Amsterdam, Netherlands: 1999), 38(4):497–515, March 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/35/33/abstract.html.

Chan:2002:DES

[360] P. M. L. Chan, R. E. Sheriff, Y. F. Hu, P. Conforto, and C. Tocci. Design and evaluation of signaling protocols for mobility management in an integrated IP environment. Computer Networks (Amsterdam, Netherlands: 1999), 38 (4):517–530, March 15, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/10/15/22/
76/35/34/abstract.html.

Xu:2002:RPM

[361] Shugong Xu and Tarek Saadawi. Revealing the problems with 802.11 medium access control protocol in multi-hop wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 38(4): 531–548, March 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
35/35/abstract.html.

Anonymous:2002:AR

[362] Anonymous. Acknowledgement to reviewers. Computer Networks (Amsterdam, Netherlands: 1999), 38(4): 549–552, March 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.nl/gej-ng/10/15/22/76/
35/36/abstract.html.

Rhee:2002:MST

[363] Injong Rhee, Nallathambi Balaguru, and George N. Rouskas. MTCP: scalable TCP-like congestion control for reliable multicast. Computer Networks (Amsterdam, Netherlands: 1999), 38 (5):553–575, April 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/76/
36/27/abstract.html.

Yu:2002:MBP

[364] Fei Yu and Victor Leung. Mobilitybased predictive call admission control and bandwidth reservation in wireless cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 38 (5):577–589, April 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/76/
36/28/abstract.html.

Kolberg:2002:PAS

[365] Mario Kolberg and Evan H. Magill. A pragmatic approach to service interaction filtering between call control services. Computer Networks (Amsterdam, Netherlands: 1999), 38 (5):591–602, April 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/76/
36/29/abstract.html.

Lee:2002:MVU

[366] Ki-Dong Lee and Sehun Kim. Modeling variable user mobility with stochastic correlation concept. Computer Networks (Amsterdam, Netherlands: 1999), 38(5):603–612, April 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/gej-ng/10/15/
22/76/36/30/abstract.html.

Kalyanasundaram:2002:ORA

[367] Suresh Kalyanasundaram, Edwin K. P. Chong, and Ness B. Shroff. Optimal resource allocation in multi-class networks with user-specified utility functions. Computer Networks (Amsterdam, Netherlands: 1999), 38(5): 613–630, April 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/76/
36/31/abstract.html.

Lee:2002:OAB

[368] Ki-Dong Lee and Sehun Kim. Optimization for adaptive bandwidth reservation in wireless multimedia networks. Computer Networks (Amsterdam, Netherlands: 1999), 38(5): 631–643, April 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/76/
36/32/abstract.html.

Chan:2002:ETC

[369] Chia-Tai Chan, Pi-Chung Wang, and Yaw-Chung Chen. An efficient traffic control scheme for TCP over ATM GFR services. Computer Networks (Amsterdam, Netherlands: 1999), 38 (5):645–662, April 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/76/
                                        36/33/abstract.html.

Xiong:2002:CTP

[370] Yijun Xiong and Lorne Mason. Comparison of two path restoration schemes in self-healing networks. Computer Networks (Amsterdam, Netherlands: 1999), 38(5):663–674, April 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/15/
                                        22/76/36/34/abstract.html.

Yeh:2002:ALD

[371] Yi-Shiung Yeh, Wei-Shen Lai, and Chung-Jaye Cheng. Applying lightweight directory access protocol service on session certification authority. Computer Networks (Amsterdam, Netherlands: 1999), 38(5):675–692, April 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/15/
                                        22/76/36/35/abstract.html.

Gavalas:2002:HNM

[372] Damianos Gavalas, Dominic Greenwood, Mohammed Ghanbari, and Mike O’Mahony. Hierarchical network management: a scalable and dynamic mobile agent-based approach. Computer Networks (Amsterdam, Netherlands: 1999), 38(6):693–711, April 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/15/
                                        22/76/37/27/abstract.html.

Choi:2002:TEE

[373] Seonho Choi. Temporally enhanced erasure codes for reliable communication protocols. Computer Networks (Amsterdam, Netherlands: 1999), 38 (6):713–730, April 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/76/
37/28/abstract.html.

Kumar:2002:TFT

[374] K. R. Renjish Kumar, A. L. Ananda, and Lillykutty Jacob. TCP-friendly traffic conditioning in DiffServ networks: a memory-based approach. Computer Networks (Amsterdam, Netherlands: 1999), 38(6):731–743, April 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/76/
37/29/abstract.html.

Ahn:2002:ARD

[375] Sungjoon Ahn and A. Udaya Shankar. Adapting to route-demand and mobility in ad hoc network routing. Computer Networks (Amsterdam, Netherlands: 1999), 38(6):745–764, April 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/76/37/30/abstract.html.

Kwok:2002:OAC

[376] Yu-Kwong Kwok and Vincent K. N. Lau. Optimal admission control algorithms for scheduling burst data in CDMA multimedia systems. Computer Networks (Amsterdam, Netherlands: 1999), 38(6):765–777, April 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/gej-ng/10/15/
22/76/37/31/abstract.html.

Busari:2002:PSW

[377] Mudashiru Busari and Carey Williamson. ProWGen: a synthetic workload generation tool for simulation evaluation of web proxy caches. Computer Networks (Amsterdam, Netherlands: 1999), 38 (6):779–794, April 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/76/
                                        37/32/abstract.html.

Cohen:2002:RPW

[378] Edith Cohen and Haim Kaplan. Refreshment policies for Web content caches. Computer Networks (Amsterdam, Netherlands: 1999), 38(6): 795–808, April 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/76/
                                        37/33/abstract.html.

Naldi:2002:IAT

[379] M. Naldi. Internet access traffic sharing in a multi-operator environment. Computer Networks (Amsterdam, Netherlands: 1999), 38(6):809–824, April 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/76/37/34/abstract.html.

Anonymous:2002:AIVa

[380] Anonymous. Author Index Volume 38. Computer Networks (Amsterdam, Netherlands: 1999), 38(6): 825–827, April 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/76/
                                        37/35/abstract.html.

Anonymous:2002:SIVa

[381] Anonymous. Subject index volume 38. Computer Networks (Amsterdam, Netherlands: 1999), 38(6): 829–831, April 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/76/
37/36/abstract.html.

Morabito:2002:GEB

[382] G. Morabito, S. Palazzo, and C. Rosenberg. Guest editorial: Broadband satellite networks: a networking perspective. Computer Networks (Amsterdam, Netherlands: 1999), 39(1):1–3, May 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/96/27/27/abstract.html.

Acar:2002:WFB

[383] Güray Açar and Catherine Rosenberg. Weighted fair bandwidth-on-demand (WFBoD) for geostationary satellite networks with on-board processing. Computer Networks (Amsterdam, Netherlands: 1999), 39 (1):5–20, May 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
27/28/abstract.html.

Chan:2002:PAM

[384] P. M. L. Chan et al. Performance analysis of mobility procedures in a hybrid space terrestrial IP environment. Computer Networks (Amsterdam, Netherlands: 1999), 39(1):21–41, May 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/96/27/29/abstract.html.

Priscoli:2002:RMA

[385] Francesco Delli Priscoli and Antonio Pietrabissa. Resource management for ATM-based geostationary satellite networks with on-board processing. Computer Networks (Amsterdam, Netherlands: 1999), 39(1):43–60, May 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/96/27/30/abstract.html.

Ercetin:2002:PGS

[386] Özgür Erçetin, Srikanth Krishnamurthy, Son Dao, and Leandros Tassiulas. Provision of guaranteed services in broadband LEO satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 39(1): 61–77, May 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/96/
                                        27/31/abstract.html.

Franck:2002:SIS

[387] Laurent Franck and Gérard Maral. Signaling for inter-satellite link routing in broadband non-GEO satellite systems. Computer Networks (Amsterdam, Netherlands: 1999), 39 (1):79–92, May 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/96/
                                        27/32/abstract.html.

Iheagwara:2002:EPI

[388] Charles Iheagwara and Andrew Blyth. Evaluation of the performance of ID systems in a switched and distributed environment: the RealSecure case study. Computer Networks (Amsterdam, Netherlands: 1999), 39 (2):93–112, June 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
49/27/abstract.html.

Wieselthier:2002:RME

[389] Jeffrey E. Wieselthier, Gam D. Nguyen, and Anthony Ephremides. Resource management in energy-limited, bandwidth-limited, transceiver-limited wireless networks for session-based multicasting. Computer Networks (Amsterdam, Netherlands: 1999), 39 (2):113–131, June 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
49/28/abstract.html.

Barakat:2002:BTB

[390] Chadi Barakat and Eitan Altman. Bandwidth tradeoff between TCP and link-level FEC. Computer Networks (Amsterdam, Netherlands: 1999), 39 (2):133–150, June 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
49/29/abstract.html.

Yu:2002:WSM

[391] Jintae Yu et al. WDM/SCM multiple access protocol with high throughput and low packet delay for passive double star networks. Computer Networks (Amsterdam, Netherlands: 1999), 39 (2):151–164, June 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/96/
                                        49/30/abstract.html.

Anjali:2002:OPL

[392] T. Anjali, C. Scoglio, J. C. de Oliveira, I. F. Akyildiz, and G. Uhl. Optimal policy for label switched path setup in MPLS networks. Computer Networks (Amsterdam, Netherlands: 1999), 39 (2):165–183, June 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/96/
                                        49/31/abstract.html.

Altman:2002:QAS

[393] Eitan Altman, Chadi Barakat, and Víctor M. Ramos R. Queuing analysis of simple FEC schemes for voice over IP. Computer Networks (Amsterdam, Netherlands: 1999), 39(2): 185–206, June 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/96/
                                        49/32/abstract.html.

Lee:2002:ACC

[394] B. P. Lee, R. K. Balan, L. Jacob, W. K. G. Seah, and A. L. Ananda. Avoiding congestion collapse on the Internet using TCP tunnels. Computer Networks (Amsterdam, Netherlands: 1999), 39(2):207–219, June 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/15/
                                        22/96/49/33/abstract.html.

Levene:2002:EWD

[395] Mark Levene and Alexandra Poulovassilis. Editorial: Web dynamics. Computer Networks (Amsterdam, Netherlands: 1999), 39(3):221–223, June 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/96/50/27/abstract.html;
 http://www.elsevier.com/gej-ng/
10/15/22/96/50/27/article.pdf.

Abiteboul:2002:XP

[396] S. Abiteboul, S. Cluet, G. Ferran, and M.-C. Rousset. The Xyleme project. Computer Networks (Amsterdam, Netherlands: 1999), 39 (3):225–238, June 21, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/96/50/28/abstract.html;
 http://www.elsevier.com/gej-ng/
10/15/22/96/50/28/article.pdf.

Bailey:2002:AOE

[397] James Bailey, Alexandra Poulovassilis, and Peter T. Wood. Analysis and optimisation of event-condition-action rules on XML. Computer Networks (Amsterdam, Netherlands: 1999), 39(3):239–259, June 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/96/50/29/abstract.html;
 http://www.elsevier.com/gej-ng/
10/15/22/96/50/29/article.pdf.

Krapivsky:2002:SPP

[398] P. L. Krapivsky and S. Redner. A statistical physics perspective on Web growth. Computer Networks (Amsterdam, Netherlands: 1999), 39 (3):261–276, June 21, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/

                                        10/15/22/96/50/30/abstract.html;
                                         http://www.elsevier.com/gej-ng/
                                        10/15/22/96/50/30/article.pdf.

Levene:2002:SME

[399] Mark Levene, Trevor Fenner, George Loizou, and Richard Wheeldon. A stochastic model for the evolution of the Web. Computer Networks (Amsterdam, Netherlands: 1999), 39 (3):277–287, June 21, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/96/50/31/abstract.html;
                                         http://www.elsevier.com/gej-ng/
                                        10/15/22/96/50/31/article.pdf.

Risvik:2002:SEW

[400] Knut Magne Risvik and Rolf Michelsen. Search engines and Web dynamics. Computer Networks (Amsterdam, Netherlands: 1999), 39(3): 289–302, June 21, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/96/50/32/abstract.html;
                                         http://www.elsevier.com/gej-ng/
                                        10/15/22/96/50/32/article.pdf.

Zhang:2002:NWU

[401] Dell Zhang and Yisheng Dong. A novel Web usage mining approach for search engines. Computer Networks (Amsterdam, Netherlands: 1999), 39 (3):303–310, June 21, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/96/50/33/abstract.html;
                                         http://www.elsevier.com/gej-ng/
                                        10/15/22/96/50/33/article.pdf.

Lemahieu:2002:CBN

[402] Wilfried Lemahieu. Context-based navigation in the Web by means of dynamically generated guided tours. Computer Networks (Amsterdam, Netherlands: 1999), 39 (3):311–328, June 21, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/96/50/34/abstract.html;
http://www.elsevier.com/gej-ng/
10/15/22/96/50/34/article.pdf.

Sang:2002:PAN

[403] Aimin Sang and San qi Li. A predictability analysis of network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 39(4): 329–345, July 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
51/27/abstract.html.

Balan:2002:THM

[404] R. K. Balan et al. TCP HACK: a mechanism to improve performance over lossy links. Computer Networks (Amsterdam, Netherlands: 1999), 39 (4):347–361, July 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
51/28/abstract.html.

Lenzini:2002:PTT

[405] L. Lenzini, E. Mingozzi, and G. Stea. Packet timed token service discipline: a scheduling algorithm based on the dual-class paradigm for providing QoS in integrated services networks. Computer Networks (Amsterdam, Netherlands: 1999), 39(4):363–384, July 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/96/51/29/abstract.html.

Bhagwat:2002:MAT

[406] Pravin Bhagwat, David A. Maltz, and Adrian Segall. MSOCKS+: an architecture for transport layer mobility. Computer Networks (Amsterdam, Netherlands: 1999), 39(4): 385–403, July 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
51/30/abstract.html.

Lee:2002:IMR

[407] Jangwon Lee and Gustavo de Veciana. IP multicast resource and topology discovery using a fan-out decrement mechanism. Computer Networks (Amsterdam, Netherlands: 1999), 39(4): 405–422, July 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
51/31/abstract.html.

Chahed:2002:MPM

[408] Tijani Chahed, Gérard Hébuterne, and Guy Pujolle. Mapping of performance metrics between IP and ATM. Computer Networks (Amsterdam, Netherlands: 1999), 39(4):423–436, July 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/96/51/32/abstract.html.

Cohen:2002:PMD

[409] Edith Cohen and Haim Kaplan. Prefetching the means for document transfer: a new approach for reducing Web latency. Computer Networks (Amsterdam, Netherlands: 1999), 39 (4):437–455, July 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
51/33/abstract.html.

Schneider:2002:SSO

[410] G. Michael Schneider and Tamas Nemeth. A simulation study of the OSPF-OMP routing algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 39(4):457–468, July 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/96/51/34/abstract.html.

Marchiori:2002:E

[411] Massimo Marchiori, Mary Ellen Zurko, and Michael R. Lyu. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):469–471, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/96/52/27/abstract.html.

Buneman:2002:KX

[412] Peter Buneman, Susan Davidson, Wenfei Fan, Carmem Hara, and Wang-Chiew Tan. Keys for XML. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):473–487, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/gej-ng/10/15/
22/96/52/28/abstract.html.

Fan:2002:UCM

[413] Wenfei Fan, Gabriel M. Kuper, and Jérôme Siméon. A unified constraint model for XML. Computer Networks (Amsterdam, Netherlands: 1999), 39 (5):489–505, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
52/29/abstract.html.

Brown:2002:MMW

[414] Allen Brown, Matthew Fuchs, Jonathan Robie, and Philip Wadler. MSL: a model for W3C XML Schema. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):507–521, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/96/52/30/abstract.html.

Agrawal:2002:VSO

[415] Rakesh Agrawal, Roberto J. Bayardo, Jr., Daniel Gruhl, and Spiros Papadimitriou. Vinci: a service-oriented architecture for rapid development of Web applications. Computer Networks (Amsterdam, Netherlands: 1999), 39(5): 523–539, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
52/31/abstract.html.

Petropoulos:2002:BXQ

[416] Michalis Petropoulos, Yannis Papakonstantinou, and Vasilis Vassalos. Building XML query forms and reports with XQForms. Computer Networks (Amsterdam, Netherlands: 1999), 39(5): 541–558, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
52/32/abstract.html.

Beck:2002:EFS

[417] Micah Beck, Terry Moore, Leif Abrahamsson, Chistophe Achouiantz, and Patrik Johansson. Enabling full service surrogates using the portable channel representation. Computer Networks (Amsterdam, Netherlands: 1999), 39 (5):559–576, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
52/33/abstract.html.

Beckett:2002:DIR

[418] David Beckett. The design and implementation of the Redland RDF application framework. Computer Networks (Amsterdam, Netherlands: 1999), 39 (5):577–588, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
52/34/abstract.html.

Kahan:2002:AOR

[419] J. Kahan, M.-R. Koivunen, E. Prud’Hommeaux, and R. R. Swick. Annotea: an open RDF infrastructure for shared Web annotations. Computer Networks (Amsterdam, Netherlands: 1999), 39(5): 589–608, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
52/35/abstract.html.

Broekstra:2002:EKR

[420] Jeen Broekstra et al. Enabling knowledge representation on the Web by extending RDF Schema. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):609–634, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/15/
                                        22/96/52/36/abstract.html.

Myllymaki:2002:EWD

[421] Jussi Myllymaki. Effective Web data extraction with standard XML technologies. Computer Networks (Amsterdam, Netherlands: 1999), 39(5): 635–644, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/96/
                                        52/37/abstract.html.

Bonifati:2002:PRS

[422] Angela Bonifati, Stefano Ceri, and Stefano Paraboschi. Pushing reactive services to XML repositories using active rules. Computer Networks (Amsterdam, Netherlands: 1999), 39(5): 645–660, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/96/
                                        52/38/abstract.html.

Strobel:2002:XSR

[423] Michael Ströbel. An XML schema representation for the communication design of electronic negotiations. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):661–680, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/96/52/39/abstract.html.

Sundaresan:2002:APM

[424] Neel Sundaresan and Reshad Moussa. Algorithms and programming models for efficient representation of XML for Internet applications. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):681–697, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/gej-ng/10/15/
22/96/52/40/abstract.html.

Judge:2002:WWM

[425] Paul Judge and Mostafa Ammar. WHIM: watermarking multicast video with a hierarchy of intermediaries. Computer Networks (Amsterdam, Netherlands: 1999), 39(6):699–712, August 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
53/27/abstract.html.

Costa:2002:DSP

[426] Luís Henrique M. K. Costa, Serge Fdida, and Otto Carlos M. B. Duarte. Developing scalable protocols for threemetric QoS routing. Computer Networks (Amsterdam, Netherlands: 1999), 39(6):713–727, August 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/gej-ng/10/15/
22/96/53/28/abstract.html.

Yang:2002:TSB

[427] Wen-Lin Yang. A tabu-search based algorithm for the multicast-streams distribution problem. Computer Networks (Amsterdam, Netherlands: 1999), 39 (6):729–747, August 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
53/29/abstract.html.

Onur:2002:MBR

[428] Ertan Onur, Hakan Deliç, Cem Ersoy, and M. Ufuk Çaǧlayan. Measurementbased replanning of cell capacities in GSM networks. Computer Networks (Amsterdam, Netherlands: 1999), 39 (6):749–767, August 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/96/
                                        53/30/abstract.html.

Cho:2002:DBM

[429] Jeong woo Cho and Dong ho Cho. Dynamic buffer management scheme based on rate estimation in packetswitched networks. Computer Networks (Amsterdam, Netherlands: 1999), 39(6):769–787, August 21, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/15/
                                        22/96/53/31/abstract.html.

Altman:2002:SDM

[430] Eitan Altman, Kostya Avrachenkov, Chadi Barakat, and Rudesindo Núñez-Queija. State-dependent M/G/1 type queueing analysis for congestion control in data networks. Computer Networks (Amsterdam, Netherlands: 1999), 39(6):789–808, August 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/15/
                                        22/96/53/32/abstract.html.

Ahmed:2002:TCM

[431] Mohiuddin Ahmed, Srikanth Krishnamurthy, Randy Katz, and Son Dao. Trajectory control of mobile gateways for range extension in ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 39(6): 809–825, August 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
53/33/abstract.html.

Iheagwara:2002:ISL

[432] Charles Iheagwara and Andrew Blyth. The impact of security layering on endto-end latency and system performance in switched and distributed e-business environments. Computer Networks (Amsterdam, Netherlands: 1999), 39 (6):827–840, August 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
53/34/abstract.html.

Anonymous:2002:AIVb

[433] Anonymous. Author index volume 39. Computer Networks (Amsterdam, Netherlands: 1999), 39(6):841– 843, August 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
53/35/abstract.html.

Anonymous:2002:SIVb

[434] Anonymous. Subject index volume 39. Computer Networks (Amsterdam, Netherlands: 1999), 39(6):845– 847, August 21, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/96/
53/36/abstract.html.

Casaca:2002:GET

[435] Augusto Casaca and Amardeo C. Sarma. Guest editorial: Towards a new Internet architecture. Computer Networks (Amsterdam, Netherlands: 1999), 40(1):1–4, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/gej-ng/10/15/
22/97/27/27/abstract.html.

Baxevanidis:2002:GRN

[436] Kyriakos Baxevanidis, Howard Davies, Ian Foster, and Fabrizio Gagliardi. Grids and research networks as drivers and enablers of future Internet architectures. Computer Networks (Amsterdam, Netherlands: 1999), 40(1): 5–17, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
27/28/abstract.html.

Noureddine:2002:IPI

[437] Waël Noureddine and Fouad Tobagi. Improving the performance of interactive TCP applications using service differentiation. Computer Networks (Amsterdam, Netherlands: 1999), 40(1): 19–43, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
27/29/abstract.html.

Miao:2002:TDM

[438] Yu-Ben Miao, Wen-Shyang Hwang, and Ce-Kuen Shieh. A transparent deployment method of RSVP-aware applications on UNIX. Computer Networks (Amsterdam, Netherlands: 1999), 40(1):45–56, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/97/27/30/abstract.html.

Benameur:2002:QSF

[439] N. Benameur, S. Ben Fredj, S. Oueslati-Boulahia, and J. W. Roberts. Quality of service and flow level admission control in the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 40(1):57–71, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/97/27/31/abstract.html.

Bianchi:2002:PFQ

[440] G. Bianchi, N. Blefari-Melazzi, and M. Femminella. Per-flow QoS support over a stateless Differentiated Services IP domain. Computer Networks (Amsterdam, Netherlands: 1999), 40(1): 73–87, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
27/32/abstract.html.

Liebeherr:2002:RAB

[441] Jörg Liebeherr and Nicolas Christin. Rate allocation and buffer management for differentiated services. Computer Networks (Amsterdam, Netherlands: 1999), 40(1):89–110, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/97/27/33/abstract.html.

Awduche:2002:ITE

[442] Daniel O. Awduche and Bijan Jabbari. Internet traffic engineering using multi-protocol label switching (MPLS). Computer Networks (Amsterdam, Netherlands: 1999), 40(1): 111–129, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/97/
                                        27/34/abstract.html.

Filsfils:2002:EMI

[443] Clarence Filsfils and John Evans. Engineering a multiservice IP backbone to support tight SLAs. Computer Networks (Amsterdam, Netherlands: 1999), 40(1):131–148, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/97/27/35/abstract.html.

Banerjee:2002:CAP

[444] Gargi Banerjee and Deepinder Sidhu. Comparative analysis of path computation techniques for MPLS traffic engineering. Computer Networks (Amsterdam, Netherlands: 1999), 40(1): 149–165, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/97/
                                        27/36/abstract.html.

Ho:2002:FSC

[445] Pin-Han Ho and H. T. Mouftah. Framework of spare capacity reallocation with S-SLSP for mesh WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 40(1): 167–179, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/97/
                                        27/37/abstract.html.

Sanmateu:2002:SMA

[446] A. Sanmateu et al. Seamless mobility across IP networks using Mobile IP. Computer Networks (Amsterdam, Netherlands: 1999), 40(1):181– 190, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
27/38/abstract.html.

Costa:2002:SSP

[447] Xavier Pérez Costa and Hannes Hartenstein. A simulation study on the performance of Mobile IPv6 in a WLAN-based cellular network. Computer Networks (Amsterdam, Netherlands: 1999), 40(1):191– 204, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
27/39/abstract.html.

Cidon:2002:OAE

[448] Israel Cidon, Shay Kutten, and Ran Soffer. Optimal allocation of electronic content. Computer Networks (Amsterdam, Netherlands: 1999), 40(2): 205–218, October 7, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
49/28/abstract.html.

Selcuk:2002:POT

[449] Ali Aydin Selçuk and Deepinder Sidhu. Probabilistic optimization techniques for multicast key management. Computer Networks (Amsterdam, Netherlands: 1999), 40(2):219–234, October 7, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/97/49/29/abstract.html.

Shan:2002:MPA

[450] Zhiguang Shan, Chuang Lin, Dan C. Marinescu, and Yang Yang. Modeling and performance analysis of QoSaware load balancing of Web-server clusters. Computer Networks (Amsterdam, Netherlands: 1999), 40(2): 235–256, October 7, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/97/
                                        49/30/abstract.html.

Wang:2002:NME

[451] S. Y. Wang and H. T. Kung. A new methodology for easily constructing extensible and high-fidelity TCP/IP network simulators. Computer Networks (Amsterdam, Netherlands: 1999), 40 (2):257–278, October 7, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/97/
                                        49/31/abstract.html.

Bordbar:2002:UUS

[452] Behzad Bordbar, John Derrick, and Gill Waters. Using UML to specify QoS constraints in ODP. Computer Networks (Amsterdam, Netherlands: 1999), 40(2):279–304, October 7, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/15/
                                        22/97/49/32/abstract.html.

Mir:2002:ESF

[453] Nader F. Mir. An efficient switching fabric for next-generation largescale computer networking. Computer Networks (Amsterdam, Netherlands: 1999), 40(2):305–315, October 7, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/gej-ng/10/15/
22/97/49/33/abstract.html.

Anonymous:2002:EBa

[454] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 40(2):CO2, October 7, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/97/49/27/abstract.html.

Devetsikiotis:2002:AME

[455] Mike Devetsikiotis and Nelson Fonseca. Guest editorial: Advances in modeling and engineering of long-range dependent traffic. Computer Networks (Amsterdam, Netherlands: 1999), 40(3): 317–318, October 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
50/28/abstract.html.

Nuzman:2002:CMT

[456] Carl Nuzman, Iraj Saniee, Wim Sweldens, and Alan Weiss. A compound model for TCP connection arrivals for LAN and WAN applications. Computer Networks (Amsterdam, Netherlands: 1999), 40(3):319– 337, October 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
50/29/abstract.html.

Figueiredo:2002:AST

[457] Daniel R. Figueiredo, Benyuan Liu, Vishal Misra, and Don Towsley. On the autocorrelation structure of TCP traffic. Computer Networks (Amsterdam, Netherlands: 1999), 40(3):

339–361, October 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/97/
                                        50/30/abstract.html.

Laskin:2002:FLM

[458] N. Laskin, I. Lambadaris, F. C. Harmantzis, and M. Devetsikiotis. Fractional Lévy motion and its application to network traffic modeling. Computer Networks (Amsterdam, Netherlands: 1999), 40(3):363–375, October 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/97/50/31/abstract.html.

Addie:2002:PEQ

[459] Ronald G. Addie, Timothy D. Neame, and Moshe Zukerman. Performance evaluation of a queue fed by a Poisson Pareto burst process. Computer Networks (Amsterdam, Netherlands: 1999), 40(3):377–397, October 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/97/50/32/abstract.html.

Mannersalo:2002:MPP

[460] Petteri Mannersalo and Ilkka Norros. A most probable path approach to queueing systems with general Gaussian input. Computer Networks (Amsterdam, Netherlands: 1999), 40(3): 399–412, October 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/97/
                                        50/33/abstract.html.

deMPereira:2002:PGP

[461] Flavio de M. Pereira, Nelson L. S. da Fonseca, and Dalton S. Arantes. On the performance of generalized processor sharing servers under long-range dependent traffic. Computer Networks (Amsterdam, Netherlands: 1999), 40 (3):413–431, October 22, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/gej-ng/10/15/
22/97/50/34/abstract.html.

Jelenkovic:2002:FBQ

[462] Predrag Jelenković and Petar Momčilović. Finite buffer queue with generalized processor sharing and heavy-tailed input processes. Computer Networks (Amsterdam, Netherlands: 1999), 40 (3):433–443, October 22, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.elsevier.com/gej-ng/10/15/
22/97/50/35/abstract.html.

Krunz:2002:AIB

[463] Marwan Krunz and Ibrahim Matta. Analytical investigation of the bias effect in variance-type estimators for inference of long-range dependence. Computer Networks (Amsterdam, Netherlands: 1999), 40(3): 445–458, October 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
50/36/abstract.html.

Markovitch:2002:EHT

[464] Natalia M. Markovitch and Udo R. Krieger. The estimation of heavytailed probability density functions, their mixtures and quantiles. Computer Networks (Amsterdam, Netherlands: 1999), 40(3):459–474, October 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/97/50/37/abstract.html.

Anonymous:2002:EBb

[465] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 40(3):CO2, October 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/97/50/27/abstract.html.

Anonymous:2002:CNJ

[466] Anonymous. Computer networks, the journal, and computer networks, the technology. Computer Networks (Amsterdam, Netherlands: 1999), 40(4): 475, November 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/97/
                                        51/26/abstract.html.

Tsaoussidis:2002:TRR

[467] V. Tsaoussidis and C. Zhang. TCPreal: receiver-oriented congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 40(4):477– 497, November 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/97/
                                        51/27/abstract.html.

Blanpain:2002:IDA

[468] Yannick Blanpain and Raghupathy Sivakumar. An incrementally deployable approach for achieving fair rate allocations. Computer Networks (Amsterdam, Netherlands: 1999), 40(4):

499–513, November 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
51/28/abstract.html.

Avancha:2002:ETP

[469] Sasikanth Avancha, Vladimir Korolev, Anupam Joshi, Timothy Finin, and Yelena Yesha. On experiments with a transport protocol for pervasive computing environments. Computer Networks (Amsterdam, Netherlands: 1999), 40(4):515–535, November 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/97/51/29/abstract.html.

Abouzeid:2002:MRE

[470] Alhussein A. Abouzeid and Sumit Roy. Modeling random early detection in a differentiated services network. Computer Networks (Amsterdam, Netherlands: 1999), 40(4):537– 556, November 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
51/30/abstract.html.

Misra:2002:PBB

[471] Archan Misra, Teunis Ott, and John Baras. Predicting bottleneck bandwidth sharing by generalized TCP flows. Computer Networks (Amsterdam, Netherlands: 1999), 40(4):557– 576, November 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
51/31/abstract.html.

Tham:2002:ADS

[472] Chen-Khong Tham, Qi Yao, and Yuming Jiang. Achieving differentiated services through multi-class probabilistic priority scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 40(4):577–593, November 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/97/51/32/abstract.html.

Anonymous:2002:EBc

[473] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 40(4):CO2, November 15, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/97/51/25/abstract.html.

Turner:2002:GEP

[474] K. J. Turner. Guest editorial: Protocol animation. Computer Networks (Amsterdam, Netherlands: 1999), 40(5): 595–598, December 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/97/
                                        52/26/abstract.html.

Combes:2002:OAT

[475] Pierre Combes, Fabrice Dubois, and Béatrice Renard. An open animation tool: application to telecommunication systems. Computer Networks (Amsterdam, Netherlands: 1999), 40(5): 599–620, December 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/15/22/97/
                                        52/27/abstract.html.

Schaible:2002:VBA

[476] Philipp Schaible and Reinhard Gotzhein. View-based animation of communication protocols in design and in operation. Computer Networks (Amsterdam, Netherlands: 1999), 40(5): 621–638, December 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
52/28/abstract.html.

Yasumoto:2002:PAB

[477] Keiichi Yasumoto, Takaaki Umedu, Hirozumi Yamaguchi, Akio Nakata, and Teruo Higashino. Protocol animation based on event-driven visualization scenarios in real-time LO-TOS. Computer Networks (Amsterdam, Netherlands: 1999), 40(5):639– 663, December 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

elsevier.com/gej-ng/10/15/22/97/
52/29/abstract.html.

Stepien:2002:GVA

[478] Bernard Stepien and Luigi Logrippo. Graphic visualization and animation of LOTOS execution traces. Computer Networks (Amsterdam, Netherlands: 1999), 40(5):665–681, December 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/gej-ng/
10/15/22/97/52/30/abstract.html.

Anonymous:2002:EBd

[479] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 40(5):CO2, December 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/15/22/97/52/25/abstract.html.

Jiang:2002:DBN

[480] Yuming Jiang. Delay bounds for a network of guaranteed rate servers with FIFO aggregation. Computer Networks (Amsterdam, Netherlands: 1999), 40 (6):683–694, December 20, 2002. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Elwalid:2002:MMA

[481] Anwar Elwalid, Cheng Jin, Steven Low, and Indra Widjaja. MATE: multipath adaptive traffic engineering. Computer Networks (Amsterdam, Netherlands: 1999), 40(6):695– 709, December 20, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Karam:2002:ADD

[482] Mansour J. Karam and Fouad A. Tobagi. Analysis of delay and delay jitter of voice traffic in the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 40(6):711–726, December 20, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ngin:2002:CTA

[483] Hoon-Tong Ngin and Chen-Khong Tham. A control-theoretical approach for achieving fair bandwidth allocations in core-stateless networks. Computer Networks (Amsterdam, Netherlands: 1999), 40(6):727–741, December 20, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2002:AIVc

[484] Anonymous. Author index volume 40. Computer Networks (Amsterdam, Netherlands: 1999), 40(6):743– 745, December 20, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2002:SIVc

[485] Anonymous. Subject index volume 40. Computer Networks (Amsterdam, Netherlands: 1999), 40(6):747– 748, December 20, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2002:AIVd

[486] Anonymous. Author index volumes 31–40. Computer Networks (Amsterdam, Netherlands: 1999), 40(6):749– 777, December 20, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2002:SIVd

[487] Anonymous. Subject index volumes 31–40. Computer Networks (Amsterdam, Netherlands: 1999), 40(6):779– 797, December 20, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2002:EB

[488] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 40(6):??, December 20, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lindemann:2003:PAG

[489] Christoph Lindemann and Axel Thümmler. Performance analysis of the general packet radio service. Computer Networks (Amsterdam, Netherlands: 1999), 41(1):1–17, January 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2003:EPA

[490] N. Zhang, Q. Shi, and M. Merabti. An efficient protocol for anonymous and fair document exchange. Computer Networks (Amsterdam, Netherlands: 1999), 41(1):19–28, January 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

DAmbrogio:2003:STA

[491] A. D’Ambrogio and G. Iazeolla. Steps towards the automatic production of performance models of web applications. Computer Networks (Amsterdam, Netherlands: 1999), 41(1): 29–39, January 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tan:2003:ABR

[492] S. K. Tan, G. Mohan, and K. C. Chua. Algorithms for burst rescheduling in WDM optical burst switching networks. Computer Networks (Amsterdam, Netherlands: 1999), 41(1): 41–55, January 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gouda:2003:SAR

[493] Mohamed G. Gouda and Chin-Tser Huang. A secure address resolution protocol. Computer Networks (Amsterdam, Netherlands: 1999), 41(1): 57–71, January 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guo:2003:SSR

[494] Liang Guo and Ibrahim Matta. Search space reduction in QoS routing. Computer Networks (Amsterdam, Netherlands: 1999), 41(1):73–88, January 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yan:2003:RSS

[495] Shuguang Yan, Jitender S. Deogun, and Maher Ali. Routing in sparse splitting optical networks with multicast traffic. Computer Networks (Amsterdam, Netherlands: 1999), 41(1): 89–113, January 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Calder:2003:FIC

[496] Muffy Calder, Mario Kolberg, Evan H. Magill, and Stephan Reiff-Marganiec. Feature interaction: a critical review and considered forecast. Computer Networks (Amsterdam, Netherlands: 1999), 41(1):115–141, January 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBa

[497] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 41(1):??, January 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xu:2003:SSO

[498] Lisong Xu, Harry G. Perros, and George N. Rouskas. A simulation study of optical burst switching and access protocols for WDM ring networks. Computer Networks (Amsterdam, Netherlands: 1999), 41(2):

143–160, February 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gan:2003:LDM

[499] Yung-Sze Gan and Chen-Khong Tham. Loss differentiated multicast congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 41(2): 161–176, February 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2003:DDM

[500] Jiangchuan Liu, Xinyan Zhang, Bo Li, Qian Zhang, and Wenwu Zhu. Distributed distance measurement for large-scale networks. Computer Networks (Amsterdam, Netherlands: 1999), 41(2):177–192, February 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2003:TBT

[501] Y. Richard Yang, Min Sik Kim, and Simon S. Lam. Transient behaviors of TCP-friendly congestion control protocols. Computer Networks (Amsterdam, Netherlands: 1999), 41(2): 193–210, February 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sailer:2003:GCC

[502] Anca Dracinschi Sailer and Serge Fdida. Generic congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 41(2):211– 225, February 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xu:2003:SNG

[503] Yong Xu, Shen-Chu Xu, and Bo-Xi Wu. Strictly nonblocking grooming of dynamic traffic in unidirectional SONET/WDM rings using genetic algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 41(2): 227–245, February 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Song:2003:CDM

[504] Yueyue Song, Avishai Wool, and Bülent Yener. Combinatorial design of multi-ring networks with combined routing and flow control. Computer Networks (Amsterdam, Netherlands: 1999), 41(2):247–267, February 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Choi:2003:CSP

[505] Sumi Choi, Jonathan Turner, and Tilman Wolf. Configuring sessions in programmable networks. Computer Networks (Amsterdam, Netherlands: 1999), 41(2):269–284, February 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBb

[506] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 41(2):??, February 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hassanein:2003:MAH

[507] Hossam Hassanein and Ioanis Nikolaidis. Mobile and ad-hoc local networks. Computer Networks (Amsterdam, Netherlands: 1999), 41(3):285– 287, February 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xue:2003:BCB

[508] Jingxin Xue and Terence D. Todd. Basestation collaboration in Bluetooth voice networks. Computer Networks (Amsterdam, Netherlands: 1999), 41 (3):289–301, February 21, 2003. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kraemer:2003:BBW

[509] Rolf Kraemer and Peter Schwander. Bluetooth based wireless Internet applications for indoor hot spots: experience of a successful experiment during CeBIT 2001. Computer Networks (Amsterdam, Netherlands: 1999), 41 (3):303–312, February 21, 2003. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Monks:2003:ESC

[510] Jeffrey Monks, Jean-Pierre Ebert, Wen-Mei W. Hwu, and Adam Wolisz. Energy saving and capacity improvement potential of power control in multi-hop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 41(3):313–330, February 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Safwat:2003:PAV

[511] Ahmed Safwat, Hossam Hassanein, and Hussein Mouftah. Power-aware virtual base stations for wireless mobile ad hoc communications. Computer Networks (Amsterdam, Netherlands: 1999), 41 (3):331–346, February 21, 2003. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Boukerche:2003:SHM

[512] Azzedine Boukerche, Sungbum Hong, and Tom Jacob. Synchronization and handoff management schemes for wireless multimedia systems. Computer Networks (Amsterdam, Netherlands: 1999), 41(3):347–362, February 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBc

[513] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 41(3):??, February 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wen:2003:LEA

[514] Wushao Wen, Biswanath Mukherjee, S.-H. Gary Chan, and Dipak Ghosal. LVMSR: an efficient algorithm to multicast layered video. Computer Networks (Amsterdam, Netherlands: 1999), 41(4):363–383, March 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Corchuelo:2003:AOI

[515] Rafael Corchuelo, José A. Pérez, and Antonio Ruiz-Cortés. Aspect-oriented interaction in multi-organisational Web-based systems. Computer Networks (Amsterdam, Netherlands: 1999), 41(4):385–406, March 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Maier:2003:HMP

[516] Martin Maier, Martin Reisslein, and Adam Wolisz. A hybrid MAC protocol for a metro WDM network using multiple free spectral ranges of an arrayedwaveguide grating. Computer Networks (Amsterdam, Netherlands: 1999), 41 (4):407–433, March 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

El-Gendy:2003:AFF

[517] Mohamed A. El-Gendy and Kang G. Shin. Assured forwarding fairness using equation-based packet marking and packet separation. Computer Networks (Amsterdam, Netherlands: 1999), 41 (4):435–450, March 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2003:FOS

[518] Xin Liu, Edwin K. P. Chong, and Ness B. Shroff. A framework for opportunistic scheduling in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 41(4): 451–474, March 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Casetti:2003:NCQ

[519] C. Casetti, R. Lo Cigno, M. Mellia, M. Munafò, and Z. Zsóka. A new class of QoS routing strategies based on network graph reduction. Computer Networks (Amsterdam, Netherlands: 1999), 41(4):475–487, March 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mandjes:2003:SMN

[520] Michel Mandjes, Debasis Mitra, and Werner Scheinhardt. Simple models of network access, with applications to the design of joint rate and admission control. Computer Networks (Amsterdam, Netherlands: 1999), 41 (4):489–504, March 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Aad:2003:PW

[521] Imad Aad and Claude Castelluccia. Priorities in WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 41(4):505–526, March 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Naor:2003:ADA

[522] Zohar Naor and Hanoch Levy. Announced dynamic access probability protocol for next generation wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 41(4): 527–544, March 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBd

[523] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 41(4):??, March 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Engbersen:2003:NP

[524] Ton Engbersen. Network processors. Computer Networks (Amsterdam, Netherlands: 1999), 41(5):545– 547, April 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Arts:2003:NPR

[525] F. Arts, P. Barri, I. Clemminck, A. Niemegeers, B. Pauwels, G. Taildeman, and M. Vrana. Network processor requirements and benchmarking. Computer Networks (Amsterdam, Netherlands: 1999), 41(5):549–562, April 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Venkatachalam:2003:HFD

[526] Muthu Venkatachalam, Prashant Chandra, and Raj Yavatkar. A highly flexible, distributed multiprocessor architecture for network processing. Computer Networks (Amsterdam, Netherlands: 1999), 41(5):563– 586, April 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Peyravian:2003:FAC

[527] Mohammad Peyravian and Jean Calvignac. Fundamental architectural considerations for network processors. Computer Networks (Amsterdam, Netherlands: 1999), 41(5):587– 600, April 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wolf:2003:PSN

[528] Tilman Wolf, Prashanth Pappu, and Mark A. Franklin. Predictive scheduling of network processors. Computer Networks (Amsterdam, Netherlands: 1999), 41(5):601–621, April 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gabrani:2003:DMM

[529] Maria Gabrani, Gero Dittmann, Andreas Döring, Andreas Herkersdorf, Patricia Sagmeister, and Jan van Lunteren. Design methodology for a modular service-driven network processor architecture. Computer Networks (Amsterdam, Netherlands: 1999), 41(5): 623–640, April 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chakraborty:2003:PEN

[530] Samarjit Chakraborty, Simon Künzli, Lothar Thiele, Andreas Herkersdorf, and Patricia Sagmeister. Performance evaluation of network processor architectures: combining simulation with analytical estimation. Computer Networks (Amsterdam, Netherlands: 1999), 41(5):641–665, April 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Theelen:2003:PMN

[531] B. D. Theelen, J. P. M. Voeten, and R. D. J. Kramer. Performance modelling of a network processor using POOSL. Computer Networks (Amsterdam, Netherlands: 1999), 41(5): 667–684, April 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBe

[532] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 41(5):??, April 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2003:E

[533] Ian F. Akyildiz and Harry Rudin. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 41(6): 685–686, April 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sivakumar:2003:BBS

[534] Raghupathy Sivakumar, Prasun Sinha, and Vaduvur Bharghavan. Braving the broadcast storm: infrastructural support for ad hoc routing. Computer Networks (Amsterdam, Netherlands: 1999), 41(6):687–706, April 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cohen:2003:PCD

[535] Edith Cohen and Haim Kaplan. Proactive caching of DNS records: addressing a performance bottleneck. Computer Networks (Amsterdam, Netherlands: 1999), 41(6):707–726, April 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bianco:2003:SAM

[536] A. Bianco, G. Galante, E. Leonardi, F. Neri, and A. Nucci. Scheduling algorithms for multicast traffic in TDM/WDM networks with arbitrary tuning latencies. Computer Networks (Amsterdam, Netherlands: 1999), 41 (6):727–742, April 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ko:2003:ABP

[537] Young-Bae Ko and Nitin H. Vaidya. Anycasting-based protocol for geocast service in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 41(6): 743–760, April 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anjali:2003:NST

[538] T. Anjali, C. Scoglio, and G. Uhl. A new scheme for traffic estimation and resource allocation for bandwidth brokers. Computer Networks (Amsterdam, Netherlands: 1999), 41(6):

761–777, April 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2003:PTS

[539] Xuan Chen and John Heidemann. Preferential treatment for short flows to reduce web latency. Computer Networks (Amsterdam, Netherlands: 1999), 41 (6):779–794, April 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Galmes:2003:CAD

[540] Sebastià Galmés and Ramon Puigjaner. Correlation analysis of a discrete-time flexible arrival process. Computer Networks (Amsterdam, Netherlands: 1999), 41(6):795– 814, April 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:AR

[541] Anonymous. Acknowledgement to reviewers. Computer Networks (Amsterdam, Netherlands: 1999), 41(6): 815–818, April 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:AI

[542] Anonymous. Author index. Computer Networks (Amsterdam, Netherlands: 1999), 41(6):819–821, April 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:SI

[543] Anonymous. Subject index. Computer Networks (Amsterdam, Netherlands: 1999), 41(6):823–824, April 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBf

[544] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 41(6):??, April 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Montpetit:2003:SHN

[545] Marie-José Montpetit and David Starobinski. Small and home networks. Computer Networks (Amsterdam, Netherlands: 1999), 42(1):1–5, May 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ganz:2003:QSS

[546] Aura Ganz, Kitti Wongthavarawat, and Anan Phonphoem. Q-Soft: software framework for QoS support in home networks. Computer Networks (Amsterdam, Netherlands: 1999), 42 (1):7–22, May 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Manzoni:2003:PIB

[547] Pietro Manzoni and Juan Carlos Cano. Providing interoperability between IEEE 802.11 and Bluetooth protocols for Home Area Networks. Computer Networks (Amsterdam, Netherlands: 1999), 42(1):23–37, May 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Qiao:2003:EEP

[548] Daji Qiao, Sunghyun Choi, Amjad Soomro, and Kang G. Shin. Energyefficient PCF operation of IEEE 802.11a WLANs via transmit power control. Computer Networks (Amsterdam, Netherlands: 1999), 42(1):39–54, May 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nomikos:2003:SMN

[549] Christos Nomikos, Aris Pagourtzis, and Stathis Zachos. Satisfying a maximum number of pre-routed requests in all-optical rings. Computer Networks (Amsterdam, Netherlands: 1999), 42 (1):55–63, May 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Su:2003:PRE

[550] Xun Su and Gustavo de Veciana. Predictive routing to enhance QoS for stream-based flows sharing excess bandwidth. Computer Networks (Amsterdam, Netherlands: 1999), 42 (1):65–80, May 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mellia:2003:TAP

[551] Marco Mellia, Ion Stoica, and Hui Zhang. TCP-aware packet marking in networks with DiffServ support. Computer Networks (Amsterdam, Netherlands: 1999), 42(1):81–100, May 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2003:REC

[552] S. Y. Wang. Reducing the energy consumption caused by flooding messages in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 42(1):101–118, May 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2003:OCQ

[553] Huamin Chen and Prasant Mohapatra. Overload control in QoS-aware web servers. Computer Networks (Amsterdam, Netherlands: 1999), 42(1): 119–133, May 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rhee:2003:CRF

[554] Seung Hyong Rhee, Takis Konstantopoulos, Hyukjoon Lee, and Kwangsue Chung. Competitive routing and flow control in communication networks of parallel links. Computer Networks (Amsterdam, Netherlands: 1999), 42(1):135–140, May 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBg

[555] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 42(1):??, May 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Misic:2003:UCL

[556] Jelena Mišić, Vojislav B. Mišić, and Yat Kwan Tang. On uplink call level QoS in DS-CDMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 42(2):141–173, June 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2003:DIN

[557] S. Y. Wang, C. L. Chou, C. H. Huang, C. C. Hwang, Z. M. Yang, C. C. Chiou, and C. C. Lin. The design and implementation of the NCTUns 1.0 network simulator. Computer Networks (Amsterdam, Netherlands: 1999), 42 (2):175–197, June 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jozsa:2003:SRS

[558] Balázs Gábor Józsa and Márton Makai. On the solution of reroute sequence planning problem in MPLS networks. Computer Networks (Amsterdam, Netherlands: 1999), 42(2):199– 210, June 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Stafford:2003:CCR

[559] Matthew Stafford, Xiangying Yang, and Gustavo de Veciana. Connection caching to reduce signaling loads with applications to softswitch telephony. Computer Networks (Amsterdam, Netherlands: 1999), 42(2): 211–229, June 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mandjes:2003:PSU

[560] Michel Mandjes. Pricing strategies under heterogeneous service requirements. Computer Networks (Amsterdam, Netherlands: 1999), 42(2): 231–249, June 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Komolafe:2003:ONP

[561] Olufemi Komolafe and David Harle. Optimal node placement in an optical packet switching Manhattan street network. Computer Networks (Amsterdam, Netherlands: 1999), 42(2): 251–260, June 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Eddy:2003:CRB

[562] Wesley M. Eddy and Mark Allman. A comparison of RED’s byte and packet modes. Computer Networks (Amsterdam, Netherlands: 1999), 42(2): 261–280, June 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:CPa

[563] Anonymous. Call for papers. Computer Networks (Amsterdam, Netherlands: 1999), 42(2):281–282, June 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBh

[564] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 42(2):??, June 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Reed:2003:ISD

[565] Rick Reed. ITU-T system design languages (SDL). Computer Networks (Amsterdam, Netherlands: 1999), 42 (3):283–284, June 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Amyot:2003:IUR

[566] Daniel Amyot. Introduction to the User Requirements Notation: learning by example. Computer Networks (Amsterdam, Netherlands: 1999), 42 (3):285–301, June 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zheng:2003:TCM

[567] Tong Zheng and Ferhat Khendek. Time consistency of MSC-2000 specifications. Computer Networks (Amsterdam, Netherlands: 1999), 42(3): 303–322, June 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Graubmann:2003:DIB

[568] Peter Graubmann. Describing interactions between MSC components: the MSC connectors. Computer Networks (Amsterdam, Netherlands: 1999), 42 (3):323–342, June 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Glasser:2003:FSS

[569] U. Glässer, R. Gotzhein, and A. Prinz. The formal semantics of SDL-2000: Status and perspectives. Computer Networks (Amsterdam, Netherlands: 1999), 42(3):343–358, June 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wong:2003:CTS

[570] W. Eric Wong, Tatiana Sugeta, J. Jenny Li, and José C. Maldonado. Coverage testing software architectural design in SDL. Computer Networks (Amsterdam, Netherlands: 1999), 42 (3):359–374, June 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Grabowski:2003:ITT

[571] Jens Grabowski, Dieter Hogrefe, György Réthy, Ina Schieferdecker, Anthony Wiles, and Colin Willcock. An introduction to the testing and test control notation (TTCN-3). Computer Networks (Amsterdam, Netherlands: 1999), 42(3):375–403, June 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2003:GTS

[572] Nam Hee Lee and Sung Deok Cha. Generating test sequences from a set of MSCs. Computer Networks (Amsterdam, Netherlands: 1999), 42(3): 405–417, June 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBi

[573] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 42(3):??, June 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hu:2003:AMA

[574] Fei Hu, Neeraj K. Sharma, and Jim Ziobro. An accurate model for analyzing wireless TCP performance with the coexistence of non-TCP traffic. Computer Networks (Amsterdam, Netherlands: 1999), 42(4):419–439, July 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Brown:2003:LIQ

[575] Timothy X. Brown and Harold N. Gabow. The limits of inputqueued switch performance with future packet arrival information. Computer Networks (Amsterdam, Netherlands: 1999), 42(4):441–460, July 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Magana:2003:PAM

[576] E. Magaña, J. Aracil, and J. Villadangos. A protocol-adaptive monitoring tree for efficient design of traffic monitoring probes. Computer Networks (Amsterdam, Netherlands: 1999), 42 (4):461–479, July 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lotker:2003:NOF

[577] Zvi Lotker and Boaz Patt-Shamir. Nearly optimal FIFO buffer management for two packet classes. Computer Networks (Amsterdam, Netherlands: 1999), 42(4):481–492, July 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2003:SIW

[578] Chi-Hsiang Wu and Rong-Hong Jan. System integration of WAP and SMS for home network system. Computer Networks (Amsterdam, Netherlands: 1999), 42(4):493–502, July 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2003:BDC

[579] Yi Yang, Lei Zhang, Jogesh K. Muppala, and Samuel T. Chanson. Bandwidth-delay constrained routing algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 42(4): 503–520, July 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bonald:2003:CFL

[580] T. Bonald and J. W. Roberts. Congestion at flow level and the impact of user behaviour. Computer Networks (Amsterdam, Netherlands: 1999), 42 (4):521–536, July 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Boucherie:2003:THB

[581] Richard J. Boucherie and Jan van der Wal. Transient handover blocking probabilities in road covering cellular mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 42 (4):537–550, July 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBj

[582] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 42(4):??, July 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Goble:2003:SWE

[583] Carole Goble. The Semantic Web: an evolution for a revolution. Computer Networks (Amsterdam, Netherlands: 1999), 42(5):551–556, August 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guha:2003:TSW

[584] R. Guha and Rob McCool. TAP: a Semantic Web platform. Computer Networks (Amsterdam, Netherlands: 1999), 42(5):557–577, August 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Handschuh:2003:CCM

[585] Siegfried Handschuh and Steffen Staab. CREAM: CREAting Metadata for the Semantic Web. Computer Networks (Amsterdam, Netherlands: 1999), 42 (5):579–598, August 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fillies:2003:PAS

[586] Christian Fillies, Gay Wood-Albrecht, and Frauke Weichhardt. Pragmatic applications of the Semantic Web using SemTalk. Computer Networks (Amsterdam, Netherlands: 1999), 42(5): 599–615, August 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Karvounarakis:2003:QSW

[587] G. Karvounarakis, A. Magganaraki, S. Alexaki, V. Christophides, D. Plexousakis, M. Scholl, and K. Tolle. Querying the Semantic Web with RQL. Computer Networks (Amsterdam, Netherlands: 1999), 42(5): 617–640, August 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Florescu:2003:XXP

[588] Daniela Florescu, Andreas Grünhagen, and Donald Kossmann. XL: an XML programming language for Web service specification and composition. Computer Networks (Amsterdam, Netherlands: 1999), 42(5):641–660, August 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Trastour:2003:SWS

[589] David Trastour, Claudio Bartolini, and Chris Preist. Semantic Web support for the business-to-business e-commerce pre-contractual lifecycle. Computer Networks (Amsterdam, Netherlands: 1999), 42(5):661–673, August 5, 2003.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Narayanan:2003:ASW

[590] Srini Narayanan and Sheila McIlraith. Analysis and simulation of Web services. Computer Networks (Amsterdam, Netherlands: 1999), 42(5): 675–693, August 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:CPb

[591] Anonymous. Call for papers. Computer Networks (Amsterdam, Netherlands: 1999), 42(5):695, August 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBk

[592] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 42(5):??, August 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kalpakis:2003:EAM

[593] Konstantinos Kalpakis, Koustuv Dasgupta, and Parag Namjoshi. Efficient algorithms for maximum lifetime data gathering and aggregation in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 42 (6):697–716, August 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Baboescu:2003:FSC

[594] F. Baboescu and G. Varghese. Fast and scalable conflict detection for packet classifiers. Computer Networks (Amsterdam, Netherlands: 1999), 42(6): 717–735, August 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jmaiel:2003:DCP

[595] Mohamed Jmaiel and Peter Pepper. Development of communication protocols using algebraic and temporal specifications. Computer Networks (Amsterdam, Netherlands: 1999), 42(6): 737–764, August 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cui:2003:VPN

[596] Wei Cui and Mostafa A. Bassiouni. Virtual private network bandwidth management with traffic prediction. Computer Networks (Amsterdam, Netherlands: 1999), 42(6):765–778, August 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2003:BTC

[597] Junshan Zhang, Ming Hu, and Ness B. Shroff. Bursty traffic over CDMA: predictive MAI temporal structure, rate control and admission control. Computer Networks (Amsterdam, Netherlands: 1999), 42(6):779–795, August 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wills:2003:UBW

[598] Craig E. Wills, Gregory Trott, and Mikhail Mikhailov. Using bundles for Web content delivery. Computer Networks (Amsterdam, Netherlands: 1999), 42(6):797–817, August 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:AIVa

[599] Anonymous. Author index volume 42. Computer Networks (Amsterdam, Netherlands: 1999), 42(6):819– 821, August 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:SIVa

[600] Anonymous. Subject index volume 42. Computer Networks (Amsterdam, Netherlands: 1999), 42(6):822– 824, August 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBl

[601] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 42(6):??, August 21, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Feridun:2003:MIS

[602] Metin Feridun and Gabi Dreo Rodosek. Management of IT services. Computer Networks (Amsterdam, Netherlands: 1999), 43(1):1–2, September 16, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Reichl:2003:CPM

[603] Peter Reichl, David Hausheer, and Burkhard Stiller. The Cumulus Pricing model as an adaptive framework for feasible, efficient, and user-friendly tariffing of Internet services. Computer Networks (Amsterdam, Netherlands: 1999), 43(1):3–24, September 16, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Brunner:2003:DIT

[604] Marcus Brunner, Alberto Gonzalez, and Paloma Martinez. From dynamic IP transport service ordering to Diff-Serv network configuration. Computer Networks (Amsterdam, Netherlands: 1999), 43(1):25–41, September 16, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Clemm:2003:GPH

[605] Alexander Clemm, Fong Shen, and Victor Lee. Generic provisioning of heterogeneous services — a close encounter with service profiles. Computer Networks (Amsterdam, Netherlands: 1999), 43(1):43–57, September 16, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Subramanian:2003:QBM

[606] Mani Subramanian and Lundy Lewis. QoS and bandwidth management in broadband cable access network. Computer Networks (Amsterdam, Netherlands: 1999), 43(1):59–73, September 16, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBm

[607] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 43(1):??, September 16, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2003:IIS

[608] Ian F. Akyildiz, Özgür B. Akan, Chao Chen, Jian Fang, and Weilian Su. InterPlaNetary Internet: state-of-the-art and research challenges. Computer Networks (Amsterdam, Netherlands: 1999), 43(2):75–112, October 7, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mohorcic:2003:DWT

[609] M. Mohorcic, A. Svigelj, G. Kandus, Y. F. Hu, and R. E. Sheriff. Demographically weighted traffic flow models for adaptive routing in packet-switched non-geostationary satellite meshed networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(2): 113–131, October 7, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altman:2003:APM

[610] Eitan Altman, Rachid El Azouzi, and Odile Pourtallier. Avoiding paradoxes in multi-agent competitive routing. Computer Networks (Amsterdam, Netherlands: 1999), 43(2):133– 146, October 7, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2003:DMT

[611] Chun-Chuan Yang and Kwin-Yee Lin. Distributed mobile tracking: a novel location management scheme for routing improvement in cellular IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(2): 147–161, October 7, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zeng:2003:AGC

[612] Helen Zeng and Imrich Chlamtac. Adaptive guard channel allocation and blocking probability estimation in PCS networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(2):

163–176, October 7, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yeung:2003:CPD

[613] Kwan L. Yeung and Ho-Lun T. Wong. Caching policy design and cache allocation in active reliable multicast. Computer Networks (Amsterdam, Netherlands: 1999), 43(2):177–193, October 7, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Giladi:2003:PNR

[614] Ran Giladi, Ephraim Korach, and Rony Ohayon. Placement of network resources in communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(2): 195–209, October 7, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Balamash:2003:MWR

[615] Abdullah Balamash and Marwan Krunz. Modeling web requests: a multifractal approach. Computer Networks (Amsterdam, Netherlands: 1999), 43 (2):211–226, October 7, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lahanas:2003:EEF

[616] Adrian Lahanas and Vassilis Tsaoussidis. Exploiting the efficiency and fairness potential of AIMD-based congestion avoidance and control. Computer Networks (Amsterdam, Netherlands: 1999), 43(2):227–245, October 7, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBn

[617] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 43(2):??, October 7, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Quintero:2003:SMP

[618] Alejandro Quintero and Samuel Pierre. Sequential and multipopulation memetic algorithms for assigning cells to switches in mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(3): 247–261, October 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hu:2003:MCA

[619] Fei Hu and Neeraj K. Sharma. Multimedia call admission control in mobile networks: a dynamical reservationpool approach. Computer Networks (Amsterdam, Netherlands: 1999), 43 (3):263–288, October 22, 2003. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zheng:2003:DPU

[620] Q. Zheng and G. Mohan. Dynamic protection using integratedrouting approach in IP-over-WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(3):289– 305, October 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jiang:2003:RBG

[621] Yuming Jiang. Relationship between guaranteed rate server and latency rate server. Computer Networks (Amsterdam, Netherlands: 1999), 43(3): 307–315, October 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tseng:2003:PSP

[622] Yu-Chee Tseng, Chih-Shun Hsu, and Ten-Yueng Hsieh. Power-saving protocols for IEEE 802.11-based multi-hop ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 43 (3):317–337, October 22, 2003. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bicakci:2003:HCO

[623] Kemal Bicakci, Gene Tsudik, and Brian Tung. How to construct optimal one-time signatures. Computer Networks (Amsterdam, Netherlands: 1999), 43(3):339–349, October 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Siachalou:2003:EQR

[624] Stavroula Siachalou and Leonidas Georgiadis. Efficient QoS routing. Computer Networks (Amsterdam, Netherlands: 1999), 43(3):351– 367, October 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Burns:2003:EID

[625] Catherine M. Burns, Johnson Kuo, and Sylvia Ng. Ecological interface design: a new approach for visualizing network management. Computer Networks (Amsterdam, Netherlands: 1999), 43(3):369–388, October 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Varadarajan:2003:RRR

[626] Srinidhi Varadarajan, Naren Ramakrishnan, and Muthukumar Thirunavukkarasu. Reinforcing reachable routes. Computer Networks (Amsterdam, Netherlands: 1999), 43(3):389–416, October 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBo

[627] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 43(3):??, October 22, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cayirci:2003:WSN

[628] Erdal Cayirci, Ramesh Govindan, Taieb Znati, and Mani Srivastava. Wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(4):417–419, November 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Avancha:2003:SSN

[629] Sasikanth Avancha, Jeffrey Undercoffer, Anupam Joshi, and John Pinkston. Secure sensor networks for perimeter protection. Computer Networks (Amsterdam, Netherlands: 1999), 43(4): 421–435, November 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Helmy:2003:MAR

[630] Ahmed Helmy. Mobility-assisted resolution of queries in large-scale mobile sensor networks (MARQ). Computer Networks (Amsterdam, Netherlands: 1999), 43(4):437–458, November 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2003:GSS

[631] Hongwei Zhang and Anish Arora. GS 3: scalable self-configuration and self-healing in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(4):459– 480, November 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

De:2003:MMR

[632] Swades De, Chunming Qiao, and Hongyi Wu. Meshed multipath routing with selective forwarding: an efficient strategy in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(4):481– 497, November 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Langendoen:2003:DLW

[633] Koen Langendoen and Niels Reijers. Distributed localization in wireless sensor networks: a quantitative comparison. Computer Networks (Amsterdam, Netherlands: 1999), 43(4):499– 518, November 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Duarte-Melo:2003:DGW

[634] Enrique J. Duarte-Melo and Mingyan Liu. Data-gathering wireless sensor networks: organization and capacity. Computer Networks (Amsterdam, Netherlands: 1999), 43(4):519– 537, November 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBp

[635] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 43(4):??, November 15, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2003:SSA

[636] Ning Wang and George Pavlou. Scalable sender access control for bidirectional multicast routing. Computer Networks (Amsterdam, Netherlands: 1999), 43(5):539–555, December 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Orincsay:2003:PPP

[637] Dániel Orincsay, Balázs Szviatovszki, and Géza Böhm. Prompt partial path optimization in MPLS networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(5):557–572, December 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Estevez-Tapiador:2003:NCN

[638] Juan M. Estévez-Tapiador, Pedro García-Teodoro, and Jesús E. Díaz-Verdejo. NSDF: a computer network system description framework and its application to network security. Computer Networks (Amsterdam, Netherlands: 1999), 43(5):573–600, December 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Coffey:2003:FVI

[639] Tom Coffey, Reiner Dojen, and Tomas Flanagan. Formal verification: an imperative step in the design of security protocols. Computer Networks (Amsterdam, Netherlands: 1999), 43(5):

601–618, December 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2003:QDE

[640] Kyungsup Kim and Chong-Ho Choi. Queue delay estimation and its application to TCP Vegas. Computer Networks (Amsterdam, Netherlands: 1999), 43(5):619–631, December 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Low:2003:LST

[641] Steven H. Low, Fernando Paganini, Jiantao Wang, and John C. Doyle. Linear stability of TCP/RED and a scalable control. Computer Networks (Amsterdam, Netherlands: 1999), 43(5): 633–647, December 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Younis:2003:EAM

[642] Mohamed Younis, Moustafa Youssef, and Khaled Arisha. Energy-aware management for cluster-based sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(5): 649–668, December 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pyun:2003:GHF

[643] Kihyun Pyun, Junehwa Song, and Heung-Kyu Lee. A generalized hierarchical fair service curve algorithm for high network utilization and linksharing. Computer Networks (Amsterdam, Netherlands: 1999), 43(5): 669–694, December 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EBq

[644] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 43(5):??, December 5, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pan:2003:ODB

[645] Jianping Pan, Y. Thomas Hou, and Bo Li. An overview of DNS-based server selections in content distribution networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(6): 695–711, December 20, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chang:2003:AAM

[646] Chao hsu Chang and Eric Hsiao kuang Wu. AMRST: adaptive multicast routing protocol for satellite–terrestrial networks. Computer Networks (Amsterdam, Netherlands: 1999), 43(6):713– 734, December 20, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Seol:2003:FAI

[647] Soonuk Seol, Myungchul Kim, Sungwon Kang, and Jiwon Ryu. Fully automated interoperability test suite derivation for communication protocols. Computer Networks (Amsterdam, Netherlands: 1999), 43(6):735– 759, December 20, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dag:2003:PCC

[648] Tamer Dağ and Ioannis Stavrakakis. Proactive congestion control for high speed networks. Computer Networks (Amsterdam, Netherlands: 1999), 43 (6):761–775, December 20, 2003. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Barbancho:2003:CAC

[649] A. M. Barbancho and A. Peinado. Cryptanalysis of anonymous channel protocol for large-scale area in wireless communications. Computer Networks (Amsterdam, Netherlands: 1999), 43 (6):777–785, December 20, 2003. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Phatak:2003:IIT

[650] Dhananjay S. Phatak, Tom Goff, and Jim Plusquellic. IP-in-IP tunneling to enable the simultaneous use of multiple IP interfaces for network level connection striping. Computer Networks (Amsterdam, Netherlands: 1999), 43 (6):787–804, December 20, 2003. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Koskela:2003:WCO

[651] Timo Koskela, Jukka Heikkonen, and Kimmo Kaski. Web cache optimization with nonlinear model using object features. Computer Networks (Amsterdam, Netherlands: 1999), 43(6): 805–817, December 20, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:AIVb

[652] Anonymous. Author index volume 43. Computer Networks (Amsterdam, Netherlands: 1999), 43(6):819– 821, December 20, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:SIVb

[653] Anonymous. Subject index volume 43. Computer Networks (Amsterdam, Netherlands: 1999), 43(6):823– 824, December 20, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2003:EB

[654] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 43(6):CO2, December 20, 2003. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bandyopadhyay:2004:MCC

[655] Seema Bandyopadhyay and Edward J. Coyle. Minimizing communication costs in hierarchically-clustered networks of wireless sensors. Computer Networks (Amsterdam, Netherlands: 1999), 44(1):1–16, January 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Park:2004:ADV

[656] Eun-Chan Park, Hyuk Lim, Kyung-Joon Park, and Chong-Ho Choi. Analysis and design of the virtual rate control algorithm for stabilizing queues in TCP networks. Computer Networks (Amsterdam, Netherlands: 1999), 44 (1):17–41, January 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2004:EQR

[657] Bin Wang and Jennifer C. Hou. An efficient QoS routing algorithm for quorumcast communication. Computer Networks (Amsterdam, Netherlands: 1999), 44(1):43–61, January 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rubenstein:2004:IRM

[658] Dan Rubenstein, Sneha Kasera, Don Towsley, and Jim Kurose. Improving reliable multicast using active parity encoding services. Computer Networks (Amsterdam, Netherlands: 1999), 44 (1):63–78, January 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nelakuditi:2004:SCP

[659] Srihari Nelakuditi, Zhi-Li Zhang, and David H. C. Du. On selection of candidate paths for proportional routing. Computer Networks (Amsterdam, Netherlands: 1999), 44(1):79– 102, January 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Seibert:2004:AOS

[660] Cristina Hristea Seibert and Fouad A. Tobagi. Assessment and optimization of schemes for tracking and routing to mobile users in packet-based networks. Computer Networks (Amsterdam, Netherlands: 1999), 44(1): 103–133, January 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBa

[661] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 44(1):CO2, January 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kamra:2004:FAB

[662] Abhinav Kamra, Huzur Saran, Sandeep Sen, and Rajeev Shorey. Fair adaptive bandwidth allocation: a rate control based active queue management discipline. Computer Networks (Amsterdam, Netherlands: 1999), 44(2): 135–152, February 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Garetto:2004:MSL

[663] M. Garetto, R. Lo Cigno, M. Meo, and M. Ajmone Marsan. Modeling short-lived TCP connections with open multiclass queuing networks. Computer Networks (Amsterdam, Netherlands: 1999), 44(2):153–176, February 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sarkar:2004:SLD

[664] Uttam K. Sarkar, S. Ramakrishnan, and Dilip Sarkar. Study of long-duration MPEG-trace segmentation methods for developing framesize-based traffic models. Computer Networks (Amsterdam, Netherlands: 1999), 44(2):177–188, February 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rezvan:2004:DCA

[665] M. Rezvan, K. Pawlikowski, and H. Sirisena. A distributed cache architecture with snooping for QoS routing in large networks. Computer Networks (Amsterdam, Netherlands: 1999), 44 (2):189–209, February 5, 2004. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Habib:2004:EEM

[666] Ahsan Habib, Maleq Khan, and Bharat Bhargava. Edge-to-edge measurementbased distributed network monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 44(2):211– 233, February 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pan:2004:RFT

[667] Jianping Pan, Y. Thomas Hou, and Bo Li. Retrieval and freshness thresholds in hierarchical caching systems. Computer Networks (Amsterdam, Netherlands: 1999), 44(2):235– 246, February 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bengi:2004:APE

[668] Kemal Bengi. Access protocols for an efficient and fair packet-switched IP-over-WDM metro network. Computer Networks (Amsterdam, Netherlands: 1999), 44(2):247–265, February 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Park:2004:APP

[669] Chang-Seop Park. Authentication protocol providing user anonymity and untraceability in wireless mobile communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 44 (2):267–273, February 5, 2004. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBb

[670] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 44(2):CO2, February 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Barbera:2004:FBM

[671] Mario Barbera, Alfio Lombardo, and Giovanni Schembra. A fluid-based model of time-limited TCP flows. Computer Networks (Amsterdam, Netherlands: 1999), 44(3):275–288, February 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Warkhede:2004:MRT

[672] Priyank Warkhede, Subhash Suri, and George Varghese. Multiway range trees: scalable IP lookup with fast updates. Computer Networks (Amsterdam, Netherlands: 1999), 44(3): 289–303, February 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Leligou:2004:MPE

[673] Helen-Catherine Leligou, John D. Angelopoulos, Charalambos Linardakis, and Alexandros Stavdas. A MAC protocol for efficient multiplexing QoS-sensitive and best-effort traffic in dynamically configurable WDM rings. Computer Networks (Amsterdam, Netherlands: 1999), 44(3):305– 317, February 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chou:2004:TEM

[674] Chun Tung Chou. Traffic engineering for MPLS-based virtual private networks. Computer Networks (Amsterdam, Netherlands: 1999), 44(3):319– 333, February 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Salvador:2004:MIT

[675] Paulo Salvador, António Pacheco, and Rui Valadas. Modeling IP traffic: joint characterization of packet arrivals and packet sizes using BMAPs. Computer Networks (Amsterdam, Netherlands: 1999), 44(3):335–352, February 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hefeeda:2004:HAC

[676] Mohamed M. Hefeeda, Bharat K. Bhargava, and David K. Y. Yau. A hybrid architecture for cost-effective on-demand media streaming. Computer Networks (Amsterdam, Netherlands: 1999), 44(3):353–382, February 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Plasser:2004:RFD

[677] Erich Plasser and Thomas Ziegler. A RED function design targeting link utilization and stable queue size behavior. Computer Networks (Amsterdam, Netherlands: 1999), 44(3):383– 410, February 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBc

[678] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 44(3):CO2, February 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Marsan:2004:SPS

[679] Marco Ajmone Marsan, Michela Meo, and Maurizio M. Munafò. Selected papers from the Second International Workshop on QoS in Multiservice IP Networks (QoS-IP 2003). Computer Networks (Amsterdam, Netherlands: 1999), 44(4):411–413, March 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Balk:2004:AVS

[680] A. Balk, M. Gerla, D. Maggiorini, and M. Sanadidi. Adaptive video streaming: pre-encoded MPEG-4 with bandwidth scaling. Computer Networks (Amsterdam, Netherlands: 1999), 44 (4):415–439, March 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lin:2004:MAU

[681] Tsung-Ching Lin, Yeali S. Sun, Shi-Chung Chang, Shao-I Chu, Yi-Ting Chou, and Mei-Wen Li. Management of abusive and unfair Internet access by quota-based priority control. Computer Networks (Amsterdam, Netherlands: 1999), 44(4):441–462, March 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fidler:2004:PBA

[682] Markus Fidler and Volker Sander. A parameter based admission control for differentiated services networks. Computer Networks (Amsterdam, Netherlands: 1999), 44(4):463–479, March 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2004:OFP

[683] Ning Wang and George Pavlou. An overlay framework for provisioning differentiated services in Source Specific Multicast. Computer Networks (Amsterdam, Netherlands: 1999), 44(4): 481–497, March 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Herreria-Alonso:2004:IAF

[684] Sergio Herrería-Alonso, Andrés Suárez-González, Manuel Fernández-Veiga, Raúl F. Rodríguez-Rubio, and Cándido López-García. Improving aggregate flow control in differentiated services networks. Computer Networks (Amsterdam, Netherlands: 1999), 44(4): 499–512, March 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Giordano:2004:DRC

[685] S. Giordano, M. Listanti, F. Mustacchio, S. Niccolini, S. Salsano, and L. Veltri. Dynamic resource configuration in DiffServ networks: control plane mechanisms and performance evaluation of a traffic control API. Computer Networks (Amsterdam, Netherlands: 1999), 44(4):513–527, March 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Toth:2004:CLL

[686] Gábor Tóth and Csaba Antal. Comparison of link-layer segmentation methods for UMTS terrestrial radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 44(4): 529–544, March 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chiasserini:2004:APM

[687] Carla-Fabiana Chiasserini, Francesca Cuomo, Leonardo Piacentini, Michele Rossi, Ilenia Tinirello, and Francesco Vacirca. Architectures and protocols for mobile computing applications: a reconfigurable approach. Computer Networks (Amsterdam, Netherlands: 1999), 44(4):545–567, March 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Callegati:2004:WTD

[688] F. Callegati, W. Cerroni, C. Raffaelli, and P. Zaffoni. Wavelength and time domain exploitation for QoS management in optical packet switches. Computer Networks (Amsterdam, Netherlands: 1999), 44(4):569–582, March 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBd

[689] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 44(4):CO2, March 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Giuliano:2004:CUS

[690] Romeo Giuliano, Gianluca Guidoni, Ibrahim Habib, and Franco Mazzenga. Coexistence of an ultrawideband spread spectrum system with fixed wireless access systems. Computer Networks (Amsterdam, Netherlands: 1999), 44(5):583–598, April 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2004:CST

[691] Yu-Chul Kim and Dong-Ho Cho. Considering spurious timeout in proxy for improving TCP performance in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 44 (5):599–616, April 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2004:SBP

[692] Jikai Li and Chunming Qiao. Schedule burst proactively for optical burst switched networks. Computer Networks (Amsterdam, Netherlands: 1999), 44(5):617–629, April 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altman:2004:SAR

[693] Eitan Altman and Tania Jiménez. Simulation analysis of RED with short lived TCP connections. Computer Networks (Amsterdam, Netherlands: 1999), 44(5):631–641, April 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Douligeris:2004:DAD

[694] Christos Douligeris and Aikaterini Mitrokotsa. DDoS attacks and defense mechanisms: classification and state-of-the-art. Computer Networks (Amsterdam, Netherlands: 1999), 44 (5):643–666, April 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pun:2004:DSC

[695] Konghong Pun and Mounir Hamdi. Dispatching schemes for Clos-network switches. Computer Networks (Amsterdam, Netherlands: 1999), 44(5): 667–679, April 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hierons:2004:TSS

[696] R. M. Hierons, T.-H. Kim, and H. Ural. On the testability of SDL specifications. Computer Networks (Amsterdam, Netherlands: 1999), 44(5):

681–700, April 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akan:2004:TAR

[697] Özgür B. Akan. On the throughput analysis of rate-based and windowbased congestion control schemes. Computer Networks (Amsterdam, Netherlands: 1999), 44(5):701–711, April 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBe

[698] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 44(5):CO2, April 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Striegel:2004:DSA

[699] A. Striegel and G. Manimaran. DSM-Cast: a scalable approach for Diff-Serv multicasting. Computer Networks (Amsterdam, Netherlands: 1999), 44 (6):713–735, April 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chang:2004:TCR

[700] Hyunseok Chang, Ramesh Govindan, Sugih Jamin, Scott J. Shenker, and Walter Willinger. Towards capturing representative AS-level Internet topologies. Computer Networks (Amsterdam, Netherlands: 1999), 44(6): 737–755, April 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ram:2004:ACI

[701] Abhishek Ram, Luiz A. DaSilva, and Srinidhi Varadarajan. Admission control by implicit signaling in support of voice over IP over ADSL. Computer Networks (Amsterdam, Netherlands: 1999), 44(6):757–772, April 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ma:2004:ASI

[702] Maode Ma and Xiaohong Huang. Adaptive scheduling for integrated traffic on WDM optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 44(6):773–788, April 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vida:2004:MHH

[703] Rolland Vida, Luís Henrique M. K. Costa, and Serge Fdida. Mobile hop-by-hop multicast routing. Computer Networks (Amsterdam, Netherlands: 1999), 44(6):789–812, April 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nandagopal:2004:SSD

[704] Thyagarajan Nandagopal, Kang-Won Lee, Jia-Ru Li, and Vaduvur Bharghavan. Scalable service differentiation using purely end-to-end mechanisms: features and limitations. Computer Networks (Amsterdam, Netherlands: 1999), 44(6):813–833, April 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Savagaonkar:2004:OPB

[705] Uday Savagaonkar, Edwin K. P. Chong, and Robert L. Givan. Online pricing for bandwidth provisioning in multi-class networks. Computer Networks (Amsterdam, Netherlands: 1999), 44(6):835–853, April 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2004:GRL

[706] X. Brian Zhang, Simon S. Lam, and Dong-Young Lee. Group rekeying with limited unicast recovery. Computer Networks (Amsterdam, Netherlands: 1999), 44(6):855–870, April 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:AIVa

[707] Anonymous. Author index volume 44. Computer Networks (Amsterdam, Netherlands: 1999), 44(6):871–873, April 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:SIVa

[708] Anonymous. Subject index volume 44. Computer Networks (Amsterdam, Netherlands: 1999), 44(6):875–876, April 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBf

[709] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 44(6):CO2, April 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hofmann:2004:GI

[710] Markus Hofmann, Robin Kravets, and Joseph D. Touch. The Global Internet. Computer Networks (Amsterdam, Netherlands: 1999), 45(1):1–3, May 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Banerjee:2004:EPL

[711] Suman Banerjee, Chris Kommareddy, and Bobby Bhattacharjee. Efficient peer location on the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 45(1):5–17, May 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fujita:2004:CGR

[712] Norihito Fujita, Yuichi Ishikawa, Atsushi Iwata, and Rauf Izmailov. Coarse-grain replica management strategies for dynamic replication of Web contents. Computer Networks (Amsterdam, Netherlands: 1999), 45 (1):19–34, May 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gast:2004:RIS

[713] Jim Gast and Paul Barford. Representing the Internet as a succinct forest. Computer Networks (Amsterdam, Netherlands: 1999), 45(1):35–44, May 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bu:2004:CBR

[714] Tian Bu, Lixin Gao, and Don Towsley. On characterizing BGP routing table growth. Computer Networks (Amsterdam, Netherlands: 1999), 45(1):45–54, May 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zappala:2004:MAA

[715] Daniel Zappala, Virginia Lo, and Chris GauthierDickey. The multicast address allocation problem: theory and practice. Computer Networks (Amsterdam, Netherlands: 1999), 45(1):55–73, May 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBg

[716] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 45(1):CO2, May 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2004:E

[717] Ian F. Akyildiz and Harry Rudin. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 45(2):75–76, June 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2004:MVI

[718] Ligang Wang, Anjali Agarwal, and J. William Atwood. Modelling and verification of interworking between SIP and H.323. Computer Networks (Amsterdam, Netherlands: 1999), 45 (2):77–98, June 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shek:2004:IAS

[719] Liza Lai-Yee Shek and Yu-Kwong Kwok. An integrated approach to scatternet traffic management in Bluetooth ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 45 (2):99–118, June 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cordeiro:2004:RPC

[720] Carlos de Morais Cordeiro, Sachin Abhyankar, and Dharma P. Agrawal. Reducing power consumption and enhancing performance by direct slave-to-slave and group communication in Bluetooth WPANs. Computer Networks (Amsterdam, Netherlands: 1999), 45 (2):119–141, June 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2004:SAE

[721] Chang N. Zhang and Chunren Lai. A systematic approach for encryption and authentication with fault tolerance. Computer Networks (Amsterdam, Netherlands: 1999), 45(2): 143–154, June 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bar-Noy:2004:EAP

[722] Amotz Bar-Noy, Vladimir Dreizin, and Boaz Patt-Shamir. Efficient algorithms for periodic scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 45(2):155–173, June 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Estevez-Tapiador:2004:MNH

[723] Juan M. Estévez-Tapiador, Pedro García-Teodoro, and Jesús E. Díaz-Verdejo. Measuring normality in HTTP traffic for anomaly-based intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 45 (2):175–193, June 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vlachos:2004:SAP

[724] Vasileios Vlachos, Stephanos Androutsellis-Theotokis, and Diomidis Spinellis. Security applications of peer-to-peer networks. Computer Networks (Amsterdam, Netherlands: 1999), 45(2):195– 205, June 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Galluccio:2004:TTP

[725] L. Galluccio, A. Leonardi, and G. Morabito. Tuning transmission power for TCP fairness in next generation wireless networks: an analytical paradigm. Computer Networks (Amsterdam, Netherlands: 1999), 45 (2):207–219, June 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Madhyastha:2004:EDT

[726] Harsha V. Madhyastha and C. Siva Ram Murthy. Efficient dynamic traffic grooming in service-differentiated WDM mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 45(2):221–235, June 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:AR

[727] Anonymous. Acknowledgement to reviewers. Computer Networks (Amsterdam, Netherlands: 1999), 45(2): 237–240, June 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBh

[728] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 45(2):CO2, June 5, 2004.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Blondia:2004:MOC

[729] Chris Blondia and Ioannis Stavrakakis. In memory of Olga Casals. Computer Networks (Amsterdam, Netherlands: 1999), 45(3):241–242, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EOC

[730] Anonymous. A eulogy for Olga Casals. Computer Networks (Amsterdam, Netherlands: 1999), 45(3):243, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vuran:2004:STC

[731] Mehmet C. Vuran, Özgür B. Akan, and Ian F. Akyildiz. Spatio-temporal correlation: theory and applications for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 45(3):245–259, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Casares-Giner:2004:ECI

[732] Vicente Casares-Giner, Pablo García-Escalle, and Vicent Pla. Evaluation of CELLULAR IP mobility tracking procedures. Computer Networks (Amsterdam, Netherlands: 1999), 45(3): 261–279, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Washington:2004:CBP

[733] Alicia Nicki Washington and Harry Perros. Call blocking probabilities in a traffic-groomed tandem optical network. Computer Networks (Amsterdam, Netherlands: 1999), 45(3): 281–294, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anastasi:2004:PSP

[734] G. Anastasi, M. Conti, E. Gregori, and A. Passarella. A performance study of power-saving polices for Wi-Fi hotspots. Computer Networks (Amsterdam, Netherlands: 1999), 45(3): 295–318, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Roberts:2004:SSB

[735] J. W. Roberts. A survey on statistical bandwidth sharing. Computer Networks (Amsterdam, Netherlands: 1999), 45(3):319–332, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Barcelo:2004:SIA

[736] José M. Barceló, Juan I. Nieto-Hipólito, and Jorge García-Vidal. Study of Internet autonomous system interconnectivity from BGP routing tables. Computer Networks (Amsterdam, Netherlands: 1999), 45(3): 333–344, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

DeCleyn:2004:SHS

[737] Peter De Cleyn, Nik Van den Wijngaert, Llorenç Cerdá, and Chris Blondia. A smooth handoff scheme using IEEE802.11 triggers—design and implementation. Computer Networks (Amsterdam, Netherlands: 1999), 45 (3):345–361, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Balaouras:2004:CSL

[738] P. Balaouras and I. Stavrakakis. Controlling smoothness and loss rate for elastic continuous media flows. Computer Networks (Amsterdam, Netherlands: 1999), 45(3):363–378, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBi

[739] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 45(3):CO2, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:POC

[740] Anonymous. Photo of Olga Casals. Computer Networks (Amsterdam, Netherlands: 1999), 45(3):iv, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Psounis:2004:MCW

[741] Konstantinos Psounis, An Zhu, Balaji Prabhakar, and Rajeev Motwani. Modeling correlations in web traces and implications for designing replacement policies. Computer Networks (Amsterdam, Netherlands: 1999), 45 (4):379–398, July 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shuai:2004:NGC

[742] Dianxun Shuai and Hongbin Zhao. A new generalized cellular automata approach to optimization of fast packet switching. Computer Networks (Amsterdam, Netherlands: 1999), 45(4):

399–419, July 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dikaiakos:2004:IIW

[743] Marios D. Dikaiakos. Intermediary infrastructures for the World Wide Web. Computer Networks (Amsterdam, Netherlands: 1999), 45(4):421– 447, July 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ngin:2004:STP

[744] Hoon-Tong Ngin and Chen-Khong Tham. Scaled time priority: an efficient approximation to waiting time priority. Computer Networks (Amsterdam, Netherlands: 1999), 45(4): 449–462, July 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shin:2004:SDM

[745] Haw-Yun Shin and Jean-Lien C. Wu. The study of dynamic multi-channel scheme with channel de-allocation in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 45 (4):463–482, July 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Critchlow:2004:SEA

[746] D. Critchlow and N. Zhang. Security enhanced accountable anonymous PKI certificates for mobile ecommerce. Computer Networks (Amsterdam, Netherlands: 1999), 45(4): 483–503, July 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

He:2004:TUP

[747] Jingyi He and S.-H. Gary Chan. TCP and UDP performance for Internet over optical packet-switched networks. Computer Networks (Amsterdam, Netherlands: 1999), 45(4): 505–521, July 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2004:DTC

[748] Liming Liu and Yumao Lu. Dynamic traffic controls for Web-server networks. Computer Networks (Amsterdam, Netherlands: 1999), 45(4): 523–536, July 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Steinder:2004:PFD

[749] M. Steinder and A. S. Sethi. Probabilistic fault diagnosis in communication systems through incremental hypothesis updating. Computer Networks (Amsterdam, Netherlands: 1999), 45 (4):537–562, July 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBj

[750] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 45(4):CO2, July 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Amyot:2004:DFI

[751] Daniel Amyot and Luigi Logrippo. Directions in feature interaction research. Computer Networks (Amsterdam, Netherlands: 1999), 45(5): 563–567, August 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Reiff-Marganiec:2004:FIP

[752] Stephan Reiff-Marganiec and Kenneth J. Turner. Feature interaction in policies. Computer Networks (Amsterdam, Netherlands: 1999), 45(5): 569–584, August 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dini:2004:PEM

[753] Petre Dini, Alexander Clemm, Tom Gray, Fuchun Joseph Lin, Luigi Logrippo, and Stephan Reiff-Marganiec. Policy-enabled mechanisms for feature interactions: reality, expectations, challenges. Computer Networks (Amsterdam, Netherlands: 1999), 45(5): 585–603, August 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nakamura:2004:DFI

[754] Masahide Nakamura, Pattara Leelaprute, Ken ichi Matsumoto, and Tohru Kikuno. On detecting feature interactions in the programmable service environment of Internet telephony. Computer Networks (Amsterdam, Netherlands: 1999), 45(5): 605–624, August 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Metzger:2004:FIE

[755] Andreas Metzger. Feature interactions in embedded control systems. Computer Networks (Amsterdam, Netherlands: 1999), 45(5):625–644, August 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zave:2004:CCT

[756] Pamela Zave, Healfdene H. Goguen, and Thomas M. Smith. Component coordination: a telecommunication case study. Computer Networks (Amsterdam, Netherlands: 1999), 45 (5):645–664, August 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Turner:2004:AIV

[757] Kenneth J. Turner. Analysing interactive voice services. Computer Networks (Amsterdam, Netherlands: 1999), 45 (5):665–685, August 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBk

[758] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 45(5):CO2, August 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ren:2004:HRT

[759] Kui Ren, Tieyan Li, Zhiguo Wan, Feng Bao, Robert H. Deng, and Kwangjo Kim. Highly reliable trust establishment scheme in ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 45(6):687–699, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altman:2004:SAG

[760] Eitan Altman, Rachid El Azouzi, and Tania Jiménez. Slotted Aloha as a game with partial information. Computer Networks (Amsterdam, Netherlands: 1999), 45(6):701–713, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lo:2004:ELA

[761] Shi-Wu Lo, Tei-Wei Kuo, Kam-Yiu Lam, and Guo-Hui Li. Efficient location area planning for cellular networks with hierarchical location databases. Computer Networks (Amsterdam, Netherlands: 1999), 45(6): 715–730, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2004:CSI

[762] Junfeng Wang, Mingtian Zhou, and Hongxia Zhou. Clock synchronization for Internet measurements: a clustering algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 45(6): 731–741, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

He:2004:CES

[763] Guanghui He, Yuan Gao, Jennifer C. Hou, and Kihong Park. A case for exploiting self-similarity of network traffic in TCP congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 45(6):743–766, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Marino:2004:ULS

[764] Perfecto Mariño, Miguel Ángel Domínguez, Francisco Poza, and Fernando Vázquez. Using LOTOS in the specification of industrial bus communication protocols. Computer Networks (Amsterdam, Netherlands: 1999), 45(6):767– 799, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Harrison:2004:ABC

[765] David Harrison, Yong Xia, Shivkumar Kalyanaraman, and Arvind Venkatesan. An accumulation-based, closedloop scheme for expected minimum rate and weighted rate services. Computer Networks (Amsterdam, Netherlands: 1999), 45(6):801–818, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tan:2004:LSS

[766] S. K. Tan, G. Mohan, and K. C. Chua. Link scheduling state information based offset management for fairness improvement in WDM optical burst switching networks. Computer Networks (Amsterdam, Netherlands: 1999), 45(6):819–834, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gevros:2004:DRM

[767] Panos Gevros and Jon Crowcroft. Distributed resource management with heterogeneous linear controls. Computer Networks (Amsterdam, Netherlands: 1999), 45(6):835–858, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:AIVb

[768] Anonymous. Author index volume 45. Computer Networks (Amsterdam, Netherlands: 1999), 45(6):859– 861, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:SIVb

[769] Anonymous. Subject index volume 45. Computer Networks (Amsterdam, Netherlands: 1999), 45(6):863– 865, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBl

[770] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 45(6):CO2, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Stiller:2004:IEP

[771] Burkhard Stiller, Lee McKnight, Martin Karsten, and Peter Reichl. Internet economics: pricing and policies. Computer Networks (Amsterdam, Netherlands: 1999), 46(1):1–2, September 16, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pongpaibool:2004:PEE

[772] Panita Pongpaibool and Hyong S. Kim. Providing end-to-end service level agreements across multiple ISP networks. Computer Networks (Amsterdam, Netherlands: 1999), 46(1): 3–18, September 16, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Heckmann:2004:OIP

[773] Oliver Heckmann, Jens Schmitt, and Ralf Steinmetz. Optimizing interconnection policies. Computer Networks (Amsterdam, Netherlands: 1999), 46 (1):19–39, September 16, 2004. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fulp:2004:BPP

[774] Errin W. Fulp and Douglas S. Reeves. Bandwidth provisioning and pricing for networks with multiple classes of service. Computer Networks (Amsterdam, Netherlands: 1999), 46(1):41– 52, September 16, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jin:2004:EBB

[775] Nan Jin and Scott Jordan. The effect of bandwidth and buffer pricing on resource allocation and QoS. Computer Networks (Amsterdam, Netherlands: 1999), 46(1):53–71, September 16, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ros:2004:MMP

[776] David Ros and Bruno Tuffin. A mathematical model of the Paris Metro Pricing scheme for charging packet networks. Computer Networks (Amsterdam, Netherlands: 1999), 46(1):73– 85, September 16, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2004:PAC

[777] Tianshu Li, Youssef Iraqi, and Raouf Boutaba. Pricing and admission control for QoS-enabled Internet. Computer Networks (Amsterdam, Netherlands: 1999), 46(1):87–110, September 16, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2004:CST

[778] Xin Wang and Henning Schulzrinne. Comparative study of two congestion pricing schemes: auction and tâtonnement. Computer Networks (Amsterdam, Netherlands: 1999), 46 (1):111–131, September 16, 2004. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Antoniadis:2004:CEI

[779] Panayotis Antoniadis, Costas Courcoubetis, and Robin Mason. Comparing economic incentives in peer-to-peer networks. Computer Networks (Amsterdam, Netherlands: 1999), 46(1): 133–146, September 16, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBm

[780] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 46(1):CO2, September 16, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Duan:2004:RAQ

[781] Qiang Duan and John N. Daigle. Resource allocation for quality of service provision in multistage buffered crossbar switches. Computer Networks (Amsterdam, Netherlands: 1999), 46(2): 147–168, October 7, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kamal:2004:PEP

[782] Ahmed E. Kamal and Hossam S. Hassanein. Performance evaluation of prioritized scheduling with buffer management for differentiated services architectures. Computer Networks (Amsterdam, Netherlands: 1999), 46(2): 169–180, October 7, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2004:NAI

[783] Kun hong Liu and Yong Xu. A new approach to improving the grooming performance with dynamic traffic in SONET rings. Computer Networks (Amsterdam, Netherlands: 1999), 46 (2):181–195, October 7, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Norden:2004:RBG

[784] Samphel Norden, Milind M. Buddhikot, Marcel Waldvogel, and Subhash Suri. Routing bandwidthguaranteed paths with restoration in label-switched networks. Computer Networks (Amsterdam, Netherlands: 1999), 46(2):197–218, October 7, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Libman:2004:OSW

[785] Lavy Libman and Ariel Orda. Optimal sliding-window strategies in networks with long round-trip delays. Computer Networks (Amsterdam, Netherlands: 1999), 46(2):219–235, October 7, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2004:RDT

[786] Jongmin Lee and Hojung Cha. A receiver-driven TCP flow control in CDMA2000 wireless networks with constrained mobile resources. Computer Networks (Amsterdam, Netherlands: 1999), 46(2):237–252, October 7, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fan:2004:PPI

[787] Jinliang Fan, Jun Xu, Mostafa H. Ammar, and Sue B. Moon. Prefixpreserving IP address anonymization: measurement-based security evaluation and a new cryptography-based scheme. Computer Networks (Amsterdam, Netherlands: 1999), 46(2): 253–272, October 7, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Korkmaz:2004:OBH

[788] Turgay Korkmaz, Marwan Krunz, and Jyothi Guntaka. OSPF-based hybrid approach for scalable dissemination of QoS parameters. Computer Networks (Amsterdam, Netherlands: 1999), 46 (2):273–293, October 7, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBn

[789] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 46(2):CO2, October 7, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Johnson:2004:NES

[790] Marjory Johnson. Networking for the earth sciences. Computer Networks (Amsterdam, Netherlands: 1999), 46 (3):295–298, October 22, 2004. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Loiacono:2004:NPM

[791] Joe Loiacono, Andy Germain, and Jeff Smith. Network performance measurements for NASA’s Earth Observation System. Computer Networks (Amsterdam, Netherlands: 1999), 46(3): 299–320, October 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hadjitheodosiou:2004:CSF

[792] Michael Hadjitheodosiou and Yingyong Chen. Communication support for future earth science space missions. Computer Networks (Amsterdam, Netherlands: 1999), 46(3):321–342, October 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Krishnan:2004:ETE

[793] Rajesh Krishnan, James P. G. Sterbenz, Wesley M. Eddy, Craig Partridge, and Mark Allman. Explicit transport error notification (ETEN) for error-prone wireless and satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 46(3): 343–362, October 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Horan:2004:TMS

[794] Stephen Horan, Anirban Chakraborti, Sandeep Muddasani, and Srinivasulu Narina. Testing MDP in a simulated space channel environment. Computer Networks (Amsterdam, Netherlands: 1999), 46(3):363–374, October 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nguyen:2004:BCS

[795] Alex T. Nguyen, Michael Hadjitheodosiou, and John S. Baras. Broadband communication support for earth observation from the International Space Station. Computer Networks (Amsterdam, Netherlands: 1999), 46(3): 375–388, October 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shen:2004:FRA

[796] Chien-Chung Shen, Sundaram Rajagopalan, Girish Borkar, and Chaiporn Jaikaeo. A flexible routing architecture for ad hoc space networks. Computer Networks (Amsterdam, Netherlands: 1999), 46(3):389–410, October 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Grossman:2004:ESD

[797] Robert L. Grossman, Yunhong Gu, David Hanley, Xinwei Hong, and Babu Krishnaswamy. Experimental studies of data transport and data access of earth-science data over networks with high bandwidth delay products. Computer Networks (Amsterdam, Netherlands: 1999), 46(3):411–421, October 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Johnson:2004:NTE

[798] Marjory Johnson, Kenneth Freeman, Raymond Gilstrap, and Richard Beck. Networking technologies enable advances in Earth Science. Computer Networks (Amsterdam, Netherlands: 1999), 46(3):423–435, October 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBo

[799] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 46(3):CO2, October 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Abdu:2004:FDE

[800] Hasina Abdu, Hanan Lutfiyya, and Michael A. Bauer. A framework for determining efficient management configurations. Computer Networks (Amsterdam, Netherlands: 1999), 46(4): 437–463, November 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Oyman:2004:OEC

[801] E. Ilker Oyman and Cem Ersoy. Overhead energy considerations for efficient routing in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 46(4):465– 478, November 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hussain:2004:DBS

[802] Alefiya Hussain, John Heidemann, and Christos Papadopoulos. Distinguishing between single and multi-source attacks using signal processing. Computer Networks (Amsterdam, Netherlands: 1999), 46(4):479–503, November 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hu:2004:TAO

[803] Ming Hu, Junshan Zhang, and John Sadowsky. Traffic aided opportunistic scheduling for wireless networks: algorithms and performance bounds. Computer Networks (Amsterdam, Netherlands: 1999), 46(4):505–518, November 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sang:2004:MBV

[804] Aimin Sang and San qi Li. Measurementbased virtual queue (VQ): to estimate the real-time bandwidth demand under loss constraint. Computer Networks (Amsterdam, Netherlands: 1999), 46 (4):519–539, November 15, 2004. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yan:2004:HCH

[805] Zhang Yan and Soong Boon-Hee. Handoff counting in hierarchical cellular system with overflow scheme. Computer Networks (Amsterdam, Netherlands: 1999), 46(4):541–554, November 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2004:USW

[806] Hui Zhang, Ashish Goel, and Ramesh Govindan. Using the small-world model to improve Freenet performance. Computer Networks (Amsterdam, Netherlands: 1999), 46(4):555– 574, November 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBp

[807] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 46(4):CO2, November 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Parr:2004:MCS

[808] G. Parr. Military communications systems and technologies. Computer Networks (Amsterdam, Netherlands: 1999), 46(5):575–579, December 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Strassner:2004:LMD

[809] John Strassner and Craig Dupler. LANS: a model-driven environment for implementing location-aware network services. Computer Networks (Amsterdam, Netherlands: 1999), 46(5): 581–603, December 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Arora:2004:LSW

[810] A. Arora, P. Dutta, S. Bapat, V. Kulathumani, H. Zhang, V. Naik, V. Mittal, H. Cao, M. Demirbas, M. Gouda, et al. A line in the sand: a wireless sensor network for target detection, classification, and tracking. Computer Networks (Amsterdam, Netherlands: 1999), 46(5):605–634, December 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Al:2004:ITR

[811] Ahmed Abd El Al, Tarek Saadawi, and Myung Lee. Improving throughput and reliability in mobile wireless networks via transport layer bandwidth aggregation. Computer Networks (Amsterdam, Netherlands: 1999), 46(5): 635–649, December 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Saffre:2004:RSN

[812] Fabrice Saffre, Richard Tateson, and Robert Ghanea-Hercock. Reliable sensor networks using decentralised channel selection. Computer Networks (Amsterdam, Netherlands: 1999), 46(5): 651–663, December 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Syrotiuk:2004:DSU

[813] V. R. Syrotiuk, M. Cui, S. Ramkumar, and C. J. Colbourn. Dynamic spectrum utilization in ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 46(5):665–678, December 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2004:PPL

[814] Yuecheng Zhang and Liang Cheng. PLACE: Protocol for Location And Coordinate Estimation—a wireless sensor network approach. Computer Networks (Amsterdam, Netherlands: 1999), 46(5):679–693, December 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Onel:2004:MHD

[815] Tolga Onel, Cem Ersoy, Erdal Cayirci, and Gerard Parr. A multicriteria handoff decision scheme for the next generation tactical communications systems. Computer Networks (Amsterdam, Netherlands: 1999), 46(5):695– 708, December 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bigham:2004:UIA

[816] John Bigham, Laurie Cuthbert, Xu Yang, Ning Lu, and Damian Ryan. Using intelligent agents for managing resources in military communications. Computer Networks (Amsterdam, Netherlands: 1999), 46(5):709– 721, December 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chivers:2004:SDF

[817] Howard Chivers and John A. Clark. Smart dust, friend or foe?—replacing identity with configuration trust. Computer Networks (Amsterdam, Netherlands: 1999), 46(5):723–740, December 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fagoonee:2004:ATC

[818] Lina Fagoonee and Bahram Honary. Application of turbo codes to tactical communications. Computer Networks (Amsterdam, Netherlands: 1999), 46 (5):741–749, December 5, 2004. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBq

[819] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 46(5):CO2, December 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2004:AMA

[820] Ning Zhang, Qi Shi, Madjid Merabti, and Robert Askwith. Autonomous mobile agent based fair exchange. Computer Networks (Amsterdam, Netherlands: 1999), 46(6):751–770, December 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

McKinnon:2004:CCS

[821] A. David McKinnon, David E. Bakken, and John C. Shovic. A configurable cryptography subsystem in a middleware framework for embedded systems. Computer Networks (Amsterdam, Netherlands: 1999), 46(6):771– 795, December 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ekici:2004:SPB

[822] Eylem Ekici. On signaling performance bounds of location management in Next Generation Wireless Networks. Computer Networks (Amsterdam, Netherlands: 1999), 46(6):797– 816, December 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jan:2004:IAC

[823] Rong-Hong Jan, Hung-Chi Chu, and Yi-Fang Lee. Improving the accuracy of cell-based positioning for wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 46(6):817– 827, December 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2004:DRI

[824] J. Li, G. Mohan, E. Cheng Tien, and K. C. Chua. Dynamic routing with inaccurate link state information in integrated IP-over-WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 46(6):829–851, December 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

DAntonio:2004:MSL

[825] S. D’Antonio, M. D’Arienzo, M. Esposito, S. P. Romano, and G. Ventre. Managing service level agreements in Premium IP networks: a businessoriented approach. Computer Networks (Amsterdam, Netherlands: 1999), 46 (6):853–866, December 20, 2004. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2004:QAR

[826] Huan Chen, Sunil Kumar, and C.- C. Jay Kuo. QoS-aware radio resource management scheme for CDMA cellular networks based on dynamic interference guard margin (IGM). Computer Networks (Amsterdam, Netherlands: 1999), 46(6):867–879, December 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:AIVc

[827] Anonymous. Author index volume 46. Computer Networks (Amsterdam, Netherlands: 1999), 46(6):881– 883, December 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:SIVc

[828] Anonymous. Subject index volume 46. Computer Networks (Amsterdam, Netherlands: 1999), 46(6):884– 886, December 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2004:EBr

[829] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 46(6):CO2, December 20, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mellia:2005:MIT

[830] M. Mellia, R. Lo Cigno, and F. Neri. Measuring IP and TCP behavior on edge nodes with Tstat. Computer Networks (Amsterdam, Netherlands: 1999), 47(1):1–21, January 14, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kwon:2005:PAO

[831] Minseok Kwon and Sonia Fahmy. Pathaware overlay multicast. Computer Networks (Amsterdam, Netherlands: 1999), 47(1):23–45, January 14, 2005.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ricciato:2005:RDT

[832] Fabio Ricciato and Ugo Monaco. Routing demands with time-varying bandwidth profiles on a MPLS network. Computer Networks (Amsterdam, Netherlands: 1999), 47(1): 47–61, January 14, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2005:DPM

[833] Shih-Lin Wu, Pao-Chu Tseng, and Zi-Tsan Chou. Distributed power management protocols for multi-hop mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 47 (1):63–85, January 14, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2005:EQS

[834] Wei Liu, Wenjing Lou, and Yuguang Fang. An efficient quality of service routing algorithm for delay-sensitive applications. Computer Networks (Amsterdam, Netherlands: 1999), 47(1): 87–104, January 14, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kumar:2005:RMN

[835] K. J. Kumar, B. S. Manoj, and C. Siva Ram Murthy. RT-MuPAC: A new multi-power architecture for voice cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 47 (1):105–128, January 14, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tsai:2005:RDS

[836] Hsien-Ming Tsai, Ai-Chun Pang, Yung-Chun Lin, and Yi-Bing Lin. Repacking on demand for speed-sensitive channel assignment. Computer Networks (Amsterdam, Netherlands: 1999), 47 (1):129–146, January 14, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBa

[837] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 47(1):CO2, January 14, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jamalipour:2005:WIT

[838] Abbas Jamalipour. Wireless IP through integration of wireless LAN and cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 47(2):147–150, February 4, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ruggeri:2005:BWL

[839] Giuseppe Ruggeri, Antonio Iera, and Sergio Polito. 802.11-based wireless-LAN and UMTS interworking: requirements, proposed solutions and open issues. Computer Networks (Amsterdam, Netherlands: 1999), 47(2): 151–166, February 4, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2005:AQF

[840] Xin Gang Wang, Geyong Min, John E. Mellor, Khalid Al-Begain, and Lin Guan. An adaptive QoS framework for integrated cellular and WLAN networks. Computer Networks (Amsterdam, Netherlands: 1999), 47(2): 167–183, February 4, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ylianttila:2005:AHL

[841] M. Ylianttila, J. Mäkelä, and K. Pahlavan. Analysis of handoff in a location-aware vertical multi-access network. Computer Networks (Amsterdam, Netherlands: 1999), 47(2): 185–201, February 4, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Calvagna:2005:MQS

[842] Andrea Calvagna, Aurelio La Corte, and Sabrina Sicari. Mobility and quality of service across heterogeneous wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 47 (2):203–217, February 4, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xu:2005:ITP

[843] Kai Xu, Ye Tian, and Nirwan Ansari. Improving TCP performance in integrated wireless communications networks. Computer Networks (Amsterdam, Netherlands: 1999), 47(2): 219–237, February 4, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Korakis:2005:PQS

[844] Thanasis Korakis and Leandros Tassiulas. Providing quality of service guarantees in wireless LANs compliant with 802.11e. Computer Networks (Amsterdam, Netherlands: 1999), 47 (2):239–255, February 4, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Palmieri:2005:MBA

[845] Francesco Palmieri. An MPLS-based architecture for scalable QoS and traffic engineering in converged multiservice mobile IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 47(2):257–269, February 4, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2005:LMC

[846] Kang-Won Lee, Young-Bae Ko, and Thyaga Nandagopal. Load mitigation in cellular data networks by peer data sharing over WLAN channels. Computer Networks (Amsterdam, Netherlands: 1999), 47(2):271–286, February 4, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Taleb:2005:DMS

[847] Tarik Taleb, Nei Kato, and Yoshiaki Nemoto. On-demand media streaming to hybrid wired/wireless networks over quasi-geostationary satellite systems. Computer Networks (Amsterdam, Netherlands: 1999), 47(2):287– 306, February 4, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBb

[848] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 47(2):CO2, February 4, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altman:2005:PAA

[849] Eitan Altman, Konstantin Avrachenkov, Chadi Barakat, and Parijat Dube. Performance analysis of AIMD mechanisms over a multi-state Markovian path. Computer Networks (Amsterdam, Netherlands: 1999), 47(3):307– 326, February 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Norden:2005:APD

[850] Samphel Norden. Analyzing the performance of deferred reservations. Computer Networks (Amsterdam, Netherlands: 1999), 47(3):327–349, February 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bicakci:2005:ISA

[851] Kemal Bicakci and Nazife Baykal. Improved server assisted signatures. Computer Networks (Amsterdam, Netherlands: 1999), 47(3):351–366, February 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cao:2005:RFQ

[852] Zhiruo Cao, Ellen Zegura, and Zheng Wang. Rainbow fair queueing: theory and applications. Computer Networks (Amsterdam, Netherlands: 1999), 47 (3):367–392, February 21, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2005:DLB

[853] J. Li, G. Mohan, and K. C. Chua. Dynamic load balancing in IP-over-WDM optical burst switching networks. Computer Networks (Amsterdam, Netherlands: 1999), 47(3):393–408, February 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Laoutaris:2005:OSC

[854] Nikolaos Laoutaris, Vassilios Zissimopoulos, and Ioannis Stavrakakis. On the optimization of storage capacity allocation for content distribution. Computer Networks (Amsterdam, Netherlands: 1999), 47(3):409–428, February 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Snoeyink:2005:LBM

[855] Jack Snoeyink, Subhash Suri, and George Varghese. A lower bound for multicast key distribution. Computer Networks (Amsterdam, Netherlands: 1999), 47(3):429–441, February 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBc

[856] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 47(3):CO2, February 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2005:E

[857] Ian F. Akyildiz and Harry Rudin. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 47(4): 443–444, March 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2005:WMN

[858] Ian F. Akyildiz, Xudong Wang, and Weilin Wang. Wireless mesh networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 47(4): 445–487, March 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lev-Tov:2005:PTA

[859] Nissan Lev-Tov and David Peleg. Polynomial time approximation schemes for base station coverage with minimum total radii. Computer Networks (Amsterdam, Netherlands: 1999), 47(4): 489–501, March 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ziviani:2005:IAM

[860] Artur Ziviani, Serge Fdida, José F. de Rezende, and Otto Carlos M. B. Duarte. Improving the accuracy of measurement-based geographic location of Internet hosts. Computer Networks (Amsterdam, Netherlands: 1999), 47(4):503–523, March 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Long:2005:YAQ

[861] Chengnian Long, Bin Zhao, Xinping Guan, and Jun Yang. The Yellow active queue management algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 47(4):525–550, March 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bouras:2005:SHS

[862] Ch. Bouras and A. Gkamas. SRAMT: a hybrid sender and receiver-based adaptation scheme for TCP friendly multicast transmission. Computer Networks (Amsterdam, Netherlands: 1999), 47 (4):551–575, March 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altuntas:2005:MMA

[863] Süleyman Altuntaş and Buyurman Baykal. Mobile multi-access IP: a proposal for mobile multi-access management in future wireless IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 47(4): 577–592, March 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:AR

[864] Anonymous. Acknowledgement to reviewers. Computer Networks (Amsterdam, Netherlands: 1999), 47(4): 593–598, March 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBd

[865] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 47(4):CO2, March 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bhasin:2005:II

[866] K. B. Bhasin. Interplanetary Internet. Computer Networks (Amsterdam, Netherlands: 1999), 47(5): 599–601, April 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hogie:2005:USI

[867] Keith Hogie, Ed Criscuolo, and Ron Parise. Using standard Internet Protocols and applications in space. Computer Networks (Amsterdam, Netherlands: 1999), 47(5):603–650, April 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rash:2005:ITF

[868] James Rash, Keith Hogie, and Ralph Casasanta. Internet technology for future space missions. Computer Networks (Amsterdam, Netherlands: 1999), 47(5):651–659, April 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2005:PAI

[869] Chunmei Liu and Eytan Modiano. On the performance of additive increase multiplicative decrease (AIMD) protocols in hybrid space-terrestrial networks. Computer Networks (Amsterdam, Netherlands: 1999), 47(5): 661–678, April 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hadjitheodosiou:2005:FAS

[870] Michael Hadjitheodosiou, Hui Zeng, Alex Nguyen, and Brenda L. Ellis. Flexible access for a space communications network with IP functionality. Computer Networks (Amsterdam, Netherlands: 1999), 47(5):679– 700, April 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Clare:2005:SBM

[871] Loren P. Clare, Jay L. Gao, Esther H. Jennings, and Clayton Okino. Spacebased multi-hop networking. Computer Networks (Amsterdam, Netherlands: 1999), 47(5):701–724, April 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bergamo:2005:HTD

[872] Marcos A. Bergamo. High-throughput Distributed Spacecraft Network: architecture and multiple access technologies. Computer Networks (Amsterdam, Netherlands: 1999), 47(5): 725–749, April 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chukkala:2005:RFC

[873] Vishwanath Chukkala, Phillip DeLeon, Stephen Horan, and Vijayakumar Velusamy. Radio frequency channel modeling for proximity networks on the martian surface. Computer Networks (Amsterdam, Netherlands: 1999), 47 (5):751–763, April 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Beck:2005:SBE

[874] Richard A. Beck, Robert K. Vincent, Doyle W. Watts, Marc A. Seibert, David P. Pleva, Michael A. Cauley, Calvin T. Ramos, Theresa M. Scott, Dean W. Harter, Mary Vickerman, et al. A space-based end-to-end prototype geographic information network for lunar and planetary exploration and emergency response (2002 and 2003 field experiments). Computer Networks (Amsterdam, Netherlands: 1999), 47 (5):765–783, April 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBe

[875] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 47(5):CO2, April 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rieck:2005:DRA

[876] Michael Q. Rieck, Sukesh Pai, and Subhankar Dhar. Distributed routing algorithms for multi-hop ad hoc networks using d-hop connected ddominating sets. Computer Networks (Amsterdam, Netherlands: 1999), 47 (6):785–799, April 22, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Beard:2005:PDB

[877] Cory Beard. Preemptive and delaybased mechanisms to provide preference to emergency traffic. Computer Networks (Amsterdam, Netherlands: 1999), 47(6):801–824, April 22, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Brunato:2005:SLT

[878] Mauro Brunato and Roberto Battiti. Statistical learning theory for location fingerprinting in wireless LANs. Computer Networks (Amsterdam, Netherlands: 1999), 47(6):825–845, April 22, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yaghmaee:2005:DPE

[879] Mohammad Hossien Yaghmaee, Mohammad Bagher Menhaj, and Hale Amintoosi. Design and performance evaluation of a fuzzy based traffic conditioner for differentiated services. Computer Networks (Amsterdam, Netherlands: 1999), 47(6): 847–869, April 22, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shyu:2005:APC

[880] Woei-Luen Shyu, Cheng-Shong Wu, and Ting-Chao Hou. Aligned prefix caching based on singleton information. Computer Networks (Amsterdam, Netherlands: 1999), 47(6): 871–884, April 22, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2005:CAN

[881] Yong Liu, Chen-Khong Tham, and Yuming Jiang. Conformance analysis in networks with service level agreements. Computer Networks (Amsterdam, Netherlands: 1999), 47(6): 885–906, April 22, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hayel:2005:MAC

[882] Y. Hayel and B. Tuffin. A mathematical analysis of the cumulus pricing scheme. Computer Networks (Amsterdam, Netherlands: 1999), 47(6): 907–921, April 22, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cui:2005:PID

[883] Yong Cui, Jianping Wu, and Ke Xu. Precomputation for intra-domain QoS routing. Computer Networks (Amsterdam, Netherlands: 1999), 47(6): 923–937, April 22, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:AIV

[884] Anonymous. Author index volume 47. Computer Networks (Amsterdam, Netherlands: 1999), 47(6):939–941, April 22, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:SIV

[885] Anonymous. Subject index volume 47. Computer Networks (Amsterdam, Netherlands: 1999), 47(6):943–944, April 22, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBf

[886] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 47(6):CO2, April 22, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fairhurst:2005:IMT

[887] Gorry Fairhurst and Marie-José Montpetit. Internet over MPEG-2 transmission networks. Computer Networks (Amsterdam, Netherlands: 1999), 48 (1):1–3, May 16, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sooriyabandara:2005:LEP

[888] Mahesh Sooriyabandara, Gorry Fairhurst, Anthony Ang, Bernhard Collini-Nocker, Hilmar Linder, and Wolfram Stering. A lightweight encapsulation protocol for IP over MPEG-2 networks: design, implementation and analysis. Computer Networks (Amsterdam, Netherlands: 1999), 48(1):5–19, May 16, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ebert:2005:LOE

[889] Johannes Ebert and Michael Schmidt. A low overhead encapsulation method for MF-TDMA. Computer Networks (Amsterdam, Netherlands: 1999), 48 (1):21–27, May 16, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Arnal:2005:CLR

[890] F. Arnal, L. Dairaine, J. Lacan, and G. Maral. Cross-layer reliability management for multicast over satellite. Computer Networks (Amsterdam, Netherlands: 1999), 48(1):29–43, May 16, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xilouris:2005:MTD

[891] G. Xilouris, A. Kourtis, and G. Stefanou. A mesh topology DVB-S network architecture for node interconnection, featuring QoS capabilities. Computer Networks (Amsterdam, Netherlands: 1999), 48(1):45–56, May 16, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kouis:2005:EDS

[892] Dimitris Kouis, Dimitris Loukatos, Kimon Kontovasilis, George Kormentzas, and Charalabos Skianis. On the effectiveness of DVB-T for the support of IP-based services in heterogeneous wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 48 (1):57–73, May 16, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hlavacs:2005:CCD

[893] Helmut Hlavacs, Michael Haddad, Christophe Lafouge, Delphine Kaplan, and Jean Ribeiro. The CODIS Content Delivery Network. Computer Networks (Amsterdam, Netherlands: 1999), 48 (1):75–89, May 16, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBg

[894] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 48(1):CO2, May 16, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yuan:2005:DLP

[895] Shengli Yuan and Jason P. Jue. Dynamic lightpath protection in WDM mesh networks under wavelengthcontinuity and risk-disjoint constraints. Computer Networks (Amsterdam, Netherlands: 1999), 48(2):91– 112, June 6, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mahanti:2005:IMC

[896] Anirban Mahanti, Derek L. Eager, and Mary K. Vernon. Improving multirate congestion control using a TCP Vegas throughput model. Computer Networks (Amsterdam, Netherlands: 1999), 48 (2):113–136, June 6, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Quan:2005:SAC

[897] Zhi Quan and Jong-Moon Chung. Statistical admission control for real-time services under earliest deadline first scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 48(2): 137–154, June 6, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Deb:2005:NDQ

[898] Swati Sinha Deb and M. E. Woodward. A new distributed QoS routing algorithm based on Fano’s method. Computer Networks (Amsterdam, Netherlands: 1999), 48(2):155–174, June 6, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pei:2005:BRI

[899] Dan Pei, Matt Azuma, Dan Massey, and Lixia Zhang. BGP-RCN: improving BGP convergence through root cause notification. Computer Networks (Amsterdam, Netherlands: 1999), 48 (2):175–194, June 6, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sarac:2005:ALR

[900] Kamil Sarac and Kevin C. Almeroth. Application layer reachability monitoring for IP multicast. Computer Networks (Amsterdam, Netherlands: 1999), 48(2):195–213, June 6, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2005:PGA

[901] Junfeng Wang, Lei Li, Fuchun Sun, and Mingtian Zhou. A probabilityguaranteed adaptive timeout algorithm for high-speed network flow detection. Computer Networks (Amsterdam, Netherlands: 1999), 48(2):215–233, June 6, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gavrilis:2005:RTD

[902] Dimitris Gavrilis and Evangelos Dermatas. Real-time detection of distributed denial-of-service attacks using RBF networks and statistical features. Computer Networks (Amsterdam, Netherlands: 1999), 48(2):235– 245, June 6, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bayse:2005:PTA

[903] Emmanuel Bayse, Ana Cavalli, Manuel Núñez, and Fatiha Zaïdi. A passive testing approach based on invariants: application to the WAP. Computer Networks (Amsterdam, Netherlands: 1999), 48(2):247–266, June 6, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liang:2005:PAC

[904] Wei Liang and Wenye Wang. On performance analysis of challenge/ response based authentication in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 48 (2):267–288, June 6, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBh

[905] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 48(2):CO2, June 6, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Devetsikiotis:2005:MNT

[906] Michael Devetsikiotis and Nelson L. S. da Fonseca. Modeling network traffic with long range dependence: characterization, visualization and tools. Computer Networks (Amsterdam, Netherlands: 1999), 48(3):289–291, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Veitch:2005:MTI

[907] Darryl Veitch, Nicolas Hohn, and Patrice Abry. Multifractality in TCP/ IP traffic: the case against. Computer Networks (Amsterdam, Netherlands: 1999), 48(3):293–313, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ribeiro:2005:STS

[908] Vinay J. Ribeiro, Zhi-Li Zhang, Sue Moon, and Christophe Diot. Smalltime scaling behavior of Internet backbone traffic. Computer Networks (Amsterdam, Netherlands: 1999), 48(3): 315–334, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sarvotham:2005:NUD

[909] Shriram Sarvotham, Rudolf Riedi, and Richard Baraniuk. Network and user driven alpha-beta on–off source model for network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 48 (3):335–350, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Melo:2005:EPC

[910] Cesar A. V. Melo and Nelson L. S. da Fonseca. Envelope process and computation of the equivalent bandwidth of multifractal flows. Computer Networks (Amsterdam, Netherlands: 1999), 48 (3):351–375, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gong:2005:SSL

[911] Wei-Bo Gong, Yong Liu, Vishal Misra, and Don Towsley. Self-similarity and long range dependence on the Internet: a second look at the evidence, origins and implications. Computer Networks (Amsterdam, Netherlands: 1999), 48 (3):377–399, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Park:2005:LRD

[912] Cheolwoo Park, Félix Hernández-Campos, J. S. Marron, and F. Donelson Smith. Long-range dependence in a changing Internet traffic mix. Computer Networks (Amsterdam, Netherlands: 1999), 48(3):401–422, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Stoev:2005:WSD

[913] Stilian Stoev, Murad S. Taqqu, Cheolwoo Park, and J. S. Marron. On the wavelet spectrum diagnostic for Hurst parameter estimation in the analysis of Internet traffic. Computer Networks (Amsterdam, Netherlands: 1999), 48 (3):423–445, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rolls:2005:QAN

[914] D. A. Rolls, G. Michailidis, and F. Hernández-Campos. Queueing analysis of network traffic: methodology and visualization tools. Computer Networks (Amsterdam, Netherlands: 1999), 48(3):447–473, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2005:NEE

[915] Han S. Kim and Ness B. Shroff. The notion of end-to-end capacity and its application to the estimation of end-to-end network delays. Computer Networks (Amsterdam, Netherlands: 1999), 48(3):475–488, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBi

[916] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 48(3):CO2, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gurses:2005:SEM

[917] Eren Gürses, Gozde Bozdagi Akar, and Nail Akar. A simple and effective mechanism for stored video streaming with TCP transport and server-side adaptive frame discard. Computer Networks (Amsterdam, Netherlands: 1999), 48 (4):489–501, July 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bao:2005:PHU

[918] Feng Bao and Robert H. Deng. Protocols that hide user’s preferences in electronic transactions. Computer Networks (Amsterdam, Netherlands: 1999), 48(4):503–515, July 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2005:FSM

[919] Yu-Yi Chen, Jinn-Ke Jan, and Chin-Ling Chen. A fair and secure mobile billing system. Computer Networks (Amsterdam, Netherlands: 1999), 48 (4):517–524, July 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hnatyshin:2005:EAR

[920] Vasil Hnatyshin and Adarshpal S. Sethi. Estimating aggregate resource reservation for dynamic, scalable, and fair distribution of bandwidth. Computer Networks (Amsterdam, Netherlands: 1999), 48(4):525–547, July 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ye:2005:SOT

[921] Qing Ye, Yuecheng Zhang, and Liang Cheng. A study on the optimal time synchronization accuracy in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 48 (4):549–566, July 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vrontis:2005:EDS

[922] Stavros Vrontis and Efstathios Sykas. Extending differentiated services architecture for multicasting provisioning. Computer Networks (Amsterdam, Netherlands: 1999), 48(4):567– 584, July 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rao:2005:DDQ

[923] S. Dharma Rao and C. Siva Ram Murthy. Distributed dynamic QoSaware routing in WDM optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 48(4): 585–604, July 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2005:EBT

[924] Jun Wang, Yaling Yang, Li Xiao, and Klara Nahrstedt. Edge-based traffic engineering for OSPF networks. Computer Networks (Amsterdam, Netherlands: 1999), 48(4):605–625, July 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2005:PPB

[925] Min-You Wu, Yan Zhu, and Wei Shu. Placement of proxy-based multicast overlays. Computer Networks (Amsterdam, Netherlands: 1999), 48(4): 627–655, July 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhu:2005:ERK

[926] Bo Zhu, Feng Bao, Robert H. Deng, Mohan S. Kankanhalli, and Guilin Wang. Efficient and robust key management for large mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 48(4): 657–682, July 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chakraborty:2005:AAG

[927] S. S. Chakraborty, M. Liinaharja, and P. Lindroos. Analysis of adaptive GBN schemes in a Gilbert–Elliott channel and optimisation of system parameters. Computer Networks (Amsterdam, Netherlands: 1999), 48(4): 683–695, July 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBj

[928] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 48(4):CO2, July 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

McDaniel:2005:WS

[929] P. McDaniel and Aviel D. Rubin. Web security. Computer Networks (Amsterdam, Netherlands: 1999), 48(5): 697–699, August 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Brumley:2005:RTA

[930] David Brumley and Dan Boneh. Remote timing attacks are practical. Computer Networks (Amsterdam, Netherlands: 1999), 48(5):701–716, August 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kruegel:2005:MMA

[931] Christopher Kruegel, Giovanni Vigna, and William Robertson. A multi-model approach to the detection of Web-based attacks. Computer Networks (Amsterdam, Netherlands: 1999), 48(5): 717–738, August 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Huang:2005:TFW

[932] Yao-Wen Huang, Chung-Hung Tsai, Tsung-Po Lin, Shih-Kun Huang, D. T. Lee, and Sy-Yen Kuo. A testing framework for Web application security assessment. Computer Networks (Amsterdam, Netherlands: 1999), 48(5): 739–761, August 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lesniewski-Laas:2005:SSS

[933] Chris Lesniewski-Laas and M. Frans Kaashoek. SSL splitting: Securely serving data from untrusted caches. Computer Networks (Amsterdam, Netherlands: 1999), 48(5): 763–779, August 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Stavrou:2005:WOB

[934] Angelos Stavrou, Debra L. Cook, William G. Morein, Angelos D. Keromytis, Vishal Misra, and Dan Rubenstein. WebSOS: an overlaybased system for protecting web servers from denial of service attacks. Computer Networks (Amsterdam, Netherlands: 1999), 48(5):781–807, August 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Agrawal:2005:XPL

[935] Rakesh Agrawal, Jerry Kiernan, Ramakrishnan Srikant, and Yirong Xu. XPref: a preference language for P3P. Computer Networks (Amsterdam, Netherlands: 1999), 48(5): 809–827, August 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBk

[936] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 48(5):CO2, August 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Balaouras:2005:SAR

[937] P. Balaouras and I. Stavrakakis. A self-adjusting rate adaptation scheme with good fairness and smoothness properties. Computer Networks (Amsterdam, Netherlands: 1999), 48(6): 829–855, August 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bahri:2005:WLA

[938] Abderraouf Bahri and Steven Chamberland. On the wireless local area network design problem with performance guarantees. Computer Networks (Amsterdam, Netherlands: 1999), 48 (6):856–866, August 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

He:2005:FMA

[939] Jun He and Hung Keng Pung. Fairness of medium access control protocols for multi-hop ad hoc wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 48(6): 867–890, August 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Phung:2005:OSO

[940] M. H. Phùng, K. C. Chua, G. Mohan, M. Motani, T. C. Wong, and P. Y. Kong. On ordered scheduling for optical burst switching. Computer Networks (Amsterdam, Netherlands: 1999), 48(6):891–909, August 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nemeth:2005:NIT

[941] F. Németh, P. Barta, R. Szabó, and J. Bíró. Network internal traffic characterization and end-to-end delay bound calculus for generalized processor sharing scheduling discipline. Computer Networks (Amsterdam, Netherlands: 1999), 48(6):910–940, August 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Whitaker:2005:IEC

[942] Roger M. Whitaker, Larry Raisanen, and Steve Hurley. The infrastructure efficiency of cellular wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 48(6): 941–959, August 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altman:2005:AAP

[943] Eitan Altman, Chadi Barakat, and Víctor M. Ramos R. Analysis of AIMD protocols over paths with variable delay. Computer Networks (Amsterdam, Netherlands: 1999), 48(6): 960–971, August 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altman:2005:AMC

[944] E. Altman, K. Avrachenkov, C. Barakat, A. A. Kherani, and B. J. Prabhu. Analysis of MIMD congestion control algorithm for high speed networks. Computer Networks (Amsterdam, Netherlands: 1999), 48(6):972–989, August 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:AI

[945] Anonymous. Author index. Computer Networks (Amsterdam, Netherlands: 1999), 48(6):990–992, August 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:SIa

[946] Anonymous. Subject index. Computer Networks (Amsterdam, Netherlands: 1999), 48(6):993–995, August 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBl

[947] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 48(6):CO2, August 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wolf:2005:NIE

[948] Lars Wolf, Ryohei Nakatsu, and Jesse Schell. Networking issues in entertainment computing. Computer Networks (Amsterdam, Netherlands: 1999), 49 (1):1–3, September 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ghini:2005:HGD

[949] V. Ghini, G. Pau, M. Roccetti, P. Salomoni, and M. Gerla. For here or to go? Downloading music on the move with an ultra reliable wireless Internet application. Computer Networks (Amsterdam, Netherlands: 1999), 49 (1):4–26, September 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Smed:2005:RBT

[950] Jouni Smed, Henrik Niinisalo, and Harri Hakonen. Realizing the bullet time effect in multiplayer games with local perception filters. Computer Networks (Amsterdam, Netherlands: 1999), 49(1):27–37, September 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Singh:2005:MNG

[951] Aameek Singh and Arup Acharya. Multiplayer networked gaming with the session initiation protocol. Computer Networks (Amsterdam, Netherlands: 1999), 49(1):38–51, September 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Claypool:2005:ELU

[952] Mark Claypool. The effect of latency on user performance in Real-Time Strategy games. Computer Networks (Amsterdam, Netherlands: 1999), 49(1): 52–70, September 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sinha:2005:NGT

[953] Amit Sinha, Kenneth Mitchell, and Deep Medhi. Network game traffic: A broadband access perspective. Computer Networks (Amsterdam, Netherlands: 1999), 49(1):71–83, September 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2005:ASS

[954] Kang-Won Lee, Bong-Jun Ko, and Seraphin Calo. Adaptive server selection for large scale interactive online games. Computer Networks (Amsterdam, Netherlands: 1999), 49(1):84– 102, September 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nguyen:2005:POD

[955] Thanh Vinh Nguyen, Farzad Safaei, Paul Boustead, and Chun Tung Chou. Provisioning overlay distribution networks. Computer Networks (Amsterdam, Netherlands: 1999), 49(1):103– 118, September 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBm

[956] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 49(1):CO2, September 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

McNair:2005:SCL

[957] Janise McNair, Tuna Tugcu, Wenye Wang, and Jiang (Linda) Xie. A survey of cross-layer performance enhancements for Mobile IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 49(2):119–146, October 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ioannidis:2005:LCD

[958] Ioannis Ioannidis and Ananth Grama. Level compressed DAGs for lookup tables. Computer Networks (Amsterdam, Netherlands: 1999), 49(2): 147–160, October 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Alharbi:2005:DBA

[959] Fahd Alharbi and Nirwan Ansari. Distributed bandwidth allocation for resilient packet ring networks. Computer Networks (Amsterdam, Netherlands: 1999), 49(2):161–171, October 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yeh:2005:DSS

[960] Jyh haw Yeh, Wei Zhang, Wen chen Hu, and Chung wei Lee. Design and simulation of a supplemental protocol for BGP. Computer Networks (Amsterdam, Netherlands: 1999), 49(2): 172–200, October 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Long:2005:CSD

[961] Fei Long, Gang Feng, and Chee Kheong Siew. Channel states dependent fair service: a new packet scheduling algorithm for CDMA. Computer Networks (Amsterdam, Netherlands: 1999), 49 (2):201–216, October 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tak:2005:MPS

[962] Sungwoo Tak and Eun Kyo Park. Modeling and performance study of restoration framework in WDM optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 49(2): 217–242, October 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bean:2005:AMA

[963] Nigel Bean and Andre Costa. An analytic modelling approach for network routing algorithms that use “ant-like” mobile agents. Computer Networks (Amsterdam, Netherlands: 1999), 49 (2):243–268, October 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Savoric:2005:APE

[964] Michael Savorić, Holger Karl, Morten Schläger, Tobias Poschwatta, and Adam Wolisz. Analysis and performance evaluation of the EFCM common congestion controller for TCP connections. Computer Networks (Amsterdam, Netherlands: 1999), 49(2): 269–294, October 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBn

[965] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 49(2):??, October 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Walke:2005:GE

[966] Bernhard Walke, Luciano Lenzini, and Bijan Jabbari. Guest editorial. Computer Networks (Amsterdam, Netherlands: 1999), 49(3):295–298, October 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Herhold:2005:CMH

[967] Patrick Herhold, Ernesto Zimmermann, and Gerhard Fettweis. Cooperative multi-hop transmission in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 49(3): 299–324, October 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Forkel:2005:HSD

[968] Ingo Forkel, Hartmut Klenner, and Andreas Kemper. High Speed Downlink Packet Access (HSDPA)—enhanced data rates for UMTS evolution. Computer Networks (Amsterdam, Netherlands: 1999), 49(3):325–340, October 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hoymann:2005:APE

[969] Christian Hoymann. Analysis and performance evaluation of the OFDMbased metropolitan area network IEEE 802.16. Computer Networks (Amsterdam, Netherlands: 1999), 49(3): 341–363, October 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Herwono:2005:PPM

[970] Ian Herwono, Joachim Sachs, and Ralf Keller. Provisioning and performance of mobility-aware personalized push services in wireless broadband hotspots. Computer Networks (Amsterdam, Netherlands: 1999), 49 (3):364–384, October 19, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pesch:2005:PES

[971] Dirk Pesch, Maria Isabel Pous, and Gerry Foster. Performance evaluation of SIP-based multimedia services in UMTS. Computer Networks (Amsterdam, Netherlands: 1999), 49(3): 385–403, October 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gozalvez:2005:SPA

[972] Javier Gozalvez and John Dunlop. System performance and adaptive configuration of link adaptation techniques in packet-switched cellular radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 49(3):404– 426, October 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liebsch:2005:GIP

[973] Marco Liebsch and Bernd Lamparter. A generic IP paging architecture and protocol. Computer Networks (Amsterdam, Netherlands: 1999), 49(3):

427–448, October 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2005:PAD

[974] Tianwei Chen, Michel Sortais, Günter Schäfer, Stefan Adams, Changpeng Fan, and Adam Wolisz. Performance analysis of a Denial of Service protection scheme for optimized and QoSaware handover. Computer Networks (Amsterdam, Netherlands: 1999), 49 (3):449–464, October 19, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBo

[975] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 49(3):??, October 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chim:2005:TDE

[976] Tat Wing Chim, Kwan L. Yeung, and King-Shan Lui. Traffic distribution over equal-cost-multi-paths. Computer Networks (Amsterdam, Netherlands: 1999), 49(4):465–475, November 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lu:2005:NEI

[977] Chenghuai Lu and Andre L. M. Dos Santos. A note on efficient implementation of prime generation algorithms in small portable devices. Computer Networks (Amsterdam, Netherlands: 1999), 49(4):476–491, November 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lancellotti:2005:HCS

[978] Riccardo Lancellotti, Francesca Mazzoni, and Michele Colajanni. Hybrid cooperative schemes for scalable and stable performance of Web content delivery. Computer Networks (Amsterdam, Netherlands: 1999), 49(4):492– 511, November 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akkaya:2005:SRE

[979] Kemal Akkaya, Mohamed Younis, and Meenakshi Bangad. Sink repositioning for enhanced performance in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 49 (4):512–534, November 15, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lu:2005:ERU

[980] Rongxing Lu and Zhenfu Cao. Efficient remote user authentication scheme using smart card. Computer Networks (Amsterdam, Netherlands: 1999), 49 (4):535–540, November 15, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Casetti:2005:LMP

[981] Claudio Casetti and Marco Mellia. A lightweight marker with partial state information for DiffServ networks. Computer Networks (Amsterdam, Netherlands: 1999), 49(4):541– 560, November 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Madhyastha:2005:GMS

[982] Harsha V. Madhyastha, Girish V. Chowdhary, N. Srinivas, and C. Siva Ram Murthy. Grooming of multicast sessions in metropolitan WDM ring networks. Computer Networks (Amsterdam, Netherlands: 1999), 49(4):561– 579, November 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ren:2005:NCT

[983] Fengyuan Ren, Chuang Lin, and Bo Wei. A nonlinear control theoretic analysis to TCP–RED system. Computer Networks (Amsterdam, Netherlands: 1999), 49(4):580–592, November 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Norden:2005:IDR

[984] Samphel Norden. Inter-domain routing: Algorithms for QoS guarantees. Computer Networks (Amsterdam, Netherlands: 1999), 49(4):593– 619, November 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBp

[985] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 49(4):??, November 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:C

[986] Anonymous. Contents. Computer Networks (Amsterdam, Netherlands: 1999), 49(5):621, December 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hogrefe:2005:TUL

[987] Dieter Hogrefe and Rick Reed. Telecommunications and UML languages. Computer Networks (Amsterdam, Netherlands: 1999), 49(5): 622–626, December 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

deWet:2005:UUM

[988] Nico de Wet and Pieter Kritzinger. Using UML models for the performance analysis of network systems. Computer Networks (Amsterdam, Netherlands: 1999), 49(5):627–642, December 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Amyot:2005:GTP

[989] Daniel Amyot, Luigi Logrippo, and Michael Weiss. Generation of test purposes from Use Case Maps. Computer Networks (Amsterdam, Netherlands: 1999), 49(5):643–660, December 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Letichevsky:2005:BPM

[990] A. Letichevsky, J. Kapitonova, A. Letichevsky, Jr., V. Volkov, S. Baranov, and T. Weigert. Basic protocols, message sequence charts, and the verification of requirements specifications. Computer Networks (Amsterdam, Netherlands: 1999), 49(5):661–675, December 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fischer:2005:TBL

[991] Joachim Fischer, Eckhardt Holz, Andreas Prinz, and Markus Scheidgen. Tool-based language development. Computer Networks (Amsterdam, Netherlands: 1999), 49(5):676– 688, December 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fliege:2005:DSC

[992] Ingmar Fliege, Alexander Geraldy, Reinhard Gotzhein, Thomas Kuhn, and Christian Webel. Developing safety-critical real-time systems with SDL design patterns and components. Computer Networks (Amsterdam, Netherlands: 1999), 49(5):689– 706, December 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:EBq

[993] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 49(5):??, December 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Motyckova:2005:CTP

[994] Lenka Motycková and David A. Carr. A cluster-tree protocol for reliable multicasting. Computer Networks (Amsterdam, Netherlands: 1999), 49(6): 707–726, December 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Alyatama:2005:WDA

[995] Anwar Alyatama. Wavelength decomposition approach for computing blocking probabilities in WDM optical networks without wavelength conversions. Computer Networks (Amsterdam, Netherlands: 1999), 49(6):727– 742, December 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Baggio:2005:DRW

[996] Aline Baggio and Maarten van Steen. Distributed redirection for the World-Wide Web. Computer Networks (Amsterdam, Netherlands: 1999), 49(6): 743–765, December 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nafaa:2005:JLP

[997] Abdelhamid Nafaa and Ahmed Mehaoua. Joint loss pattern characterization and unequal interleaved FEC protection for robust H.264 video distribution over wireless LAN. Computer Networks (Amsterdam, Netherlands: 1999), 49 (6):766–786, December 19, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Salamah:2005:DAC

[998] Muhammed Salamah and Hashem Lababidi. Dynamically adaptive channel reservation scheme for cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 49(6):787– 796, December 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pietrabissa:2005:VQA

[999] Antonio Pietrabissa, Tiziano Inzerilli, Olivier Alphand, Pascal Berthou, Thierry Gayraud, Michel Mazzella, Eddy Fromentin, and Fabrice Lucas. Validation of a QoS architecture for DVB-RCS satellite networks via the SATIP6 demonstration platform. Computer Networks (Amsterdam, Netherlands: 1999), 49(6):797–815, December 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chang:2005:IBA

[1000] XiaoLin Chang and Jogesh K. Muppala. On improving bandwidth assurance in AF-based DiffServ networks using a control theoretic approach. Computer Networks (Amsterdam, Netherlands: 1999), 49(6):816–839, December 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ercetin:2005:PSD

[1001] Ozgur Ercetin and Leandros Tassiulas. Pricing strategies for differentiated services content delivery networks. Computer Networks (Amsterdam, Netherlands: 1999), 49(6):840–855, December 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Salles:2005:FED

[1002] Ronaldo M. Salles and Javier A. Barria. Fair and efficient dynamic bandwidth allocation for multi-application networks. Computer Networks (Amsterdam, Netherlands: 1999), 49(6): 856–877, December 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bag-Mohammadi:2005:NBM

[1003] Mozafar Bag-Mohammadi and Nasser Yazdani. Next Branch Multicast (NBM) routing protocol. Computer Networks (Amsterdam, Netherlands: 1999), 49(6):878–897, December 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2005:SIb

[1004] Anonymous. Subject index. Computer Networks (Amsterdam, Netherlands: 1999), 49(6):898–900, December 19, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nomikos:2006:RWA

[1005] Christos Nomikos, Aris Pagourtzis, Katerina Potika, and Stathis Zachos. Routing and wavelength assignment in multifiber WDM networks with nonuniform fiber cost. Computer Networks (Amsterdam, Netherlands: 1999), 50 (1):1–14, January 16, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jiang:2006:DBP

[1006] Yuming Jiang. Delay bound and packet scale rate guarantee for some expedited forwarding networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(1):15–28, January 16, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lambrecht:2006:OMT

[1007] Thijs Lambrecht, Bart Duysburgh, Tim Wauters, Filip De Turck, Bart Dhoedt, and Piet Demeester. Optimizing multimedia transcoding multicast trees. Computer Networks (Amsterdam, Netherlands: 1999), 50(1): 29–45, January 16, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lan:2006:MSC

[1008] Kun chan Lan and John Heidemann.

A measurement study of correlations of Internet flow characteristics. Computer Networks (Amsterdam, Netherlands: 1999), 50(1):46–62, January 16, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Miorandi:2006:QMH

[1009] Daniele Miorandi, Arzad A. Kherani, and Eitan Altman. A queueing model for HTTP traffic over IEEE 802.11 WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 50(1): 63–79, January 16, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cui:2006:ASQ

[1010] Jun-Hong Cui, Li Lao, Michalis Faloutsos, and Mario Gerla. AQoSM: Scalable QoS multicast provisioning in Diff-Serv networks. Computer Networks (Amsterdam, Netherlands: 1999), 50 (1):80–105, January 16, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cheng:2006:PFI

[1011] An-Hsun Cheng and Yuh-Jzer Joung. Probabilistic file indexing and searching in unstructured peer-to-peer networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(1):106– 127, January 16, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bartolini:2006:PAC

[1012] N. Bartolini and E. Casalicchio. A performance analysis of context transfer protocols for QoS enabled Internet services. Computer Networks (Amsterdam, Netherlands: 1999), 50(1): 128–144, January 16, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBa

[1013] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(1):??, January 16, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Perros:2006:ON

[1014] Harry Perros. Optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(2):145–148, February 8, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2006:UVN

[1015] Li-Wei Chen, Poompat Saengudomlert, and Eytan Modiano. Uniform versus non-uniform band switching in WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(2): 149–167, February 8, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Haque:2006:GSP

[1016] Anwar Haque, Pin-Han Ho, and Raouf Boutaba. Group shared protection for spare capacity reconfiguration in optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 50 (2):168–180, February 8, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Frederick:2006:SGR

[1017] Michael T. Frederick, Pallab Datta, and Arun K. Somani. Sub-graph routing: A generalized fault-tolerant strategy for link failures in WDM optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(2): 181–199, February 8, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Singhal:2006:CSV

[1018] Narendra K. Singhal, Canhui Ou, and Biswanath Mukherjee. Cross-sharing vs. self-sharing trees for protecting multicast sessions in mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(2):200–206, February 8, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pattavina:2006:PDR

[1019] Achille Pattavina. Performance of deflection routing algorithms in IP optical transport networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(2):207–218, February 8, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Battestilli:2006:PSO

[1020] Tzvetelina Battestilli and Harry Perros. A performance study of an optical burst switched network with dynamic simultaneous link possession. Computer Networks (Amsterdam, Netherlands: 1999), 50(2):219–236, February 8, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bouabdallah:2006:EBS

[1021] Nizar Bouabdallah, Andre-Luc Beylot, and Guy Pujolle. Efficient bandwidth sharing and fair access in single-hub WDM rings. Computer Networks (Amsterdam, Netherlands: 1999), 50(2): 237–250, February 8, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Marin-Tordera:2006:HRA

[1022] Eva Marín-Tordera, Xavier Masip-Bruin, Sergio Sánchez-Lopez, Josep Solé-Pareta, and Jordi Domingo-Pascual. A hierarchical routing approach for optical transport networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(2): 251–267, February 8, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mason:2006:TDD

[1023] Lorne Mason, Anton Vinokurov, Ning Zhao, and David Plant. Topological design and dimensioning of Agile All-Photonic Networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(2):268–287, February 8, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBb

[1024] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(2):??, February 8, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2006:DCR

[1025] Sun-Myeng Kim and Young-Jong Cho. A distributed collision resolution scheme for improving the performance in wireless LANs. Computer Networks (Amsterdam, Netherlands: 1999), 50 (3):289–300, February 22, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zheng:2006:MLP

[1026] Qin Zheng and Gurusamy Mohan.

Multi-layer protection in IP-over-WDM networks with and with no backup lightpath sharing. Computer Networks (Amsterdam, Netherlands: 1999), 50(3):301–316, February 22, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bhatia:2006:FNR

[1027] Randeep Bhatia, Murali Kodialam, and T. V. Lakshman. Fast network re-optimization schemes for MPLS and optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 50 (3):317–331, February 22, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Angrisani:2006:TAB

[1028] Leopoldo Angrisani, Salvatore D’Antonio, Marcello Esposito, and Michele Vadursi. Techniques for available bandwidth measurement in IP networks: A performance comparison. Computer Networks (Amsterdam, Netherlands: 1999), 50(3):332–349, February 22, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhu:2006:HSS

[1029] Yongqing Zhu, Maode Ma, and Tee Hiang Cheng. Hierarchical scheduling to support differentiated services in Ethernet passive optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(3):350–366, February 22, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

DiSorte:2006:PEP

[1030] Dario Di Sorte, Mauro Femminella, Leonardo Piacentini, and Gianluca Reali. Performance evaluation of the push-mode-multicast based candidate access router discovery (PMM CARD). Computer Networks (Amsterdam, Netherlands: 1999), 50(3): 367–397, February 22, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pei:2006:ACD

[1031] Dan Pei, Beichuan Zhang, Daniel Massey, and Lixia Zhang. An analysis of convergence delay in path vector routing protocols. Computer Networks (Amsterdam, Netherlands: 1999), 50 (3):398–421, February 22, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chandrayana:2006:RST

[1032] Kartikeya Chandrayana, Sthanunathan Ramakrishnan, Biplab Sikdar, and Shivkumar Kalyanaraman. On randomizing the sending times in TCP and other window based algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 50(3):422–447, February 22, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bonmariage:2006:SON

[1033] Nicolas Bonmariage and Guy Leduc. A survey of optimal network congestion control for unicast and multicast transmission. Computer Networks (Amsterdam, Netherlands: 1999), 50(3): 448–468, February 22, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBc

[1034] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(3):??, February 22, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Boutaba:2006:MPP

[1035] Raouf Boutaba and Alan Marshall. Management in peer-to-peer systems: Trust, reputation and security. Computer Networks (Amsterdam, Netherlands: 1999), 50(4):469–471, March 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Marti:2006:TTC

[1036] Sergio Marti and Hector Garcia-Molina. Taxonomy of trust: Categorizing P2P reputation systems. Computer Networks (Amsterdam, Netherlands: 1999), 50(4):472–484, March 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Despotovic:2006:PRM

[1037] Zoran Despotovic and Karl Aberer. P2P reputation management: Probabilistic estimation vs. social networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(4): 485–500, March 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gupta:2006:TOB

[1038] Minaxi Gupta, Mostafa H. Ammar, and Mustaque Ahamad. Trade-offs between reliability and overheads in peer-to-peer reputation tracking. Computer Networks (Amsterdam, Netherlands: 1999), 50(4):501–522, March 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sherwood:2006:CPG

[1039] Rob Sherwood, Seungjoon Lee, and Bobby Bhattacharjee. Cooperative peer groups in NICE. Computer Networks (Amsterdam, Netherlands: 1999), 50(4):523–544, March 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mekouar:2006:PPM

[1040] Loubna Mekouar, Youssef Iraqi, and Raouf Boutaba. Peer-to-peer’s most wanted: Malicious peers. Computer Networks (Amsterdam, Netherlands: 1999), 50(4):545–562, March 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Papaioannou:2006:RBP

[1041] Thanasis G. Papaioannou and George D.

Stamoulis. Reputation-based policies that provide the right incentives in peer-to-peer environments. Computer Networks (Amsterdam, Netherlands: 1999), 50(4):563–578, March 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pathak:2006:BFT

[1042] Vivek Pathak and Liviu Iftode. Byzantine fault tolerant public key authentication in peer-to-peer systems. Computer Networks (Amsterdam, Netherlands: 1999), 50(4):579–596, March 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBd

[1043] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(4):??, March 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2006:E

[1044] Ian F. Akyildiz and Harry Rudin.

Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 50(5): 597–598, April 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tamasi:2006:COD

[1045] Levente Tamási, Dániel Orincsay, and Balázs Gábor Józsa. Cost-optimal design of VoIP networks using the VPN concept. Computer Networks (Amsterdam, Netherlands: 1999), 50(5): 599–614, April 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cohen:2006:SWB

[1046] Edith Cohen and Balachander Krishnamurthy. A short walk in the Blogistan. Computer Networks (Amsterdam, Netherlands: 1999), 50(5): 615–630, April 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

vandenBerg:2006:QAB

[1047] Hans van den Berg, Michel Mandjes, Remco van de Meent, Aiko Pras, Frank Roijers, and Pieter Venemans. QoSaware bandwidth provisioning for IP network links. Computer Networks (Amsterdam, Netherlands: 1999), 50 (5):631–647, April 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jin:2006:SWC

[1048] Shudong Jin and Azer Bestavros. Small-world characteristics of Internet topologies and implications on multicast scaling. Computer Networks (Amsterdam, Netherlands: 1999), 50 (5):648–666, April 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhu:2006:IUM

[1049] Weiying Zhu and Min Song. Integration of unicast and multicast scheduling in input-queued packet switches. Computer Networks (Amsterdam, Netherlands: 1999), 50(5):667–687, April 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Risso:2006:NEX

[1050] Fulvio Risso and Mario Baldi. NetPDL:

An extensible XML-based language for packet header description. Computer Networks (Amsterdam, Netherlands: 1999), 50(5):688–706, April 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2006:PAS

[1051] Bo Li, Roberto Battiti, and Akira Fukuda. Performance analysis of a service-dependent handoff scheme in voice/data integrated cellular mobile systems. Computer Networks (Amsterdam, Netherlands: 1999), 50(5): 707–730, April 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBe

[1052] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(5):??, April 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mathy:2006:ODS

[1053] L. Mathy, D. Hutchison, T. Plagemann, and P. Steenkiste. Overlay distribution structures and their applications. Computer Networks (Amsterdam, Netherlands: 1999), 50(6): 739–741, April 13, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhu:2006:DOR

[1054] Yong Zhu, Constantinos Dovrolis, and Mostafa Ammar. Dynamic overlay routing based on available bandwidth estimation: A simulation study. Computer Networks (Amsterdam, Netherlands: 1999), 50(6):742–762, April 13, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2006:UWL

[1055] Ju Wang and Andrew A. Chien. Understanding when location-hiding using overlay networks is feasible. Computer Networks (Amsterdam, Netherlands: 1999), 50(6):763–780, April 13, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2006:UIM

[1056] Beichuan Zhang, Wenjie Wang, Sugih Jamin, Daniel Massey, and Lixia Zhang. Universal IP multicast delivery. Computer Networks (Amsterdam, Netherlands: 1999), 50(6): 781–806, April 13, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lakshminarayanan:2006:EHC

[1057] Karthik Lakshminarayanan, Ananth Rao, Ion Stoica, and Scott Shenker. End-host controlled multicast routing. Computer Networks (Amsterdam, Netherlands: 1999), 50(6):807– 825, April 13, 2006. CODEN ????

ISSN 1389-1286 (print), 1872-7069 (electronic).

Banerjee:2006:OEO

[1058] Suman Banerjee, Christopher Kommareddy, Koushik Kar, Bobby Bhattacharjee, and Samir Khuller. OMNI: An efficient overlay multicast infrastructure for real-time applications. Computer Networks (Amsterdam, Netherlands: 1999), 50(6):826– 841, April 13, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liang:2006:FOM

[1059] Jian Liang, Rakesh Kumar, and Keith W. Ross. The FastTrack overlay: A measurement study. Computer Networks (Amsterdam, Netherlands: 1999), 50(6):842–858, April 13, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2006:CMD

[1060] Xiaotao Liu, Jiang Lan, Prashant Shenoy, and Krithi Ramaritham. Consistency maintenance in dynamic peerto-peer overlay networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(6):859–876, April 13, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBf

[1061] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(6):??, April 13, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gungor:2006:SCN

[1062] V. C. Gungor and F. C. Lambert. A survey on communication networks for electric system automation. Computer Networks (Amsterdam, Netherlands: 1999), 50(7):877–897, May 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fazio:2006:IAA

[1063] M. Fazio, M. Villari, and A. Puliafito. IP address autoconfiguration in ad hoc networks: Design, implementation and measurements. Computer Networks (Amsterdam, Netherlands: 1999), 50 (7):898–920, May 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Luo:2006:ORU

[1064] Hongli Luo, Mei-Ling Shyu, and Shu-Ching Chen. An optimal resource utilization scheme with end-to-end congestion control for continuous media stream transmission. Computer Networks (Amsterdam, Netherlands: 1999), 50(7):921–937, May 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cayirci:2006:QSN

[1065] Erdal Cayirci, Vedat Coskun, and Caghan Cimen. Querying sensor networks by using dynamic task sets. Computer Networks (Amsterdam, Netherlands: 1999), 50(7):938– 952, May 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jan:2006:OMW

[1066] Rong-Hong Jan, Ching-Peng Lin, and Maw-Sheng Chern. An optimization model for Web content adaptation. Computer Networks (Amsterdam, Netherlands: 1999), 50(7): 953–965, May 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Capone:2006:DOQ

[1067] Antonio Capone, Luigi Fratta, and Fabio Martignon. Dynamic online QoS routing schemes: Performance and bounds. Computer Networks (Amsterdam, Netherlands: 1999), 50(7): 966–981, May 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altman:2006:PDS

[1068] Eitan Altman, Dhiman Barman, Rachid El Azouzi, David Ros, and Bruno Tuffin. Pricing differentiated services: A game-theoretic approach. Computer Networks (Amsterdam, Netherlands: 1999), 50(7): 982–1002, May 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2006:CAP

[1069] Xudong Wang. A capacity acquisition protocol for channel reservation in CDMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 50 (7):1003–1021, May 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBg

[1070] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(7):??, May 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bianchi:2006:SPI

[1071] Giuseppe Bianchi, Marco Listanti, Michela Meo, and Maurizio M. Munafò. Selected papers from the 3rd international workshop on QoS in multiservice IP networks (QoS-IP 2005). Computer Networks (Amsterdam, Netherlands: 1999), 50(8):1023–1025, June 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fidler:2006:CNC

[1072] Markus Fidler and Stephan Recker. Conjugate network calculus: A dual approach applying the Legendre transform. Computer Networks (Amsterdam, Netherlands: 1999), 50(8): 1026–1039, June 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sayenko:2006:CAR

[1073] Alexander Sayenko, Timo Hämäläinen, Jyrki Joutsensalo, and Lari Kannisto. Comparison and analysis of the revenue-based adaptive queuing models. Computer Networks (Amsterdam, Netherlands: 1999), 50(8):1040–1058, June 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Leino:2006:ILB

[1074] J. Leino and J. Virtamo. Insensitive load balancing in data networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(8):1059–1068, June 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bolla:2006:CPI

[1075] Raffaele Bolla, Roberto Bruschi, and Franco Davoli. Capacity planning in IP Virtual Private Networks under mixed traffic. Computer Networks (Amsterdam, Netherlands: 1999), 50 (8):1069–1085, June 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wille:2006:AIN

[1076] Emilio C. G. Wille, Marco Mellia, Emilio Leonardi, and Marco Ajmone Marsan. Algorithms for IP network design with end-to-end QoS constraints. Computer Networks (Amsterdam, Netherlands: 1999), 50(8): 1086–1103, June 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dan:2006:EPS

[1077] György Dán, Viktória Fodor, and Gunnar Karlsson. On the effects of the packet size distribution on FEC performance. Computer Networks (Amsterdam, Netherlands: 1999), 50(8): 1104–1129, June 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gamage:2006:CON

[1078] Manodha Gamage, Mitsuo Hayasaka, and Tetsuya Miki. A connectionoriented network architecture with guaranteed QoS for future real-time applications over the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 50(8):1130–1144, June 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

deLaunois:2006:LNP

[1079] Cédric de Launois, Bruno Quoitin, and Olivier Bonaventure. Leveraging network performance with IPv6 multihoming and multiple providerdependent aggregatable prefixes. Computer Networks (Amsterdam, Netherlands: 1999), 50(8):1145–1157, June 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Peters:2006:QMM

[1080] Liesbeth Peters, Ingrid Moerman, Bart Dhoedt, and Piet Demeester. Q-MEHROM: Mobility support and resource reservations for mobile senders and receivers. Computer Networks (Amsterdam, Netherlands: 1999), 50 (8):1158–1175, June 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Neglia:2006:AMN

[1081] Giovanni Neglia, Vincenzo Falletta, and Giuseppe Bianchi. An analytical model of a new packet marking algorithm for TCP flows. Computer Networks (Amsterdam, Netherlands: 1999), 50(8):1176–1191, June 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBh

[1082] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(8):??, June 6, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ozkasap:2006:TCT

[1083] Öznur Özkasap and Mine Çağlar. Traffic characterization of transport level reliable multicasting: Comparison of epidemic and feedback controlled loss recovery. Computer Networks (Amsterdam, Netherlands: 1999), 50(9): 1193–1218, June 20, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shuai:2006:PON

[1084] Dianxun Shuai and Xiang Feng. The parallel optimization of network bandwidth allocation based on generalized particle model. Computer Networks (Amsterdam, Netherlands: 1999), 50 (9):1219–1246, June 20, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ssu:2006:DDD

[1085] Kuo-Feng Ssu, Chih-Hsun Chou, Hewijin Christine Jiau, and Wei-Te Hu. Detection and diagnosis of data inconsistency failures in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(9): 1247–1260, June 20, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

DeAngelis:2006:IRA

[1086] Flavio De Angelis, Ibrahim Habib, Fabrizio Davide, and Mahmoud Naghshineh. Increasing revenues via adaptive filtering in wired/wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(9): 1261–1280, June 20, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bu:2006:SRO

[1087] Tian Bu, Samphel Norden, and Thomas Woo. A survivable DoSresistant overlay network. Computer Networks (Amsterdam, Netherlands: 1999), 50(9):1281–1301, June 20, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2006:FFA

[1088] Jiwei Chen, Fernando Paganini, M. Y. Sanadidi, Ren Wang, and Mario Gerla. Fluid-flow analysis of TCP Westwood with RED. Computer Networks (Amsterdam, Netherlands: 1999), 50(9): 1302–1326, June 20, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lafrance:2006:UEC

[1089] Stéphane Lafrance. Using equivalencechecking to verify robustness to denial of service. Computer Networks (Amsterdam, Netherlands: 1999), 50(9): 1327–1348, June 20, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Reddy:2006:MMP

[1090] T. Bheemarjuna Reddy, S. Sriram, B. S. Manoj, and C. Siva Ram Murthy. MuSeQoR: Multi-path failure-tolerant security-aware QoS routing in Ad hoc wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 50 (9):1349–1383, June 20, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2006:FIM

[1091] Hailong Li, Malvin Wei Liak Tan, and Ian Li-Jin Thng. Fairness issue and monitor-based algorithm in optical burst switching networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(9):1384–1405, June 20, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2006:PCU

[1092] Fu-Yuan Lee and Shiuhpyng Shieh. Packet classification using diagonalbased tuple space search. Computer Networks (Amsterdam, Netherlands: 1999), 50(9):1406–1423, June 20, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBi

[1093] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(9):??, June 20, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Levene:2006:SIW

[1094] Mark Levene and Alexandra Poulovassilis. Special issue on Web dynamics. Computer Networks (Amsterdam, Netherlands: 1999), 50(10): 1425–1429, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ke:2006:WDT

[1095] Yiping Ke, Lin Deng, Wilfred Ng, and Dik-Lun Lee. Web dynamics and their ramifications for the development of Web search engines. Computer Networks (Amsterdam, Netherlands: 1999), 50(10):1430–1447, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bar-Ilan:2006:MCR

[1096] Judit Bar-Ilan, Mazlita Mat-Hassan, and Mark Levene. Methods for comparing rankings of search engine results. Computer Networks (Amsterdam, Netherlands: 1999), 50(10): 1448–1463, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Baeza-Yates:2006:DCW

[1097] Ricardo Baeza-Yates and Barbara Poblete. Dynamics of the Chilean Web structure. Computer Networks (Amsterdam, Netherlands: 1999), 50(10): 1464–1473, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2006:DPS

[1098] Hua-Fu Li, Suh-Yin Lee, and Man-Kwan Shan. DSM-PLW: Single-pass mining of path traversal patterns over streaming Web click-sequences. Computer Networks (Amsterdam, Netherlands: 1999), 50(10):1474–1487, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nasraoui:2006:FME

[1099] Olfa Nasraoui, Carlos Rojas, and Cesar Cardona. A framework for mining evolving trends in Web data streams using dynamic learning and retrospective validation. Computer Networks (Amsterdam, Netherlands: 1999), 50 (10):1488–1512, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Papamarkos:2006:ECA

[1100] George Papamarkos, Alexandra Poulovassilis, and Peter T. Wood. Eventcondition-action rules on RDF metadata in P2P environments. Computer Networks (Amsterdam, Netherlands: 1999), 50(10):1513–1532, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ceri:2006:MWA

[1101] Stefano Ceri, Florian Daniel, and Federico M. Facca. Modeling web applications reacting to user behaviors. Computer Networks (Amsterdam, Netherlands: 1999), 50(10):1533–1546, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sanka:2006:DAE

[1102] Anoop Sanka, Shravan Chamakura, and Sharma Chakravarthy. A dataflow approach to efficient change detection of HTML/XML documents in Web-VigiL. Computer Networks (Amsterdam, Netherlands: 1999), 50(10): 1547–1563, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Malandrino:2006:TWD

[1103] Delfina Malandrino and Vittorio Scarano. Tackling Web dynamics by programmable proxies. Computer Networks (Amsterdam, Netherlands: 1999), 50(10):1564–1580, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

DePrisco:2006:APD

[1104] Roberto De Prisco and Sergio Rajsbaum. Algorithmic problems in distributed systems. Computer Networks (Amsterdam, Netherlands: 1999), 50 (10):1581–1582, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sun:2006:MRW

[1105] Qixiang Sun, Neil Daswani, and Hector Garcia-Molina. Maximizing remote work in flooding-based peer-to-peer systems. Computer Networks (Amsterdam, Netherlands: 1999), 50 (10):1583–1598, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Garcia-Luna-Aceves:2006:NAD

[1106] J. J. Garcia-Luna-Aceves, Marc Mosko, and Charles E. Perkins. A new approach to on-demand loop-free routing in networks using sequence numbers. Computer Networks (Amsterdam, Netherlands: 1999), 50(10):1599–1615, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Abraham:2006:ARD

[1107] Ittai Abraham and Danny Dolev. Asynchronous resource discovery. Computer Networks (Amsterdam, Netherlands: 1999), 50(10):1616–1629, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fraigniaud:2006:HSL

[1108] Pierre Fraigniaud and Cyril Gavoille. Header-size lower bounds for end-to-end communication in memoryless networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(10): 1630–1638, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Castelluccia:2006:SAA

[1109] Claude Castelluccia, Stanisław Jarecki, Jihye Kim, and Gene Tsudik. Secure acknowledgment aggregation and multisignatures with limited robustness. Computer Networks (Amsterdam, Netherlands: 1999), 50(10):1639–1652, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBj

[1110] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(10):??, July 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhu:2006:ACB

[1111] Hua Zhu and Imrich Chlamtac. Admission control and bandwidth reservation in multi-hop ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(11):1653–1674, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sharma:2006:IAC

[1112] Puneet Sharma, Jack Brassil, Sung-Ju Lee, and Kang G. Shin. Improving aggregated channel performance through decentralized channel monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 50(11): 1675–1691, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tansupasiri:2006:UAN

[1113] T. Tansupasiri, K. Kanchanasut, C. Barakat, and P. Jacquet. Using active networks technology for dynamic QoS. Computer Networks (Amsterdam, Netherlands: 1999), 50(11): 1692–1709, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Figueiredo:2006:MIR

[1114] Gustavo B. Figueiredo, Nelson L. S.

da Fonseca, and José A. S. Monteiro.

A minimum interference routing algorithm with reduced computational complexity. Computer Networks (Amsterdam, Netherlands: 1999), 50(11): 1710–1732, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shang:2006:PRD

[1115] Hao Shang and Craig E. Wills. Piggybacking related domain names to improve DNS performance. Computer Networks (Amsterdam, Netherlands: 1999), 50(11):1733–1748, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Banchs:2006:TAO

[1116] Albert Banchs and Luca Vollero. Throughput analysis and optimal configuration of 802.11e EDCA. Computer Networks (Amsterdam, Netherlands: 1999), 50(11):1749–1768, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Krashakov:2006:URD

[1117] Serge A. Krashakov, Anton B. Teslyuk, and Lev N. Shchur. On the universality of rank distributions of website popularity. Computer Networks (Amsterdam, Netherlands: 1999), 50(11): 1769–1780, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jung:2006:KBP

[1118] Eunjin Jung, Alex X. Liu, and Mohamed G. Gouda. Key bundles and parcels: Secure communication in many groups. Computer Networks (Amsterdam, Netherlands: 1999), 50 (11):1781–1798, August 10, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altman:2006:LSC

[1119] Eitan Altman, Rachid El Azouzi, David Ros, and Bruno Tuffin. Loss strategies for competing AIMD flows. Computer Networks (Amsterdam, Netherlands: 1999), 50(11):1799–1815, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2006:DUM

[1120] Hyang-Won Lee and Song Chong. A distributed utility max–min flow control algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 50 (11):1816–1830, August 10, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Baboescu:2006:FPC

[1121] Florin Baboescu, Priyank Warkhede, Subhash Suri, and George Varghese. Fast packet classification for twodimensional conflict-free filters. Computer Networks (Amsterdam, Netherlands: 1999), 50(11):1831–1842, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kambourakis:2006:SSC

[1122] Georgios Kambourakis, Angelos Rouskas, Stefanos Gritzalis, and Dimitrios Geneiatakis. Support of subscribers’ certificates in a hybrid WLAN-3G environment. Computer Networks (Amsterdam, Netherlands: 1999), 50(11): 1843–1859, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Argyriou:2006:SHT

[1123] Antonios Argyriou and Vijay Madisetti. A soft-handoff transport protocol for media flows in heterogeneous mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(11): 1860–1871, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Khlifi:2006:LCO

[1124] Hechmi Khlifi and Jean-Charles Grégoire. Low-complexity offline and online clock skew estimation and removal. Computer Networks (Amsterdam, Netherlands: 1999), 50(11): 1872–1884, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBk

[1125] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(11):??, August 10, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hou:2006:NMS

[1126] Jennifer C. Hou and P. R. Kumar. Network modelling and simulation. Computer Networks (Amsterdam, Netherlands: 1999), 50(12):1885–1886, August 24, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Han:2006:DPD

[1127] Yijie Han, Richard J. La, Armand M. Makowski, and Seungjoon Lee. Distribution of path durations in mobile ad-hoc networks—Palm’s Theorem to the rescue. Computer Networks (Amsterdam, Netherlands: 1999), 50(12): 1887–1900, August 24, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Park:2006:SAP

[1128] Kyung-Joon Park, Hyuk Lim, and Chong-Ho Choi. Stochastic analysis of packet-pair probing for network bandwidth estimation. Computer Networks (Amsterdam, Netherlands: 1999), 50 (12):1901–1915, August 24, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hei:2006:MBE

[1129] Xiaojun Hei, Brahim Bensaou, and Danny H. K. Tsang. Model-based endto-end available bandwidth inference using queueing analysis. Computer Networks (Amsterdam, Netherlands: 1999), 50(12):1916–1937, August 24, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guven:2006:MBO

[1130] Tuna Güven, Richard J. La, Mark A.

Shayman, and Bobby Bhattacharjee. Measurement-based optimal routing on overlay architectures for unicast sessions. Computer Networks (Amsterdam, Netherlands: 1999), 50(12):1938– 1951, August 24, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Khayari:2006:DEW

[1131] Rachid El Abdouni Khayari. Design and evaluation of web proxies by leveraging self-similarity of web traffic. Computer Networks (Amsterdam, Netherlands: 1999), 50(12): 1952–1973, August 24, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Figueiredo:2006:EFS

[1132] Daniel R. Figueiredo, Benyuan Liu, Yang Guo, Jim Kurose, and Don Towsley. On the efficiency of fluid simulation of networks. Computer Networks (Amsterdam, Netherlands: 1999), 50 (12):1974–1994, August 24, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2006:INC

[1133] Hwangnam Kim. Integrating networkcalculus-based simulation with packetlevel simulation for TCP-operated networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(12): 1995–2012, August 24, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dimitropoulos:2006:ELS

[1134] Xenofontas A. Dimitropoulos and George F. Riley. Efficient large-scale BGP simulations. Computer Networks (Amsterdam, Netherlands: 1999), 50 (12):2013–2027, August 24, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2006:GSD

[1135] Yu Liu, Boleslaw K. Szymanski, and Adnan Saifee. Genesis: A scalable distributed system for large-scale parallel network simulation. Computer Networks (Amsterdam, Netherlands: 1999), 50(12):2028–2053, August 24, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lao:2006:FRS

[1136] Li Lao, Jun-Hong Cui, and Mario Gerla. A framework for realistic and systematic multicast performance evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 50(12): 2054–2070, August 24, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Golaup:2006:MTM

[1137] Assen Golaup and Hamid Aghvami. A multimedia traffic modeling framework for simulation-based performance evaluation studies. Computer Networks (Amsterdam, Netherlands: 1999), 50 (12):2071–2087, August 24, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Marchese:2006:LBC

[1138] Mario Marchese and Maurizio Mongelli. On-line bandwidth control for quality of service mapping over satellite independent service access points. Computer Networks (Amsterdam, Netherlands: 1999), 50(12): 2088–2111, August 24, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2006:CDQ

[1139] Frank Y. Li, Norvald Stol, and Bjarne E. Helvik. Conformance deterioration in QoS networks: Phenomenon, analysis and elimination. Computer Networks (Amsterdam, Netherlands: 1999), 50(12):2112–2126, August 24, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBl

[1140] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(12):??, August 24, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2006:NGD

[1141] Ian F. Akyildiz, Won-Yeol Lee, Mehmet C. Vuran, and Shantidev Mohanty. NeXt generation/dynamic spectrum access/cognitive radio wireless networks: A survey. Computer Networks (Amsterdam, Netherlands: 1999), 50(13):2127–2159, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zou:2006:TMS

[1142] Jianping Zou, M. Ümit Uyar, Mariusz A. Fecko, and Sunil Samtani. Throughput models for SCTP with parallel subflows. Computer Networks (Amsterdam, Netherlands: 1999), 50 (13):2160–2182, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gong:2006:RUW

[1143] Mingwei Gong and Carey Williamson. Revisiting unfairness in Web server scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 50(13): 2183–2203, September 15, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chang:2006:SQB

[1144] Xiaolin Chang and Jogesh K. Muppala. A stable queue-based adaptive controller for improving AQM performance. Computer Networks (Amsterdam, Netherlands: 1999), 50(13):2204– 2224, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Obanaik:2006:SPE

[1145] Venkatesh Obanaik, Lillykutty Jacob, and A. L. Ananda. Secure performance enhancing proxy: To ensure end-to-end security and enhance TCP performance over IPv6 wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(13):2225– 2238, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhou:2006:AFJ

[1146] Yunkai Zhou and Harish Sethu. On achieving fairness in the joint allocation of buffer and bandwidth resources: Principles and algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 50(13):2239–2254, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Han:2006:QSP

[1147] Seung Chul Han, Puneet Zaroo, David K. Y. Yau, Yu Dong, Prem Gopalan, and John C. S. Lui. Quality of service provisioning for composable routing elements. Computer Networks (Amsterdam, Netherlands: 1999), 50(13):2255– 2270, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mhamdi:2006:HPS

[1148] Lotfi Mhamdi, Mounir Hamdi, Christopher Kachris, Stephan Wong, and Stamatis Vassiliadis. High-performance switching based on buffered crossbar fabrics. Computer Networks (Amsterdam, Netherlands: 1999), 50(13):2271– 2285, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Atteo:2006:LWA

[1149] E. Atteo, S. Avallone, and S. P. Romano. A link weight assignment algorithm for traffic-engineered networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(13):2286– 2294, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Heikkinen:2006:PGA

[1150] T. Heikkinen. A potential game approach to distributed power control and scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 50(13): 2295–2311, September 15, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sun:2006:DMD

[1151] Haibin Sun, John C. S. Lui, and David K. Y. Yau. Distributed mechanism in detecting and defending against the low-rate TCP attack. Computer Networks (Amsterdam, Netherlands: 1999), 50(13):2312–2330, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guirguis:2006:ELA

[1152] Mina Guirguis, Azer Bestavros, and Ibrahim Matta. Exogenous-loss aware traffic management in overlay networks: Toward global fairness. Computer Networks (Amsterdam, Netherlands: 1999), 50(13):2331– 2348, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ahuja:2006:OPS

[1153] Satyajeet Ahuja, Marwan Krunz, and Turgay Korkmaz. Optimal path selection for minimizing the differential delay in Ethernet-over-SONET. Computer Networks (Amsterdam, Netherlands: 1999), 50(13): 2349–2363, September 15, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shen:2006:BAW

[1154] Shu Shen, Gaoxi Xiao, and Tee Hiang Cheng. Benefits of advertising wavelength availability in distributed lightpath establishment. Computer Networks (Amsterdam, Netherlands: 1999), 50(13):2364–2379, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cooper:2006:SMA

[1155] Brian F. Cooper and Hector Garcia-Molina. SIL: A model for analyzing scalable peer-to-peer search networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(13):2380– 2400, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ren:2006:ROS

[1156] Kui Ren, Wenjing Lou, Kai Zeng, Feng Bao, Jianying Zhou, and Robert H. Deng. Routing optimization security in mobile IPv6. Computer Networks (Amsterdam, Netherlands: 1999), 50(13): 2401–2419, September 15, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBm

[1157] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(13):??, September 15, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Plattner:2006:GE

[1158] Bernhard Plattner and James P. G. Sterbenz. Guest Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 50(14):2421–2425, October 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shalaby:2006:SSL

[1159] Nadia Shalaby, Yitzchak Gottlieb, Mike Wawrzoniak, and Larry Peterson. Snow on silk: Layering NodeOS functionality on conventional operating systems. Computer Networks (Amsterdam, Netherlands: 1999), 50(14): 2426–2443, October 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bos:2006:DEC

[1160] Herbert Bos, Bart Samwel, and Ilja Booij. Dynamically extending the Corral with native code for high-speed packet processing. Computer Networks (Amsterdam, Netherlands: 1999), 50 (14):2444–2461, October 5, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Donnelly:2006:LTT

[1161] Austin Donnelly. Lightweight thread tunnelling in network applications. Computer Networks (Amsterdam, Netherlands: 1999), 50(14):2462–2474, October 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dhillon:2006:BLA

[1162] Chetan Singh Dhillon, Mary Bond, James Griffioen, and Kenneth L. Calvert. Building layered active services. Computer Networks (Amsterdam, Netherlands: 1999), 50(14):2475– 2487, October 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Schmid:2006:HFS

[1163] S. Schmid, T. Chart, M. Sifalakis, and A. C. Scott. A highly flexible service composition framework for real-life networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(14): 2488–2505, October 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Solarski:2006:DMC

[1164] Marcin Solarski, Matthias Bossardt, and Thomas Becker. Deployment and management of component-based services in active networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(14):2506–2518, October 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Smith:2006:LRP

[1165] Paul Smith, Steven Simpson, and David Hutchison. Locating resources in a programmable networking environment. Computer Networks (Amsterdam, Netherlands: 1999), 50(14): 2519–2531, October 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sehgal:2006:FCB

[1166] Amit Sehgal, Kenneth L. Calvert, and James Griffioen. A flexible concastbased grouping service. Computer Networks (Amsterdam, Netherlands: 1999), 50(14):2532–2547, October 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anagnostakis:2006:FNM

[1167] K. G. Anagnostakis, M. B. Greenwald, S. Ioannidis, D. Li, and J. M. Smith. Flexible network monitoring with FLAME. Computer Networks (Amsterdam, Netherlands: 1999), 50 (14):2548–2563, October 5, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rogers:2006:AAN

[1168] Craig Milo Rogers. ANQL – An Active Networks Query Language. Computer Networks (Amsterdam, Netherlands: 1999), 50(14):2564–2576, October 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBn

[1169] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(14):??, October 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pack:2006:CEA

[1170] Sangheon Pack, Kyoungae Kim, and Yanghee Choi. A cost-effective approach to selective IP paging scheme using explicit multicast. Computer Networks (Amsterdam, Netherlands: 1999), 50(15):2577–2594, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vidhyashankar:2006:SAS

[1171] V. Vidhyashankar, B. S. Manoj, and C. Siva Ram Murthy. Slot allocation schemes for delay sensitive traffic support in asynchronous wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(15): 2595–2613, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2006:ERM

[1172] Jiang Li, Murat Yuksel, and Shivkumar Kalyanaraman. Explicit rate multicast congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 50 (15):2614–2640, October 18, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rajendran:2006:OQS

[1173] Raj Kumar Rajendran and Dan Rubenstein. Optimizing the quality of scalable video streams on P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(15):2641– 2658, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Meddahi:2006:PMM

[1174] Ahmed Meddahi and Hossam Afifi.

“Packet-E-Model”: E-Model for VoIP quality evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 50 (15):2659–2675, October 18, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cho:2006:FSM

[1175] Chia Yuan Cho, Winston Khoon Guan Seah, and Yong Huat Chew. A framework and source model for design and evaluation of Robust Header Compression. Computer Networks (Amsterdam, Netherlands: 1999), 50(15):2676– 2712, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2006:CMA

[1176] Wenye Wang and Guoliang Xue. A cost-minimization algorithm for fast location tracking in mobile wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(15):2713– 2726, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hui:2006:SWO

[1177] Ken Y. K. Hui, John C. S. Lui, and David K. Y. Yau. Small-world overlay P2P networks: Construction, management and handling of dynamic flash crowds. Computer Networks (Amsterdam, Netherlands: 1999), 50(15): 2727–2746, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

St-Hilaire:2006:UUN

[1178] Marc St-Hilaire, Steven Chamberland, and Samuel Pierre. Uplink UMTS network design—an integrated approach. Computer Networks (Amsterdam, Netherlands: 1999), 50(15): 2747–2761, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2006:ITP

[1179] Jung-Shian Li and Chuan-Gang Liu. Improve TCP performance with a novel adaptive queue management scheme in IEEE 802.17 RPR optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(15):2762– 2786, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gupta:2006:PRL

[1180] Aparna Gupta, Shivkumar Kalyanaraman, and Lingyi Zhang. Pricing of risk for loss guaranteed intra-domain Internet service contracts. Computer Networks (Amsterdam, Netherlands: 1999), 50(15):2787–2804, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kowalik:2006:RTF

[1181] Karol Kowalik and Martin Collier. Reducing traffic fluctuations of link state QoS routing algorithms in virtual circuit networks. Computer Networks (Amsterdam, Netherlands: 1999), 50 (15):2805–2819, October 18, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Meng:2006:CMM

[1182] Xiaoqiao Meng, Thyaga Nandagopal, Li Li, and Songwu Lu. Contour maps: Monitoring and diagnosis in sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(15):2820– 2838, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2006:MLS

[1183] Hai Liu, Pengjun Wan, and Xiaohua Jia. Maximal lifetime scheduling for K to 1 sensor–target surveillance networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(15):2839– 2854, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wei:2006:CAH

[1184] Qing Wei, Károly Farkas, Christian Prehofer, Paulo Mendes, and Bernhard Plattner. Context-aware handover using active network technology. Computer Networks (Amsterdam, Netherlands: 1999), 50(15):2855–2872, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2006:ERI

[1185] Huaming Wu and Alhussein A. Abouzeid. Error resilient image transport in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(15):2873–2887, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chai:2006:LMG

[1186] Ning Chai, Boon Sain Yeo, and Yong Huat Chew. Location management for GPRS. Computer Networks (Amsterdam, Netherlands: 1999), 50 (15):2888–2901, October 18, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kwon:2006:GRP

[1187] Sungoh Kwon and Ness B. Shroff. Geographic routing in the presence of location errors. Computer Networks (Amsterdam, Netherlands: 1999), 50(15): 2902–2917, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBo

[1188] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(15):??, October 18, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

He:2006:SSS

[1189] Guanghui He and Jennifer C. Hou.

On sampling self-similar Internet traffic. Computer Networks (Amsterdam, Netherlands: 1999), 50(16):2919– 2936, November 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gavalas:2006:LCI

[1190] Damianos Gavalas and Christina Tanya Politi. Low-cost itineraries for multihop agents designed for scalable monitoring of multiple subnets. Computer Networks (Amsterdam, Netherlands: 1999), 50(16):2937–2952, November 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

McDaniel:2006:OAI

[1191] Patrick McDaniel, William Aiello, Kevin Butler, and John Ioannidis. Origin authentication in interdomain routing. Computer Networks (Amsterdam, Netherlands: 1999), 50(16):2953– 2980, November 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vacirca:2006:ADT

[1192] Francesco Vacirca, Thomas Ziegler, and Eduard Hasenleithner. An algorithm to detect TCP spurious timeouts and its application to operational UMTS/GPRS networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(16):2981–3001, November 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2006:GTA

[1193] Kuan-Ta Chen, Polly Huang, and Chin-Laung Lei. Game traffic analysis: An MMORPG perspective. Computer Networks (Amsterdam, Netherlands: 1999), 50(16):3002–3023, November 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yu:2006:PSM

[1194] Ming Yu, Wenjui Li, and Li jin W. Chung. A practical scheme for MPLS fault monitoring and alarm correlation in backbone networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(16):3024– 3042, November 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Balon:2006:SDF

[1195] Simon Balon, Laurent Mélon, and Guy Leduc. A scalable and decentralized fast-rerouting scheme with efficient bandwidth sharing. Computer Networks (Amsterdam, Netherlands: 1999), 50(16):3043–3063, November 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zeinalipour-Yazti:2006:STA

[1196] Demetrios Zeinalipour-Yazti and Vana Kalogeraki. Structuring topologically aware overlay networks using domain names. Computer Networks (Amsterdam, Netherlands: 1999), 50(16):3064– 3082, November 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lam:2006:FRS

[1197] Simon S. Lam and Huaiyu Liu. Failure recovery for structured P2P networks: Protocol design and performance under churn. Computer Networks (Amsterdam, Netherlands: 1999), 50(16): 3083–3104, November 14, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hamad:2006:MWR

[1198] Ashraf Hamad, Tao Wu, Ahmed E. Kamal, and Arun K. Somani. On multicasting in wavelength-routing mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(16): 3105–3164, November 14, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gao:2006:CEG

[1199] Zhenguo Gao, Ling Wang, Mei Yang, and Xiaozong Yang. CNPGSDP: An efficient group-based service discovery protocol for MANETs. Computer Networks (Amsterdam, Netherlands: 1999), 50(16):3165–3182, November 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dolev:2006:IRA

[1200] Danny Dolev, Sugih Jamin, Osnat (Ossi) Mokryn, and Yuval Shavitt. Internet resiliency to attacks and failures under BGP policy routing. Computer Networks (Amsterdam, Netherlands: 1999), 50(16):3183– 3196, November 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guillaume:2006:RMD

[1201] Jean-Loup Guillaume, Matthieu Latapy, and Damien Magoni. Relevance of massively distributed explorations of the Internet topology: Qualitative results. Computer Networks (Amsterdam, Netherlands: 1999), 50(16):3197– 3224, November 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBp

[1202] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(16):??, November 14, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xenakis:2006:GCO

[1203] Christos Xenakis, Nikolaos Laoutaris, Lazaros Merakos, and Ioannis Stavrakakis. A generic characterization of the overheads imposed by IPsec and associated cryptographic algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 50(17):3225–3241, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Touati:2006:GNB

[1204] Corinne Touati, Eitan Altman, and Jérôme Galtier. Generalized Nash Bargaining Solution for bandwidth allocation. Computer Networks (Amsterdam, Netherlands: 1999), 50(17):3242– 3263, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Campo:2006:PLD

[1205] Celeste Campo, Carlos García-Rubio, Andrés Marín López, and Florina Almenárez. PDP: A lightweight discovery protocol for local-scope interactions in wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(17):3264–3283, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bauer:2006:TAC

[1206] Daniel Bauer, John N. Daigle, Ilias Iliadis, and Paolo Scotton. Topology aggregation for combined additive and restrictive metrics. Computer Networks (Amsterdam, Netherlands: 1999), 50 (17):3284–3299, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cheng:2006:RDB

[1207] Gang Cheng and Nirwan Ansari.

Rate-distortion based link state update. Computer Networks (Amsterdam, Netherlands: 1999), 50(17):3300– 3314, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xiao:2006:BSL

[1208] Li Xiao, Guanghui He, and Klara Nahrstedt. BGP session lifetime modeling in congested networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(17):3315–3333, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ge:2006:MEI

[1209] Yu Ge and Winston Seah. A method to efficiently integrate Internet Telephony call signaling with dynamic resource negotiation. Computer Networks (Amsterdam, Netherlands: 1999), 50(17): 3334–3352, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Christin:2006:QQS

[1210] Nicolas Christin and Jörg Liebeherr. The QoSbox: Quantitative service differentiation in BSD routers. Computer Networks (Amsterdam, Netherlands: 1999), 50(17):3353–3374, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Er:2006:PAM

[1211] Inn Inn Er and Winston K. G. Seah. Performance analysis of mobility-based d-hop (MobDHop) clustering algorithm for mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(17):3375–3399, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xiao:2006:NFA

[1212] Shi Xiao, Gaoxi Xiao, and Yiu-Wing Leung. A network flow approach for static and dynamic traffic grooming in WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 50 (17):3400–3415, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fu:2006:BQS

[1213] Xiaoming Fu, Hannes Tschofenig, and Dieter Hogrefe. Beyond QoS signaling: A new generic IP signaling framework. Computer Networks (Amsterdam, Netherlands: 1999), 50(17):3416– 3433, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Na:2006:GNG

[1214] Jongkeun Na and Chong kwon Kim.

GLR: A novel geographic routing scheme for large wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(17):3434– 3448, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gu:2006:DHM

[1215] Yaoyao Gu, Doruk Bozdağ, Robert W.

Brewer, and Eylem Ekici. Data harvesting with mobile elements in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 50 (17):3449–3465, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tugcu:2006:AMC

[1216] Tuna Tugcu, H. Birkan Yilmaz, and Feodor Vainstein. Analytical modeling of CAC in next generation wireless systems. Computer Networks (Amsterdam, Netherlands: 1999), 50(17): 3466–3484, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Risson:2006:SRT

[1217] John Risson and Tim Moors. Survey of research towards robust peer-to-peer networks: Search methods. Computer Networks (Amsterdam, Netherlands: 1999), 50(17):3485–3521, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:EBq

[1218] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 50(17):??, December 5, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pao:2006:EPC

[1219] Derek Pao, Yiu Keung Li, and Peng Zhou. Efficient packet classification using TCAMs. Computer Networks (Amsterdam, Netherlands: 1999), 50(18): 3523–3535, December 21, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ma:2006:TMS

[1220] Miao Ma. Tabu marking scheme to speedup IP traceback. Computer Networks (Amsterdam, Netherlands: 1999), 50(18):3536–3549, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sahasrabudhe:2006:POB

[1221] Amol Sahasrabudhe and D. Manjunath. Performance of optical burst switched networks: A two moment analysis. Computer Networks (Amsterdam, Netherlands: 1999), 50(18):3550– 3563, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Monaco:2006:UOD

[1222] U. Monaco, F. Cuomo, T. Melodia, F. Ricciato, and M. Borghini. Understanding optimal data gathering in the energy and latency domains of a wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 50(18):3564–3584, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2006:GFC

[1223] Hongwei Zhang and Anish Arora.

Guaranteed fault containment and local stabilization in routing. Computer Networks (Amsterdam, Netherlands: 1999), 50(18):3585–3607, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mundur:2006:OSA

[1224] Padmavathi Mundur and Poorva Arankalle. Optimal server allocations for streaming multimedia applications on the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 50 (18):3608–3621, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dong:2006:LLP

[1225] Yingfei Dong, Changho Choi, and Zhi-Li Zhang. LIPS: A lightweight permit system for packet source origin accountability. Computer Networks (Amsterdam, Netherlands: 1999), 50(18): 3622–3641, December 21, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rawat:2006:PPO

[1226] Anuj Rawat and Mark A. Shayman.

Preventing persistent oscillations and loops in IBGP configuration with route reflection. Computer Networks (Amsterdam, Netherlands: 1999), 50(18): 3642–3665, December 21, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rammohan:2006:LMR

[1227] N. Rammohan and C. Siva Ram Murthy. On-line multicast routing with QoS constraints in WDM networks with no wavelength converters. Computer Networks (Amsterdam, Netherlands: 1999), 50(18):3666– 3685, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Huang:2006:FTH

[1228] Yin-Fu Huang and Min-Hsiu Chuang. Fault tolerance for home agents in mobile IP. Computer Networks (Amsterdam, Netherlands: 1999), 50(18):3686– 3700, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Delicato:2006:EHS

[1229] Flávia Delicato, Fábio Protti, Luci Pirmez, and José Ferreira de Rezende. An efficient heuristic for selecting active nodes in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 50(18):3701– 3720, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2006:PBC

[1230] Gang Wu, Edwin K. P. Chong, and Robert Givan. Predictive buffer control in delivering remotely stored video using proxy servers. Computer Networks (Amsterdam, Netherlands: 1999), 50 (18):3721–3742, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zheng:2006:HRE

[1231] Yanxing Zheng, Turgay Korkmaz, Wenhua Dou, and Jing Tian. Highly responsive and efficient QoS routing using pre- and on-demand computations along with a new normal measure. Computer Networks (Amsterdam, Netherlands: 1999), 50(18):3743– 3762, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bartolini:2006:SBA

[1232] Novella Bartolini. Session based access control in geographically replicated Internet services. Computer Networks (Amsterdam, Netherlands: 1999), 50 (18):3763–3783, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhou:2006:NRS

[1233] Kaiyu Zhou, Kwan L. Yeung, and Victor O. K. Li. Nonlinear RED: A simple yet efficient active queue management scheme. Computer Networks (Amsterdam, Netherlands: 1999), 50(18):3784– 3794, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2006:SIV

[1234] Anonymous. Subject index volume 50. Computer Networks (Amsterdam, Netherlands: 1999), 50(18):3795– 3803, December 21, 2006. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sebe:2007:SSB

[1235] Francesc Sebé and Josep Domingo-Ferrer. Scalability and security in biased many-to-one communication. Computer Networks (Amsterdam, Netherlands: 1999), 51(1):1–13, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Careglio:2007:SCW

[1236] D. Careglio, J. Solé Pareta, and S. Spadaro. Service category-to-wavelength selection technique for QoS support in connection-oriented optical packet switching. Computer Networks (Amsterdam, Netherlands: 1999), 51 (1):14–30, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

To:2007:POH

[1237] K. K. To and Jack Y. B. Lee. Parallel overlays for high data-rate multicast data transfer. Computer Networks (Amsterdam, Netherlands: 1999), 51 (1):31–42, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2007:TDC

[1238] Junfeng Wang, Lei Li, and Mingtian Zhou. Topological dynamics characterization for LEO satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(1): 43–53, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2007:IOS

[1239] Zhi Li and Prasant Mohapatra. On investigating overlay service topologies. Computer Networks (Amsterdam, Netherlands: 1999), 51(1):54–68, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zheng:2007:CCE

[1240] Shanyu Zheng, David Manz, and Jim Alves-Foss. A communication– computation efficient group key algorithm for large and dynamic groups. Computer Networks (Amsterdam, Netherlands: 1999), 51(1): 69–93, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tran:2007:DFT

[1241] Hung Tuan Tran and Thomas Ziegler.

A design framework towards the profitable operation of service overlay networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(1): 94–113, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Carofiglio:2007:BFM

[1242] G. Carofiglio, M. Garetto, E. Leonardi, A. Tarello, and M. Ajmone Marsan. Beyond fluid models: Modelling TCP mice in IP networks under nonstationary random traffic. Computer Networks (Amsterdam, Netherlands: 1999), 51(1):114–133, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Feng:2007:LMB

[1243] Li Feng, Qinglin Zhao, and Hanwen Zhang. Location management based on distance and direction for PCS networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(1): 134–152, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Reddy:2007:PMQ

[1244] T. Bheemarjuna Reddy, John P. John, and C. Siva Ram Murthy. Providing MAC QoS for multimedia traffic in 802.11e based multi-hop ad hoc wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (1):153–176, January 17, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Meng:2007:PPW

[1245] Tao Meng and Hong-Fei Yan. On the peninsula phenomenon in web graph and its implications on web search. Computer Networks (Amsterdam, Netherlands: 1999), 51(1): 177–189, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kutten:2007:ARD

[1246] Shay Kutten and David Peleg. Asynchronous resource discovery in peerto-peer networks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (1):190–206, January 17, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Spadaro:2007:TAA

[1247] Salvatore Spadaro, Alessandro D’Alessandro, Antonio Manzalini, and Josep Solé-Pareta. TRIDENT: An automated approach to traffic engineering in IP/MPLS over ASON/GMPLS networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(1): 207–223, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hierons:2007:RCA

[1248] R. M. Hierons and H. Ural. Reducing the cost of applying adaptive test cases. Computer Networks (Amsterdam, Netherlands: 1999), 51(1): 224–238, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Khan:2007:SRA

[1249] Javed I. Khan and Raid Y. Zaghal.

Symbiotic rate adaptation for time sensitive elastic traffic with interactive transport. Computer Networks (Amsterdam, Netherlands: 1999), 51(1): 239–257, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yamaguchi:2007:DPS

[1250] Hirozumi Yamaguchi, Khaled El-Fakih, Gregor v. Bochmann, and Teruo Higashino. Deriving protocol specifications from service specifications written as Predicate/Transition-nets. Computer Networks (Amsterdam, Netherlands: 1999), 51(1):258–284, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hu:2007:SSW

[1251] Fei Hu, Waqaas Siddiqui, and Krishna Sankar. Scalable security in Wireless Sensor and Actuator Networks (WSANs): Integration re-keying with routing. Computer Networks (Amsterdam, Netherlands: 1999), 51(1): 285–308, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Postigo-Boix:2007:CEM

[1252] Marcos Postigo-Boix, Joan Garcia-Haro, and Jose L. Melús-Moreno. A cost-efficient method for streaming stored content in a guaranteed QoS Internet. Computer Networks (Amsterdam, Netherlands: 1999), 51(1): 309–335, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tang:2007:MGR

[1253] Wenting Tang, Yun Fu, Ludmila Cherkasova, and Amin Vahdat. Modeling and generating realistic streaming media server workloads. Computer Networks (Amsterdam, Netherlands: 1999), 51(1):336–356, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBa

[1254] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(1):??, January 17, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Reiff-Marganiec:2007:GE

[1255] Stephan Reiff-Marganiec and Mark D. Ryan. Guest Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 51(2):357–358, February 7, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Weiss:2007:TCW

[1256] M. Weiss, B. Esfandiari, and Y. Luo. Towards a classification of web service feature interactions. Computer Networks (Amsterdam, Netherlands: 1999), 51(2):359–381, February 7, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Crespo:2007:DRF

[1257] Rui Gustavo Crespo, Miguel Carvalho, and Luigi Logrippo. Distributed resolution of feature interactions for Internet applications. Computer Networks (Amsterdam, Netherlands: 1999), 51 (2):382–397, February 7, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shehata:2007:TIR

[1258] Mohamed Shehata, Armin Eberlein, and Abraham O. Fapojuwo. A taxonomy for identifying requirement interactions in software systems. Computer Networks (Amsterdam, Netherlands: 1999), 51(2):398–425, February 7, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chi:2007:TGI

[1259] Caixia Chi and Ruibing Hao. Test generation for interaction detection in feature-rich communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 51(2):426– 438, February 7, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Miller:2007:TBA

[1260] A. Miller, M. Calder, and A. F. Donaldson. A template-based approach for the generation of abstractable and reducible models of featured networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(2): 439–455, February 7, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Schobbens:2007:GSF

[1261] Pierre-Yves Schobbens, Patrick Heymans, Jean-Christophe Trigaux, and Yves Bontemps. Generic semantics of feature diagrams. Computer Networks (Amsterdam, Netherlands: 1999), 51 (2):456–479, February 7, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Leung:2007:PEF

[1262] Wu-Hon F. Leung. Program entanglement, feature interaction and the feature language extensions. Computer Networks (Amsterdam, Netherlands: 1999), 51(2):480–495, February 7, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Turner:2007:PCC

[1263] Kenneth J. Turner and Lynne Blair. Policies and conflicts in call control. Computer Networks (Amsterdam, Netherlands: 1999), 51(2):496– 514, February 7, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2007:HFI

[1264] Xiaotao Wu and Henning Schulzrinne. Handling feature interactions in the language for end system services. Computer Networks (Amsterdam, Netherlands: 1999), 51(2):515–535, February 7, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kolberg:2007:MFI

[1265] Mario Kolberg and Evan H. Magill. Managing feature interactions between distributed SIP call control services. Computer Networks (Amsterdam, Netherlands: 1999), 51(2):536– 557, February 7, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBb

[1266] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(2):??, February 7, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Praveen:2007:UFE

[1267] B. Praveen, J. Praveen, and C. Siva Ram Murthy. On using forward error correction for loss recovery in optical burst switched networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):559–568, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rodrigues:2007:BUD

[1268] Carlo K. da S. Rodrigues and Rosa M. M. Leão. Bandwidth usage distribution of multimedia servers using Patching. Computer Networks (Amsterdam, Netherlands: 1999), 51(3): 569–587, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chang:2007:FBM

[1269] Yeim-Kuan Chang. Fast binary and multiway prefix searches for packet forwarding. Computer Networks (Amsterdam, Netherlands: 1999), 51(3): 588–605, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2007:ICC

[1270] Sungchang Kim, Young-Chon Kim, Bin-Yeong Yoon, and Minho Kang. An integrated congestion control mechanism for optimized performance using two-step rate controller in optical burst switching networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):606–620, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Choi:2007:ULS

[1271] Young-June Choi, Jin-Ghoo Choi, and Saewoong Bahk. Upper-level scheduling supporting multimedia traffic in cellular data networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):621–631, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yu:2007:IQA

[1272] Dong Yu and Deborah Frincke. Improving the quality of alerts and predicting intruder’s next goal with Hidden Colored Petri-Net. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):632–654, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wan:2007:ACP

[1273] Zhiguo Wan, Robert H. Deng, Feng Bao, and Akkihebbal L. Ananda. Access control protocols with twolayer architecture for wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):655– 670, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2007:PAL

[1274] Hailong Li and Ian Li-Jin Thng. Performance analysis of limited number of wavelength converters by share per node in optical switching network. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):671– 682, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Haga:2007:GMP

[1275] Péter Hága, Krisztián Diriczi, Gábor Vattay, and István Csabai. Granular model of packet pair separation in Poissonian traffic. Computer Networks (Amsterdam, Netherlands: 1999), 51 (3):683–698, February 21, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kelif:2007:AGC

[1276] Jean Marc Kelif, Eitan Altman, and Ioannis Koukoutsidis. Admission and GoS control in a multiservice WCDMA system. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):699– 711, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Joung:2007:CTL

[1277] Yuh-Jzer Joung and Jiaw-Chang Wang. Chord^{2}: A two-layer Chord for reducing maintenance overhead via heterogeneity. Computer Networks (Amsterdam, Netherlands: 1999), 51 (3):712–731, February 21, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gao:2007:PRI

[1278] Zhiqiang Gao and Nirwan Ansari. A practical and robust inter-domain marking scheme for IP traceback. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):732–750, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Benson:2007:PST

[1279] Glenn Benson. Portable security transaction protocol. Computer Networks (Amsterdam, Netherlands: 1999), 51 (3):751–766, February 21, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2007:NPC

[1280] Meng Chang Chen, Li-Ping Tung, Yeali S. Sun, and Wei-Kuan Shih. A new per-class flow fixed proportional differentiated service for multi-service wireless LAN. Computer Networks (Amsterdam, Netherlands: 1999), 51 (3):767–780, February 21, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mandjes:2007:GTQ

[1281] Michel Mandjes, Petteri Mannersalo, and Ilkka Norros. Gaussian tandem queues with an application to dimensioning of switch fabric interfaces. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):781– 797, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ye:2007:PCR

[1282] Zhenqiang Ye, Lap Kong Law, Srikanth V. Krishnamurthy, Zhong Xu, Suvidhean Dhirakaosal, Satish K. Tripathi, and Mart Molle. Predictive channel reservation for handoff prioritization in wireless cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):798– 822, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Law:2007:NAP

[1283] Lap Kong Law, Srikanth V. Krishnamurthy, and Michalis Faloutsos. A novel adaptive protocol for lightweight efficient multicasting in ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):823– 834, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sanchez-Lopez:2007:FSE

[1284] Sergio Sánchez-Lopez, Xavier Masip-Bruin, Josep Sole-Pareta, and Jordi Domingo-Pascual. Fast setup of end-to-end paths for bandwidth constrained applications in an IP/MPLS– ATM integrated environment. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):835–852, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pan:2007:IBS

[1285] Jianping Pan, Lin Cai, Xuemin (Sherman) Shen, and Jon W. Mark. Identitybased secure collaboration in wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (3):853–865, February 21, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2007:DPP

[1286] Jenshiuh Liu, Zhi-Jian Lee, and Yeh-Ching Chung. Dynamic probabilistic packet marking for efficient IP traceback. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):866– 882, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tsai:2007:TTA

[1287] Yuh-Ren Tsai and Shiuh-Jeng Wang. Two-tier authentication for cluster and individual sets in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):883– 900, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Biersack:2007:OAF

[1288] E. W. Biersack, D. Carra, R. Lo Cigno, P. Rodriguez, and P. Felber. Overlay architectures for file distribution: Fundamental performance analysis for homogeneous and heterogeneous cases. Computer Networks (Amsterdam, Netherlands: 1999), 51(3): 901–917, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBc

[1289] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(3):??, February 21, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2007:E

[1290] Ian. F. Akyildiz and Harry Rudin. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 51(4):

919–920, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2007:SWM

[1291] Ian F. Akyildiz, Tommaso Melodia, and Kaushik R. Chowdhury. A survey on wireless multimedia sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(4): 921–960, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kredo:2007:MAC

[1292] Kurtis Kredo II and Prasant Mohapatra. Medium access control in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (4):961–994, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2007:PEM

[1293] Yi Li, Gaoxi Xiao, and H. Ghafouri-Shiraz. Performance evaluation of multi-fiber optical packet switches. Computer Networks (Amsterdam, Netherlands: 1999), 51(4):995–1012, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Macia-Fernandez:2007:ELR

[1294] Gabriel Maciá-Fernández, Jesús E.

Díaz-Verdejo, and Pedro García-Teodoro. Evaluation of a low-rate DoS attack against iterative servers. Computer Networks (Amsterdam, Netherlands: 1999), 51(4):1013–1030, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ferentinos:2007:ADO

[1295] Konstantinos P. Ferentinos and Theodore A. Tsiligiridis. Adaptive design optimization of wireless sensor networks using genetic algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 51(4):1031–1051, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Willke:2007:CIU

[1296] Theodore L. Willke and Nicholas F. Maxemchuk. Coordinated interaction using reliable broadcast in mobile wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (4):1052–1059, March 14, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ghaderi:2007:JCP

[1297] Majid Ghaderi, Raouf Boutaba, and Gary W. Kenward. Joint call and packet QoS in cellular packet networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(4): 1060–1071, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2007:CEO

[1298] Jinwon Lee, Hyonik Lee, Seungwoo Kang, Su Myeon Kim, and Junehwa Song. CISS: An efficient object clustering framework for DHT-based peer-to-peer applications. Computer Networks (Amsterdam, Netherlands: 1999), 51 (4):1072–1094, March 14, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mohanty:2007:PAN

[1299] Shantidev Mohanty and Jiang Xie. Performance analysis of a novel architecture to integrate heterogeneous wireless systems. Computer Networks (Amsterdam, Netherlands: 1999), 51(4): 1095–1105, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gouda:2007:SFD

[1300] Mohamed G. Gouda and Alex X.

Liu. Structured firewall design. Computer Networks (Amsterdam, Netherlands: 1999), 51(4):1106–1120, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jayachandran:2007:PEQ

[1301] Praveen Jayachandran, Praveen Bhamidipati, and C. Siva Ram Murthy. On providing elastic QoS in optical burst switched networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(4):1121–1132, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Huang:2007:SFB

[1302] Yaqing Huang and Roch Guérin. A simple FIFO-based scheme for differentiated loss guarantees. Computer Networks (Amsterdam, Netherlands: 1999), 51(4):1133–1150, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Su:2007:EEE

[1303] Wei-Tsung Su, Ko-Ming Chang, and Yau-Hwang Kuo. eHIP: An energyefficient hybrid intrusion prohibition system for cluster-based wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(4): 1151–1168, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Castelluccia:2007:RSK

[1304] Claude Castelluccia, Nitesh Saxena, and Jeong Hyun Yi. Robust selfkeying mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(4):1169–1182, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zarifzadeh:2007:RFL

[1305] Sajjad Zarifzadeh, Nasser Yazdani, and Hamed Khanmirza. A routing framework for load balancing of bandwidth sensitive traffic in differentiated service networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(4): 1183–1204, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ricciato:2007:DCB

[1306] Fabio Ricciato, Francesco Vacirca, and Philipp Svoboda. Diagnosis of capacity bottlenecks via passive monitoring in 3G networks: An empirical analysis. Computer Networks (Amsterdam, Netherlands: 1999), 51(4): 1205–1231, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBd

[1307] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(4):??, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:LR

[1308] Anonymous. List of Reviewers. Computer Networks (Amsterdam, Netherlands: 1999), 51(4):??, March 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Frincke:2007:IDS

[1309] Deborah Frincke, Andreas Wespi, and Diego Zamboni. From intrusion detection to self-protection. Computer Networks (Amsterdam, Netherlands: 1999), 51(5):1233–1238, April 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ingham:2007:LDR

[1310] Kenneth L. Ingham, Anil Somayaji, John Burge, and Stephanie Forrest. Learning DFA representations of HTTP for protecting web applications. Computer Networks (Amsterdam, Netherlands: 1999), 51(5): 1239–1255, April 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Portokalidis:2007:SZH

[1311] Georgios Portokalidis and Herbert Bos.

SweetBait: Zero-hour worm detection and containment using low- and high-interaction honeypots. Computer Networks (Amsterdam, Netherlands: 1999), 51(5):1256–1274, April 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Laureano:2007:PHB

[1312] M. Laureano, C. Maziero, and E. Jamhour. Protecting host-based intrusion detectors through virtual machines. Computer Networks (Amsterdam, Netherlands: 1999), 51(5):

1275–1283, April 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Williams:2007:CEH

[1313] Paul D. Williams and Eugene H. Spafford. CuPIDS: An exploration of highly focused, co-processor-based information system protection. Computer Networks (Amsterdam, Netherlands: 1999), 51(5):1284–1298, April 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gelenbe:2007:SAA

[1314] Erol Gelenbe and George Loukas. A self-aware approach to denial of service defence. Computer Networks (Amsterdam, Netherlands: 1999), 51(5): 1299–1314, April 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Swimmer:2007:UDM

[1315] Morton Swimmer. Using the danger model of immune systems for distributed defense in modern data networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(5): 1315–1333, April 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2007:AAI

[1316] Yu-Sung Wu, Bingrui Foo, Yu-Chun Mao, Saurabh Bagchi, and Eugene H. Spafford. Automated adaptive intrusion containment in systems of interacting services. Computer Networks (Amsterdam, Netherlands: 1999), 51 (5):1334–1360, April 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Goel:2007:AHP

[1317] Ashvin Goel, Wu chang Feng, Wu chi Feng, and David Maier. Automatic high-performance reconstruction and recovery. Computer Networks (Amsterdam, Netherlands: 1999), 51(5): 1361–1377, April 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBe

[1318] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(5):??, April 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2007:MMS

[1319] Hwangnam Kim and Jennifer C.

Hou. Mixed-mode simulation for IEEE 802.11-operated WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 51(6):1379–1402, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2007:WQM

[1320] Jeng Farn Lee, Meng Chang Chen, and Yeali Sun. WF^{2} Q-M: Worst-case fair weighted fair queueing with maximum rate control. Computer Networks (Amsterdam, Netherlands: 1999), 51(6): 1403–1420, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2007:GMC

[1321] Jiang Li, Murat Yuksel, Xingzhe Fan, and Shivkumar Kalyanaraman. Generalized multicast congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 51(6):1421–1443, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2007:PDR

[1322] Lan Wang, Daniel Massey, and Lixia Zhang. Persistent detection and recovery of state inconsistencies. Computer Networks (Amsterdam, Netherlands: 1999), 51(6):1444–1458, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tarkoma:2007:CSH

[1323] Sasu Tarkoma and Jaakko Kangasharju. On the cost and safety of handoffs in content-based routing systems. Computer Networks (Amsterdam, Netherlands: 1999), 51(6): 1459–1482, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vales-Alonso:2007:PEM

[1324] Javier Vales-Alonso, Esteban Egea-López, Alejandro Martínez-Sala, Pablo Pavón-Mariño, M. Victoria Bueno-Delgado, and Joan García-Haro. Performance evaluation of MAC transmission power control in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(6): 1483–1498, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bisnik:2007:ORW

[1325] Nabhendra Bisnik and Alhussein A. Abouzeid. Optimizing random walk search algorithms in P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(6): 1499–1514, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rangarajan:2007:EUR

[1326] Hari Rangarajan and J. J. Garcia-Luna-Aceves. Efficient use of route requests for loop-free on-demand routing in ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(6):1515–1529, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Manoj:2007:MNM

[1327] B. S. Manoj, K. J. Kumar, and C. Siva Ram Murthy. MuPAC: A new multi-power architecture for capacity enhancement in cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(6): 1530–1554, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chew:2007:SSL

[1328] Y. H. Chew, B. S. Yeo, and D. C. M.

Kuan. Sensitivity study of location management area partitioning in cellular communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 51(6):1555–1573, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2007:AOP

[1329] Bo Li and Roberto Battiti. Achieving optimal performance in IEEE 802.11 wireless LANs with the combination of link adaptation and adaptive backoff. Computer Networks (Amsterdam, Netherlands: 1999), 51(6):1574–1600, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Khlifi:2007:ABA

[1330] Hechmi Khlifi and Jean-Charles Grégoire. ARTP: A buffer-aware rate control protocol for media streaming. Computer Networks (Amsterdam, Netherlands: 1999), 51(6):1601–1615, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Argiriou:2007:CSM

[1331] Nikos Argiriou and Leonidas Georgiadis. Channel sharing by multiclass rate adaptive streams: Performance region and optimization. Computer Networks (Amsterdam, Netherlands: 1999), 51(6):1616–1629, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pack:2007:PCM

[1332] Sangheon Pack, Taekyoung Kwon, and Yanghee Choi. A performance comparison of mobility anchor point selection schemes in Hierarchical Mobile IPv6 networks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (6):1630–1642, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jin:2007:AOM

[1333] Mushi Jin and Oliver W. W. Yang. APOSN: Operation, modeling and performance evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 51(6):1643–1659, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Daoud:2007:LCP

[1334] Ashraf Al Daoud and Murat Alanyali.

Loss-cognizant pricing in networks with greedy users. Computer Networks (Amsterdam, Netherlands: 1999), 51(6): 1660–1683, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zheng:2007:ACC

[1335] Feng Zheng and John Nelson. An H_{∞} approach to congestion control design for AQM routers supporting TCP flows in wireless access networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(6):1684–1704, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gao:2007:HTS

[1336] Feng Gao and M. E. Woodward. A hybrid topological–stochastic partitioning method for scaling QoS routing algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 51(6): 1705–1724, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Poo:2007:MPR

[1337] Gee-Swee Poo and Haibo Wang. Multipath routing versus tree routing for VPN bandwidth provisioning in the hose model. Computer Networks (Amsterdam, Netherlands: 1999), 51(6): 1725–1743, April 25, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBf

[1338] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(6):??, April 25, 2007.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kobayashi:2007:PFL

[1339] Katsushi Kobayashi, Pascale Vicat-Blanc Primet, and Joe Touch. Protocols for fast, long-distance networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(7): 1745–1747, May 16, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ha:2007:IBT

[1340] Sangtae Ha, Long Le, Injong Rhee, and Lisong Xu. Impact of background traffic on performance of high-speed TCP variant protocols. Computer Networks (Amsterdam, Netherlands: 1999), 51 (7):1748–1762, May 16, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2007:CHS

[1341] Shan Chen and Brahim Bensaou. Can high-speed networks survive with DropTail queues management? Computer Networks (Amsterdam, Netherlands: 1999), 51(7):1763–1776, May 16, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gu:2007:UUB

[1342] Yunhong Gu and Robert L. Grossman. UDT: UDP-based data transfer for high-speed wide area networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(7):1777–1799, May 16, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Huang:2007:NHS

[1343] Xiaomeng Huang, Chuang Lin, and Fengyuan Ren. A novel high speed transport protocol based on explicit virtual load feedback. Computer Networks (Amsterdam, Netherlands: 1999), 51(7):1800–1814, May 16, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sarolahti:2007:DAS

[1344] Pasi Sarolahti, Mark Allman, and Sally Floyd. Determining an appropriate sending rate over an underutilized network path. Computer Networks (Amsterdam, Netherlands: 1999), 51(7): 1815–1832, May 16, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nakauchi:2007:ERF

[1345] Kiyohide Nakauchi and Katsushi Kobayashi. An explicit router feedback framework for high bandwidth-delay product networks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (7):1833–1846, May 16, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xu:2007:EEB

[1346] Lisong Xu. Extending equation-based congestion control to high-speed and long-distance networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(7):1847–1859, May 16, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBg

[1347] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(7):??, May 16, 2007.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cohen:2007:ASP

[1348] Edith Cohen, Amos Fiat, and Haim Kaplan. Associative search in peer to peer networks: Harnessing latent semantics. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 1861–1881, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Raj:2007:SIM

[1349] Alex Raj and Oliver C. Ibe. A survey of IP and multiprotocol label switching fast reroute schemes. Computer Networks (Amsterdam, Netherlands: 1999), 51(8):1882–1907, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cohen:2007:GIP

[1350] Reuven Cohen and Amnon Shochot. The “Global-ISP” paradigm. Computer Networks (Amsterdam, Netherlands: 1999), 51(8):1908–1921, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2007:QPI

[1351] Hwangnam Kim, Jennifer C. Hou, Chunyu Hu, and Ye Ge. QoS provisioning in IEEE 802.11-compliant networks: Past, present, and future. Computer Networks (Amsterdam, Netherlands: 1999), 51(8):1922–1941, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Choi:2007:SPS

[1352] Jin-Hee Choi, Young-Pil Kim, and Chuck Yoo. Self-prevention of socket buffer overflow. Computer Networks (Amsterdam, Netherlands: 1999), 51 (8):1942–1954, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ge:2007:AST

[1353] Ye Ge, Jennifer C. Hou, and Sunghyun Choi. An analytic study of tuning systems parameters in IEEE 802.11e enhanced distributed channel access. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 1955–1980, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Weber:2007:FLQ

[1354] Steven Weber and Gustavo de Veciana.

Flow-level QoS for a dynamic load of rate adaptive sessions sharing a bottleneck link. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 1981–1997, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Galluccio:2007:GLA

[1355] Laura Galluccio, Giacomo Morabito, Sergio Palazzo, Marco Pellegrini, M. Elena Renda, and Paolo Santi. Georoy: A location-aware enhancement to Viceroy peer-to-peer algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 1998–2014, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wei:2007:CPD

[1356] Jianbin Wei and Cheng-Zhong Xu.

Consistent proportional delay differentiation: A fuzzy control approach. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 2015–2032, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Psaras:2007:WTT

[1357] Ioannis Psaras and Vassilis Tsaoussidis. Why TCP timers (still) don’t work well. Computer Networks (Amsterdam, Netherlands: 1999), 51(8):2033–2048, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2007:DMM

[1358] Chun-Chuan Yang, Chun-Shian Tsai, Junn-Yen Hu, and Tzu-Chien Chuang. On the design of mobility management scheme for 802.16-based network environment. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 2049–2066, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Leguay:2007:DSI

[1359] Jérémie Leguay, Matthieu Latapy, Timur Friedman, and Kavé Salamatian. Describing and simulating Internet routes. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 2067–2085, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chek:2007:DEP

[1360] Michael Cho-Hoi Chek and Yu-Kwong Kwok. Design and evaluation of practical coexistence management schemes for Bluetooth and IEEE 802.11b systems. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 2086–2103, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dhillon:2007:PAA

[1361] S. S. Dhillon and P. Van Mieghem. Performance analysis of the AntNet algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 2104–2125, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Manoj:2007:ULA

[1362] B. S. Manoj, Archana Sekhar, and C. Siva Ram Murthy. On the use of limited autonomous mobility for dynamic coverage maintenance in sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 2126–2143, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

MacLarty:2007:TPW

[1363] Glen MacLarty and Michael Fry. Towards a platform for wide-area overlay network deployment and management. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 2144–2162, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ramasubramanian:2007:DMR

[1364] Srinivasan Ramasubramanian, Harish Krishnamoorthy, and Marwan Krunz. Disjoint multipath routing using colored trees. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 2163–2180, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cetinkaya:2007:IEM

[1365] Coskun Cetinkaya. Improving the efficiency of multipath traffic via opportunistic traffic scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 51(8):2181–2197, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dolcy:2007:FPE

[1366] Dawn S. L. Dolcy, Costas C. Constantinou, and Steven F. Quigley. A Fokker–Planck equation method predicting Buffer occupancy in a single queue. Computer Networks (Amsterdam, Netherlands: 1999), 51(8): 2198–2216, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBh

[1367] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(8):??, June 6, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Domingo-Ferrer:2007:ASC

[1368] Josep Domingo-Ferrer, Joachim Posegga, Francesc Sebé, and Vicenç Torra. Advances in smart cards. Computer Networks (Amsterdam, Netherlands: 1999), 51(9):2219–2222, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Preneel:2007:SRD

[1369] Bart Preneel. A survey of recent developments in cryptographic algorithms for smart cards. Computer Networks (Amsterdam, Netherlands: 1999), 51 (9):2223–2233, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lu:2007:NSC

[1370] HongQian Karen Lu. Network smart card review and analysis. Computer Networks (Amsterdam, Netherlands: 1999), 51(9):2234–2248, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Torres:2007:ANS

[1371] Joaquin Torres, Antonio Izquierdo, and Jose Maria Sierra. Advances in network smart cards authentication. Computer Networks (Amsterdam, Netherlands: 1999), 51(9):2249–2261, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Munilla:2007:HMF

[1372] J. Munilla and A. Peinado. HB-MP: A further step in the HB-family of lightweight authentication protocols. Computer Networks (Amsterdam, Netherlands: 1999), 51(9):2262–2267, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Solanas:2007:DAS

[1373] Agusti Solanas, Josep Domingo-Ferrer, Antoni Martínez-Ballesté, and Vanesa Daza. A distributed architecture for scalable private RFID tag identification. Computer Networks (Amsterdam, Netherlands: 1999), 51(9): 2268–2279, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cheng:2007:SEI

[1374] Ting-Fang Cheng, Jung-San Lee, and Chin-Chen Chang. Security enhancement of an IC-card-based remote login mechanism. Computer Networks (Amsterdam, Netherlands: 1999), 51 (9):2280–2287, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lopez:2007:SCB

[1375] Andrés Marín López, Daniel Díaz Sánchez, Florina Almenárez, Carlos García Rubio, and Celeste Campo. Smart card-based agents for fair nonrepudiation. Computer Networks (Amsterdam, Netherlands: 1999), 51(9): 2288–2298, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sebe:2007:SMO

[1376] Francesc Sebé, Alexandre Viejo, and Josep Domingo-Ferrer. Secure manyto-one symbol transmission for implementation on smart cards. Computer Networks (Amsterdam, Netherlands: 1999), 51(9):2299–2307, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Karetsos:2007:TWB

[1377] George Karetsos, Angelos Rouskas, Bijan Jabbari, and Bernhard Walke. Topics in wireless broadband systems. Computer Networks (Amsterdam, Netherlands: 1999), 51(9): 2308–2310, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cicconetti:2007:DPA

[1378] Claudio Cicconetti, Luciano Lenzini, Enzo Mingozzi, and Giovanni Stea. Design and performance analysis of the Real-Time HCCA scheduler for IEEE 802.11e WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 51 (9):2311–2325, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Perez-Costa:2007:DPS

[1379] Xavier Pérez-Costa, Daniel Camps-Mur, and Albert Vidal. On distributed power saving mechanisms of wireless LANs 802.11e U-APSD vs 802.11 power save mode. Computer Networks (Amsterdam, Netherlands: 1999), 51 (9):2326–2344, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vassis:2007:PAI

[1380] D. Vassis and G. Kormentzas. Performance analysis of IEEE 802.11 ad hoc networks in the presence of hidden terminals. Computer Networks (Amsterdam, Netherlands: 1999), 51(9): 2345–2352, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Max:2007:SSI

[1381] Sebastian Max, Guido R. Hiertz, Erik Weiss, Dee Denteneer, and Bernhard H. Walke. Spectrum sharing in IEEE 802.11s wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(9): 2353–2367, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hosein:2007:QSU

[1382] Patrick Hosein. QoS scheduling of uplink resources in OFDMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(9): 2368–2378, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Leoleis:2007:FME

[1383] Georgios A. Leoleis and Iakovos S. Venieris. Fast MIPv6 extensions supporting seamless multicast handovers. Computer Networks (Amsterdam, Netherlands: 1999), 51(9): 2379–2396, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Karyotis:2007:RBA

[1384] Vasileios Karyotis and Symeon Papavassiliou. Risk-based attack strategies for mobile ad hoc networks under probabilistic attack modeling framework. Computer Networks (Amsterdam, Netherlands: 1999), 51(9): 2397–2410, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBi

[1385] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(9):??, June 20, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Misic:2007:IMS

[1386] Jelena Mišić and Carol J. Fung. The impact of master–slave bridge access mode on the performance of multi-cluster 802.15.4 network. Computer Networks (Amsterdam, Netherlands: 1999), 51(10):2411–2449, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bejerano:2007:ELB

[1387] Yigal Bejerano, Seung-Jae Han, and Amit Kumar. Efficient loadbalancing routing for wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(10): 2450–2466, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mao:2007:PLE

[1388] Guoqiang Mao, Brian D. O. Anderson, and Barış Fidan. Path loss exponent estimation for wireless sensor network localization. Computer Networks (Amsterdam, Netherlands: 1999), 51(10): 2467–2483, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Popescu:2007:DIS

[1389] Bogdan C. Popescu, Bruno Crispo, Andrew S. Tanenbaum, and Arno Bakker. Design and implementation of a secure wide-area object middleware. Computer Networks (Amsterdam, Netherlands: 1999), 51(10):2484–2513, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Leng:2007:MAM

[1390] Supeng Leng, Liren Zhang, Huirong Fu, and Jianjun Yang. Mobility analysis of mobile hosts with random walking in ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(10):2514–2528, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mao:2007:WSN

[1391] Guoqiang Mao, Barış Fidan, and Brian D. O. Anderson. Wireless sensor network localization techniques. Computer Networks (Amsterdam, Netherlands: 1999), 51(10):2529–2553, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lu:2007:UFB

[1392] Ning Lu and John Bigham. On utilityfair bandwidth adaptation for multiclass traffic QoS provisioning in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (10):2554–2564, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Todtmann:2007:ADP

[1393] Birger Tödtmann and Erwin P. Rathgeb. Anticipatory distributed packet filter configurations for carriergrade IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (10):2565–2579, July 11, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Geneiatakis:2007:FPS

[1394] Dimitris Geneiatakis, Georgios Kambourakis, Costas Lambrinoudakis, Tasos Dagiuklas, and Stefanos Gritzalis. A framework for protecting a SIPbased infrastructure against malformed message attacks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (10):2580–2593, July 11, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shah:2007:HAP

[1395] Vasudev Shah, Ece Gelal, and Srikanth V. Krishnamurthy. Handling asymmetry in power heterogeneous ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (10):2594–2615, July 11, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Duan:2007:DCS

[1396] Zhenhai Duan, Yingfei Dong, and Kartik Gopalan. DMTP: Controlling spam through message delivery differentiation. Computer Networks (Amsterdam, Netherlands: 1999), 51(10): 2616–2630, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Khayat:2007:MLV

[1397] I. El Khayat, P. Geurts, and G. Leduc.

Machine-learnt versus analytical models of TCP throughput. Computer Networks (Amsterdam, Netherlands: 1999), 51(10):2631–2644, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Blefari-Melazzi:2007:ACP

[1398] N. Blefari-Melazzi, D. Di Sorte, M. Femminella, and G. Reali. Autonomic control and personalization of a wireless access network. Computer Networks (Amsterdam, Netherlands: 1999), 51(10):2645–2676, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Belenky:2007:DPM

[1399] Andrey Belenky and Nirwan Ansari.

On deterministic packet marking.

Computer Networks (Amsterdam, Netherlands: 1999), 51(10):2677–2700, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Choi:2007:QSP

[1400] Baek-Young Choi, Sue Moon, Rene Cruz, Zhi-Li Zhang, and Christophe Diot. Quantile sampling for practical delay monitoring in Internet backbone networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(10): 2701–2716, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kuzmanovic:2007:RCC

[1401] A. Kuzmanovic and E. W. Knightly. Receiver-centric congestion control with a misbehaving receiver: Vulnerabilities and end-point solutions. Computer Networks (Amsterdam, Netherlands: 1999), 51(10):2717–2737, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Marsan:2007:FAC

[1402] M. Ajmone Marsan, C. Casetti, G. Mardente, and M. Mellia. A framework for admission control and path allocation in DiffServ networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(10):2738–2752, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2007:IRC

[1403] Kang-Won Lee, Suresh Chari, Anees Shaikh, Sambit Sahu, and Pau-Chen Cheng. Improving the resilience of content distribution networks to large scale distributed denial of service attacks. Computer Networks (Amsterdam, Netherlands: 1999), 51(10): 2753–2770, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2007:CTA

[1404] Sun-Myeng Kim and Young-Jong Cho.

Channel time allocation scheme based on feedback information in IEEE 802.11e wireless LANs. Computer Networks (Amsterdam, Netherlands: 1999), 51(10):2771–2787, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhu:2007:ETN

[1405] Yujie Zhu, Aravind Velayutham, Oyebamiji Oladeji, and Raghupathy Sivakumar. Enhancing TCP for networks with guaranteed bandwidth services. Computer Networks (Amsterdam, Netherlands: 1999), 51(10): 2788–2804, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Huang:2007:SMD

[1406] Chun-Ying Huang, Yun-Peng Chiu, Kuan-Ta Chen, and Chin-Laung Lei. Secure multicast in dynamic environments. Computer Networks (Amsterdam, Netherlands: 1999), 51(10): 2805–2817, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jin:2007:NAE

[1407] Shudong Jin and Hongbo Jiang. Novel approaches to efficient flooding search in peer-to-peer networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(10):2818–2832, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Elias:2007:NAD

[1408] J. Elias, F. Martignon, A. Capone, and G. Pujolle. A new approach to dynamic bandwidth allocation in Quality of Service networks: Performance and bounds. Computer Networks (Amsterdam, Netherlands: 1999), 51(10): 2833–2853, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ramasubramanian:2007:LTD

[1409] Srinivasan Ramasubramanian, Mithun Harkara, and Marwan Krunz. Linear time distributed construction of colored trees for disjoint multipath routing. Computer Networks (Amsterdam, Netherlands: 1999), 51(10): 2854–2866, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2007:PME

[1410] Xiaolan Zhang, Giovanni Neglia, Jim Kurose, and Don Towsley. Performance modeling of epidemic routing. Computer Networks (Amsterdam, Netherlands: 1999), 51(10):2867–2891, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vidales:2007:MEE

[1411] Pablo Vidales, Carlos J. Bernardos, Ignacio Soto, David Cottingham, Javier Baliosian, and Jon Crowcroft. MIPv6 experimental evaluation using overlay networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(10): 2892–2915, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBj

[1412] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(10):??, July 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Smaragdakis:2007:MRD

[1413] Georgios Smaragdakis, Nikolaos Laoutaris, Azer Bestavros, Ibrahim Matta, and Ioannis Stavrakakis. Mistreatmentresilient distributed caching. Computer Networks (Amsterdam, Netherlands: 1999), 51(11):2917–2937, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chiu:2007:FTC

[1414] Dah Ming Chiu and Adrian Sai-Wah Tam. Fairness of traffic controls for inelastic flows in the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 51(11):2938–2957, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wong:2007:NMA

[1415] Eric W. M. Wong, Andrew Zalesky, Zvi Rosberg, and Moshe Zukerman. A new method for approximating blocking probability in overflow loss networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(11): 2958–2975, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gupta:2007:SUW

[1416] Ashima Gupta and Prasant Mohapatra. A survey on ultra wide band medium access control schemes. Computer Networks (Amsterdam, Netherlands: 1999), 51(11):2976–2993, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Alves:2007:RTC

[1417] Mário Alves and Eduardo Tovar. Real-time communications over wired/ wireless PROFIBUS networks supporting inter-cell mobility. Computer Networks (Amsterdam, Netherlands: 1999), 51(11):2994–3012, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kuran:2007:SEB

[1418] Mehmet S. Kuran and Tuna Tugcu. A survey on emerging broadband wireless access technologies. Computer Networks (Amsterdam, Netherlands: 1999), 51(11):3013–3046, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xiong:2007:STA

[1419] Lixiang Xiong and Guoqiang Mao. Saturated throughput analysis of IEEE 802.11e EDCA. Computer Networks (Amsterdam, Netherlands: 1999), 51 (11):3047–3068, August 8, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lao:2007:TGT

[1420] Li Lao, Jun-Hong Cui, and Mario Gerla. Tackling group-to-tree matching in large scale group communications. Computer Networks (Amsterdam, Netherlands: 1999), 51(11): 3069–3089, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2007:ETC

[1421] Hualiang Chen, Zhongxin Liu, Zengqiang Chen, and Zhuzhi Yuan. Extending TCP congestion control to multicast. Computer Networks (Amsterdam, Netherlands: 1999), 51(11): 3090–3109, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vanhatupa:2007:ETE

[1422] Timo Vanhatupa, Marko Hännikäinen, and Timo D. Hämäläinen. Evaluation of throughput estimation models and algorithms for WLAN frequency planning. Computer Networks (Amsterdam, Netherlands: 1999), 51(11): 3110–3124, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lufei:2007:EAQ

[1423] Hanping Lufei and Weisong Shi.

Energy-aware QoS for application sessions across multiple protocol domains in mobile computing. Computer Networks (Amsterdam, Netherlands: 1999), 51(11):3125–3141, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Munaretto:2007:RQS

[1424] Anelise Munaretto and Mauro Fonseca.

Routing and quality of service support for mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(11):3142–3156, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2007:CPA

[1425] Xiaolong Li and Aaron D. Striegel.

A case for Passive Application Layer Multicast. Computer Networks (Amsterdam, Netherlands: 1999), 51(11): 3157–3171, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lucio:2007:GLS

[1426] Gilberto Flores Lucio, Martin J. Reed, and Ian D. Henning. Guided local search as a network planning algorithm that incorporates uncertain traffic demands. Computer Networks (Amsterdam, Netherlands: 1999), 51(11): 3172–3196, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hassen:2007:KMC

[1427] H. Ragab Hassen, A. Bouabdallah, H. Bettahar, and Y. Challal. Key management for content access control in a hierarchy. Computer Networks (Amsterdam, Netherlands: 1999), 51(11): 3197–3219, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lai:2007:TSP

[1428] Yuan-Cheng Lai. Two schedulers to provide delay proportion and reduce queueing delay simultaneously. Computer Networks (Amsterdam, Netherlands: 1999), 51(11):3220–3231, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kiesel:2007:MPE

[1429] Sebastian Kiesel and Michael Scharf. Modeling and performance evaluation of transport protocols for firewall control. Computer Networks (Amsterdam, Netherlands: 1999), 51(11):3232– 3251, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Misic:2007:AEC

[1430] Jelena Mišić. Algorithm for equalization of cluster lifetimes in a multilevel Beacon enabled 802.15.4 sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 51(11): 3252–3264, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cai:2007:CAM

[1431] Lin X. Cai, Xuemin (Sherman) Shen, Jon W. Mark, and Lin Cai. Capacity analysis and MAC enhancement for UWB broadband wireless access networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(11): 3265–3277, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2007:LFC

[1432] Zhenjiang Li and J. J. Garcia-Luna-Aceves. Loop-free constrained path computation for hop-by-hop QoS routing. Computer Networks (Amsterdam, Netherlands: 1999), 51(11):3278– 3293, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shelford:2007:AOR

[1433] Steven Shelford, Gholamali C. Shoja, and Eric G. Manning. Achieving optimal revenues in dynamically priced network services with QoS guarantees. Computer Networks (Amsterdam, Netherlands: 1999), 51(11):3294– 3304, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Niyato:2007:QAB

[1434] Dusit Niyato and Ekram Hossain. QoSaware bandwidth allocation and admission control in IEEE 802.16 broadband wireless access networks: A non-cooperative game theoretic approach. Computer Networks (Amsterdam, Netherlands: 1999), 51(11): 3305–3321, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBk

[1435] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(11):??, August 8, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Habib:2007:IAQ

[1436] Ahsan Habib and John Chuang. Improving application QoS with residential multihoming. Computer Networks (Amsterdam, Netherlands: 1999), 51 (12):3323–3337, August 22, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shin:2007:MIA

[1437] Soo Young Shin, Hong Seong Park, and Wook Hyun Kwon. Mutual interference analysis of IEEE 802.15.4 and IEEE 802.11b. Computer Networks (Amsterdam, Netherlands: 1999), 51(12): 3338–3353, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wuu:2007:LPF

[1438] Lih-Chyau Wuu, Tzong-Jye Liu, and Kuo-Ming Chen. A longest prefix first search tree for IP lookup. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3354–3367, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2007:QMH

[1439] Jiann-Liang Chen, Ming-Chiao Chen, and Yi-Ru Chian. QoS management in heterogeneous home networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3368–3379, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Naiksatam:2007:EBU

[1440] Sumit Naiksatam, Stephen A. Chiappari, and Silvia Figueira. Efficient bandwidth utilization in LambdaGrids using pricing incentives. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3380–3391, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sridhara:2007:RPS

[1441] Vinay Sridhara and Stephan Bohacek.

Realistic propagation simulation of urban mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3392–3412, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lopez:2007:GTC

[1442] Luis López, Antonio Fernández, and Vicent Cholvi. A game theoretic comparison of TCP and digital fountain based protocols. Computer Networks (Amsterdam, Netherlands: 1999), 51 (12):3413–3426, August 22, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Balasubramanian:2007:AEE

[1443] Saravanan Balasubramanian and Demet Aksoy. Adaptive energy-efficient registration and online scheduling for asymmetric wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3427–3447, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Patcha:2007:OAD

[1444] Animesh Patcha and Jung-Min Park. An overview of anomaly detection techniques: Existing solutions and latest technological trends. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3448–3470, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Antonatos:2007:DAH

[1445] S. Antonatos, P. Akritidis, E. P. Markatos, and K. G. Anagnostakis. Defending against hitlist worms using network address space randomization. Computer Networks (Amsterdam, Netherlands: 1999), 51(12): 3471–3490, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Karapantazis:2007:CAC

[1446] Stylianos Karapantazis and Fotini-Niovi Pavlidou. Call admission control in multiservice high altitude platform (HAP) W-CDMA cellular systems. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3491– 3506, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zachariadis:2007:DMT

[1447] Grigorios Zachariadis and Javier A.

Barria. Demand management for telecommunications services. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3507–3524, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mukherjee:2007:SSS

[1448] Ritesh Mukherjee and J. William Atwood. Scalable solutions for secure group communications. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3525–3548, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bouras:2007:AQS

[1449] Christos Bouras and Afrodite Sevasti.

An analytical QoS service model for delay-based differentiation. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3549–3563, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2007:QAD

[1450] Yang Wang, Chuang Lin, Quan-Lin Li, and Yuguang Fang. A queueing analysis for the denial of service (DoS) attacks in computer networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(12): 3564–3573, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tam:2007:QSR

[1451] W.-Y. Tam, K.-S. Lui, S. Uludag, and K. Nahrstedt. Quality-of-service routing with path information aggregation. Computer Networks (Amsterdam, Netherlands: 1999), 51(12): 3574–3594, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pucha:2007:PIT

[1452] Himabindu Pucha, Saumitra M. Das, and Y. Charlie Hu. The performance impact of traffic patterns on routing protocols in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(12): 3595–3616, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lam:2007:CUI

[1453] Patrick P. Lam, Soung C. Liew, and Jack Y. B. Lee. Cellular universal IP for nested network mobility. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3617–3631, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Saxena:2007:TCP

[1454] Nitesh Saxena, Gene Tsudik, and Jeong Hyun Yi. Threshold cryptography in P2P and MANETs: The case of access control. Computer Networks (Amsterdam, Netherlands: 1999), 51 (12):3632–3649, August 22, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nadarajah:2007:CPG

[1455] Saralees Nadarajah and Samuel Kotz.

On the convolution of Pareto and gamma distributions. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):3650–3654, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Eun:2007:PAT

[1456] Do Young Eun and Xinbing Wang. Performance analysis of TCP/AQM with generalized AIMD under intermediate buffer sizes. Computer Networks (Amsterdam, Netherlands: 1999), 51(12): 3655–3671, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBl

[1457] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(12):??, August 22, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Balamash:2007:PAC

[1458] Abdullah Balamash, Marwan Krunz, and Philippe Nain. Performance analysis of a client-side caching/prefetching system for Web traffic. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3673–3692, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Das:2007:SRB

[1459] Saumitra M. Das, Himabindu Pucha, and Y. Charlie Hu. On the scalability of rendezvous-based location services for geographic wireless ad hoc routing. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3693– 3714, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gouda:2007:SAP

[1460] Mohamed G. Gouda, Alex X. Liu, Lok M. Leung, and Mohamed A. Alam. SPP: An anti-phishing single password protocol. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3715– 3726, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vural:2007:PDM

[1461] Serdar Vural and Eylem Ekici. Probability distribution of multi-hop-distance in one-dimensional sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3727– 3749, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Khalil:2007:LWD

[1462] Issa Khalil, Saurabh Bagchi, and Ness B. Shroff. Lite Worp : Detection and isolation of the wormhole attack in static multihop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3750– 3772, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lestas:2007:ACP

[1463] Marios Lestas, Andreas Pitsillides, Petros Ioannou, and George Hadjipollas. Adaptive congestion protocol: A congestion control protocol with learning capability. Computer Networks (Amsterdam, Netherlands: 1999), 51(13): 3773–3798, September 12, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mondal:2007:ITP

[1464] Sakib A. Mondal and Faisal B. Luqman. Improving TCP performance over wired–wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3799–3811, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Choi:2007:APP

[1465] Baek-Young Choi, Sue Moon, Zhi-Li Zhang, Konstantina Papagiannaki, and Christophe Diot. Analysis of point-to-point packet delay in an operational network. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3812– 3827, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2007:AOF

[1466] Zhi Li, Lihua Yuan, Prasant Mohapatra, and Chen-Nee Chuah. On the analysis of overlay failure detection and recovery. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3828– 3843, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Downey:2007:TSC

[1467] Allen B. Downey. TCP self-clocking and bandwidth sharing. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3844–3863, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bremler-Barr:2007:PBA

[1468] Anat Bremler-Barr, Nir Halachmi, and Hanoch Levy. Protecting bursty applications against traffic aggressiveness. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3864– 3877, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bouabdallah:2007:CES

[1469] Nizar Bouabdallah and Harry Perros. Cost-effective single-hub WDM ring networks: A proposal and analysis. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3878– 3901, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mas:2007:PBA

[1470] Ignacio Más and Gunnar Karlsson. Probe-based admission control for a differentiated-services Internet. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3902– 3918, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhao:2007:ROR

[1471] Yong-Xiang Zhao and Chang-Jia Chen. A redundant overbooking reservation algorithm for OBS/OPS networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3919– 3934, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Patcha:2007:NAD

[1472] Animesh Patcha and Jung-Min Park. Network anomaly detection with incomplete audit data. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):3935–3955, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBm

[1473] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(13):??, September 12, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

He:2007:PLT

[1474] Qi He, Constantinos Dovrolis, and Mostafa Ammar. On the predictability of large transfer TCP throughput. Computer Networks (Amsterdam, Netherlands: 1999), 51(14):3959– 3977, October 10, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2007:DSM

[1475] Yang Yang, Feiyi Huang, Xiaohu Ge, Xiaodong Zhang, Xuanye Gu, Mohsen Guizani, and Hsiao-Hwa Chen. Double sense multiple access for wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(14): 3978–3988, October 10, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Randaccio:2007:GMR

[1476] L. Sanna Randaccio and L. Atzori. Group multicast routing problem: A genetic algorithms based approach. Computer Networks (Amsterdam, Netherlands: 1999), 51(14): 3989–4004, October 10, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sarkar:2007:MLP

[1477] Mahasweta Sarkar and Rene L. Cruz. A MAC layer power management scheme for efficient energy delay tradeoff in a WLAN. Computer Networks (Amsterdam, Netherlands: 1999), 51(14): 4005–4031, October 10, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

deCola:2007:PAC

[1478] Tomaso de Cola, Harald Ernst, and Mario Marchese. Performance analysis of CCSDS File Delivery Protocol and erasure coding techniques in deep space environments. Computer Networks (Amsterdam, Netherlands: 1999), 51 (14):4032–4049, October 10, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2007:IVF

[1479] Wenji Wu and Matt Crawford. Interactivity vs. fairness in networked Linux systems. Computer Networks (Amsterdam, Netherlands: 1999), 51(14): 4050–4069, October 10, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chang:2007:TTA

[1480] Chih-Yung Chang and Chao-Tsun Chang. TARP: A traffic-aware restructuring protocol for Bluetooth radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(14):4070– 4091, October 10, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Baldi:2007:SSE

[1481] Mario Baldi, Guido Marchetto, and Yoram Ofek. A scalable solution for engineering streaming traffic in the future Internet. Computer Networks (Amsterdam, Netherlands: 1999), 51(14): 4092–4111, October 10, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hanbali:2007:IMP

[1482] A. Al Hanbali, A. A. Kherani, R. Groenevelt, P. Nain, and E. Altman. Impact of mobility on the performance of relaying in ad hoc networks – Extended version. Computer Networks (Amsterdam, Netherlands: 1999), 51 (14):4112–4130, October 10, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ta:2007:TPA

[1483] Duong Nguyen Binh Ta and Suiping Zhou. A two-phase approach to interactivity enhancement for large-scale distributed virtual environments. Computer Networks (Amsterdam, Netherlands: 1999), 51(14):4131–4152, October 10, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gunes:2007:MSO

[1484] Mehmet H. Gunes, Sevcan Bilir, Kamil Sarac, and Turgay Korkmaz. A measurement study on overhead distribution of value-added Internet services. Computer Networks (Amsterdam, Netherlands: 1999), 51(14):4153– 4173, October 10, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bar:2007:GDP

[1485] Sagy Bar, Mira Gonen, and Avishai Wool. A geographic directed preferential Internet topology model. Computer Networks (Amsterdam, Netherlands: 1999), 51(14):4174–4188, October 10, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rakotoarivelo:2007:SDM

[1486] Thierry Rakotoarivelo, Patrick Sénac, Aruna Seneviratne, and Michel Diaz. SPAD: A distributed middleware architecture for QoS enhanced alternate path discovery. Computer Networks (Amsterdam, Netherlands: 1999), 51 (14):4189–4211, October 10, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Norden:2007:SRP

[1487] Samphel Norden and Katherine Guo. Support for resilient Peer-to-Peer gaming. Computer Networks (Amsterdam, Netherlands: 1999), 51(14):4212– 4233, October 10, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBn

[1488] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(14):??, October 10, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kashyap:2007:ITC

[1489] Abhishek Kashyap, Kwangil Lee, Mehdi Kalantari, Samir Khuller, and Mark Shayman. Integrated topology control and routing in wireless optical mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (15):4237–4251, October 24, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2007:AGA

[1490] Shigang Chen, Yibei Ling, Randy Chow, and Ye Xia. AID: A global anti-DoS service. Computer Networks (Amsterdam, Netherlands: 1999), 51(15): 4252–4269, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pavon-Marino:2007:PDB

[1491] P. Pavon-Mariño, J. Garcia-Haro, and A. Jajszczyk. Parallel desynchronized block matching: A feasible scheduling algorithm for the input-buffered wavelength-routed switch. Computer Networks (Amsterdam, Netherlands: 1999), 51(15):4270–4283, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Krishnamurthy:2007:SLI

[1492] Vaishnavi Krishnamurthy, Michalis Faloutsos, Marek Chrobak, Jun-Hong Cui, Li Lao, and Allon G. Percus. Sampling large Internet topologies for simulation purposes. Computer Networks (Amsterdam, Netherlands: 1999), 51 (15):4284–4302, October 24, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2007:SKD

[1493] Weichao Wang and Tylor Stransky.

Stateless key distribution for secure intra and inter-group multicast in mobile wireless network. Computer Networks (Amsterdam, Netherlands: 1999), 51 (15):4303–4321, October 24, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lootah:2007:TTB

[1494] Wesam Lootah, William Enck, and Patrick McDaniel. TARP: Ticket-based address resolution protocol. Computer Networks (Amsterdam, Netherlands: 1999), 51(15):4322–4337, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2007:AMD

[1495] Dongwook Lee, Gayathri Chandrasekaran, Mukundan Sridharan, and Prasun Sinha. Association management for data dissemination over wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(15):4338–4355, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Papapetrou:2007:DDR

[1496] E. Papapetrou, S. Karapantazis, and F.-N. Pavlidou. Distributed ondemand routing for LEO satellite systems. Computer Networks (Amsterdam, Netherlands: 1999), 51(15):4356– 4376, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Papadimitriou:2007:SCC

[1497] Panagiotis Papadimitriou and Vassilis Tsaoussidis. SSVP: A congestion control scheme for real-time video streaming. Computer Networks (Amsterdam, Netherlands: 1999), 51(15):4377– 4395, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kliazovich:2007:PIW

[1498] Dzmitry Kliazovich, Fabrizio Granelli, and Mario Gerla. Performance improvement in wireless networks using cross-layer ARQ. Computer Networks (Amsterdam, Netherlands: 1999), 51 (15):4396–4411, October 24, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cusani:2007:NMP

[1499] Roberto Cusani, Tiziano Inzerilli, and Luca Valentini. Network monitoring and performance evaluation in a 3.5G network. Computer Networks (Amsterdam, Netherlands: 1999), 51(15): 4412–4420, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Turner:2007:RAO

[1500] Kenneth J. Turner and Koon Leai Larry Tan. A rigorous approach to orchestrating grid services. Computer Networks (Amsterdam, Netherlands: 1999), 51(15):4421–4441, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lim:2007:ILL

[1501] Hyuk Lim and Jennifer C. Hou. Identifying lossy links in wired/wireless networks by exploiting sparse characteristics. Computer Networks (Amsterdam, Netherlands: 1999), 51(15):4442– 4459, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cholvi:2007:SFN

[1502] Vicent Cholvi and Juan Echagüe. Stability of FIFO networks under adversarial models: State of the art. Computer Networks (Amsterdam, Netherlands: 1999), 51(15):4460–4474, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2007:STF

[1503] Lijun Wang, Lin Cai, Xinzhi Liu, and Xuemin (Sherman) Shen. Stability and TCP-friendliness of AIMD/RED systems with feedback delays. Computer Networks (Amsterdam, Netherlands: 1999), 51(15):4475–4491, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guitart:2007:DOC

[1504] Jordi Guitart, David Carrera, Vicenç Beltran, Jordi Torres, and Eduard Ayguadé. Designing an overload control strategy for secure e-commerce applications. Computer Networks (Amsterdam, Netherlands: 1999), 51(15): 4492–4510, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBo

[1505] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(15):??, October 24, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Courtenage:2007:IWC

[1506] Simon Courtenage, David Lewis, and Thanassis Tiropanis. Innovations in Web communications infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4511– 4513, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Michlmayr:2007:UTC

[1507] Elke Michlmayr, Arno Pany, and Gerti Kappel. Using taxonomies for contentbased routing with ants. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4514–4528, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vazquez:2007:MHB

[1508] Juan Ignacio Vazquez and Diego López de Ipiña. mRDP: An HTTP-based lightweight semantic discovery protocol. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4529– 4542, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gu:2007:IRS

[1509] Tao Gu, Hung Keng Pung, and Daqing Zhang. Information retrieval in schema-based P2P systems using onedimensional semantic space. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4543–4560, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2007:GUD

[1510] Frank Zhigang Wang, Na Helian, Sining Wu, Yuhui Deng, Vineet Khare, and Michael Parker. GridJet: An underlying data-transporting protocol for accelerating Web communications. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4561– 4573, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Benz:2007:SCH

[1511] Dominik Benz, Karen H. L. Tso, and Lars Schmidt-Thieme. Supporting collaborative hierarchical classification: Bookmarks as an example. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4574– 4585, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fenner:2007:SEG

[1512] Trevor Fenner, Mark Levene, George Loizou, and George Roussos. A stochastic evolutionary growth model for social networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4586–4595, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kormentzas:2007:MCN

[1513] George Kormentzas and Thomas Magedanz. Middleware challenges for next generation networks and services. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4596– 4598, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Boonma:2007:BBI

[1514] Pruet Boonma and Junichi Suzuki. BiSNET: A biologically-inspired middleware architecture for self-managing wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4599–4616, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Aoul:2007:FLB

[1515] Yassine Hadjadj Aoul, Ahmed Mehaoua, and Charalabos Skianis. A fuzzy logicbased AQM for real-time traffic over Internet. Computer Networks (Amsterdam, Netherlands: 1999), 51(16): 4617–4633, November 14, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kangasharju:2007:XMM

[1516] Jaakko Kangasharju, Tancred Lindholm, and Sasu Tarkoma. XML messaging for mobile devices: From requirements to implementation. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4634–4654, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jorstad:2007:PUF

[1517] Ivar Jørstad and Do Thanh van.

Personalised ubiquitous file access with XML Web Services. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4655–4668, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Borcoci:2007:SIA

[1518] Eugen Borcoci, Georgios Kormentzas, Abolghasem (Hamid) Asgari, and Toufik Ahmed. Service invocation admission control algorithm for multidomain IP environments. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4669–4678, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lioudakis:2007:MAP

[1519] Georgios V. Lioudakis, Eleftherios A. Koutsoloukas, Nikolaos L. Dellas, Nikolaos Tselikas, Sofia Kapellaki, George N. Prezerakos, Dimitra I. Kaklamani, and Iakovos S. Venieris. A middleware architecture for privacy protection. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4679– 4696, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sher:2007:VAC

[1520] Muhammad Sher and Thomas Magedanz. A vulnerabilities analysis and corresponding middleware security extensions for securing NGN applications. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):4697– 4709, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBp

[1521] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(16):??, November 14, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jayashree:2007:TEL

[1522] S. Jayashree and C. Siva Ram Murthy. Towards estimating lifetime of ad hoc wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 51 (17):4711–4726, December 5, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pour:2007:HGK

[1523] Alireza Nemaney Pour, Kazuya Kumekawa, Toshihiko Kato, and Shuichi Itoh. A hierarchical group key management scheme for secure multicast increasing efficiency of key distribution in leave operation. Computer Networks (Amsterdam, Netherlands: 1999), 51(17):4727–4743, December 5, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xu:2007:MST

[1524] Lisong Xu and Josh Helzer. Media streaming via TFRC: An analytical study of the impact of TFRC on userperceived media quality. Computer Networks (Amsterdam, Netherlands: 1999), 51(17):4744–4764, December 5, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Correia:2007:TPC

[1525] Luiz H. A. Correia, Daniel F. Macedo, Aldri L. dos Santos, Antonio A. F. Loureiro, and José Marcos S. Nogueira. Transmission power control techniques for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(17):4765–4779, December 5, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xia:2007:MSP

[1526] Jianhong Xia, Lixin Gao, and Teng Fei. A measurement study of persistent forwarding loops on the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 51(17):4780– 4796, December 5, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gaspary:2007:FSP

[1527] Luciano Paschoal Gaspary, Marinho P. Barcellos, André Detsch, and Rodolfo S. Antunes. Flexible security in peer-to-peer applications: Enabling new opportunities beyond file sharing. Computer Networks (Amsterdam, Netherlands: 1999), 51(17): 4797–4815, December 5, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ilie:2007:SMG

[1528] Dragos Ilie and Adrian Popescu. Statistical models for Gnutella signaling traffic. Computer Networks (Amsterdam, Netherlands: 1999), 51(17):4816– 4835, December 5, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chu:2007:RMB

[1529] Jian Chu and Chin-Tau Lea. A restorable MPLS-based hose-model VPN network. Computer Networks (Amsterdam, Netherlands: 1999), 51 (17):4836–4848, December 5, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Melia:2007:TIC

[1530] Telemaco Melia, Antonio de la Oliva, Albert Vidal, Ignacio Soto, Daniel Corujo, and Rui Aguiar. Toward IP converged heterogeneous mobility: A network controlled approach. Computer Networks (Amsterdam, Netherlands: 1999), 51(17):4849–4866, December 5, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Huynh:2007:MEN

[1531] Minh Huynh and Prasant Mohapatra. Metropolitan Ethernet Network: A move from LAN to MAN. Computer Networks (Amsterdam, Netherlands: 1999), 51(17):4867–4894, December 5, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2007:EBq

[1532] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 51(17):??, December 5, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pitsillides:2007:MHC

[1533] Andreas Pitsillides and Christophoros Christophorou. MBMS Handover control: A new approach for efficient handover in MBMS enabled 3G cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(18): 4897–4918, December 19, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yi:2007:FHI

[1534] Yung Yi and Sanjay Shakkottai.

FluNet: A hybrid Internet simulator for fast queue regimes. Computer Networks (Amsterdam, Netherlands: 1999), 51(18):4919–4937, December 19, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Daza:2007:CTM

[1535] Vanesa Daza, Javier Herranz, Paz Morillo, and Carla Ràfols. Cryptographic techniques for mobile ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 51(18): 4938–4950, December 19, 2007. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kounavis:2007:TSP

[1536] M. E. Kounavis, A. Kumar, R. Yavatkar, and H. Vin. Two stage packet classification using most specific filter matching and transport level sharing. Computer Networks (Amsterdam, Netherlands: 1999), 51(18):4951– 4978, December 19, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dramitinos:2007:AMA

[1537] Manos Dramitinos, George D. Stamoulis, and Costas Courcoubetis. An auction mechanism for allocating the bandwidth of networks to their users. Computer Networks (Amsterdam, Netherlands: 1999), 51(18):4979– 4996, December 19, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ssu:2007:IDT

[1538] Kuo-Feng Ssu, Chih-Hsun Chou, Wei-Tong Wang, and Shin-Hung Chung. Improving data transmission with helping nodes for geographical ad hoc routing. Computer Networks (Amsterdam, Netherlands: 1999), 51(18):4997– 5010, December 19, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Huang:2007:SCP

[1539] Shu Huang and Rudra Dutta. Spare capacity provisioning for quasi-static traffic. Computer Networks (Amsterdam, Netherlands: 1999), 51(18):5011– 5035, December 19, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lu:2007:RED

[1540] Kejie Lu, Dapeng Wu, Jieyan Fan, Sinisa Todorovic, and Antonio Nucci. Robust and efficient detection of DDoS attacks for large-scale Internet. Computer Networks (Amsterdam, Netherlands: 1999), 51(18):5036–5056, December 19, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Abu-Ghazaleh:2008:GE

[1541] Nael Abu-Ghazaleh, Enrique Alba, Carla-Fabiana Chiasserini, and Renato LoCigno. Guest Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 52(1):1–3, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Charlet:2008:EEM

[1542] Damien Charlet, Valérie Issarny, and Rafik Chibout. Energy-efficient middleware-layer multi-radio networking: An assessment in the area of service discovery. Computer Networks (Amsterdam, Netherlands: 1999), 52 (1):4–24, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhu:2008:RLT

[1543] Junhua Zhu, Ka-Lok Hung, Brahim Bensaou, and Farid Nait-Abdesselam. Rate-lifetime tradeoff for reliable communication in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(1): 25–43, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Avin:2008:PCR

[1544] Chen Avin and Bhaskar Krishnamachari. The power of choice in random walks: An empirical study. Computer Networks (Amsterdam, Netherlands: 1999), 52(1):44–60, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hu:2008:STN

[1545] Chunyu Hu, Hwangnam Kim, and Jennifer C. Hou. Short-term non-uniform access in IEEE 802.11-compliant WLANs: A study on its impact on the saturation performance. Computer Networks (Amsterdam, Netherlands: 1999), 52(1):61–76, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Razafindralambo:2008:DDP

[1546] Tahiry Razafindralambo, Isabelle Guérin-lassous, Luigi Iannone, and Serge Fdida. Dynamic and distributed packet aggregation to solve the performance anomaly in 802.11 wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(1): 77–95, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sayenko:2008:SSI

[1547] Alexander Sayenko, Olli Alanen, and Timo Hämäläinen. Scheduling solution for the IEEE 802.16 base station. Computer Networks (Amsterdam, Netherlands: 1999), 52(1):96–115, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Stuedi:2008:MCT

[1548] Patrick Stuedi and Gustavo Alonso.

Modeling and computing throughput capacity of wireless multihop networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(1):116– 129, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bamis:2008:MAP

[1549] Athanasios Bamis, Azzedine Boukerche, Ioannis Chatzigiannakis, and Sotiris Nikoletseas. A mobility aware protocol synthesis for efficient routing in ad hoc mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(1):130–154, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bononi:2008:MFP

[1550] Luciano Bononi, Marco Di Felice, Gabriele D’Angelo, Michele Bracuto, and Lorenzo Donatiello. MoVES: A framework for parallel and distributed simulation of wireless vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(1): 155–179, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

ElRakabawy:2008:TGA

[1551] Sherif M. ElRakabawy, Alexander Klemm, and Christoph Lindemann. TCP with gateway adaptive pacing for multihop wireless networks with Internet connectivity. Computer Networks (Amsterdam, Netherlands: 1999), 52 (1):180–198, January 18, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bruno:2008:ACF

[1552] Raffaele Bruno, Marco Conti, and Enrico Gregori. An accurate closed-form formula for the throughput of longlived TCP connections in IEEE 802.11 WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 52(1): 199–212, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Skianis:2008:ESI

[1553] Charalambos Skianis. Editorial for the special issue on synergy of telecommunication and broadcasting networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(1): 213–214, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tompros:2008:IHA

[1554] Spyridon L. Tompros and Spyridon Denazis. Interworking of heterogeneous access networks and QoS provisioning via IP multimedia core networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(1):215–227, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Santos:2008:MBN

[1555] Justino Santos, Diogo Gomes, Susana Sargento, Rui L. Aguiar, Nigel Baker, Madiha Zafar, and Ahsan Ikram. Multicast/broadcast network convergence in next generation mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(1): 228–247, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Charilas:2008:PSN

[1556] Dimitris Charilas, Ourania Markaki, Dimitris Nikitopoulos, and Michael Theologou. Packet-switched network selection with the highest QoS in 4G networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(1): 248–258, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Park:2008:LBV

[1557] Joon-Sang Park and Seung-Jae Han.

Load balancing for video streaming services in hierarchical wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(1):259– 274, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bisio:2008:PEB

[1558] Igor Bisio and Mario Marchese. Performance evaluation of bandwidth allocation methods in a geostationary satellite channel in the presence of Internet traffic. Computer Networks (Amsterdam, Netherlands: 1999), 52(1): 275–291, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gardikis:2008:DIC

[1559] G. Gardikis, S. Orfanos, G. Kormentzas, E. Pallis, and A. Kourtis. Dynamic IP configuration of terminals in broadcasting networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(1):292–302, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBa

[1560] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(1):??, January 18, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kasnavi:2008:CBI

[1561] Soraya Kasnavi, Paul Berube, Vincent Gaudet, and José Nelson Amaral. A cache-based Internet protocol address lookup architecture. Computer Networks (Amsterdam, Netherlands: 1999), 52(2):303–326, February 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Al-Karaki:2008:EVB

[1562] Jamal N. Al-Karaki and Ahmed E. Kamal. Efficient virtual-backbone routing in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(2):327–350, February 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Oida:2008:IQS

[1563] Kazumasa Oida. On the impact of QoS-sensitivity on correlation structures. Computer Networks (Amsterdam, Netherlands: 1999), 52(2):351– 359, February 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Toumpis:2008:MNK

[1564] S. Toumpis. Mother nature knows best: A survey of recent results on wireless networks based on analogies with physics. Computer Networks (Amsterdam, Netherlands: 1999), 52(2): 360–383, February 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2008:ABE

[1565] Changhan Kim, Younghoon Kim, Jae-Heon Yang, and Ikjun Yeom. Analysis of bandwidth efficiency in overlay multicasting. Computer Networks (Amsterdam, Netherlands: 1999), 52 (2):384–398, February 8, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2008:LVI

[1566] Jun Li, Jelena Mirkovic, Toby Ehrenkranz, Mengqiu Wang, Peter Reiher, and Lixia Zhang. Learning the valid incoming direction of IP packets. Computer Networks (Amsterdam, Netherlands: 1999), 52(2): 399–417, February 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chou:2008:DRM

[1567] Chih-Hsun Chou, Kuo-Feng Ssu, and Hewijin Christine Jiau. Dynamic route maintenance for geographic forwarding in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(2):418–431, February 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Merayo:2008:FTT

[1568] Mercedes G. Merayo, Manuel Núñez, and Ismael Rodríguez. Formal testing from timed finite state machines. Computer Networks (Amsterdam, Netherlands: 1999), 52(2):432–460, February 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shihada:2008:NTD

[1569] Basem Shihada and Pin-Han Ho. A novel TCP with dynamic Burst-Contention Loss notification over OBS networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(2): 461–471, February 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBb

[1570] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(2):??, February 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2008:E

[1571] Ian F. Akyildiz and Harry Rudin. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 52(3):473– 474, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mandjes:2008:BTU

[1572] M. Mandjes and M. Ramakrishnan. Bandwidth trading under misaligned objectives: Decentralized measurement-based control. Computer Networks (Amsterdam, Netherlands: 1999), 52(3):475–492, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nayyeri:2008:LST

[1573] Amir Nayyeri, Sajjad Zarifzadeh, Nasser Yazdani, and Mohammad Mahmoody-Ghidary. Load sensitive topology control: Towards minimum energy consumption in dense ad hoc sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (3):493–513, February 22, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2008:DBS

[1574] Yanni Ellen Liu and Johnny W. Wong.

Deadline-based scheduling in support of real-time data delivery. Computer Networks (Amsterdam, Netherlands: 1999), 52(3):514–530, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chang:2008:CTB

[1575] Tai-Jung Chang, Kuochen Wang, and Yi-Ling Hsieh. A color-theory-based energy efficient routing algorithm for mobile wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(3):531–541, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jin:2008:EEE

[1576] Yan Jin, Ling Wang, Yoohwan Kim, and Xiaozong Yang. EEMC: An energy-efficient multi-level clustering algorithm for large-scale wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(3): 542–562, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bolla:2008:HOQ

[1577] Raffaele Bolla, Roberto Bruschi, Franco Davoli, and Matteo Repetto. Hybrid optimization for QoS control in IP Virtual Private Networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(3):563–580, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2008:ANN

[1578] Chiapin Wang and Tsungnan Lin. Application of neural networks for achieving 802.11 QoS in heterogeneous channels. Computer Networks (Amsterdam, Netherlands: 1999), 52(3):581– 592, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nelakuditi:2008:MTL

[1579] Srihari Nelakuditi, Zifei Zhong, Junling Wang, Ram Keralapura, and Chen-Nee Chuah. Mitigating transient loops through interface-specific forwarding. Computer Networks (Amsterdam, Netherlands: 1999), 52(3): 593–609, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gandham:2008:DTO

[1580] Shashidhar Gandham, Ying Zhang, and Qingfeng Huang. Distributed time-optimal scheduling for convergecast in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(3):610–629, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ibanez:2008:ASS

[1581] G. Ibáñez, A. García-Martínez, A. Azcorra, and I. Soto. ABridges: Scalable, self-configuring Ethernet campus networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(3): 630–649, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hossfeld:2008:ASV

[1582] Tobias Hoßfeld and Andreas Binzenhöfer.

Analysis of Skype VoIP traffic in UMTS: End-to-end QoS and QoE measurements. Computer Networks (Amsterdam, Netherlands: 1999), 52(3): 650–666, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2008:MOZ

[1583] Kyung Hee Kim, Jang Hyun Baek, and Che Soong Kim. Modeling and optimization of zone-based registration considering cell-by-cell location area for mobile communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(3):667–674, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Karakaya:2008:CFR

[1584] Murat Karakaya, İbrahim Körpeoğlu, and Özgür Ulusoy. Counteracting free riding in Peer-to-Peer networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(3):675–694, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2008:GTD

[1585] Chunxi Li and Changjia Chen. On Gnutella topology dynamics by studying leaf and ultra connection jointly in phase space. Computer Networks (Amsterdam, Netherlands: 1999), 52 (3):695–719, February 22, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lin:2008:CCD

[1586] Kate Ching-Ju Lin and Cheng-Fu Chou. Cooperative content dissemination in multi-channel WLAN hotspots. Computer Networks (Amsterdam, Netherlands: 1999), 52(3): 720–738, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tian:2008:JRF

[1587] Hongtao Tian, Sanjay K. Bose, Choi Look Law, and Wendong Xiao. Joint routing and flow rate optimization in multi-rate ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(3):739– 764, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBc

[1588] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(3):??, February 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2008:CNE

[1589] Ye (Geoffrey) Li, Petri Mähönen, Milind M. Buddhikot, and Ying-Chang Liang. Computer Networks (Elsevier) Special Issue on Cognitive Wireless Networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(4): 775–777, March 14, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

deBaynast:2008:ABC

[1590] Alexandre de Baynast, Petri Mähönen, and Marina Petrova. ARQ-based crosslayer optimization for wireless multicarrier transmission on cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(4): 778–794, March 14, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Han:2008:ACC

[1591] Chuan Han, Jun Wang, Yaling Yang, and Shaoqian Li. Addressing the control channel design problem: OFDMbased transform domain communication system in cognitive radio. Computer Networks (Amsterdam, Netherlands: 1999), 52(4):795–815, March 14, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Attar:2008:CRT

[1592] A. Attar, S. A. Ghorashi, M. Sooriyabandara, and A. H. Aghvami. Challenges of real-time secondary usage of spectrum. Computer Networks (Amsterdam, Netherlands: 1999), 52(4): 816–830, March 14, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Krishnamurthy:2008:TED

[1593] Srinivasan Krishnamurthy, Mansi Thoppian, Srikant Kuppa, R. Chandrasekaran, Neeraj Mittal, S. Venkatesan, and Ravi Prakash. Time-efficient distributed layer-2 auto-configuration for cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(4):831–849, March 14, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2008:EAS

[1594] Jie Chen, Chulin Liao, Chuan Han, and Shaoqian Li. Extending available spectrum in cognitive radio: Hierarchical spectrum sharing network. Computer Networks (Amsterdam, Netherlands: 1999), 52(4):850–863, March 14, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zou:2008:QAD

[1595] Chao Zou, Tao Jin, Chunxiao Chigan, and Zhi Tian. QoS-aware distributed spectrum sharing for heterogeneous wireless cognitive networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(4):864–878, March 14, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gandhi:2008:TRT

[1596] Sorabh Gandhi, Chiranjeeb Buragohain, Lili Cao, Haitao Zheng, and Subhash Suri. Towards real-time dynamic spectrum auctions. Computer Networks (Amsterdam, Netherlands: 1999), 52(4):879–897, March 14, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBd

[1597] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(4):??, March 14, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bansod:2008:MEB

[1598] Neelesh Bansod, Ashish Malgi, Byung K. Choi, and Jean Mayo. MuON: Epidemic based mutual anonymity in unstructured P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(5):915–934, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Deng:2008:PAD

[1599] Leiwen Deng, Yan Gao, Yan Chen, and Aleksandar Kuzmanovic. Pollution attacks and defenses for Internet caching systems. Computer Networks (Amsterdam, Netherlands: 1999), 52 (5):935–956, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shevtekar:2008:RBT

[1600] Amey Shevtekar and Nirwan Ansari. A router-based technique to mitigate reduction of quality (RoQ) attacks. Computer Networks (Amsterdam, Netherlands: 1999), 52(5):957–970, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tu:2008:EBS

[1601] Yung-Cheng Tu, Meng Chang Chen, Yeali S. Sun, and Wei-Kuan Shih. Enhanced bulk scheduling for supporting delay sensitive streaming applications. Computer Networks (Amsterdam, Netherlands: 1999), 52(5): 971–987, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Junhai:2008:RMR

[1602] Luo Junhai, Xue Liu, and Ye Danxia.

Research on multicast routing protocols for mobile ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(5):988–997, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Viger:2008:DUP

[1603] Fabien Viger, Brice Augustin, Xavier Cuvellier, Clémence Magnien, Matthieu Latapy, Timur Friedman, and Renata Teixeira. Detection, understanding, and prevention of traceroute measurement artifacts. Computer Networks (Amsterdam, Netherlands: 1999), 52 (5):998–1018, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wouhaybi:2008:BRL

[1604] Rita H. Wouhaybi and Andrew T. Campbell. Building resilient lowdiameter peer-to-peer topologies. Computer Networks (Amsterdam, Netherlands: 1999), 52(5):1019–1039, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hahm:2008:MMD

[1605] Seong il Hahm, Jong won Lee, and Chong kwon Kim. Maximizing multiuser diversity gains in IEEE 802.11 WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 52(5): 1040–1057, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sarangan:2008:FFR

[1606] V. Sarangan, M. R. Devarapalli, and S. Radhakrishnan. A framework for fast RFID tag reading in static and mobile environments. Computer Networks (Amsterdam, Netherlands: 1999), 52 (5):1058–1073, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bin:2008:NBF

[1607] Liu Bin, Lin Chuang, Qiao Jian, He Jianping, and Peter Ungsunan. A NetFlow based flow analysis and monitoring system in enterprise networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(5): 1074–1092, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ma:2008:QCG

[1608] Yue Ma and Chengwen Zhang. Quick convergence of genetic algorithm for QoS-driven web service selection. Computer Networks (Amsterdam, Netherlands: 1999), 52(5):1093–1104, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rajah:2008:SBF

[1609] Kannan Rajah, Sanjay Ranka, and Ye Xia. Scheduling bulk file transfers with start and end times. Computer Networks (Amsterdam, Netherlands: 1999), 52(5):1105–1122, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBe

[1610] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(5):??, April 10, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Exposito:2008:ICL

[1611] Ernesto Exposito, Nicolas Van Wambeke, Christophe Chassot, and Khalil Drira. Introducing a cross-layer interpreter for multimedia streams. Computer Networks (Amsterdam, Netherlands:

1999), 52(6):1125–1141, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Panagakis:2008:AAL

[1612] Antonis Panagakis, Athanasios Vaios, and Ioannis Stavrakakis. Approximate analysis of LRU in the case of short term correlations. Computer Networks (Amsterdam, Netherlands: 1999), 52 (6):1142–1152, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Popova:2008:CMM

[1613] Larissa Popova, Thomas Herpel, Wolfgang Gerstacker, and Wolfgang Koch. Cooperative mobile-to-mobile file dissemination in cellular networks within a unified radio interface. Computer Networks (Amsterdam, Netherlands: 1999), 52(6):1153–1165, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2008:EPP

[1614] Ji Li, Kwan L. Yeung, and Chun-Kit Chan. Efficient path protection in bi-directional WDM systems. Computer Networks (Amsterdam, Netherlands: 1999), 52(6):1166–1178, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

CoutodaSilva:2008:QAI

[1615] Ana Paula Couto da Silva, Martín Varela, Edmundo de Souza e Silva, Rosa M. M. Leão, and Gerardo Rubino. Quality assessment of interactive voice applications. Computer Networks (Amsterdam, Netherlands: 1999), 52 (6):1179–1192, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2008:JSM

[1616] Yueping Zhang, Derek Leonard, and Dmitri Loguinov. Jetmax: Scalable max–min congestion control for highspeed heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(6):1193–1219, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chobanyan:2008:EEA

[1617] Alexander Chobanyan, Matt Mutka, Vidiadhar Mandrekar, and Ning Xi. End-to-end available bandwidth as a random autocorrelated QoS-relevant time-series. Computer Networks (Amsterdam, Netherlands: 1999), 52(6): 1220–1237, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Veeraraghavan:2008:FTD

[1618] Vilas Veeraraghavan and Steven Weber. Fundamental tradeoffs in distributed algorithms for rate adaptive multimedia streams. Computer Networks (Amsterdam, Netherlands: 1999), 52(6):1238–1251, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fodor:2008:PAS

[1619] Gábor Fodor, Miklós Telek, and Chrysostomos Koutsimanis. Performance analysis of scheduling and interference coordination policies for OFDMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(6):1252–1271, April 24, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Banerjee:2008:PWS

[1620] Anirban Banerjee, Michalis Faloutsos, and Laxmi Bhuyan. The P2P war: Someone is monitoring your activities. Computer Networks (Amsterdam, Netherlands: 1999), 52(6): 1272–1280, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Harhira:2008:NAC

[1621] Hichem Ayed Harhira and Samuel Pierre. A novel admission control mechanism in GMPLS-based IP over optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (6):1281–1290, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Belotti:2008:MLM

[1622] Pietro Belotti, Antonio Capone, Giuliana Carello, and Federico Malucelli. Multi-layer MPLS network design: The impact of statistical multiplexing. Computer Networks (Amsterdam, Netherlands: 1999), 52(6): 1291–1307, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ho:2008:FRF

[1623] Cheng-Yuan Ho, Yaw-Chung Chen, Yi-Cheng Chan, and Cheng-Yun Ho. Fast retransmit and fast recovery schemes of transport protocols: A survey and taxonomy. Computer Networks (Amsterdam, Netherlands: 1999), 52(6): 1308–1327, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Huang:2008:SSC

[1624] Shih-Chang Huang, Rong-Hong Jan, and Wuu Yang. SCPS: A selfconfiguring power-saving protocol for wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(6):1328–1342, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBf

[1625] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(6):??, April 24, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Szymaniak:2008:PLS

[1626] Michal Szymaniak, David Presotto, Guillaume Pierre, and Maarten van Steen. Practical large-scale latency estimation. Computer Networks (Amsterdam, Netherlands: 1999), 52(7): 1343–1364, May 15, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kherani:2008:STT

[1627] A. Kherani, R. El-Khoury, R. El-Azouzi, and E. Altman. Stability– throughput tradeoff and routing in multi-hop wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(7): 1365–1389, May 15, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guitart:2008:DCP

[1628] Jordi Guitart, David Carrera, Vicenç Beltran, Jordi Torres, and Eduard Ayguadé. Dynamic CPU provisioning for self-managed secure web applications in SMP hosting platforms. Computer Networks (Amsterdam, Netherlands: 1999), 52(7):1390–1409, May 15, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bai:2008:CAM

[1629] Baochun Bai, Janelle Harms, and Yuxi Li. Configurable active multicast congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 52 (7):1410–1432, May 15, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ozianyi:2008:NPA

[1630] Vitalis G. Ozianyi, Neco Ventura, and Eugene Golovins. A novel pricing approach to support QoS in 3G networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(7): 1433–1450, May 15, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Joung:2008:ANP

[1631] Yuh-Jzer Joung. Approaching neighbor proximity and load balance for range query in P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(7):1451–1472, May 15, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

DeVleeschouwer:2008:LRW

[1632] Christophe De Vleeschouwer and Pascal Frossard. Loss-resilient windowbased congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 52(7):1473–1491, May 15, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Guo:2008:TRT

[1633] Lei Guo, Xingwei Wang, Jin Cao, and Lemin Li. Traffic recovery time constrained shared sub-path protection algorithm in survivable WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(7): 1492–1505, May 15, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ahuja:2008:AAD

[1634] Sandeep K. Ahuja and Srinivasan Ramasubramanian. All-to-all disjoint multipath routing using cycle embedding. Computer Networks (Amsterdam, Netherlands: 1999), 52(7): 1506–1517, May 15, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBg

[1635] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(7):??, May 15, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vergados:2008:GRD

[1636] Dimitrios D. Vergados, Alexandros Panoutsakopoulos, and Christos Douligeris. Group registration with distributed databases for location tracking in 3G wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(8):1521–1544, June 12, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nam:2008:DSE

[1637] Seung Yeob Nam, Hyu-Dae Kim, and Hyong S. Kim. Detector SherLOCK: Enhancing TRW with Bloom filters under memory and performance constraints. Computer Networks (Amsterdam, Netherlands: 1999), 52(8): 1545–1566, June 12, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shue:2008:PFS

[1638] Craig A. Shue, Minaxi Gupta, and Matthew P. Davy. Packet forwarding with source verification. Computer Networks (Amsterdam, Netherlands: 1999), 52(8):1567–1582, June 12, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Franklin:2008:LLA

[1639] A. Antony Franklin and C. Siva Ram Murthy. A link layer adaptive pacing scheme for improving throughput of transport protocols in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(8): 1583–1602, June 12, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nair:2008:BSF

[1640] Padmaraj M. V. Nair, Suku V. S. Nair, Marco Marchetti, Girish Chiruvolu, and Maher Ali. Bandwidth sensitive fast failure recovery scheme for Metro Ethernet. Computer Networks (Amsterdam, Netherlands: 1999), 52 (8):1603–1616, June 12, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Garbacki:2008:BPL

[1641] Paweł Garbacki, Dick H. J. Epema, and Maarten van Steen. Brokerplacement in latency-aware peer-to-peer networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (8):1617–1633, June 12, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Huang:2008:SME

[1642] Rongsheng Huang, Hongqiang Zhai, Chi Zhang, and Yuguang Fang. SAM-MAC: An efficient channel assignment scheme for multi-channel ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(8): 1634–1646, June 12, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2008:DEF

[1643] Qi Wang, Robert Atkinson, and John Dunlop. Design and evaluation of flow handoff signalling for multihomed mobile nodes in wireless overlay networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(8): 1647–1674, June 12, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBh

[1644] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(8):??, June 12, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ang:2008:IBG

[1645] Chee-Wei Ang and Chen-Khong Tham.

iMST: A bandwidth-guaranteed topology control algorithm for TDMA-based ad hoc networks with sectorized antennas. Computer Networks (Amsterdam, Netherlands: 1999), 52(9): 1675–1692, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shi:2008:UAU

[1646] Minghui Shi, Xuemin (Sherman) Shen, Jon W. Mark, Dongmei Zhao, and Yixin Jiang. User authentication and undeniable billing support for agent-based roaming service in WLAN/cellular integrated mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(9): 1693–1702, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ma:2008:JTP

[1647] Hui Ma, Jing Zhu, Sumit Roy, and Soo Young Shin. Joint transmit power and physical carrier sensing adaptation based on loss differentiation for high density IEEE 802.11 WLAN. Computer Networks (Amsterdam, Netherlands: 1999), 52(9):1703–1720, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Flammini:2008:SCP

[1648] Michele Flammini, Gianpiero Monaco, Luca Moscardelli, Mordechai Shalom, and Shmuel Zaks. Selfishness, collusion and power of local search for the ADMs minimization problem. Computer Networks (Amsterdam, Netherlands: 1999), 52(9):1721–1731, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Berzin:2008:MLB

[1649] Oleg Berzin. Mobility label based network: Mobility support in label switched networks with multi-protocol BGP. Computer Networks (Amsterdam, Netherlands: 1999), 52(9): 1732–1744, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wuhib:2008:DDG

[1650] Fetahi Wuhib, Mads Dam, and Rolf Stadler. Decentralized detection of global threshold crossings using aggregation trees. Computer Networks (Amsterdam, Netherlands: 1999), 52(9): 1745–1761, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Batista:2008:SAR

[1651] Daniel M. Batista, Nelson L. S. da Fonseca, Flavio K. Miyazawa, and Fabrizio Granelli. Self-adjustment of resource allocation for grid applications. Computer Networks (Amsterdam, Netherlands: 1999), 52(9):1762–1781, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dong:2008:LBL

[1652] Qunfeng Dong, Ashutosh Shukla, Vivek Shrivastava, Dheeraj Agrawal, Suman Banerjee, and Koushik Kar. Load balancing in large-scale RFID systems. Computer Networks (Amsterdam, Netherlands: 1999), 52(9): 1782–1796, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Konstantopoulos:2008:CMA

[1653] Charalampos Konstantopoulos, Damianos Gavalas, and Grammati Pantziou. Clustering in mobile ad hoc networks through neighborhood stabilitybased mobility prediction. Computer Networks (Amsterdam, Netherlands: 1999), 52(9):1797–1824, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2008:MCN

[1654] Xue Jun Li, Boon-Chong Seet, and Peter Han Joo Chong. Multihop cellular networks: Technology and economics. Computer Networks (Amsterdam, Netherlands: 1999), 52(9): 1825–1837, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chan:2008:ESE

[1655] Mo-Che Chan, Shun-Yun Hu, and Jehn-Ruey Jiang. An efficient and secure event signature (EASES) protocol for peer-to-peer massively multiplayer online games. Computer Networks (Amsterdam, Netherlands: 1999), 52 (9):1838–1845, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Siekkinen:2008:RCA

[1656] Matti Siekkinen, Guillaume Urvoy-Keller, Ernst W. Biersack, and Denis Collange. A root cause analysis toolkit for TCP. Computer Networks (Amsterdam, Netherlands: 1999), 52 (9):1846–1858, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBi

[1657] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(9):??, June 26, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tomkos:2008:CNJ

[1658] Ioannis Tomkos, Davide Careglio, Anna Tzanakaki, and Josep Solé-Pareta. Computer Networks journal special issue on: “Challenges and opportunities in advanced optical networking”. Computer Networks (Amsterdam, Netherlands: 1999), 52(10): 1861–1863, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ichikawa:2008:UWN

[1659] Haruhisa Ichikawa, Masashi Shimizu, Kazunori Akabane, Osamu Ishida, and Mitsuo Teramoto. A Ubiquitous wireless network architecture and its impact on optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(10):1864–1872, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:PSA

[1660] Anonymous. Photonics in switching:

Architectures, systems and enabling technologies. Computer Networks (Amsterdam, Netherlands: 1999), 52(10): 1873–1890, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yao:2008:RSD

[1661] Wang Yao and Byrav Ramamurthy.

Rerouting schemes for dynamic traffic grooming in optical WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(10): 1891–1904, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pavani:2008:DAI

[1662] Gustavo Sousa Pavani, Luiz Gustavo Zuliani, Helio Waldman, and Mauricio Magalhães. Distributed approaches for impairment-aware routing and wavelength assignment algorithms in GM-PLS networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (10):1905–1915, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lopez:2008:BDT

[1663] Víctor López, José Alberto Hernández, Javier Aracil, Juan P. Fernández Palacios, and Óscar González de Dios. A Bayesian decision theory approach for the techno-economic analysis of an all-optical router (extended version). Computer Networks (Amsterdam, Netherlands: 1999), 52(10): 1916–1926, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dijkstra:2008:MLN

[1664] Freek Dijkstra, Bert Andree, Karst Koymans, Jeroen van der Ham, Paola Grosso, and Cees de Laat. A multilayer network model based on ITU-T G.805. Computer Networks (Amsterdam, Netherlands: 1999), 52(10): 1927–1937, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Escalona:2008:ARS

[1665] Eduard Escalona, Salvatore Spadaro, Jaume Comellas, and Gabriel Junyent. Advance reservations for service-aware GMPLS-based optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(10):1938–1950, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ruepp:2008:RAO

[1666] Sarah Ruepp, Nicola Andriolli, Jakob Buron, Lars Dittmann, and Lars Ellegaard. Restoration in all-optical GM-PLS networks with limited wavelength conversion. Computer Networks (Amsterdam, Netherlands: 1999), 52(10): 1951–1964, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tornatore:2008:ISS

[1667] Massimo Tornatore, Matteo Carcagní, Canhui (Sam) Ou, Biswanath Mukherjee, and Achille Pattavina. Intelligent shared-segment protection. Computer Networks (Amsterdam, Netherlands: 1999), 52(10):1965–1974, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Velasco:2008:IOP

[1668] Luis Velasco, Salvatore Spadaro, Jaume Comellas, and Gabriel Junyent. Introducing OMS protection in GMPLS-based optical ring networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(10): 1975–1987, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Awad:2008:DDR

[1669] Christian Awad, Brunilde Sansò, and André Girard. DiffServ for differentiated reliability in meshed IP/WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(10): 1988–2012, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mazloom:2008:CCF

[1670] Amin R. Mazloom, Preetam Ghosh, Kalyan Basu, and Sajal K. Das. Co-CONet : A collision-free containerbased core optical network. Computer Networks (Amsterdam, Netherlands: 1999), 52(10):2013–2032, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2008:ROO

[1671] Qian Chen, Gurusamy Mohan, and Kee Chaing Chua. Route optimization in optical burst switched networks considering the streamline effect. Computer Networks (Amsterdam, Netherlands: 1999), 52(10):2033–2044, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lin:2008:COM

[1672] Hui-Tang Lin and Wang-Rong Chang. CORNet: An OBS metro ring network with QoS support and fairness control. Computer Networks (Amsterdam, Netherlands: 1999), 52(10): 2045–2064, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zervas:2008:SEO

[1673] Georgios Zervas, Yixuan Qin, Reza Nejabati, Dimitra Simeonidou, Franco Callegati, Aldo Campi, and Walter Cerroni. SIP-enabled Optical Burst Switching architectures and protocols for application-aware optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(10): 2065–2076, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nasaruddin:2008:DSV

[1674] Nasaruddin and Tetsuo Tsujioka. Design of strict variable-weight optical orthogonal codes for differentiated quality of service in optical CDMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(10): 2077–2086, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Awaji:2008:GAO

[1675] Yoshinari Awaji, Hideaki Furukawa, Naoya Wada, Eddie Kong, Peter Chan, and Ray Man. Guidelines for amplification of optical packets in WDM environment regarding impact of transient response of erbium-doped fiber amplifier. Computer Networks (Amsterdam, Netherlands: 1999), 52(10): 2087–2093, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBj

[1676] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(10):??, July 16, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Meshkova:2008:SRD

[1677] Elena Meshkova, Janne Riihijärvi, Marina Petrova, and Petri Mähönen. A survey on resource discovery mechanisms, peer-to-peer and service discovery frameworks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (11):2097–2128, August 8, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chang:2008:WAW

[1678] Tae-Young Chang, Zhenyun Zhuang, Aravind Velayutham, and Raghupathy Sivakumar. WebAccel: Accelerating Web access for low-bandwidth hosts. Computer Networks (Amsterdam, Netherlands: 1999), 52(11): 2129–2147, August 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tang:2008:JSA

[1679] Jian Tang, Satyajayant Misra, and Guoliang Xue. Joint spectrum allocation and scheduling for fair spectrum sharing in cognitive radio wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(11): 2148–2158, August 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Amaldi:2008:OMM

[1680] E. Amaldi, A. Capone, M. Cesana, I. Filippini, and F. Malucelli. Optimization models and methods for planning wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(11):2159–2171, August 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Allen:2008:PSP

[1681] S. M. Allen, R. M. Whitaker, and S. Hurley. Personalised subscription pricing for optimised wireless mesh network deployment. Computer Networks (Amsterdam, Netherlands: 1999), 52 (11):2172–2188, August 8, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chang:2008:EAN

[1682] Chih-Yung Chang and Hsu-Ruey Chang. Energy-aware node placement, topology control and MAC scheduling for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(11):2189–2204, August 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhao:2008:CTC

[1683] Qun Zhao and Mohan Gurusamy. Connected K-target coverage problem in wireless sensor networks with different observation scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 52(11):2205–2220, August 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2008:DSF

[1684] W. J. Liu and J. Gong. Double sampling for flow measurement on high speed links. Computer Networks (Amsterdam, Netherlands: 1999), 52(11): 2221–2226, August 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Misic:2008:TEC

[1685] Jelena Mišić. Traffic and energy consumption of an IEEE 802.15.4 network in the presence of authenticated, ECC Diffie–Hellman ephemeral key exchange. Computer Networks (Amsterdam, Netherlands: 1999), 52(11): 2227–2236, August 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dasgupta:2008:DTE

[1686] Sukrit Dasgupta, Jaudelice C. de Oliveira, and J.-P. Vasseur. Dynamic traffic engineering for mixed traffic on international networks: Simulation and analysis on real network and traffic scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 52(11): 2237–2258, August 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBk

[1687] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(11):??, August 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:PN

[1688] Anonymous. Publisher’s Note. Computer Networks (Amsterdam, Netherlands: 1999), 52(12):2259, August 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2008:NNC

[1689] Ian F. Akyildiz, Fernando Brunetti, and Cristina Blázquez. Nanonetworks: A new communication paradigm. Computer Networks (Amsterdam, Netherlands: 1999), 52(12):2260–2279, August 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Qiao:2008:ECW

[1690] Daji Qiao, Kang G. Shin, and Zakhia Abichar. Energy-conservation in 802.11 WLANs via transmission-strategy-aware airtime allocation. Computer Networks (Amsterdam, Netherlands: 1999), 52(12):2280–2291, August 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yick:2008:WSN

[1691] Jennifer Yick, Biswanath Mukherjee, and Dipak Ghosal. Wireless sensor network survey. Computer Networks (Amsterdam, Netherlands: 1999), 52(12): 2292–2330, August 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chao:2008:OCA

[1692] Chih-Min Chao. OVSF code assignment strategies with minimal fragmentations for WCDMA systems. Computer Networks (Amsterdam, Netherlands: 1999), 52(12):2331–2343, August 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lopez:2008:MNC

[1693] José Gutiérrez López, Mohamed Imine, Rubén Cuevas Rumín, Jens M. Pedersen, and Ole B. Madsen. Multilevel network characterization using regular topologies. Computer Networks (Amsterdam, Netherlands: 1999), 52(12): 2344–2359, August 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Naraghi-Pour:2008:LFT

[1694] Mort Naraghi-Pour and Vinay Desai.

Loop-free traffic engineering with path protection in MPLS VPNs. Computer Networks (Amsterdam, Netherlands: 1999), 52(12):2360–2372, August 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2008:NMA

[1695] Yalin Chen, Jue-Sam Chou, and Hung-Min Sun. A novel mutual authentication scheme based on quadratic residues for RFID systems. Computer Networks (Amsterdam, Netherlands: 1999), 52(12):2373–2380, August 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Datta:2008:GTA

[1696] Pallab Datta and Arun K. Somani. Graph transformation approaches for diverse routing in shared risk resource group (SRRG) failures. Computer Networks (Amsterdam, Netherlands: 1999), 52(12):2381–2394, August 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kliazovich:2008:LWI

[1697] Dzmitry Kliazovich, Fabrizio Granelli, and Daniele Miorandi. Logarithmic window increase for TCP Westwood+ for improvement in high speed, long distance networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (12):2395–2410, August 22, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jain:2008:PSU

[1698] Manish Jain and Constantine Dovrolis. Path selection using available bandwidth estimation in overlay-based video streaming. Computer Networks (Amsterdam, Netherlands: 1999), 52 (12):2411–2418, August 22, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altinel:2008:BIP

[1699] İ. Kuban Altınel, Necati Aras, Evren Güney, and Cem Ersoy. Binary integer programming formulation and heuristics for differentiated coverage in heterogeneous sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(12):2419–2431, August 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Benini:2008:ARI

[1700] M. Benini and S. Sicari. Assessing the risk of intercepting VoIP calls. Computer Networks (Amsterdam, Netherlands: 1999), 52(12):2432–2446, August 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBl

[1701] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(12):??, August 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pavlidou:2008:RTR

[1702] Fotini-Niovi Pavlidou, Andreas Pitsillides, Henning Schulzrinne, and Dorghan Sisalem. Research and trials for reliable VoIP applications. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2447– 2449, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liao:2008:MPM

[1703] Jianxin Liao, Jingyu Wang, and Xiaomin Zhu. A multi-path mechanism for reliable VoIP transmission over wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (13):2450–2460, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kalama:2008:VDR

[1704] Maria Kalama, Guray Acar, Barry Evans, and Ariane Isoard. VoIP over DVB-RCS satellite systems: Trial results and the impact of adaptive speech coding using cross-layer design. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2461– 2472, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jelassi:2008:APS

[1705] Sofiene Jelassi and Habib Youssef. Adaptive playout scheduling algorithm tailored for real-time packet-based voice conversations over wireless adhoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (13):2473–2488, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hsieh:2008:EVS

[1706] Hung-Yun Hsieh, You-En Lin, and Hsiao-Pu Lin. Enhancing VoIP service for ubiquitous communication in a campus WLAN with partial coverage. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2489– 2504, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Estepa:2008:ARE

[1707] Antonio Estepa and Rafael Estepa.

Accurate resource estimation for homogeneous VoIP aggregated traffic. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2505– 2517, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zisiadis:2008:VD

[1708] Dimitris Zisiadis, Spyros Kopsidas, and Leandros Tassiulas. VIPSec defined. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2518– 2528, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akan:2008:SIW

[1709] Ozgur B. Akan, Pascal Frossard, Qian Zhang, and Nikil Jayant. Special issue on wireless multimedia sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2529– 2531, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Saxena:2008:DDC

[1710] Navrati Saxena, Abhishek Roy, and Jitae Shin. Dynamic duty cycle and adaptive contention window based QoS-MAC protocol for wireless multimedia sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2532–2542, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Na:2008:OGS

[1711] Chewoo Na, Yaling Yang, and Amitabh Mishra. An optimal GTS scheduling algorithm for time-sensitive transactions in IEEE 802.15.4 networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2543–2557, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tezcan:2008:SOW

[1712] Nurcan Tezcan and Wenye Wang. Self-orienting wireless multimedia sensor networks for occlusion-free viewpoints. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2558– 2567, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Suh:2008:DIE

[1713] Changsu Suh, Zeeshan Hameed Mir, and Young-Bae Ko. Design and implementation of enhanced IEEE 802.15.4 for supporting multimedia service in Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2568–2581, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Malhotra:2008:DCA

[1714] Baljeet Malhotra, Ioanis Nikolaidis, and Janelle Harms. Distributed classification of acoustic targets in wireless audio-sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2582–2593, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lu:2008:LCE

[1715] Qin Lu, Wusheng Luo, Jidong Wang, and Bo Chen. Low-complexity and energy efficient image compression scheme for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2594– 2603, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBm

[1716] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):??, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cohen:2008:PTQ

[1717] Edith Cohen, Nadav Grossaug, and Haim Kaplan. Processing top-k queries from samples. Computer Networks (Amsterdam, Netherlands: 1999), 52 (14):2605–2622, October 9, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sundaresan:2008:RAH

[1718] Karthikeyan Sundaresan and Raghupathy Sivakumar. Routing in ad-hoc networks with MIMO links: Optimization considerations and protocols. Computer Networks (Amsterdam, Netherlands: 1999), 52(14):2623–2644, October 9, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dainotti:2008:ITM

[1719] Alberto Dainotti, Antonio Pescapé, Pierluigi Salvo Rossi, Francesco Palmieri, and Giorgio Ventre. Internet traffic modeling by means of Hidden Markov Models. Computer Networks (Amsterdam, Netherlands: 1999), 52 (14):2645–2662, October 9, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mellia:2008:PAT

[1720] Marco Mellia, Michela Meo, Luca Muscariello, and Dario Rossi. Passive analysis of TCP anomalies. Computer Networks (Amsterdam, Netherlands: 1999), 52(14):2663–2676, October 9, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fernandes:2008:STS

[1721] Stênio Fernandes, Carlos Kamienski, Judith Kelner, Dênio Mariz, and Djamel Sadok. A stratified traffic sampling methodology for seeing the big picture. Computer Networks (Amsterdam, Netherlands: 1999), 52(14): 2677–2689, October 9, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Henderson:2008:CUM

[1722] Tristan Henderson, David Kotz, and Ilya Abyzov. The changing usage of a mature campus-wide wireless network. Computer Networks (Amsterdam, Netherlands: 1999), 52(14):2690– 2712, October 9, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Al-Abri:2008:IBL

[1723] Dawood Al-Abri and Janise McNair. On the interaction between localization and location verification for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (14):2713–2727, October 9, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ergin:2008:ESI

[1724] Mesut Ali Ergin, Kishore Ramachandran, and Marco Gruteser. An experimental study of inter-cell interference effects on system performance in unplanned wireless LAN deployments. Computer Networks (Amsterdam, Netherlands: 1999), 52(14): 2728–2744, October 9, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Oliviero:2008:RAA

[1725] F. Oliviero, L. Peluso, and S. P.

Romano. REFACING: An autonomic approach to network security based on multidimensional trustworthiness. Computer Networks (Amsterdam, Netherlands: 1999), 52(14): 2745–2763, October 9, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chang:2008:TTL

[1726] Chao-Tsun Chang, Chih-Yung Chang, and Sheng-Wen Chang. TMCP: Twolayer multicast communication protocol for Bluetooth radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(14):2764–2778, October 9, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akkaya:2008:MCC

[1727] K. Akkaya and S. Janapala. Maximizing connected coverage via controlled actor relocation in wireless sensor and actor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (14):2779–2796, October 9, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shi:2008:CUB

[1728] Yi Shi and Y. Thomas Hou. On the capacity of UWB-based wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(14): 2797–2804, October 9, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Menth:2008:RNA

[1729] Michael Menth, Stefan Kopf, Joachim Charzinski, and Karl Schrodi. Resilient network admission control. Computer Networks (Amsterdam, Netherlands: 1999), 52(14):2805–2815, October 9, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBn

[1730] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(14):??, October 9, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Latapy:2008:CCC

[1731] Matthieu Latapy and Walter Willinger. Complex computer and communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(15): 2817–2818, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fraiwan:2008:SAC

[1732] M. Fraiwan and G. Manimaran. Scheduling algorithms for conducting conflict-free measurements in overlay networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(15): 2819–2830, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gonen:2008:FDC

[1733] Mira Gonen, Dana Ron, Udi Weinsberg, and Avishai Wool. Finding a dense-core in Jellyfish graphs. Computer Networks (Amsterdam, Netherlands: 1999), 52(15):2831–2841, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Scherrer:2008:DSD

[1734] A. Scherrer, P. Borgnat, E. Fleury, J.- L. Guillaume, and C. Robardet. Description and simulation of dynamic mobility networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (15):2842–2858, October 23, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vilhar:2008:PRA

[1735] Andrej Vilhar and Roman Novak. Policy relationship annotations of predefined AS-level topologies. Computer Networks (Amsterdam, Netherlands: 1999), 52(15):2859–2871, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vik:2008:EST

[1736] Knut-Helge Vik, Pål Halvorsen, and Carsten Griwodz. Evaluating Steinertree heuristics and diameter variations for application layer multicast. Computer Networks (Amsterdam, Netherlands: 1999), 52(15):2872–2893, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gojmerac:2008:TLC

[1737] Ivan Gojmerac, Peter Reichl, and Lasse Jansen. Towards low-complexity Internet traffic engineering: The Adaptive Multi-Path algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 52(15):2894–2907, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Karlin:2008:ASA

[1738] Josh Karlin, Stephanie Forrest, and Jennifer Rexford. Autonomous security for autonomous systems. Computer Networks (Amsterdam, Netherlands: 1999), 52(15):2908–2923, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2008:TAS

[1739] Hwangnam Kim, Jennifer C. Hou, and Hyuk Lim. TranSim: Accelerating simulation of large-scale IP networks through preserving network invariants. Computer Networks (Amsterdam, Netherlands: 1999), 52(15): 2924–2946, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Harks:2008:CCU

[1740] Tobias Harks and Tobias Poschwatta. Congestion control in utility fair networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(15):2947– 2960, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2008:GTA

[1741] Jessie Hui Wang, Dah Ming Chiu, and John C. S. Lui. A game–theoretic analysis of the implications of overlay network traffic on ISP peering. Computer Networks (Amsterdam, Netherlands: 1999), 52(15):2961–2974, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Botta:2008:AIN

[1742] Alessio Botta, Antonio Pescapé, and Giorgio Ventre. An approach to the identification of network elements composing heterogeneous end-to-end paths. Computer Networks (Amsterdam, Netherlands: 1999), 52(15):2975– 2987, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Varvarigos:2008:RSC

[1743] Emmanouel (Manos) Varvarigos, Vasileios Sourlas, and Konstantinos Christodoulopoulos. Routing and scheduling connections in networks that support advance reservations. Computer Networks (Amsterdam, Netherlands: 1999), 52(15): 2988–3006, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Domingo-Ferrer:2008:PHS

[1744] Josep Domingo-Ferrer, Alexandre Viejo, Francesc Sebé, and Úrsula González-Nicolás. Privacy homomorphisms for social networks with private relationships. Computer Networks (Amsterdam, Netherlands: 1999), 52 (15):3007–3016, October 23, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2008:SSC

[1745] Hui Chen, Yang Xiao, and Susan V.

Vrbsky. Scalability study of cache access mechanisms in multiple-cell wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (15):3017–3027, October 23, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBo

[1746] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(15):??, October 23, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Han:2008:EEE

[1747] Junghee Han, David Watson, and Farnam Jahanian. Enhancing endto-end availability and performance via topology-aware overlay networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(16):3029–3046, November 13, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Machado:2008:SGT

[1748] Renita Machado and Sirin Tekinay. A survey of game-theoretic approaches in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(16):3047–3061, November 13, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2008:ITP

[1749] Xiuchao Wu, Mun Choon Chan, and A. L. Ananda. Improving TCP performance in heterogeneous mobile environments by exploiting the explicit cooperation between server and mobile host. Computer Networks (Amsterdam, Netherlands: 1999), 52(16): 3062–3074, November 13, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shao:2008:CBF

[1750] Zuhua Shao. Certificate-based fair exchange protocol of signatures from pairings. Computer Networks (Amsterdam, Netherlands: 1999), 52(16):3075– 3084, November 13, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lai:2008:DPR

[1751] Yuan-Cheng Lai and Ching-Neng Lai. DCCP partial reliability extension with sequence number compensation. Computer Networks (Amsterdam, Netherlands: 1999), 52(16):3085– 3100, November 13, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fang:2008:MPL

[1752] Jen-Yu Fang, Hung-Chi Chu, Rong-Hong Jan, and Wuu Yang. A multiple power-level approach for wireless sensor network positioning. Computer Networks (Amsterdam, Netherlands: 1999), 52(16):3101–3118, November 13, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2008:LBD

[1753] Cungang Yang, Celia Li, and Jie Xiao. Location-based design for secure and efficient wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(16):3119– 3129, November 13, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

MoniaGhobadi:2008:ROA

[1754] MoniaGhobadi, Sudhakar Ganti, and Gholamali C. Shoja. Resource optimization algorithms for virtual private networks using the hose model. Computer Networks (Amsterdam, Netherlands: 1999), 52(16):3130– 3147, November 13, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pan:2008:VFS

[1755] Qiang Pan, Jianming Wei, Haitao Liu, and Maolin Hu. Virtual field strategy for collaborative signal and information processing in wireless heterogeneous sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (16):3148–3168, November 13, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fallah:2008:ATT

[1756] Yaser Pourmohammadi Fallah and Hussein Alnuweiri. Analysis of temporal and throughput fair scheduling in multirate WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 52 (16):3169–3183, November 13, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBp

[1757] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(16):??, November 13, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sanchez-Artigas:2008:TCA

[1758] Marc Sànchez-Artigas, Pedro García-López, Antonio F. Gómez-Skarmeta, and José Santa. TR-clustering: Alleviating the impact of false clustering on P2P overlay networks. Computer Networks (Amsterdam, Netherlands:

1999), 52(17):3185–3204, December 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wei:2008:CAN

[1759] Wei Wei, Bing Wang, Chun Zhang, Jim Kurose, and Don Towsley. Classification of access network types: Ethernet, wireless LAN, ADSL, cable modem or dialup? Computer Networks (Amsterdam, Netherlands: 1999), 52(17): 3205–3217, December 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Verma:2008:CBS

[1760] Vijay K. Verma, Ramesh C. Joshi, Bin Xie, and Dharma P. Agrawal. Combating the bloated state problem in mobile agents based network monitoring applications. Computer Networks (Amsterdam, Netherlands: 1999), 52(17): 3218–3228, December 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Campos:2008:FAC

[1761] Rui Campos and Manuel Ricardo. A fast algorithm for computing minimum routing cost spanning trees. Computer Networks (Amsterdam, Netherlands: 1999), 52(17):3229–3247, December 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dimitropoulos:2008:ESS

[1762] Xenofontas Dimitropoulos, Marc Stoecklin, Paul Hurley, and Andreas Kind. The eternal sunshine of the sketch data structure. Computer Networks (Amsterdam, Netherlands: 1999), 52(17):3248–3257, December 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pla:2008:ORP

[1763] Vicent Pla, Jorma Virtamo, and Jorge Martínez-Bauset. Optimal robust policies for bandwidth allocation and admission control in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(17):3258– 3272, December 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pack:2008:DHA

[1764] Sangheon Pack and Wonjun Lee. Dual home agent (DHA)-based location management scheme in integrated cellular–WLAN networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(17):3273–3283, December 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shioda:2008:ENW

[1765] Shigeo Shioda, Kenji Ohtsuka, and Takehiro Sato. An efficient networkwide broadcasting based on hoplimited shortest-path trees. Computer Networks (Amsterdam, Netherlands: 1999), 52(17):3284–3295, December 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jiang:2008:EQO

[1766] Jehn-Ruey Jiang. Expected quorum overlap sizes of quorum systems for asynchronous power-saving in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 52 (17):3296–3306, December 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2008:EBq

[1767] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 52(17):??, December 8, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2008:SPR

[1768] Chih-Chiang Wang and Khaled Harfoush. Shortest-path routing in randomized DHT-based Peer-to-Peer systems. Computer Networks (Amsterdam, Netherlands: 1999), 52(18):3307– 3317, December 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jaisson:2008:IPM

[1769] Pascal Jaisson and Florian De Vuyst.

An innovating PDE model based on fluid flow paradigm for multithread systems. Computer Networks (Amsterdam, Netherlands: 1999), 52(18):3318– 3324, December 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cicic:2008:BPF

[1770] Tarik Cicić. On basic properties of fault-tolerant multi-topology routing. Computer Networks (Amsterdam, Netherlands: 1999), 52(18):3325– 3341, December 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Su:2008:QDA

[1771] Wei-Tsung Su, Yau-Hwang Kuo, and Po-Cheng Huang. A QoS-driven approach for service-oriented device anycasting in ubiquitous environments. Computer Networks (Amsterdam, Netherlands: 1999), 52(18):3342– 3357, December 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nguyen-Vuong:2008:UCC

[1772] Quoc-Thinh Nguyen-Vuong, Nazim Agoulmine, and Yacine Ghamri-Doudane. A user-centric and contextaware solution to interface management and access network selection in heterogeneous wireless environments. Computer Networks (Amsterdam, Netherlands: 1999), 52(18):3358– 3372, December 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Comuzzi:2008:PQD

[1773] Marco Comuzzi, Chiara Francalanci, and Paolo Giacomazzi. Pricing the quality of differentiated services for media-oriented real-time applications: A multi-attribute negotiation approach. Computer Networks (Amsterdam, Netherlands: 1999), 52(18): 3373–3391, December 22, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tsao:2008:AFQ

[1774] Shih-Chiang Tsao, Yuan-Cheng Lai, Le-Chi Tsao, and Ying-Dar Lin. On applying fair queuing discipline to schedule requests at access gateway for downlink differential QoS. Computer Networks (Amsterdam, Netherlands: 1999), 52(18):3392–3404, December 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tay:2008:EAT

[1775] Y. C. Tay, Dinh Nguyen Tran, Eric Yi Liu, Wei Tsang Ooi, and Robert Morris. Equilibrium analysis through separation of user and network behavior. Computer Networks (Amsterdam, Netherlands: 1999), 52(18):3405– 3420, December 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jaekel:2008:NAD

[1776] Arunita Jaekel, Subir Bandyopadhyay, and Yash Aneja. A new approach for designing fault-tolerant WDM networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(18):3421– 3432, December 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nezhad:2008:LPA

[1777] Alireza A. Nezhad, Ali Miri, and Dimitris Makrakis. Location privacy and anonymity preserving routing for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(18):3433–3452, December 22, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2009:ENC

[1778] Hwangnam Kim and Jennifer C.

Hou. Enabling network calculus-based simulation for TCP congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 53(1):1–24, January 16, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Langar:2009:MAC

[1779] Rami Langar, Nizar Bouabdallah, and Raouf Boutaba. Mobility-aware clustering algorithms with interference constraints in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(1): 25–44, January 16, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2009:EML

[1780] Ai Chen, Dongwook Lee, and Prasun Sinha. Efficient multicasting over largescale WLANs through controlled association. Computer Networks (Amsterdam, Netherlands: 1999), 53(1): 45–59, January 16, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Seetharaman:2009:IDP

[1781] Srinivasan Seetharaman and Mostafa Ammar. Inter-domain policy violations in multi-hop overlay routes: Analysis and mitigation. Computer Networks (Amsterdam, Netherlands: 1999), 53 (1):60–80, January 16, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dusi:2009:THD

[1782] M. Dusi, M. Crotti, F. Gringoli, and L. Salgarelli. Tunnel Hunter: Detecting application-layer tunnels with statistical fingerprinting. Computer Networks (Amsterdam, Netherlands: 1999), 53 (1):81–97, January 16, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2009:DCA

[1783] Jongtack Kim and Saewoong Bahk. Design of certification authority using secret redistribution and multicast routing in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(1):98–109, January 16, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xia:2009:APL

[1784] Ye Xia, Shigang Chen, Chunglae Cho, and Vivekanand Korgaonkar. Algorithms and performance of loadbalancing with multiple hash functions in massive content distribution. Computer Networks (Amsterdam, Netherlands: 1999), 53(1):110–125, January 16, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBa

[1785] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(1):??, January 16, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bianco:2009:GES

[1786] Andrea Bianco, Achille Pattavina, and Alfio Lombardo. Guest editorial for the special issue. Computer Networks (Amsterdam, Netherlands: 1999), 53 (2):127–129, February 13, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rossi:2009:USS

[1787] Dario Rossi, Marco Mellia, and Michela Meo. Understanding Skype signaling. Computer Networks (Amsterdam, Netherlands: 1999), 53(2):130– 140, February 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tinnirello:2009:SEP

[1788] Ilenia Tinnirello, Domenico Giustiniano, Luca Scalia, and Giuseppe Bianchi. On the side-effects of proprietary solutions for fading and interference mitigation in IEEE 802.11b/g outdoor links. Computer Networks (Amsterdam, Netherlands: 1999), 53(2): 141–152, February 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

He:2009:PES

[1789] Jiayue He, Augustin Chaintreau, and Christophe Diot. A performance evaluation of scalable live video streaming with nano data centers. Computer Networks (Amsterdam, Netherlands: 1999), 53(2):153–167, February 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Petracca:2009:HHS

[1790] Michele Petracca, Robert Birke, and Andrea Bianco. HERO: High-speed enhanced routing operation in Ethernet NICs for software routers. Computer Networks (Amsterdam, Netherlands: 1999), 53(2):168–179, February 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Capone:2009:RRO

[1791] Antonio Capone, Jocelyne Elias, and Fabio Martignon. Routing and resource optimization in service overlay networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(2): 180–190, February 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bermolen:2009:SVR

[1792] Paola Bermolen and Dario Rossi. Support vector regression for link load prediction. Computer Networks (Amsterdam, Netherlands: 1999), 53(2): 191–201, February 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Eramo:2009:PLA

[1793] V. Eramo, A. Germoni, C. Raffaelli, and M. Savi. Packet loss analysis of shared-per-wavelength multi-fiber alloptical switch with parallel scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 53(2):202– 216, February 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bernaschi:2009:OOS

[1794] M. Bernaschi, F. Cacace, G. Iannello, M. Vellucci, and L. Vollero. OpenCAP-WAP: An open source CAPWAP implementation for the management and configuration of WiFi hot-spots. Computer Networks (Amsterdam, Netherlands: 1999), 53(2):217–230, February 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Giacomazzi:2009:RAA

[1795] Paolo Giacomazzi and Gabriella Saddemi. Resource allocation and admission control for the provisioning of quality of service in networks of static priority schedulers. Computer Networks (Amsterdam, Netherlands: 1999), 53(2):231–243, February 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBb

[1796] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(2):??, February 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Akyildiz:2009:E

[1797] Ian F. Akyildiz and Harry Rudin. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 53(3): 245, February 27, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mountrouidou:2009:DPB

[1798] Xenia Mountrouidou and Harry Perros. On the departure process of burst aggregation algorithms in optical burst switching. Computer Networks (Amsterdam, Netherlands: 1999), 53(3): 247–264, February 27, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Stassopoulou:2009:WRD

[1799] Athena Stassopoulou and Marios D. Dikaiakos. Web robot detection: A probabilistic reasoning approach. Computer Networks (Amsterdam, Netherlands: 1999), 53(3):265–278, February 27, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

He:2009:RDB

[1800] Xinming He, Christos Papadopoulos, John Heidemann, Urbashi Mitra, and Usman Riaz. Remote detection of bottleneck links using spectral and statistical methods. Computer Networks (Amsterdam, Netherlands: 1999), 53 (3):279–298, February 27, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ibrahim:2009:AFD

[1801] Marc Ibrahim, Kinda Khawam, Abed Ellatif Samhat, and Samir Tohme. Analytical framework for dimensioning hierarchical WiMax–WiFi networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(3): 299–309, February 27, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Barlet-Ros:2009:RNM

[1802] Pere Barlet-Ros, Gianluca Iannaccone, Josep Sanjuàs-Cuxart, and Josep Solé-Pareta. Robust network monitoring in the presence of non-cooperative traffic queries. Computer Networks (Amsterdam, Netherlands: 1999), 53(3): 310–321, February 27, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kong:2009:DDT

[1803] Peng-Yong Kong and Inn-Inn Er.

DTSMA: Distributed time-spread multiple access for wireless mesh networks with IEEE 802.16d MAC protocol. Computer Networks (Amsterdam, Netherlands: 1999), 53(3):322– 337, February 27, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2009:SAM

[1804] Lijun Wang, Lin Cai, Xinzhi Liu, Xuemin (Sherman) Shen, and Junshan Zhang. Stability analysis of multiple-bottleneck networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(3):338–352, February 27, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yan:2009:KHH

[1805] Junzhi Yan, Jianfeng Ma, and Hongyue Liu. Key hierarchies for hierarchical access control in secure group communications. Computer Networks (Amsterdam, Netherlands: 1999), 53(3): 353–364, February 27, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cai:2009:SCO

[1806] Han Cai, Do Young Eun, Sangtae Ha, Injong Rhee, and Lisong Xu. Stochastic convex ordering for multiplicative decrease Internet congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 53(3):365– 381, February 27, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Demirkol:2009:IRP

[1807] Ilker Demirkol, Cem Ersoy, Fatih Alagöz, and Hakan Deliç. The impact of a realistic packet traffic model on the performance of surveillance wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (3):382–399, February 27, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fan:2009:GRT

[1808] Xing Fan, Magnus Jonsson, and Jan Jonsson. Guaranteed real-time communication in packet-switched networks with FCFS queuing. Computer Networks (Amsterdam, Netherlands: 1999), 53(3):400–417, February 27, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Derbel:2009:AAN

[1809] Hajer Derbel, Nazim Agoulmine, and Mikaël Salaün. ANEMA: Autonomic network management architecture to support self-configuration and selfoptimization in IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(3):418–430, February 27, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBc

[1810] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(3):??, February 27, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Canonico:2009:CDI

[1811] Roberto Canonico, Carmen Guerrero, and Andreas Mauthe. Content distribution infrastructures for community networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(4): 431–433, March 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Skevik:2009:ECP

[1812] Karl-André Skevik, Vera Goebel, and Thomas Plagemann. Evaluation of a comprehensive P2P video-on-demand streaming system. Computer Networks (Amsterdam, Netherlands: 1999), 53 (4):434–455, March 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Vassilakis:2009:IPS

[1813] Constantinos Vassilakis, Nikolaos Laoutaris, and Ioannis Stavrakakis. On the impact of playout scheduling on the performance of peer-to-peer live streaming. Computer Networks (Amsterdam, Netherlands: 1999), 53(4): 456–469, March 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Silverston:2009:TAP

[1814] Thomas Silverston, Olivier Fourmaux, Alessio Botta, Alberto Dainotti, Antonio Pescapé, Giorgio Ventre, and Kavé Salamatian. Traffic analysis of peer-to-peer IPTV communities. Computer Networks (Amsterdam, Netherlands: 1999), 53(4):470–484, March 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bolla:2009:MSS

[1815] R. Bolla, R. Gaeta, A. Magnetto, M. Sciuto, and M. Sereno. A measurement study supporting P2P filesharing community models. Computer Networks (Amsterdam, Netherlands: 1999), 53(4):485–500, March 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zink:2009:CYN

[1816] Michael Zink, Kyoungwon Suh, Yu Gu, and Jim Kurose. Characteristics of YouTube network traffic at a campus network – Measurements, models, and implications. Computer Networks (Amsterdam, Netherlands: 1999), 53(4): 501–514, March 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Andrade:2009:RDS

[1817] Nazareno Andrade, Elizeu Santos-Neto, Francisco Brasileiro, and Matei Ripeanu. Resource demand and supply in BitTorrent content-sharing communities. Computer Networks (Amsterdam, Netherlands: 1999), 53(4): 515–527, March 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cuevas:2009:FHP

[1818] Ruben Cuevas, Albert Cabellos-Aparicio, Angel Cuevas, Jordi Domingo-Pascual, and Arturo Azcorra. fP2P– HN: A P2P-based route optimization architecture for mobile IP-based community networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (4):528–540, March 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2009:PED

[1819] P. Yang and M. Chuah. Performance evaluations of data-centric information retrieval schemes for DTNs. Computer Networks (Amsterdam, Netherlands: 1999), 53(4):541–555, March 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Martinez-Yelmo:2009:HPI

[1820] Isaias Martinez-Yelmo, Alex Bikfalvi, Ruben Cuevas, Carmen Guerrero, and Jaime Garcia. H-P2PSIP: Interconnection of P2PSIP domains for global multimedia services based on a hierarchical DHT overlay network. Computer Networks (Amsterdam, Netherlands: 1999), 53(4):556–568, March 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBd

[1821] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(4):??, March 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Macedo:2009:DLE

[1822] Mário Macedo, António Grilo, and Mário Nunes. Distributed Latency-Energy Minimization and interference avoidance in TDMA Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(5): 569–582, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Psaras:2009:PAI

[1823] Ioannis Psaras and Vassilis Tsaoussidis. On the properties of an additive increase rate accelerator. Computer Networks (Amsterdam, Netherlands: 1999), 53(5):583–595, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2009:FMM

[1824] Y. Wang, M.Ü. Uyar, S. S. Batth, and M. A. Fecko. Fault masking by multiple timing faults in timed EFSM models. Computer Networks (Amsterdam, Netherlands: 1999), 53 (5):596–612, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nayebi:2009:ANT

[1825] A. Nayebi and H. Sarbazi-Azad. Analysis of k-Neigh topology control protocol for mobile wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(5):613–633, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ratnam:2009:DSI

[1826] Krishanthmohan Ratnam, Mohan Gurusamy, and Luying Zhou. Differentiated survivability with improved fairness in IP/MPLS-over-WDM optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (5):634–649, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chrysostomou:2009:FEM

[1827] C. Chrysostomou, A. Pitsillides, and Y. A. Sekercioglu. Fuzzy explicit marking: A unified congestion controller for Best-Effort and Diff-Serv networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(5): 650–667, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gilly:2009:ABM

[1828] Katja Gilly, Salvador Alcaraz, Carlos Juiz, and Ramon Puigjaner. Analysis of burstiness monitoring and detection in an adaptive Web system. Computer Networks (Amsterdam, Netherlands: 1999), 53(5):668–679, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hierons:2009:OCP

[1829] R. M. Hierons and H. Ural. Overcoming controllability problems with fewest channels between testers. Computer Networks (Amsterdam, Netherlands: 1999), 53(5):680–690, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sue:2009:URO

[1830] Chuan-Ching Sue and Jing-Ying Yeh. Utilize 100% redundancy to offer higher-level multiple fault restoration in WDM networks without wavelength conversion. Computer Networks (Amsterdam, Netherlands: 1999), 53(5): 691–705, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Blefari-Melazzi:2009:ESD

[1831] N. Blefari-Melazzi, J. N. Daigle, and M. Femminella. Efficient and stateless deployment of VoIP services. Computer Networks (Amsterdam, Netherlands: 1999), 53(5):706–726, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bartolini:2009:STS

[1832] Novella Bartolini, Giancarlo Bongiovanni, and Simone Silvestri. Self-* through self-learning: Overload control for distributed web systems. Computer Networks (Amsterdam, Netherlands: 1999), 53(5):727–743, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rahbar:2009:DBB

[1833] Akbar Ghaffar Pour Rahbar and Oliver W. W. Yang. Distribution-based bandwidth access scheme in slotted all-optical packet-switched networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(5): 744–758, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBe

[1834] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(5):??, April 9, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mellia:2009:TCA

[1835] Marco Mellia, Antonio Pescapè, and Luca Salgarelli. Traffic classification and its applications to modern networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(6): 759–760, April 23, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Palmieri:2009:NRB

[1836] Francesco Palmieri and Ugo Fiore. A nonlinear, recurrence-based approach to traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 53(6):761–773, April 23, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Apiletti:2009:CNT

[1837] Daniele Apiletti, Elena Baralis, Tania Cerquitelli, and Vincenzo D’Elia. Characterizing network traffic by means of the NetMine framework. Computer Networks (Amsterdam, Netherlands: 1999), 53(6):774–789, April 23, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2009:EAI

[1838] Wei Li, Marco Canini, Andrew W.

Moore, and Raffaele Bolla. Efficient application identification and the temporal and spatial stability of classification schema. Computer Networks (Amsterdam, Netherlands: 1999), 53 (6):790–809, April 23, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sass:2009:ASH

[1839] Detlef Saß, Simon Hauger, and Martin Köhn. Architecture and scalability of a high-speed traffic measurement platform with a highly flexible packet classification. Computer Networks (Amsterdam, Netherlands: 1999), 53(6): 810–820, April 23, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2009:SPC

[1840] Pi-Chung Wang. Scalable packet classification with controlled crossproducting. Computer Networks (Amsterdam, Netherlands: 1999), 53(6): 821–834, April 23, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Marsono:2009:TSC

[1841] Muhammad N. Marsono, M. Watheq El-Kharashi, and Fayez Gebali. Targeting spam control on middleboxes: Spam detection based on layer-3 email content classification. Computer Networks (Amsterdam, Netherlands: 1999), 53(6):835–848, April 23, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hu:2009:PIP

[1842] Yan Hu, Dah-Ming Chiu, and John C. S. Lui. Profiling and identification of P2P traffic. Computer Networks (Amsterdam, Netherlands: 1999), 53 (6):849–863, April 23, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Perdisci:2009:MMC

[1843] Roberto Perdisci, Davide Ariu, Prahlad Fogla, Giorgio Giacinto, and Wenke Lee. McPAD: A multiple classifier system for accurate payloadbased anomaly detection. Computer Networks (Amsterdam, Netherlands: 1999), 53(6):864–881, April 23, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBf

[1844] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(6):??, April 23, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2009:CAU

[1845] Xin Wang, J. J. Garcia-Luna-Aceves, and Hamid R. Sadjadpour. Channel access using opportunistic reservations and virtual MIMO. Computer Networks (Amsterdam, Netherlands: 1999), 53(7):883–895, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Park:2009:FAR

[1846] Kyung-Joon Park, Hyuk Lim, Jennifer C. Hou, and Chong-Ho Choi. Feedback-assisted robust estimation of available bandwidth. Computer Networks (Amsterdam, Netherlands: 1999), 53(7):896–912, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Youssef:2009:GBC

[1847] Mina Youssef and Caterina Scoglio. On graph-based characteristics of optimal overlay topologies. Computer Networks (Amsterdam, Netherlands: 1999), 53 (7):913–925, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Azodolmolky:2009:SPL

[1848] Siamak Azodolmolky, Mirosław Klinkowski, Eva Marin, Davide Careglio, Josep Solé Pareta, and Ioannis Tomkos. A survey on physical layer impairments aware routing and wavelength assignment algorithms in optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(7):926–944, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Al-Karaki:2009:DAR

[1849] Jamal N. Al-Karaki, Raza Ul-Mustafa, and Ahmed E. Kamal. Data aggregation and routing in Wireless Sensor Networks: Optimal and heuristic algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 53 (7):945–960, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Karapistoli:2009:MPL

[1850] Eirini Karapistoli, Ioannis Gragopoulos, Ioannis Tsetsinas, and Fotini-Niovi Pavlidou. A MAC protocol for lowrate UWB wireless sensor networks using directional antennas. Computer Networks (Amsterdam, Netherlands: 1999), 53(7):961–972, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Martinez-Bauset:2009:MCM

[1851] Jorge Martinez-Bauset, David Garcia-Roger, M. Jose Domenech-Benlloch, and Vicent Pla. Maximizing the capacity of mobile cellular networks with heterogeneous traffic. Computer Networks (Amsterdam, Netherlands: 1999), 53 (7):973–988, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Islam:2009:MRA

[1852] Salekul Islam and J. William Atwood.

Multicast receiver access control by IGMP-AC. Computer Networks (Amsterdam, Netherlands: 1999), 53(7): 989–1013, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Harada:2009:PCP

[1853] Junko Harada, Shigeo Shioda, and Hiroshi Saito. Path coverage properties of randomly deployed sensors with finite data-transmission ranges. Computer Networks (Amsterdam, Netherlands: 1999), 53(7):1014–1026, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2009:TEE

[1854] Yueping Zhang, Saurabh Jain, and Dmitri Loguinov. Towards experimental evaluation of explicit congestion control. Computer Networks (Amsterdam, Netherlands: 1999), 53(7): 1027–1039, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bianco:2009:MSM

[1855] Andrea Bianco and Alessandra Scicchitano. Multicast support in multi-chip centralized schedulers in Input Queued switches. Computer Networks (Amsterdam, Netherlands: 1999), 53(7): 1040–1049, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Caro:2009:ELS

[1856] Luis F. Caro, Dimitri Papadimitriou, and Jose L. Marzo. Enhancing label space usage for Ethernet VLANlabel switching. Computer Networks (Amsterdam, Netherlands: 1999), 53 (7):1050–1061, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kovacs:2009:CLO

[1857] Ákos Kovács and István Gódor. Crosslayer optimized wireless multicast for layered media. Computer Networks (Amsterdam, Netherlands: 1999), 53 (7):1062–1072, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sbai:2009:EED

[1858] Mohamed Karim Sbai and Chadi Barakat. Experiences on enhancing data collection in large networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(7):1073–1086, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Huang:2009:QAV

[1859] Chenn-Jung Huang, Kai-Wen Hu, You-Jia Chen, Chun-Hua Chen, and Yun-Cheng Luo. A QoS-aware VoD resource sharing scheme for heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (7):1087–1098, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Addie:2009:PAP

[1860] Ronald G. Addie, Timothy D. Neame, and Moshe Zukerman. Performance analysis of a Poisson–Pareto queue over the full range of system parameters. Computer Networks (Amsterdam, Netherlands: 1999), 53(7): 1099–1113, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2009:LDB

[1861] Bang Wang, Cheng Fu, and Hock Beng Lim. Layered diffusion-based coverage control in Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(7): 1114–1124, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mohorko:2009:MMI

[1862] Jože Mohorko and Matjaž Fras. Mathematical model of IRIS replication mechanism for the simulation of tactical networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (7):1125–1136, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Markovich:2009:CSB

[1863] Natalia M. Markovich, Astrid Undheim, and Peder J. Emstad. Classification of slice-based VBR video traffic and estimation of link loss by exceedance. Computer Networks (Amsterdam, Netherlands: 1999), 53(7): 1137–1153, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBg

[1864] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(7):??, May 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:PMC

[1865] Anonymous. “Performance Modeling of Computer Networks: Special Issue in Memory of Dr. Gunter Bolch” (eulogies and editorials). Computer Networks (Amsterdam, Netherlands: 1999), 53 (8):1155–1157, June 11, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gelabert:2009:OPS

[1866] Xavier Gelabert, Ian F. Akyildiz, Oriol Sallent, and Ramon Agustí. Operating point selection for primary and secondary users in cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(8): 1158–1170, June 11, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Herpel:2009:SDP

[1867] Thomas Herpel, Kai-Steffen Hielscher, Ulrich Klehmet, and Reinhard German. Stochastic and deterministic performance evaluation of automotive CAN communication. Computer Networks (Amsterdam, Netherlands: 1999), 53(8):1171–1185, June 11, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Al-Begain:2009:IMG

[1868] Khalid Al-Begain, Alexander Dudin, Arseniy Kazimirsky, and Suleiman Yerima. Investigation of the M_{2} / G_{2} / 1 / ∞, N queue with restricted admission of priority customers and its application to HSDPA mobile systems. Computer Networks (Amsterdam, Netherlands: 1999), 53(8):1186–1201, June 11, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Heindl:2009:AFP

[1869] Armin Heindl and Gilles Pokam. An analytic framework for performance modeling of software transactional memory. Computer Networks (Amsterdam, Netherlands: 1999), 53(8): 1202–1214, June 11, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Heegaard:2009:NSM

[1870] Poul E. Heegaard and Kishor S.

Trivedi. Network survivability modeling. Computer Networks (Amsterdam, Netherlands: 1999), 53(8): 1215–1234, June 11, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Horvath:2009:TBD

[1871] András Horváth, Gábor Horváth, and Miklós Telek. A traffic based decomposition of two-class queueing networks with priority service. Computer Networks (Amsterdam, Netherlands: 1999), 53(8):1235–1248, June 11, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mylosz:2009:BSS

[1872] J. Mylosz and H. Daduna. On the behavior of stable subnetworks in non-ergodic networks with unreliable nodes. Computer Networks (Amsterdam, Netherlands: 1999), 53(8): 1249–1263, June 11, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wuchner:2009:FSM

[1873] Patrick Wüchner, János Sztrik, and Hermann de Meer. Finite-source M / M / S retrial queue with search for balking and impatient customers from the orbit. Computer Networks (Amsterdam, Netherlands: 1999), 53(8): 1264–1273, June 11, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ludtke:2009:MPC

[1874] D. Lüdtke and D. Tutsch. The modeling power of CINSim: Performance evaluation of interconnection networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(8): 1274–1288, June 11, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bazan:2009:ATA

[1875] Peter Bazan and Reinhard German. Approximate transient analysis of large stochastic models with WinPEPSY-QNS. Computer Networks (Amsterdam, Netherlands: 1999), 53(8): 1289–1301, June 11, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBh

[1876] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(8):??, June 11, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhao:2009:FSE

[1877] Yao Zhao and Yan Chen. FAD and SPA: End-to-end link-level loss rate inference without infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 53(9):1303–1318, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2009:DBT

[1878] Mingyu Chen, Jinsong Zhang, Manohar N.

Murthi, and Kamal Premaratne. Delay-based TCP congestion avoidance: A network calculus interpretation and performance improvements. Computer Networks (Amsterdam, Netherlands: 1999), 53(9): 1319–1340, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Saidi:2009:USR

[1879] Mohand Yazid Saidi, Bernard Cousin, and Jean-Louis Le Roux. Using Shared Risk Link Groups to enhance backup path computation. Computer Networks (Amsterdam, Netherlands: 1999), 53 (9):1341–1353, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fortuna:2009:TDC

[1880] Carolina Fortuna and Mihael Mohorcic. Trends in the development of communication networks: Cognitive networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(9): 1354–1376, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Krishna:2009:SMP

[1881] T. Vamsi Krishna and Amitabha Das. A survey on MAC protocols in OSA networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(9): 1377–1394, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shu:2009:EEP

[1882] Tao Shu and Marwan Krunz. Energyefficient power/rate control and scheduling in hybrid TDMA/CDMA wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(9):1395–1408, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ozkasap:2009:EBR

[1883] Oznur Ozkasap, Zulkuf Genc, and Emre Atsan. Epidemic-based reliable and adaptive multicast for mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (9):1409–1430, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zeinalipour-Yazti:2009:FHR

[1884] Demetrios Zeinalipour-Yazti, Zografoula Vagena, Vana Kalogeraki, Dimitrios Gunopulos, Vassilis J. Tsotras, Michail Vlachos, Nick Koudas, and Divesh Srivastava. Finding the K highest-ranked answers in a distributed network. Computer Networks (Amsterdam, Netherlands: 1999), 53(9): 1431–1449, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rodrigo:2009:MBP

[1885] Miguel de Vega Rodrigo, Guy Latouche, and Marie-Ange Remiche. Modeling bufferless packet-switching networks with packet dependencies. Computer Networks (Amsterdam, Netherlands: 1999), 53(9):1450–1466, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Saidi:2009:PBH

[1886] Mohand Yazid Saidi, Bernard Cousin, and Jean-Louis Le Roux. PLRbased heuristic for backup path computation in MPLS networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(9):1467–1479, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Han:2009:ONS

[1887] Seung Chul Han and Ye Xia. Optimal node-selection algorithm for parallel download in overlay contentdistribution networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(9):1480–1496, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Korcak:2009:VTD

[1888] Ömer Korçak and Fatih Alagöz. Virtual topology dynamics and handover mechanisms in Earth-fixed LEO satellite systems. Computer Networks (Amsterdam, Netherlands: 1999), 53(9):

1497–1511, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2009:PAR

[1889] Haodong Wang, Bo Sheng, and Qun Li. Privacy-aware routing in sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(9): 1512–1529, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Alouneh:2009:NPP

[1890] Sahel Alouneh, Anjali Agarwal, and Abdeslam En-Nouaary. A novel path protection scheme for MPLS networks using multi-path routing. Computer Networks (Amsterdam, Netherlands: 1999), 53(9):1530–1545, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2009:SSS

[1891] Yu-Wei Chen and Chao-Yao Hsieh. SingBroad: A scheduling scheme for broadcasting continuous multimedia data over a single channel. Computer Networks (Amsterdam, Netherlands: 1999), 53(9):1546–1554, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pallas:2009:HTH

[1892] G. D. Pallas, G. I. Papadimitriou, and A. S. Pomportsis. Hybrid TDMA with heuristic traffic shaping: An efficient bandwidth allocation approach for heavy loaded LANs. Computer Networks (Amsterdam, Netherlands: 1999), 53(9):1555–1566, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBi

[1893] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(9):??, June 25, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dobson:2009:EAS

[1894] Simon Dobson, John Strassner, and Hermann de Meer. Editorialautonomic and self-organising systems. Computer Networks (Amsterdam, Netherlands: 1999), 53(10): 1567–1569, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sadjadi:2009:TAC

[1895] S. Masoud Sadjadi and Philip K.

McKinley. Transparent autonomization in CORBA. Computer Networks (Amsterdam, Netherlands: 1999), 53 (10):1570–1586, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Latre:2009:AAO

[1896] Steven Latré, Pieter Simoens, Bart De Vleeschauwer, Wim Van de Meerssche, Filip De Turck, Bart Dhoedt, Piet Demeester, Steven Van den Berghe, and Edith Gilon de Lumley. An autonomic architecture for optimizing QoE in multimedia access networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(10):1587–1602, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mukherjee:2009:SME

[1897] Tridib Mukherjee, Georgios Varsamopoulos, and Sandeep K. S. Gupta. Selfmanaging energy-efficient multicast support in MANETs under end-to-end reliability constraints. Computer Networks (Amsterdam, Netherlands: 1999), 53(10):1603–1627, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dragoni:2009:SPS

[1898] Nicola Dragoni, Fabio Massacci, and Ayda Saidane. A self-protecting and self-healing framework for negotiating services and trust in autonomic communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 53(10):1628–1648, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jiang:2009:MDS

[1899] Shanshan Jiang, Yuan Xue, and Douglas C. Schmidt. Minimum disruption service composition and recovery in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(10):1649–1665, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Balasubramaniam:2009:PCB

[1900] Sasitharan Balasubramaniam, Dmitri Botvich, Brendan Jennings, Steven Davy, William Donnelly, and John Strassner. Policy-constrained bioinspired processes for autonomic route management. Computer Networks (Amsterdam, Netherlands: 1999), 53 (10):1666–1682, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xu:2009:NTM

[1901] Hui Xu and J. J. Garcia-Luna-Aceves. Neighborhood tracking for mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (10):1683–1696, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Watteyne:2009:CVC

[1902] Thomas Watteyne, Isabelle Augé-Blum, Mischa Dohler, Stéphane Ubéda, and Dominique Barthel. Centroid virtual coordinates – A novel nearshortest path routing paradigm. Computer Networks (Amsterdam, Netherlands: 1999), 53(10):1697–1711, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Poggi:2009:SAU

[1903] Nicolas Poggi, Toni Moreno, Josep Lluis Berral, Ricard Gavaldà, and Jordi Torres. Self-adaptive utility-based web session management. Computer Networks (Amsterdam, Netherlands: 1999), 53 (10):1712–1721, July 14, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rodero-Merino:2009:SMT

[1904] Luis Rodero-Merino, Antonio Fernández Anta, Luis López, and Vicent Cholvi. Self-managed topologies in P2P networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(10): 1722–1736, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dressler:2009:RBS

[1905] Falko Dressler, Isabel Dietrich, Reinhard German, and Bettina Krüger. A rule-based system for programming self-organized sensor and actor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(10): 1737–1750, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Altman:2009:ETP

[1906] Eitan Altman, Rachid El-Azouzi, Yezekael Hayel, and Hamidou Tembine. The evolution of transport protocols: An evolutionary game perspective. Computer Networks (Amsterdam, Netherlands: 1999), 53(10): 1751–1759, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Streitberger:2009:SES

[1907] Werner Streitberger and Torsten Eymann. A simulation of an economic, self-organising resource allocation approach for application layer networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(10): 1760–1770, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBj

[1908] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(10):??, July 14, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Menth:2009:TCR

[1909] Michael Menth and Matthias Hartmann. Threshold configuration and routing optimization for PCN-based resilient admission control. Computer Networks (Amsterdam, Netherlands: 1999), 53(11):1771–1783, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xiong:2009:DAS

[1910] Naixue Xiong, Yi Pan, Xiaohua Jia, Jong Hyuk Park, and Yingshu Li. Design and analysis of a self-tuning feedback controller for the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 53(11):1784–1797, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yaghmaee:2009:PBR

[1911] Mohammad Hossein Yaghmaee and Donald A. Adjeroh. Priority-based rate control for service differentiation and congestion control in wireless multimedia sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(11):1798–1811, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zarifzadeh:2009:JRA

[1912] Sajjad Zarifzadeh, Amir Nayyeri, Nasser Yazdani, Ahmad Khonsari, and Hamid Hajabdolali Bazzaz. Joint range assignment and routing to conserve energy in wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(11):1812–1829, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Urdaneta:2009:WWA

[1913] Guido Urdaneta, Guillaume Pierre, and Maarten van Steen. Wikipedia workload analysis for decentralized hosting. Computer Networks (Amsterdam, Netherlands: 1999), 53(11): 1830–1845, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). See corrigendum

Chang:2009:PAM

[1914] Moonjeong Chang, Hyunjeong Lee, and Meejeong Lee. A per-application mobility management platform for application-specific handover decision in overlay networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(11):1846–1858, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pereira:2009:FSD

[1915] Flávio de Melo Pereira, Nelson L. S. da Fonseca, and Dalton S. Arantes. A fair scheduling discipline for Ethernet passive optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(11):1859–1878, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kim:2009:EET

[1916] Jaesub Kim and Kyu Ho Park. An energy-efficient, transport-controlled MAC protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(11): 1879–1902, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cho:2009:UCD

[1917] Soohyun Cho and Riccardo Bettati. Use of competition detection in TCP for simple topology networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(11):1903–1925, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rosenberg:2009:MHR

[1918] Eric Rosenberg. Minimizing hierarchical routing error. Computer Networks (Amsterdam, Netherlands: 1999), 53 (11):1926–1938, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Catanuto:2009:AOR

[1919] Roberto Catanuto, Stavros Toumpis, and Giacomo Morabito. On asymptotically optimal routing in large wireless networks and Geometrical Optics analogy. Computer Networks (Amsterdam, Netherlands: 1999), 53(11): 1939–1955, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gan:2009:NPS

[1920] Chai-Hien Gan, Phone Lin, and Chung-Min Chen. A novel prebuffering scheme for IPTV service. Computer Networks (Amsterdam, Netherlands: 1999), 53(11):1956–1966, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBk

[1921] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(11):??, July 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wong:2009:QQB

[1922] May Wong and Demet Aksoy. QUAD:

Quadrant-based relative location estimates for representative topologies in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(12):1967–1979, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nakamura:2009:RRA

[1923] Eduardo F. Nakamura, Heitor S. Ramos, Leandro A. Villas, Horacio A. B. F. de Oliveira, Andre L. L. de Aquino, and Antonio A. F. Loureiro. A reactive role assignment for data routing in event-based wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(12): 1980–1996, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yun:2009:BCM

[1924] Jeongkyun Yun, Jin-Ghoo Choi, and Saewoong Bahk. A backwardcompatible multiple-round collision avoidance scheme for contention based medium access control. Computer Networks (Amsterdam, Netherlands: 1999), 53(12):1997–2010, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jiang:2009:FCG

[1925] Yuming Jiang, Qinghe Yin, Yong Liu, and Shengming Jiang. Fundamental calculus on generalized stochastically bounded bursty traffic for communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (12):2011–2021, August 13, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ozdemir:2009:SDA

[1926] Suat Ozdemir and Yang Xiao. Secure data aggregation in wireless sensor networks: A comprehensive overview. Computer Networks (Amsterdam, Netherlands: 1999), 53(12): 2022–2037, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Garcia:2009:PAI

[1927] Roberto García, Xabiel G. Pañeda, David Melendi, and Victor Garcia. Probabilistic analysis and interdependence discovery in the user interactions of a video news on demand service. Computer Networks (Amsterdam, Netherlands: 1999), 53(12):2038– 2049, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Karapantazis:2009:VCS

[1928] Stylianos Karapantazis and Fotini-Niovi Pavlidou. VoIP: A comprehensive survey on a promising technology. Computer Networks (Amsterdam, Netherlands: 1999), 53(12): 2050–2090, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Belbekkouche:2009:NRL

[1929] Abdeltouab Belbekkouche, Abdelhakim Hafid, and Michel Gendreau. Novel reinforcement learning-based approaches to reduce loss probability in buffer-less OBS networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(12):2091–2105, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Demirkol:2009:EDO

[1930] Ilker Demirkol and Cem Ersoy. Energy and delay optimized contention for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(12):2106–2119, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xu:2009:ASP

[1931] Yuedong Xu, John C. S. Lui, and Dah-Ming Chiu. Analysis and scheduling of practical network coding in OFDMA relay networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (12):2120–2139, August 13, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

He:2009:RLS

[1932] Yuan He, Hao Ren, Yunhao Liu, and Baijian Yang. On the reliability of large-scale distributed systems – A topological view. Computer Networks (Amsterdam, Netherlands: 1999), 53 (12):2140–2152, August 13, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Berzin:2009:MLB

[1933] Oleg Berzin. Mobility label based network: Hierarchical mobility management and packet forwarding architecture. Computer Networks (Amsterdam, Netherlands: 1999), 53(12): 2153–2181, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2009:HSS

[1934] Chao-Chin Yang. High speed and secure optical CDMA-based passive optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (12):2182–2191, August 13, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Martignon:2009:DIM

[1935] Fabio Martignon, Stefano Paris, and Antonio Capone. Design and implementation of MobiSEC: A complete security architecture for wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(12): 2192–2207, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Franklin:2009:IMC

[1936] A. Antony Franklin, Vibhav Bukkapatanam, and C. Siva Ram Murthy. Impact of multiple channels and radios on the performance of a TDMA based wireless mesh network. Computer Networks (Amsterdam, Netherlands: 1999), 53(12):2208–2224, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jeng:2009:RCA

[1937] Andy An-Kai Jeng and Rong-Hong Jan. Role and channel assignments for wireless mesh networks using hybrid approach. Computer Networks (Amsterdam, Netherlands: 1999), 53(12): 2225–2240, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bremler-Barr:2009:BOB

[1938] Anat Bremler-Barr, Nir Chen, Jussi Kangasharju, Osnat Mokryn, and Yuval Shavitt. Bringing order to BGP: Decreasing time and message complexity. Computer Networks (Amsterdam, Netherlands: 1999), 53(12): 2241–2256, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBl

[1939] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(12):??, August 13, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kermarrec:2009:E

[1940] Anne-Marie Kermarrec and Maarten van Steen. Editorial. Computer Networks (Amsterdam, Netherlands: 1999), 53(13):2257–2258, August 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ozkasap:2009:SFS

[1941] Oznur Ozkasap, Mine Caglar, Emrah Cem, Emrah Ahi, and Emre Iskender. Stepwise fair-share buffering for gossip-based peer-to-peer data dissemination. Computer Networks (Amsterdam, Netherlands: 1999), 53(13): 2259–2274, August 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tran-Quang:2009:NGB

[1942] Vinh Tran-Quang and Takumi Miyoshi. A novel gossip-based sensing coverage algorithm for dense wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(13): 2275–2287, August 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bakhshi:2009:AMI

[1943] Rena Bakhshi, Daniela Gavidia, Wan Fokkink, and Maarten van Steen. An analytical model of information dissemination for a gossip-based protocol. Computer Networks (Amsterdam, Netherlands: 1999), 53(13):2288– 2303, August 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Patel:2009:REI

[1944] Jay A. Patel, Étienne Rivière, Indranil Gupta, and Anne-Marie Kermarrec. Rappel: Exploiting interest and network locality to improve fairness in publish-subscribe systems. Computer Networks (Amsterdam, Netherlands: 1999), 53(13):2304–2320, August 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jelasity:2009:MGB

[1945] Márk Jelasity, Alberto Montresor, and Ozalp Babaoğlu. T-Man: Gossipbased fast overlay topology construction. Computer Networks (Amsterdam, Netherlands: 1999), 53(13):2321– 2339, August 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bortnikov:2009:BBR

[1946] Edward Bortnikov, Maxim Gurevich, Idit Keidar, Gabriel Kliot, and Alexander Shraer. Brahms: Byzantine resilient random membership sampling. Computer Networks (Amsterdam, Netherlands: 1999), 53(13):2340– 2359, August 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Morales:2009:AAA

[1947] Ramsés Morales and Indranil Gupta.

AVCOL: Availability-aware information aggregation in large distributed systems under uncollaborative behavior. Computer Networks (Amsterdam, Netherlands: 1999), 53(13):2360– 2372, August 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Keidar:2009:ESM

[1948] Idit Keidar, Roie Melamed, and Ariel Orda. EquiCast: Scalable multicast with selfish users. Computer Networks (Amsterdam, Netherlands: 1999), 53 (13):2373–2386, August 28, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBm

[1949] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(13):??, August 28, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Artishchev-Zapolotsky:2009:ILO

[1950] Maria Artishchev-Zapolotsky. Improved layout of the odd–even sorting network. Computer Networks (Amsterdam, Netherlands: 1999), 53(14):2387– 2395, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Luo:2009:TMB

[1951] Junhai Luo, Xue Liu, and Mingyu Fan. A trust model based on fuzzy recommendation for mobile ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(14):2396– 2407, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tsakountakis:2009:GAS

[1952] Alexandros Tsakountakis, Georgios Kambourakis, and Stefanos Gritzalis. A generic accounting scheme for next generation networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(14):2408–2426, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gobel:2009:IBC

[1953] Johannes Göbel, Anthony Krzesinski, and Michel Mandjes. Incentive-based control of ad hoc networks: A performance study. Computer Networks (Amsterdam, Netherlands: 1999), 53 (14):2427–2443, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Behsaz:2009:TIF

[1954] B. Behsaz, P. Gburzynski, and M. Mac-Gregor. Transport-independent fairness. Computer Networks (Amsterdam, Netherlands: 1999), 53(14):2444– 2457, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xiao:2009:TTD

[1955] Yang Xiao, Yanping Zhang, Miao Peng, Hui Chen, Xiaojiang Du, Bo Sun, and Kui Wu. Two and three-dimensional intrusion object detection under randomized scheduling algorithms in sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (14):2458–2475, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Este:2009:SVM

[1956] Alice Este, Francesco Gringoli, and Luca Salgarelli. Support Vector Machines for TCP traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 53(14):2476– 2490, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zang:2009:TMD

[1957] Hui Zang and Antonio Nucci. Traffic monitor deployment in IP networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(14):2491– 2501, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rad:2009:CAC

[1958] A. Hamed Mohsenian Rad and Vincent W. S. Wong. Congestion-aware channel assignment for multi-channel wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (14):2502–2516, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sekercioglu:2009:SMB

[1959] Y. Ahmet Şekercioğlu, Milosh Ivanovich, and Alper Yeğin. A survey of MAC based QoS implementations for WiMAX networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (14):2517–2536, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shah-Heydari:2009:HAD

[1960] Shahram Shah-Heydari and Oliver Yang. Heuristic algorithms for designing self-repairing protection trees in mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (14):2537–2551, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shifrin:2009:SBN

[1961] Mark Shifrin and Isaac Keslassy. Smallbuffer networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (14):2552–2565, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chang:2009:LAB

[1962] Ben-Jye Chang, Shu-Yu Lin, and Jun-Yu Jin. LIAD: Adaptive bandwidth prediction based Logarithmic Increase Adaptive Decrease for TCP congestion control in heterogeneous wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(14):2566– 2585, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Calafate:2009:MBT

[1963] Carlos T. Calafate, P. Manzoni, Juan-Carlos Cano, and M. P. Malumbres. Markovian-based traffic modeling for mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(14):2586–2600, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBn

[1964] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(14):??, September 18, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2009:DSA

[1965] Patrick P. C. Lee, Tian Bu, and Thomas Woo. On the detection of signaling DoS attacks on 3G/WiMax wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (15):2601–2616, October 12, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bergfeldt:2009:RTA

[1966] Erik Bergfeldt, Svante Ekelin, and Johan M. Karlsson. Real-time availablebandwidth estimation using filtering and change detection. Computer Networks (Amsterdam, Netherlands: 1999), 53(15):2617–2645, October 12, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wu:2009:SRP

[1967] Wenji Wu, Phil Demar, and Matt Crawford. Sorting reordered packets with interrupt coalescing. Computer Networks (Amsterdam, Netherlands: 1999), 53(15):2646–2662, October 12, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ssu:2009:IRD

[1968] Kuo-Feng Ssu, Chun-Hao Yang, Chih-Hsun Chou, and An-Kuo Yang. Improving routing distance for geographic multicast with Fermat points in mobile ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (15):2663–2673, October 12, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2009:AMC

[1969] Ian W. C. Lee and Abraham O. Fapojuwo. Analysis and modeling of a campus wireless network TCP/IP traffic. Computer Networks (Amsterdam, Netherlands: 1999), 53(15):2674– 2687, October 12, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Srivastava:2009:TEM

[1970] Shekhar Srivastava, Appie van de Liefvoort, and Deep Medhi. Traffic engineering of MPLS backbone networks in the presence of heterogeneous streams. Computer Networks (Amsterdam, Netherlands: 1999), 53(15): 2688–2702, October 12, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2009:MFC

[1971] Sirui Yang, Hai Jin, Bo Li, and Xiaofei Liao. A modeling framework of content pollution in Peer-to-Peer video streaming systems. Computer Networks (Amsterdam, Netherlands: 1999), 53(15): 2703–2715, October 12, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Antoniou:2009:ANS

[1972] Josephina Antoniou, Ioannis Koukoutsidis, Eva Jaho, Andreas Pitsillides, and Ioannis Stavrakakis. Access network synthesis game in next generation networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(15): 2716–2726, October 12, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hernandez-Orallo:2009:WSP

[1973] Enrique Hernández-Orallo and Joan Vila-Carbó. Web server performance analysis using histogram workload models. Computer Networks (Amsterdam, Netherlands: 1999), 53(15): 2727–2739, October 12, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Martin:2009:PBI

[1974] S. Martin and G. Leduc. A practical bytecode interpreter for programmable routers on IXP network processors. Computer Networks (Amsterdam, Netherlands: 1999), 53(15): 2740–2751, October 12, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBo

[1975] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(15):??, October 12, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gine:2009:MCO

[1976] Lluís Parcerisa Giné and Ian F. Akyildiz. Molecular communication options for long range nanonetworks. Computer Networks (Amsterdam, Netherlands: 1999), 53(16):2753–2766, November 10, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lee:2009:ITF

[1977] Jongmin Lee, Hojung Cha, and Rhan Ha. Improving TCP fairness and performance with bulk transmission control over lossy wireless channel. Computer Networks (Amsterdam, Netherlands: 1999), 53(16):2767– 2781, November 10, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cordeiro:2009:CLC

[1978] Weverton Luis da Costa Cordeiro, Guilherme Sperb Machado, Fabrício Girardi Andreis, Alan Diego dos Santos, Cristiano Bonato Both, Luciano Paschoal Gaspary, Lisandro Zambenedetti Granville, Claudio Bartolini, and David Trastour. Change Ledge: Change design and planning in networked systems based on reuse of knowledge and automation. Computer Networks (Amsterdam, Netherlands: 1999), 53(16):2782–2799, November 10, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2009:FPV

[1979] Alex X. Liu. Firewall policy verification and troubleshooting. Computer Networks (Amsterdam, Netherlands: 1999), 53(16):2800–2809, November 10, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Rosberg:2009:RCM

[1980] Zvi Rosberg and Fariza Sabrina. Rate control of multi class priority flows with end-to-end delay and rate constraints for QoS networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (16):2810–2824, November 10, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cha:2009:ESP

[1981] Meeyoung Cha, W. Art Chaovalitwongse, Jennifer Yates, Aman Shaikh, and Sue Moon. Efficient and scalable provisioning of always-on multicast streaming services. Computer Networks (Amsterdam, Netherlands: 1999), 53 (16):2825–2839, November 10, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Houngbadji:2009:SDA

[1982] Therence Houngbadji and Samuel Pierre. SubCast: A distributed addressing and routing system for large scale wireless sensor and actor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(16):2840– 2854, November 10, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bhardwaj:2009:MFB

[1983] Onkar Bhardwaj, G. V. V. Sharma, Manoj Panda, and Anurag Kumar. Modeling finite buffer effects on TCP traffic over an IEEE 802.11 infrastructure WLAN. Computer Networks (Amsterdam, Netherlands: 1999), 53(16): 2855–2869, November 10, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBp

[1984] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(16):??, November 10, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kant:2009:VDC

[1985] Krishna Kant and Bhuvan Urgaonkar. Virtualized Data Centers. Computer Networks (Amsterdam, Netherlands: 1999), 53(17):2871–2872, December 3, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Iyer:2009:VMM

[1986] Ravi Iyer, Ramesh Illikkal, Omesh Tickoo, Li Zhao, Padma Apparao, and Don Newell. VM^{3}: Measuring, modeling and managing VM shared resources. Computer Networks (Amsterdam, Netherlands: 1999), 53(17): 2873–2887, December 3, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mukherjee:2009:STT

[1987] Tridib Mukherjee, Ayan Banerjee, Georgios Varsamopoulos, Sandeep K. S. Gupta, and Sanjay Rungta. Spatio-temporal thermal-aware job scheduling to minimize energy consumption in virtualized heterogeneous data centers. Computer Networks (Amsterdam, Netherlands: 1999), 53(17): 2888–2904, December 3, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gmach:2009:RPM

[1988] Daniel Gmach, Jerry Rolia, Ludmila Cherkasova, and Alfons Kemper. Resource pool management: Reactive versus proactive or let’s be friends. Computer Networks (Amsterdam, Netherlands: 1999), 53(17):2905–2922, December 3, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wood:2009:SBB

[1989] Timothy Wood, Prashant Shenoy, Arun Venkataramani, and Mazin Yousif. Sandpiper: Black-box and gray-box resource management for virtual machines. Computer Networks (Amsterdam, Netherlands: 1999), 53 (17):2923–2938, December 3, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Kant:2009:DCE

[1990] Krishna Kant. Data center evolution: A tutorial on state of the art, issues, and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 53 (17):2939–2965, December 3, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2009:EBq

[1991] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 53(17):??, December 3, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Garcia-Reinoso:2009:ZCR

[1992] Jaime García-Reinoso, Iván Vidal, Francisco Valera, and Arturo Azcorra. Zero config residential gateway experiences for next generation smart homes. Computer Networks (Amsterdam, Netherlands: 1999), 53(18):2967– 2984, December 24, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fu:2009:IDR

[1993] Jing Fu, Peter Sjödin, and Gunnar Karlsson. Intra-domain routing convergence with centralized control. Computer Networks (Amsterdam, Netherlands: 1999), 53(18):2985–2996, December 24, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Moradi:2009:RBS

[1994] Shahab Moradi, A. Hamed Mohsenian-Rad, and Vincent W. S. Wong. An RLbased scheduling algorithm for video traffic in high-rate wireless personal area networks. Computer Networks (Amsterdam, Netherlands: 1999), 53 (18):2997–3010, December 24, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mur:2009:ASW

[1995] Daniel Camps Mur, Xavier Pérez-Costa, and Sebastia Sallent Ribes. An adaptive solution for Wireless LAN distributed power saving modes. Computer Networks (Amsterdam, Netherlands: 1999), 53(18):3011–3030, December 24, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2009:OFC

[1996] Jiming Chen, Shibo He, Youxian Sun, Preetha Thulasiraman, and Xuemin (Sherman) Shen. Optimal flow control for utility-lifetime tradeoff in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(18):3031–3041, December 24, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ssu:2009:DSA

[1997] Kuo-Feng Ssu, Wei-Tong Wang, and Wen-Chung Chang. Detecting Sybil attacks in Wireless Sensor Networks using neighboring information. Computer Networks (Amsterdam, Netherlands: 1999), 53(18):3042–3056, December 24, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cuomo:2009:PAI

[1998] Francesca Cuomo, Emanuele Cipollone, and Anna Abbagnale. Performance analysis of IEEE 802.15.4 wireless sensor networks: An insight into the topology formation process. Computer Networks (Amsterdam, Netherlands: 1999), 53(18):3057–3075, December 24, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Boushaba:2009:HAL

[1999] Mustapha Boushaba, Abdelhakim Hafid, and Abderrahim Benslimane. High accuracy localization method using AoA in sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 53(18):3076–3088, December 24, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jiang:2009:SAP

[2000] Yixin Jiang, Yanfei Fan, Xuemin (Sherman) Shen, and Chuang Lin. A selfadaptive probabilistic packet filtering scheme against entropy attacks in network coding. Computer Networks (Amsterdam, Netherlands: 1999), 53(18): 3089–3101, December 24, 2009. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Urdaneta:2010:CWW

[2001] Guido Urdaneta, Guillaume Pierre, and Maarten van Steen. Corrigendum to “Wikipedia workload analysis for decentralized hosting” [Computer Networks 53 (11) (2009) 1830–1845]. Computer Networks (Amsterdam, Netherlands: 1999), 54(5):877–878, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). See [1913].

Herman:2000:NIW

[2002] Ivan Herman and Albert Vezza, editors. Ninth International World Wide Web Conference (WWW9): The Web: The Next Generation: May 15– 19, 2000, Amsterdam, The Netherlands, volume 33(1–6) of Computer Networks (Amsterdam, Netherlands: 1999). Foretec Seminars, Inc., ????, 2000. ISBN 1-930792-00-X. LCCN ????

URL http://www9.org/;   http://
www9.org/w9cdrom/start.html.